A Bibliography of Publications about the GNU/Linux Operating System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2024 Version 2.153

Title word cross-reference

+ [Ste01e]. $100 [CS95]. $24.95 [Ano95a]. $39.95 [Sch04]. $44.99 [Sta07b]. $49.95 [Jen05]. $49.99 [Hid04, Tho05]. $59.99 [Koh06]. $99 [Kro00]. = [Ste01e]. × [Hun99].

* [TYKZ07]. */ [TYKZ07]. *BSD [Den99a]. *icomment [TYKZ07].

-dienste [WF03].

/*icomment [TYKZ07]. /GNOME [Wri00, Pen99].

0 [Hid04, Koh06, Sch04, Sta07b, Tho05]. 0-13-101415-3 [Hid04]. 0-13-144853-6 [Tho05]. 0-13-167984-8 [Sta07b]. 0-596-00482-6 [Sch04]. 0-7821-4428-4 [Koh06]. ’03 [ACM03b]. 046 [Sav11]. ’05 [ACM05b, MS05].

1 [FOP06, Jen05, She03]. 1-59327-036-4 [Jen05]. 1-GHz [Ano03b]. 1.0 [Coc01]. 1.2 [Gar98]. 1.7 [ZB19]. 1.7-GHz [ZB19]. 1.x [KGG00]. 10 [DWV06]. 10-Gigabit [cFJH^{+}03]. 10th [USE96a]. 12-step [Mil01]. 12th [MS05]. 1394 [Ale00, HKP09]. 14-16 [ACM06]. 18th [KD96]. 1999 [Den99b, Tim99]. 19th [ACM03b, SS05b]. 1Z0 [Sav11]. 1Z0-046 [Sav11].

2 [Ano94c, Com00, Com03, Gab07, MK04]. 2.0 [Böl01, Car98, McN99, PF97, Swe01]. 2.0.1 [ISO05]. 2.1 [BR95, CV00]. 2.2 [Ano00b, BB99b, Bra04]. 2.4 [Cal00]. 2.6 [BS05, PTS^{+}14, TCM07]. 2000 [Bru02, Kro00, MYH00, War01]. 2003 [ACM03a, ACM03b, Ano03d]. 2004 [ACM04]. 2005 [ACM05a]. 2006 [IEE06]. 20th [Ano94b, IEE06]. 22-nm [ZB19]. 23360 [ISO05]. 2e [Str94a].

3 [Gra05, Hid04]. 3.0.F [Ano96b]. 3.X [HH05a]. 32-Bit [Ano02d, Moh97b]. 37th [IEE04]. 39 [Kie99b]. 3rd [Ano99a, Ano99b].

4 [G^{+}06a, Jen05, Koh06]. 4.0 [Ano02c, Ano03f]. 488 [PPSM96]. 4th [Ano00a, Sch04, USE00].

5 [Ano02c]. 5.2 [Bon99, Dos99b, P^{+}08]. 5th [USE01].

6 [Dan99b, Ran99, Sch04, Tho05]. 6.0 [Dun99, Gra99]. 6.1 [HS99]. 64 [Gig06]. 64-Bit [Cha06, ZB19, Ano03b, GHL^{+}04, SGP^{+}23].

7 [HS00, Neg01, SCB^{+}13, Hag00]. 7.1 [Bar01c]. 7.2 [BW02].

8 [Ano03a, Neg02, Smi99, Sob03, Sta07b]. 8th [Pop94].

9 [Gan08]. ’94 [Ano94b, DW94]. 95 [Bur97]. 95/NT [Bur97]. ’96 [KD96].

= [RB08].

A-Z [Cor97]. A300 [YKS^{+}02]. Aaron [Sch04]. Abaci [SWF^{+}22]. Abaci-finder [SWF^{+}22]. abduction [CDOY09]. abstractions [KYB^{+}07]. Accelerating [MKO^{+}17]. Acceleration [IBM05]. Accelerator [OBSB23, PS22, PGW^{+}20]. access [KS09a, LPH^{+}07, Ram96]. Accidental [Ray99c, TD01, Ray01b]. accurate [KS09b]. achieving [WTKW08].

Acid [Lew99b]. ACM [ACM03b, ACM04, ACM05a, ACM05b, MS05]. ACM/IEEE [ACM05a]. ACM/USENIX [ACM05b]. Across [MAP^{+}22, Mon00b, GR09]. Action [NR03]. ActiveX [Kro99]. activity [MB08]. Acumen [Kro99]. Ada [SB99]. Ada95 [Gar09]. Ada95/C [Gar09]. adaptable [SAN08]. Adaptation [EJMT94, ZBG^{+}05]. Adapter [ZST^{+}07]. adaptive [MR07, SJP^{+}09]. addicts [Mil01]. Adding [AR07, Ano03a]. adds [Ano03d]. administracion [Bla97]. Administration [Bau03, Cam04, CDD00, Fri03, Gag01, G^{+}00, G^{+}02, GA04, Hun00a, Hun00c, Hun00b, Hun01b, KC95, KC98, NSH02, USE96a, Wei04, AT07, Aul01, Car99, Col02, GJ08, KC00, NSH06, NSHW10, Pon05, Ron05b, Sha00, Sha01, Smi01, SS01, SS02, Tay03, Tro96b, TB05, Wel99b, WWS98, Zad01]. Administrator [Kir95a, Kir95b, Kir95c, Luc00a, Mag04, BDP05, DRP01, Kir94, KD00, MM00, Par96, Par00]. Administrators [BDD^{+}06, MYH00, Hei99, K^{+}06, Kom96, Man01, MY03]. Admins [Jan03]. Adopt [Sta05b]. Adoption [Ksh04, CM06]. Adopts [Bar01b]. Advanced [CF20, Gar99, IEE06, MOS01, Ron05a, SS05b, Smi02, ISTE08]. Advantages [Cha06]. AES [OST05]. AflIot [DCH^{+}22]. against [WJS07]. agent [LLT07]. agent-based [LLT07]. Agents [Dat14, Jon02]. aggregate [VFMM08]. Agile [PGW^{+}20, USC^{+}08, AOAH15]. Agile-SD [AOAH15]. Aglets [Jon02]. AI [FES^{+}99]. AINA [SS05b]. AIX [Kuo99]. Ajaykumar [Wer10]. Alexandria [MS05]. Algebra [uRH95, uRH96]. algorithm [AOAH15, MAT08]. algorithms [TK07]. all-in-one [Dul14]. aller [DF00]. Almost [Pau95, Ste01c, Kau96]. Alpha [ST03]. Altera [Ano03b]. alternative [ALP^{+}21, SPF^{+}07]. alternatives [CKE^{+}07]. Amazon [Gib17]. Amber [Cha13]. Ambidextrous [O’S02]. AMD [Ano01b, SuS01, Ano03b]. AMD-PCs [SuS01]. Amorphous [GZK^{+}01]. Analysis [ARJ08, Ano02d, CF20, DRN^{+}23, GD18, MR18, She03, SHB23, ZB19, Ano03c, BJM^{+}22, CDOY09, CKE^{+}07, DDA08, FFPC22, HRS^{+}10, KMP^{+}20, LLA07, LF95, LWL06, MMW96, PFH06, RKSV20, RS14, RGGL^{+}09, SAF07, Wün13, dOdO16]. Analytic [UPS^{+}07]. analytical [UPS^{+}05]. Analyzer [Ano03e, Ano03c]. analyzes [Ano03b]. Analyzing [BKB^{+}18, FD10, SCB^{+}13]. Andrew [Teo13]. Android [KMP^{+}20, MKO^{+}17, SFE10]. Android-Powered [SFE10]. Announced [DDJ98b]. Announcement [Ano99a, Ano99b, Ano00a]. Announces [Ano03e, Bar00b, Ano03c, Ano03d]. annoyances [Jan06a]. Annual [Ano99a, Ano99b, Ano00a, Den99a, USE00, USE01, USE96b]. anomaly [DFC^{+}23]. Anpassen [Gün02]. answers [RC98]. Antarctica [Ana99]. Anwenderhandbuch [Ano01b, Ron05a, Ano01b, Ron01a, Ron01b]. Anwendung [Bor01, WF03, G^{+}00, G^{+}02, GA04]. anything [SAN08]. Aonix [Kro99]. AP [SMTW96, SW98]. AP/Linux [SMTW96, SW98]. Apache [Aul01, Gab07, G^{+}06b, Hun01b, LW03, Sai01, TSC02]. API [CDM97, Car98]. AppForce [Ano03c]. Apple [Ano02c, Ano03b]. appliance [S^{+}07]. appliances [S^{+}07]. Application [AJ05, Ano02c, Dun99, Gre19, JT98, KBPV99, Roe02, ZB19, Ano03c, Gab07, JPS00, JT05, Jon05, LHB24, LZ07, MMW96, Pen99, USR09, YW07, YSCO04, ZBG^{+}05]. application-aware [YW07]. Application-Class [ZB19]. application-layer [Ano03c]. application-level [LZ07]. application-specific [LHB24]. Applications [Abb03, AR22, Ano02b, Ano02c, Asb00, Bol99, Böl01, CS01, Kro00, LW00, Loz05, Mar04, Mot02, SS05b, UJE^{+}22, Ano95b, Cha11, Gon99, Gue19, GMTW08, Har99, Cor05a, MKO^{+}17, PBM^{+}22, RS14, ST03, SMXH01, Spe00b, SKP^{+}02, UPS^{+}05, UPS^{+}07, USC^{+}08, Yeo05]. Approach [Mud97, Sta07b, ALP^{+}21, AKL^{+}96, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, GMV07, Hal06, Hal07, OMCB07, RFS05, RFS06, SMS08]. apps [KMP^{+}20]. arbeiten [Ste00a]. architectural [PL06]. Architecture [LBN14, PS22, CLDA07, DKK07, IZG^{+}07, Mau08, Vre03, Wer10, Wil98]. architectures [PBH^{+}07, RFS05, RFS06, TF04]. area [TK07]. ARM [Jor04, GNDB16, Jor04]. ARM-based [Jor04]. ARM-basierender [Jor04]. ARMLinux [WW01a, WW01b]. arrival [GMV07]. Art [Dec04, BDF^{+}03, WB06]. Asbestos [Lew99a]. Asian [CM06]. asked [Pub11]. aspects [MS07]. Assembly [Dun00, Cas17]. Assessment [BP96]. Asset [Kro00]. Associates [Ano95a]. assurance [FOP06]. ASSURE [SLP^{+}09]. Assuring [Ano14]. astronomical [Mil96]. Astronomy [Ano95g, Ano95f]. asymmetric [GCN^{+}10]. asymmetry [KS09b]. asymmetry-aware [KS09b]. asynchronous [BJL^{+}06]. Atlanta [Ano99a, Den99b, USE00, Ano00a]. ATM [Alm96, Col97, JWKZ99]. Atomic [Ano03e, BLH20]. Attack [Ahm08, Ras07]. Attacks [Han00, LCHL21, OST05, Chi02, HMK^{+}20]. attempt [Fel13]. audio [Ran06]. ausgereifte [Bra04]. Australia [Ano94f]. authentication [Gan08]. authorization [GJJ05]. authors [Hou96]. AutoBash [SAF07]. Automated [CDD^{+}07]. Automatic [GJJ05, FM02, MSZ09, SLP^{+}09]. Automatically [CGP^{+}08, LPH^{+}07]. Automating [Bau03, PLHM08]. automation [Lum13, Pop94]. automatisieren [RB06a]. Automounter [Zad01]. AUUG94 [Ano94f]. aux [RB05b]. Availability [Bra00, SHB23]. available [Daw94, Kop05, Tim99]. Avilon [Kro99]. Awards [Eri95]. Aware [XZY^{+}10, KS09b, MB08, TAS07, WW02, XSaJ08, YW07, ZYCZ10]. Away [Den99a, Sta07a]. awesome [Gri14]. Awk [Bla02, Bla04]. Axis [Dou02].

B [RBM13]. B-Tree [RBM13]. B.S. [Ran97, Ran99, Ran00]. Baan [Kro99]. Back [Ano02c, Lea12]. Back-Office [Ano02c]. backdoor [Fel13]. Backup [Pre07]. Bad [BK13, TYKZ07]. Balancing [Kro99, Mon00a]. bandwidth [BSR06]. BarbequeRTRM [BMF15]. bare [Daw94]. bare-bones [Daw94]. Base [Coc01, ISO05, Dup05, RS14, Cor05a, Yeo05]. Based [Ano02d, Chu03, CPJ^{+}98, GVK^{+}03, GZK^{+}01, HLT^{+}21, HHD^{+}16, PNK02, TSEK01, ALC20, ABB19, AKL^{+}96, AOAH15, Ano95b, Ano03c, Ano03d, Ano06, ALL06, BDH^{+}09, CMMST20, CM06, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, DCH^{+}22, FMR07, Gra06b, GFRD96, GMV07, HMK^{+}20, HNH03, Jor04, KEP07, Kro00, Lac06, LLT07, MKSB22, McC07, MFPP03, PBM^{+}22, PV02, Pon05, SPF^{+}07, Sto07, WG08, WN15, Wil98]. Bash [Bur04, May01, Bla02, Bla04]. basics [Hsi01, Mus02]. basierenden [Lac06]. basierender [Jor04]. basierte [Ric14]. Basis [GA04]. BASS [PL06]. Bastille [Bea04]. Batch [LD05]. Batteries [Coc01]. Battery [Sta04]. Battery-powered [Sta04]. Bazaar [Ray99c, Ray01b]. BDP [AOAH14]. Be [FES^{+}99, Pub11, Tim99, Gia99]. BeagleBone [BK13, Dat14, Gri13, Gri14, Lum13, Mol14, Ric13, Ric14, SP15, YK15]. bearbeiten [GGZ02]. Bedienung [KGG00]. Befehle [Gün02]. before [SBRM09]. Beginner [RGRG05, Pet01a, Sha00, Sha01].

Beginning [MSB^{+}96, MS99, Ron05a, Tho06, Wri00, Ste99]. Behave [Swa00a]. Behind [Swa98]. Being [ASW^{+}01, Ano95e]. Ben [RB08]. benchmark [PL06]. benefit [HRW09]. Benefits [CPJ^{+}98, AMS^{+}08, AMSS08]. Beowulf [Hul05, DDJ98a, GLS03, Spe00a, Ste01a, Ste02]. Best [Wal99]. Bet [Bar00c]. Beta [Geo94]. Beta-Testing [Geo94]. Betriebssystem [CK06a, CK06b, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i, CK06j, CK06k, SuS01]. Better [Coc01, Ded09, Com00, gL02]. BetterBASIC [ASW^{+}01]. Between [Woe03, NAGL10, SMM^{+}09]. beyond [Mus02]. bi [CDOY09]. bi-abduction [CDOY09]. Bible [Goe00, Neg01, PR96a, Blu08, Blu11, DWV06, Gri01a, HZ99, HH05a, Hun01a, Neg02, Neg06a, Neg07a, PWN01, WY94, W^{+}95, vH06, vH07]. biblia [Pit98a]. Big [Sta04, Gar99]. Bildbearbeitung [DF00]. Bilder [DF00]. binaries [MM06, Sun99]. Binary [Hub02, Lew99a, Lew99b, DCH^{+}22]. binary-level [DCH^{+}22]. biodata [Wün13]. biology [Wün13]. Biometric [Coc03a]. Biopolis [Ano06]. Bit [Ano02d, Cha06, ZB19, Ano03b, GHL^{+}04, Moh97b, SGP^{+}23, SKP96]. Black [BK13, Sta04, HRD^{+}07, Gri13, Gri14]. black-box [HRD^{+}07]. BlackParrot [PGW^{+}20]. Blind [But02b, Eri97]. blocking [FD10]. Blue [AUW08, Gar99]. Bolton [ACM03b]. Bolts [Ano99f]. Bone [BK13]. bones [Daw94]. Book [Ano95a, Cha13, Elb01, Hid04, Kad95c, Kad95b, Men12, Sch04, Teo13, Wer10, CDM98, HBB^{+}06, H^{+}07a, H^{+}07b, SAN08, Cha13]. Books [Sch04]. Bookshelf [GS00, Gri01b, Ros98, SWL^{+}01, Tay99b, Woe02, Woe03, FWS^{+}03, WLS^{+}01]. Boot [HHD^{+}16, Neg06a, Neg07a]. Bootable [SGP^{+}23]. bootables [Neg06c, Neg07b]. Bootstrap [Reg11]. Borders [Coc03a].

Borland [Lis02]. both [DJC07]. bottlenecks [SBRM09]. Boulevard [ACM99]. boundary [GFRD96]. boundary-scan [GFRD96]. Bourne [May01, May01]. Bourne-Shell [May01]. Boutell [Koh06]. box [HRD^{+}07, LHB24]. BPF [CF20, Gre19, WC22]. Brain [Bar01b]. Branding [Ano95l]. Bridge [Ano02d]. Bridging [ACM04]. Briefs [Gar99, Gla99]. bring [VMKB05]. Bringing [Vad00]. Broken [Ahm08]. brought [Lea12]. browsing [FFPC22]. BSD [DF00, GGZ02, Kuo99, Lin02a, May01, WF03, ASW^{+}01, Cor05b, Far94, GGZ02, gL02, Lin02a, May01, McN99]. BSD/ [Far94]. BT8233 [Col97]. BTRFS [RBM13]. Buch [KGG00]. Buffer [LBN14, DJC07]. Bug [CPJ^{+}98, Kie99a, LZAH22]. Bug-Free [CPJ^{+}98]. Buggy [CPJ^{+}98]. Bugs [AHSM21, BLH20, CCVLW23, LPH^{+}07, TYKZ07]. Build [Asb00, Gib17, Kro00, Pau95, Bur97, Gue19, Kop05, NH14]. Builder [Coc03a]. Building [AB04, Bau02, Don96a, Fel96a, HE95, ILG19, Cor05a, Luc05, SY00, Spe00b, Tho05, WSL^{+}99, Yag03, Yeo05, YW07, Yor99, ZCC00, AUW08, Boo03, Gon99, Gri14, Jan03, KAR^{+}06, Mol14, Neg06c, Neg07b, S^{+}07]. Bulgaria [Pop94]. bus [Ano03b]. Business [Ano02c, Lat02, Lei99, SH99a, Gon99]. Busters [Ano94c].

C [Ano02b, DF00, Swa00b, Ano03d, But02a, Gar09, LH00, Lla06, Pin02, Ste00b, Ste01c, Ste01d, Ste01b, Ste01e, Ste01f, Ste01g]. C/C [Ano03d]. CA [Ano95a, USE01]. Cache [LBN14, OST05, DJC07, GB07, IZG^{+}07]. cache/memory [IZG^{+}07]. Calculated [Coc01]. Calculations [TSEK01]. Caldera [Ano95b, Ano95c, SRBB99]. California [ACM93, USE96b]. call [CKR08, XCS^{+}22]. calls [BKB^{+}18, Lov07]. Camera [Dou02]. Can [Bar00c, Ste01d, Ano02d]. Canada [ACM06]. capabilities [BP96]. Capability [JKS^{+}15]. Capable [OBSB23, VTS^{+}23]. Capacity [Coc01, FSW02]. Capo [MHKT09]. care [HBB13]. cas [Cor00, GVK^{+}03]. CAS-System [GVK^{+}03]. Case [cFJH^{+}03, Kie99b, LS01, Lei99, MAP^{+}22, OST05, BTS09, FFPC22, MW08, NH14, WW02]. Catching [SJP^{+}09]. Categorization [YSCO04]. Cathedral [Ray99c, Ray01b]. causality [SAF07]. CCS [MS05]. CD [Ano01c, Ano01b, DF00, SuS01, Cha99, FWS^{+}03, GGZ02, HS99, Kom02, OR96, Ser00, SWL^{+}01, WLS^{+}01]. CD-ROM [Kom02, OR96]. CD-ROMs [Ano01c, HS99]. CDROM [HS99, HS00]. CDs [Gün02, GGZ02, Neg06c, Neg07b, SC98]. CE [Ano03d]. CE-based [Ano03d]. Center [ACM99, UJE^{+}22]. centralizada [VD01]. centralized [VD01]. Centre [Ano94f]. cert [Tom11]. Certification [Mag04, Yor99, Dea01, Eck11, Mag00, Mag01a, Mag01b, Mag01c, MS02, P^{+}06, Sai01, Sul02]. Certified [Mag04, Pri99, Sav11, DRP01, Jan04, RC98]. Certifying [SK04]. CFD [GVK^{+}03]. Cg [Ano03e]. Chains [BB99b]. Challenge [Sta06]. challenged [Kro00]. Challenges [Cha06]. Changing [Edw98]. channels [CMR^{+}21]. characterization [AJ05]. Characterizing [CMMST20]. Cheap [Far94]. Cheat [Rul07]. Checker [Kro99]. checking [BDH^{+}09, KS19, Kla93]. checkpoint [BBHL08]. checkpoint/restart [BBHL08]. Cheshire [OBSB23]. Chicago [ACM05b, USE96a]. Children [AMM^{+}18]. Chip [Bar01b, Don04, ISTE08]. Chips [Coc01]. choice [TF96]. Christopher [Sta07b]. Chrome [FFPC22, RR15a]. chuan [TGW97]. Citius [Ano00b]. Class [WQR13, ZB19]. Classes [Han00]. Classic [RB05a]. classification [AWW08, SWF^{+}22]. cleaner [Gat01]. Cleaning [Gat01]. clearly [Pfa99, Pfa01]. ClearSight [Ano03c]. Client [CS01, Lac06, ON06]. Client-Server [CS01]. clone [Ano95b, Bur97, Daw94]. cloning [LLA07]. Cloud [MKSB22, ALC20, BBB^{+}20, SCB^{+}13]. Cluster [DR01, GVK^{+}03, GSN^{+}01, GSHL03, Hul05, Jen05, KBPV99, KEB^{+}99, KKH03, Kop00, LS01, O’K00b, SGHL01, Ste01a, Wol04b, GLS03, Kop05, Ste02, Vre03, EEK^{+}04, Kop05]. Clustered [Tho05, Luc05, MAT08]. Clustering [DDJ98a, Lum01, RMZ01, Boo03, TAS07]. Clusters [DDJ98b, cFJH^{+}03, Meh00, MFPP03, PKB01, Spe00a, TSEK01, Boo03, CKE^{+}07, LWL06, PS07, PKB03, Slo05, Spe00b]. CLX [Lis02]. CMG [Ano94b]. CMP [TAS07, IZG^{+}07]. co [Yu07]. co-evolution [Yu07]. CoCreate [Ano03d]. Codagen [Ano03e]. Code [Bon01, Cou17, Hub02, CP01, CP02, FRRJ07, Kau96, Moo01b, Moo01a, PC01, SLQP07]. CodeGuide [Ano02d]. CodeWarrior [Ano02d]. coding [Cas17]. CoLiS [BJM^{+}22]. collaboration [Ano03d]. Collaborative [IEE06]. collateral [PLM06, PLHM08]. collectively [Hou96]. collector [JHT^{+}07]. com [Bar00c]. Combination [SW98]. Combining [DRL^{+}02, LZ07]. Coming [Wol04b]. command [Blu08, Blu11, Pfa00, Pur97, Sho12, Ste99]. commands [Gol95, VR00]. commentary [Max01]. comments [TYKZ07]. Commercial [BR95, Ano95b, Daw94, Lev95]. Commits [JJW^{+}24, WG18]. commodity [CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CLDA07, Kop05, MR07, SLQP07]. common [MS07, YSCO04]. Communicating [NAGL10]. Communication [CK06b, SGL^{+}00, ZST^{+}07, IEE93, Lin02a, YY96]. Communications [MS05, BP96]. communities [ACM04]. Community [WSL^{+}99]. como [RÓ01]. compaction [CDD^{+}05]. Companies [Bar00c]. Companion [OR96, Hei99, Kom96]. Company [ASW^{+}01]. Comparative [AOAH14, PSK08, SHB23, SKP^{+}02]. Compare [Sta04]. Comparing [PBH^{+}07]. comparison [LB96, PS07, SK10, TCM07, TF04]. Compatible [Ano02c, SGP^{+}23, AAB^{+}05, MM06]. Compétition [Cor00]. Competitive [LSP07]. Compiler [CS06, HKS^{+}04, Hua06, Ste01c, CBC^{+}08a, CBC^{+}08b, CBC^{+}08c, GHL^{+}04]. Compilers [Ano02b]. Compiling [Hua06]. complet [Tau96]. Complete [HS99, Kom00, Nor99, Pet98b, Pet00b, QH01, SU95a, Ano97a, Bru02, Chi02, DDA08, Gay99, Kie98, KAR^{+}06, Pet96, Pet99, Pet01b, Pet02a, Pet02b, Pur97, Ric98, Sho12, SU94b, SU96a, SU97, Tau95, Tay99a]. complex [Gri13, WYW^{+}13]. Component [Kro99, Lis02, Yu07]. Components [TEGF98, Ano03c, HRD^{+}07]. Compositional [CDOY09]. comprehensive [CKE^{+}07, Gar00, MSW06]. compression [BGM99]. Compromise [Ahm08]. Computational [Gar99, Wün13, AJD^{+}10]. Computer [Ano94a, CCO^{+}05, DW08, uRH95, uRH96, JB95, MS05, Alm96, AC11, Ano94d, AUW08, FOP06, Lla06, Shu06, CPJ^{+}98]. Computeralgebra [uRH95, uRH96]. Computers [DDJ98b, Kop00, Man99, Nie00, SNF04, Ano97b, Red98, Tro96b]. Computing [ACM04, ASW^{+}01, Bar00b, Cer03, GZK^{+}01, Hom00, Hul05, IEE06, ILG19, KIKT01, Kro00, Lum01, SSSS97, Ste01a, Thr01, Tom02, AKL^{+}96, Ano94b, AUW08, DW94, GLS03, SCB^{+}13, Ste02, Wal01b, Wol02, Zem94]. Compuware [Ano03e]. Concept [BB99b]. Conception [Dup05]. concepts [Kra05, Sul02]. Concludes [Coc03b]. Concurrency [AMM^{+}18, LZ07, LPH^{+}07]. concurrent [Gue19, LSP07]. Conference [ACM93, ACM04, ACM05b, ACM06, Ano94a, Ano94f, Ano00a, Ano02a, Ano05a, Ano06, DDJ98a, KD96, MS05, SS05b, USE96a, USE00, USE01, ACM05a, Ano94b, Ano94d, IEE93, Pop94, USE96b, Den99a]. configurable [BWH^{+}07]. configuraciones [VD01]. Configuration [Ale00, EDK00, Kof97, VRJ97, Jan01, Kof99, Kof00, Mag00, Mag01a, SU94b, SUM95, SUM96, SMM97, SAF07, VRJ95, VRJ96, VRFJ98]. Configurations [CGDB24, VD01]. Configuring [Rul07]. congestion [AOAH15]. Congress [Ano94f]. connection [SMS08]. connection-oriented [SMS08]. Cons [Sie00]. Consistency [RKSV20]. Constraints [GSHL03]. consumption [CKE^{+}07]. contained [Han97]. Container [SHB23, SPF^{+}07]. Container-based [SPF^{+}07]. containers [ALC20, BBK^{+}16, GBC^{+}18]. containers-based [ALC20]. Contention [XSaJ08]. Contention-aware [XSaJ08]. Contest [Coc03b, DDJ98a]. Contests [DDJ98b]. context [BLH20, LLA07, PFH06, WW02]. context-aware [WW02]. context-sensitive [LLA07, PFH06]. Continuously [GAH16]. contributors [HNH03]. Control [BMF15, CGDB24, Dou02, GGZ02, Tim03, AMSS08, AOAH15, Gar09, GAH16, Kar16, KYB^{+}07, RPB^{+}09, ZBG^{+}05, AMS^{+}08]. controlled [Gib17, JB95, Sto07]. Controller [ZST^{+}07]. Controlling [Kro00]. Convention [ACM99]. Conversation [Flo94]. Converse [KBPV99]. converting [MM06]. Cookbook [Kal16, Loz05, Stu01, YK15, BSB03, G^{+}06a, P^{+}08, Sch05, Sch07, Stu04]. cool [NW03, Neg06b]. Cooperation [Daw94]. Cooperative [Wie93]. Coprocessor [LS01, AV04]. Copy [KLK^{+}02, GMTW08, KS19]. copying [YY96]. Copyleft [O’S02]. Copyright [O’S02, Zic01]. Core [ISO05, Max01, SGP^{+}23, XZY^{+}10, ZB19, CAW08, FD10, Kas14, Sul02, Sob06]. Corel [Dan01, Smi99]. cores [SMM^{+}09]. Corporate [Kau96, Moh97a]. Corrections [AMS^{+}08]. correlation [PFH06]. correlations [LPH^{+}07]. correntropy [HMK^{+}20]. Coscheduling [SGHL01]. Cost [Rob95b, Sta05a, ZB19, Ano03d, CC97, GFRD96, ZS93]. Could [Wel99c]. Count [Nis01b, WJS07]. counter [SZD^{+}08a, SZD^{+}08b, SZD^{+}08c]. Countermeasures [OST05]. Counters [DRN^{+}23]. Countries [Ksh04]. coupling [YSCO04]. Course [Kom00, Kom02, QH01, BW02]. courses [MB95]. covering [HHH11, HHHT10]. covert [CMR^{+}21]. CPU [Kar16]. crack [Lea12]. Crackberries [Sta05a]. crafting [BK13]. Cram [Pri99, Sav11]. crash [BW02, SWF^{+}22]. Crayons [CCVLW23]. Create [Pau95, Gri13, Gri14]. Creating [Böl01, Wyl95, ZS93, LHB24]. creation [CK06b, GGZ02]. crew [Lea12]. Critic [Lew99a, Lew99b]. critical [Gar09]. Croatia [KD96]. cross [Bro03]. cross-platform [Bro03]. Crypt [HHAW19]. Crypto [Ahm08]. cryptographic [AV04]. Cryptography [DDJ98a]. CUBIC [HRX08]. CUPS [GGZ02, GGZ02]. curriculum [FOP06]. curve [GMV07]. customizations [Kra07]. customizing [Neg06c, Neg07b]. CVS [Kie99b]. Cyber [Wal99]. Cyberspace [Wal99]. Cyrix [Ano01b].

Dallas [Ano94d]. Damn [SAN08]. Dance [Wol04a]. Dangerous [DDJ98b]. Data [Cas14, GVK^{+}03, PNK02, UJE^{+}22, BGM99, BC08, JYX^{+}16, TK07, YY96, Lin02a]. Database [Ano02c, Bar01b, BF97, JPS00, PWN01]. dataset [DFC^{+}23]. date [Jan06b]. Dateien [DF00]. Datenverschlüsselung [Lin02a]. dating [KMP^{+}20]. d’autore [Zic01]. Dawn [Lea12]. day [HMK^{+}20]. days [LH00]. DBA [Emm11]. de-facto [Rus08]. Deadline [LSAF16]. deadly [Tox07]. Dealing [GSHL03]. Death [CGP^{+}08]. deb [VD01]. Debain [Dum05]. Debian [Ano00b, Ano01b, CK06a, CK06b, CK06h, CK06k, DF00, Her04, Ahm08, Ano01a, Ano01c, Ano01b, Ano05b, BJM^{+}22, Bel00a, BGO02, Bra04, CV00, CK06a, CK06b, Den99a, Epp00a, G^{+}00, G^{+}02, GA04, Gan04, GO99, HH05a, Hun01a, Kle00, Kra05, Mac02, Mac99, McC99a, Neg06a, Neg07a, Ron01a, Ron01b, Ron05b, Ron05a, SS05a, VD01, WN15]. Debian-GNU-Linux-Powerpack [Gan04]. debug [FD10, Reg11]. Debugger [MZG14]. Debugging [MR18, Zac01, Ano03d, Bes05]. decade [BKB^{+}18]. December [Ano94b, IEE04]. decentralized [RPB^{+}09]. Decoder [Ale00, MYH00]. Decomposing [JKS^{+}15]. découverte [Fab97]. DeCSS [Coc03b]. Dedicated [DRL^{+}02, GSHL03, SGHL01]. Deep [HLT^{+}21]. Defense [WJS07]. definitive [Emm11]. deformation [HR95]. Degunking [SD05a, SD05b]. del [Bla97]. deleted [BC08]. d’emploi [TW96]. demystified [Sut02]. denied [Chi02]. Denmark [DW94]. dependencies [FMK^{+}07]. Deploying [Hal05, MKSB22, PKB01, LHB24, PKB03]. Deployment [BBD^{+}00]. derivation [MSZ09]. Design [Ano02d, BGM99, EEK^{+}04, EM00, GD18, SGP^{+}23, Var94, WSL^{+}99, Wan02, WC22, ARJ08, AWW08, Ano03b, Ano03c, BBS96, Gab07, Gia99, Reg11, S^{+}07, Wer10, WXX08]. designed [PPSM96]. desk [Haw00, Haw02]. Desktop [AH04, Ano95d, Böl01, Bra04, Dec04, KGG00, Ano95c, Bri05a, Dea01, Hal05, Hek97, PB05, SH99b, S^{+}00, S^{+}05]. Desktop- [Bra04]. detail [XCS^{+}22]. detect [JYX^{+}16]. detecting [CMR^{+}21, HMK^{+}20, LPH^{+}07, WG08]. Detection [BLH20, DFC^{+}23, PFH06, Ras07, WC22]. determination [WB06]. determinism [LWV^{+}10]. deterministic [MHKT09]. develop [Reg11]. developed [Daw94]. developer [Inf95, Moh98, Ano02c]. Developers [Ano02a, Ano05a, Far01, WW01a, HNH03]. Developing [BR99, Coc03a, Har99, Ksh04, CS06, PBM^{+}22]. Development [Ano03e, Bro03, Cam04, CPJ^{+}98, Coo09a, Coo09b, DH02, Dun99, JT98, Kro99, Kro00, NR03, RP02, SNM22, Ste01f, SuS01, Swe01, Wie93, Ben96, BV06, CM06, Gar09, GHL^{+}04, G^{+}06b, JT05, LHB24, LW03, Lov05, Lov10, MMW96, Pen99, VMKB05, Kro99, Lac06]. develops [Ano95b]. Device [Ano02d, Coo09c, Coo09d, Don97b, Rub98, RC01, CRKH05, DCH^{+}22, GD96, Kha02, MM06, PLM06, PLHM08, TF04, Ven08]. Devices [Kro00, SFE10, Ano03d, GR09, KS09a, MKO^{+}17, PPSM96, Rus08]. devil [XCS^{+}22]. d’experts [Par99]. d’exploitation [Cor00]. DGPS [HR95]. Dialogue [Lew99a]. Dickens [FES^{+}99]. Didn’t [Wel99c]. Diego [ACM93, USE96b]. dienste [WF03]. Different [FES^{+}99]. Diffie [ASW^{+}01]. Digital [AC11, FFPC22, Kom02, Kro00, GFRD96, Ano95e]. digitale [Zic01]. DigiTerra [Ano02c]. d’installation [LR97]. Direct [KEB^{+}99, KS09a, LP10, LWL06, Ram96]. direct-access [KS09a]. directory [Gol95]. diritto [Zic01]. DIS [ISO05]. Disaster [Bar00c]. Discharged [Ros10]. Disconcerting [AMM^{+}18]. Discover [Oua97a]. Discrete [TT01]. Disk [GW94, Vad00]. display [DM19]. disruptive [ALP^{+}21]. distance [AOAH15, WN15]. Distributed [Asb00, Gag00a, GZK^{+}01, LLT07, Mes00, Tan95, Tom02, BP96, Bro03, GCN^{+}10, GAH16, LBH09]. Distribution [MSSvK08, BFS07, DL08, DLP10, PF96, PF97]. distributions [Ano96b, LZAH22, LZAH23, Neg06a, Neg07a, Wie99b]. DivX [Zic01]. DMCA [Coc03b]. DNS [Hun00a]. Do [Jen02, Wel99c]. Dobb [Eri95, Eri97]. Documentation [CPJ^{+}98, Tay99a]. Documenting [PLHM08]. Does [Wol04a]. Dog [Coc01, WBS01]. dogs [Wis03]. Doku [Ano01c]. Domain [OBSB23, LF95]. Domain-Specific [OBSB23]. Domains [JKS^{+}15, LAK09]. Domino [RMZ01]. done [Gra06b]. DongTing [DFC^{+}23]. données [Dup05]. Doors [Kau96]. Dorothea [Hid04]. DOS-Programme [Str94a]. DOS/ [EJMT94]. Dot [Bar00c]. Dot-com [Bar00c]. double [RR95]. double-edged [RR95]. down [RFS05, RFS06]. Downs [Chu03]. Dr. [Eri95, Eri97, LL94a]. Draft [ISO05]. DRAM [LZZ^{+}08]. DRDDR [JYX^{+}16]. dream [Lum13]. DREAMS [HR95]. Dress [Ahm08]. driven [FRRJ07, LP10, RS14, SZD^{+}08a, SZD^{+}08b, SZD^{+}08c]. Driver [Geo94, Col97, GD96, TF04]. Drivers [Coo09c, Coo09d, Don97b, Rub98, RC01, Ste01e, CRKH05, Kha02, MM06, MSZ09, PLM06, PLHM08, Ven08]. Drives [DDJ98a]. driving [Bri05b]. Drucken [Gün02]. Drucksysteme [GGZ02]. DRX. [Wt95a]. DST [Sta07a]. DTM [PS07]. dual [Gar09]. Duly [CKM^{+}99]. Dummies [HS99, HS00, Mea99, Bel00a, Dul14, mH99b, Hug97, Hug98, MWM00, SP15, WWH97]. d’une [Dup05]. Duplicate [Nie00]. Durchstarter [Han99]. DVD [SuS01]. Dynamic [ALC20, MR07, Oom13, PKFH02, MS07, PBM^{+}22, USC^{+}08]. Dynamically [MZG14, Don04, OMCB07].

Early [ASW^{+}01]. Easily [PKB01, PKB03]. easy [Gra05, Mus02]. Echo [Gib17]. Economics [Ksh04]. economists [YL08]. ed [Sch04]. eDesktop [Cal00]. Edge [Don99, Sta05a, Tor99]. edged [RR95].

Edgewatch [San98]. Editieren [Str94a]. Edition [Hul05, KGG00, SRBB99, TGB95a, BN01, Cha13, TG96, Teo13, Bar99a]. Editor [DM97, Ste01c]. Editorial [FOP06]. Editors [Ste01e, BB99a]. Education [DW08]. Effective [BLH20, BMF15, Sto07, ZS93]. effektiven [Kuo99]. efficiency [BJL^{+}06, SAL19]. Efficient [LBH09, WR03, WJY^{+}07, BCP08, Kar16, LWV^{+}10, NLK09]. Effort [CIC13]. Eighteen [MFBWW20]. Eilert [Hid04]. Einführung [CK06a, CK06b, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i, CK06j, CK06k]. einrichten [Ste00a]. Einsatz [Wei04, SU94c, SU95b, SE96]. Einsteiger [Ric14, Ron05a]. Einsteigerseminar [RGRG05]. Eisenhaendler [Hid04]. Electromagnetic [TSEK01]. electronic [BK13, HH04]. Elektronik [Ric14]. Elektronik-Projekte [Ric14]. Elektronische [Ano01a, Ano01c, Ano01b]. element [LWL06]. Elimination [YY96]. Ellen [Sch04]. Emacs [Str94a]. Emacspeak [Ram96]. Email [Gag00b, ZS93]. Embedded [Abb03, DDJ98a, DH02, Hal07, Hol02, Jor04, Kro99, Lom01, Nis01a, Nis01b, Nis01c, Nis03, Sut02, Weh01, Woe02, Yag03, ARJ08, Ano03c, BCP08, FRRJ07, Kha02, Mol14, PLKH08, Reg11, SJP^{+}09, SKP^{+}02, Ano02d, Coc03b, Coc03a, DH02, Fis02, Hal06, Jor04, Nis01c, Sta07b]. embedding [Ano96a, DOF96]. Empirical [MSSvK08, GAH16]. Empirix [Ano03e]. Emulation [LBN^{+}18, GD96]. Emulator [MZG14]. emWare [Ano02d]. Enabling [GBC^{+}18]. Encoding [CPJ^{+}98, Mud97, Lin02a]. Encryption [PS22, HHAW19]. encyclopedia [PR96b, PR97, Wor98]. End [SBRM09, Alm96]. End-to-end [SBRM09]. Energy [FRRJ07, Kar16, ZB19, CKE^{+}07, SAL19]. Energy-efficient [Kar16].

Energy-optimizing [FRRJ07]. enforcement [RKSV20]. Enforcing [Oom13]. Engineer [Pri99, Jan04]. Engineering [Ano95d, BFN^{+}06, EJMT94, Pop94]. English [SU96d, VRFJ97a]. Enhanced [MMC06, McC05, GHRS05]. enhancements [Don22]. Enhances [Ano03e, Ano03c]. enjeux [Cor00]. EnlightenDSM [Gag00a]. enough [PBR^{+}08]. Ensuring [RS14]. entering [Moh97a]. Enterprise [G^{+}06a, Jen05, Kop05, McC04, Sob06, Ano94b, Men12, Spe00b, Asb00]. entertainment [NW03, Neg06b]. entropy [SNM22]. Entwicklung [Lac06]. Environment [Ano02d, Ano03e, Don98, GZK^{+}01, Hub02, IEE06, KKH03, Koh06, Kro99, SGL^{+}00, Ste00a, CLDA07, Gar09, GHL^{+}04, MB05, PPSM96, TK07, Van06, VFMM08]. Environments [ACM05b, ACM06, SSSS97, CMR^{+}21, GBC^{+}18]. ERP [Bar00c]. Error [RGGL^{+}09, HRD^{+}07]. erstellen [Gün02, GGZ02]. eServer [SvBC^{+}04, AV04]. Essay [Ano99d]. Essential [Fri03, Hea97, Ven08, Kom96, Pet00c, Wal01b]. Essentials [LCJH05, Smi12]. Estimation [Kro00, CIC13]. Estudio [RÓ01]. ESX [Rul07]. etc [AC11]. Ethernet [cFJH^{+}03]. ethics [Mag01c]. EUVL [Bar01b]. Evaluating [BJL^{+}06, NTW07, SSSS97, EP05, PL06]. Evaluation [RLB08, TD96, Tem96, Ano94b, BFN^{+}06, EEK^{+}04, LZ07, PSK08, SK10, SKP^{+}02]. even [Moh97b]. Event [TT01]. events [LZ07]. every [Lov07, War04, War15]. Everyday [MM01, Stu01, Gra05, Stu04]. Everyone [Coc01]. Everything [Woe03]. Everywhere [Ana99]. Evolution [LADADL14, Kor06, PLM06, WYW^{+}13, Yu07]. evolutions [PLHM08]. evoluzione [Zic01]. Exam [Pri99, DRP01, Jan04, Sav11]. Examination [Mag04]. Examining [NL00].

Example [Gay00, Qui02, Rob04, Wal00, MMC06, Qui00]. exams [Tom11]. Excellence [Eri95]. Exception [BJ01]. exciting [Gri13]. EXE [CGP^{+}08]. Executable [Coc03b]. Execution [ACM05b, ACM06, BBB^{+}20, CLDA07, XSaJ08]. exercises [Coo09a, Coo09b, Coo09c, Coo09d]. expect [Daw94]. Experience [BFS07, SK04, ETKF07]. Experiences [AAB^{+}05, NV06]. Experimental [BSSS96, JWKZ99, AJD^{+}10]. Experiments [GZK^{+}01, WN15]. Expert [Sav11]. explained [Pfa99, Pfa01]. exploitation [Ron05b]. exploiting [DJC07]. Explorer‘s [Dum05]. Exploring [AH04, Mol14]. Expo [Ano06]. Expose [O’S02]. exposed [HLK01]. exposition [She02]. ext3 [AR07]. Extend [Ano03a]. Extended [OST05, WC22]. Extending [BF08, Hus94, TEGF98]. Extends [Ano03e]. extensible [BWH^{+}07]. extension [CBC^{+}08a, CBC^{+}08b, CBC^{+}08c]. extension-oriented [CBC^{+}08a, CBC^{+}08b, CBC^{+}08c]. Extensions [WQR13]. External [Rob95a, LWV^{+}10]. extra [Ano03a].

Face [Edw98, IEE93]. facile [Ano96b]. Facility [Thr01]. Fact [KEB^{+}99]. facto [Rus08]. factor [ALC20]. Factory [Kro99]. Fail [Coc03b]. Failover [Roe02]. Failures [MR18]. fair [CAW08, LBH09]. fair-share [CAW08]. fairness [WTKW08]. Farm [Bar00c]. Farms [Thr01]. Fast [DR01, SRTH15, SMM^{+}09, MSZ09, dBB08]. FastTrack [SC98]. Fault [FES^{+}99, LW00]. Fault-Tolerant [LW00]. Faults [PTS^{+}11, PTS^{+}14]. FDSOI [ZB19]. Feature [PQM^{+}21]. features [Daw94]. featuring [SC98]. Fedora [HH05b, Men12, McC04, Neg06a, Neg07a, Sob06, Tyl06]. feedback [ZBG^{+}05]. feedback-control [ZBG^{+}05]. Feel [Kro00]. FENRIS [Tim99].

FGMC [HMK^{+}20]. FGMC-HADS [HMK^{+}20]. Fiberoptic [Coc01]. Fiction [KEB^{+}99]. FIDES [ISTE08]. Field [TSEK01]. Figgins [Sch04]. Fighting [DDJ98b]. File [Bra00, GR09, LADADL14, MLE^{+}00, O’K00b, SRTH15, TECB03, WQR13, AR07, BGM99, Bar01a, BBS96, FMK^{+}07, Gat01, Gia99, GPK^{+}07, GGZ02, JHT^{+}07, Jan03, KAS^{+}06, RGGL^{+}09]. Filesystem [RBM13]. Filesystems [vH02, Len97]. Films [GZK^{+}01]. filosofia [Cor05b]. Filtering [LCHL21, SKP96, WJS07]. filters [NJ11]. final [Ano01b]. Find [Ste01d, Daw94]. finder [SWF^{+}22]. Finding [AHSM21, SBRM09]. Findings [CMMST20]. Fine [ETKF07, RPB^{+}09]. fine-grained [RPB^{+}09]. finite [LWL06]. Firefox [FFPC22]. Firewall [AT05, BB99b, Don96a, Mat99b, Roe02]. firewalls [Ras07, SY00, Zie00, ZC02, ZCC00]. firmware [Fra19, GHL^{+}04, SvBC^{+}04, Sut02]. First [ACM05b, DW94, HR95, Ano03c, Ano02d]. Fits [UJE^{+}22]. FL [Ano94b]. Flames [Swa98]. Flannel [Moh97a]. Flash [Ano02d, Ano03a]. Flash-Based [Ano02d]. Flexibility [DRL^{+}02]. Flexible [Ano96a, TEGF98, DKK07, DOF96]. FLOSS [CIC13]. Flow [Cas14, CGDB24, BDH^{+}09, DKK07, GHRS05, KYB^{+}07, RKSV20, RPB^{+}09]. flow-based [BDH^{+}09]. fluid [LWL06]. fly [BGM99, SZD^{+}08a, SZD^{+}08b, SZD^{+}08c]. Focus [Ano99c, Ano99d, Ano99f, Ano99e]. Folly [CKM^{+}99]. fonctionnement [CDM97, Car98]. fonts [Ano03a]. footprint [CDD^{+}07, PLKH08]. Force [Ano03e]. forecasting [SBRM09]. forensic [FFPC22, KMP^{+}20]. forensics [AC11]. form [Ano02d]. Formatierung [GGZ02]. Formatting [GGZ02]. forme [Cha99]. Fortgeschritten [Ron05a]. Fortran [Ano02b, CS06]. Forum [CKM^{+}99]. forwarding [SH13]. Fotos [DF00]. foundation [BDH^{+}09]. FPGA [PBM^{+}22]. FPGAs [Ano02d, MKSB22]. Frame [Coo98]. Framework [BMF15, Kro00, LW00, Lis02, GJJ05, KEP07, LL94b, WXX08]. Frankencamera [AJD^{+}10]. Free [Lin02c, CPJ^{+}98, Eri97, Pau95, Pom04, Sun99, Tim99, Tro96a, Wal99, Way00, Ano95b, Ano95e, Cor05b, Cus04, Gra06b, Kau96, Kop05, Moh97a, Zic01, Lin02b, NR03]. Free/Open [NR03]. FreeBSD [Luc00a, Luc00b]. freedom [YL08]. FreeDOS [WF03]. freely [Daw94]. Freeware [Edw98]. Freeze [HHAW19]. freie [Lin02b, Lin02c]. French [Ron05b]. frequency [Kar16]. Friendly [ZBG^{+}05, HRX08]. Frightening [AMM^{+}18]. full [Daw94]. Fully [PS22]. Fully-Homomorphic [PS22]. Fun [TD01, Cas17]. functional [BFS07, DL08, DLP10]. Functionality [LBN14, Ano03a]. Functions [JMF14, LCHL21]. Fundamentals [Rob04]. funded [Ano95b]. futex [BF08]. Future [ASW^{+}01, Bla99, San98, Ano94f]. Fuzzing [DCH^{+}22]. Fuzzy [HMK^{+}20]. fwsnort [Ras07].

G [Ano02c, Men12]. gadget [Gib17].

Games [LH01, SuS01, Cas17, Kos00]. gap [BYVF08]. garbage [JHT^{+}07]. Gast [WF03]. Gast-Systeme [WF03]. Gateway [CCO^{+}05, Roe02, Wyl95]. Gaussian [HMK^{+}20]. Gcompris [CK06k]. GDB [MZG14]. GDK [Har99]. ge [Gün02]. ge-packt [Gün02]. Geek [Cas17]. Geeks [Gra04, Gra06b, Jan06a]. gekonnte [Ste00a]. Gelato [Ano06]. Gene [AUW08]. Gene/P [AUW08]. General [LW00, FST^{+}09]. Generalized [FMK^{+}07]. Generate [Bar01b]. Generating [CGP^{+}08, XCS^{+}22]. Generation [AR22, HKS02, MZG14, Thi04, ISTE08, Mat99a, TF96]. generator [Don22]. generic [AUW08]. Gentoo [Thi04]. George [ACM03b]. Georgia [USE00]. German [Lin02b, Lin02c, Bra04, CK06a, CK06b, GA04, GGZ02, uRH95, uRH96, Jor04, Lac06, Lin02a, Mag04, May01, RGRG05, Ron05a, Spr07]. gestalten [DF00]. gestión [VD01]. Get [DDJ98a, Ste01c, Gra06b]. get-things-done [Gra06b]. Getting [Car96, Ric13, Wt95b, WBS01, Wel94, Wel95b]. Ghost [Ano94c]. GHz [Ano03b, ZB19]. Gigabit [cFJH^{+}03]. GILK [PKFH02]. GIMP [DF00, DF00]. GIMP- [DF00]. GIS [SNF04]. GIST [FST^{+}09]. Git [GAH16]. Given [Den99a]. Glass [CKM^{+}99]. Glibc [Gar00]. glimpse [GGZ02]. glimpse-Tools [GGZ02]. Global [Ahm08, AUW08, Lla06, O’K00b]. global-scale [AUW08]. GNOME [Pin02, Ste00a, Wri00, Pen99]. GNU [Ano00b, Ano01a, Ano01c, Ano01b, Lin02b, Lin02c, Ano05b, Bra04, CV00, DF00, G^{+}00, G^{+}02, GA04, Gan04, Gün02, HH05a, Jor04, Kle00, Mac02, Mag04, Ron01a, Ron01b, Ron05a, Ste00a, SuS01, Ygg94, Ana99, Lin02b, Lin02c, Bel00a, BGO02, Cor00, Cor05b, Cou17, DRP01, Don04, Dum05, GA04, Gar00, GHL^{+}04, GO99, GS00, Hom00, Hun01a, Jan01, Jon05, Jor04, LO97, Mac99, Mag00, Mag01a, Mag01b, Mag01c, MS02, Mag04, McC99a, NR03, Lás05, PR96a, RÓ01, Ron05b, Ron05a, Sai01, SS05a, SNF04, Sul02, Swa00b, TF96, VD01, WY94, W^{+}95, YL08, Ygg93, Zac01, Zic01, BGM99]. GNU-Linux [Kle00]. GNU-Linux-Zertifizierung [Mag04]. GNU-Softwaretools [Jor04]. GNU/ [BGM99]. GNU/Linux [Ano00b, Ano01b, Lin02b, Lin02c, Ano05b, CV00, G^{+}02, GA04, Gün02, HH05a, Ron05a, Lás05, RÓ01, Ana99, Lin02b, Bel00a, BGO02, Cor00, Cor05b, Don04, Dum05, GA04, Gar00, GO99, Hom00, Hun01a, Jon05, Mac99, Mag04, McC99a, NR03, Ron05b, Ron05a, SS05a, SNF04, VD01, YL08, Zac01, Lin02c, Cor05b, RÓ01, Zic01]. GNU/Linux-Anwenderhandbuch [Ano01b]. GNU/Linux-Systeme [Gün02]. go [GP00, Gue19]. Goal [CKM^{+}99]. goals [GHRS05]. Goes [Bar00c, Bar01b]. Going [Bar01b, Ste00b]. Golang [Gue19]. Gold [SC98]. goniometer [JB95]. Good [Ste01d, Kau96]. Google [Sta06]. Googlebombs [Sta07a]. Googling [Sta06]. gorilla [Oua97b]. gourmet [HP01]. Government [Bar01b]. Governmental [CPJ^{+}98]. GPGPU [MKO^{+}17]. GPIB [Tim03]. GPS [PV02]. Grafik [Str94a]. Grafiken [DF00]. grained [ETKF07, RPB^{+}09]. Graner [Cha13]. graphical [Ano96a, DOF96]. Graphics [Kro99, Lin00, Bar00b]. GRASS [SNF04]. Graver [Cha99]. Gray [Moh97a]. Green [CPJ^{+}98, SKP^{+}02]. Grid [Mes00, CCO^{+}05, Wol04b]. Grids [cFJH^{+}03]. Group [Moe02]. Groups [BMF15]. groupware [LL94b]. Grown [AMM^{+}18]. Grown-ups [AMM^{+}18]. Grundlagen [DF00, G^{+}00, G^{+}02, GA04, SU94c, Str94a, SU95b, SE96]. GSM [AKL^{+}96]. GTD [Har99]. GTK [DF00, Ste00a, Pen99, Wri00]. GTK-Versionen [DF00]. guarantees [GMV07]. guest [MSZ09, BB99a]. GUI [Bro00, Gar99]. Guide [BM02, BDD^{+}06, C^{+}03, DG00, Gag01, GO99, Han97, Kir95a, Kir95b, Kir95c, Lat02, McC02, PDG02, RB08, Ros98, Sob97, Sob03, Tra96, WW01a, WW01b, Bar04, BDP05, BGO02, Bri05a, Chi02, Coo09a, Coo09b, Coo09c, Coo09d, DRP01, Eck11, FT95, Gra05, Gre94, Hah09, HP01, HNV98, Jan04, Kir94, KD00, Kom96, LR97, MM00, Men12, Nor99, Par96, Par00, Pet01a, Ran97, Ran99, Ran00, Ric98, Sha00, Sha01, S^{+}07, Sob06, TR99, Tom11, Wal01b]. guidebook [Gra06b]. Guided [UJE^{+}22]. Guix [Cou17]. gut [BW04].

H [MRW99]. H-RMC [MRW99]. habitual [Mil01]. Hack [Chi02]. hacker [TR99]. Hacking [HLK01, Kra07]. Hacks [ORC06, Fli03, Kra07, PB05, Ran04, Ran06, Ran07, VJ06]. HADS [HMK^{+}20]. Hahn [Hah09]. Haiku [Lea12]. Hall [Hid04, Sta07b, Tho05]. Hallinan [Sta07b]. Handbook [Bra04, EDK00, KC98, NSH02, Ron05a, Smi00a, Ker10, KC00, NSH06, NSHW10]. Handbuch [WF03, Bra04]. Handheld [SNF04]. handler [LP10]. Handling [BJ01, GVK^{+}03, RLB08]. Hands [Ano03f, DM19, Gib17, Gue19, Koh06, SC98, Tom11, MB95, MB05, Shu06, S^{+}07]. Hands-on [Gue19, Koh06, SC98, Tom11, MB95, MB05, Shu06, S^{+}07]. happen [SBRM09]. Happn [KMP^{+}20]. Happning [KMP^{+}20]. Haptic [Ano03f]. Hard [TEGF98]. Hardening [Tur05, NJ11]. Hardware [Car96, DRL^{+}02, Hol02, KIKT01, SZD^{+}08a, SZD^{+}08b, SZD^{+}08c, Smi00a, YK15, vD06, ARJ08, Kop05, MHKT09, RHP^{+}07, VMKB05]. hardware/software [ARJ08]. Harley [Hah09]. healing [SLP^{+}09]. heap [LLA07]. Hellman [ASW^{+}01]. Helmke [Cha13, Teo13]. Here [CS95, Oua97b]. Heterogeneous [VTS^{+}23, Ano96a, DOF96, GR09, GCN^{+}10, PBM^{+}22]. Hierarchical [ABB19, HLT^{+}21]. High [ACM04, Alm96, Bar01b, Bra00, DRL^{+}02, Gig06, IEE06, JMF14, Lum01, MLE^{+}00, PKB01, SHB23, Slo05, SSSS97, AOAH14, AOAH15, BSR06, HRX08, SPF^{+}07, Way00, AOAH14]. High-Availability [Bra00]. high-bandwidth [BSR06]. high-BDP [AOAH14]. High-MCC [JMF14]. High-Performance [Gig06, IEE06, Lum01, MLE^{+}00, PKB01, SPF^{+}07]. High-speed [Alm96, AOAH14, AOAH15, HRX08]. High-Tech [Bar01b, Way00]. highly [Kop05]. History [Bou95, Bou96a, Bou96b, Cer03, Pom04]. HLA [LLT07]. HOL [BFN^{+}06]. Hole [Sta04]. Holistic [VFMM08]. Home [FES^{+}99, Lum13, NW03, Neg06b]. Homomorphic [PS22]. honeypots [ALL06]. Honors [Coc03a]. hooks [GJJ05]. hop [WJS07]. hop-count [WJS07]. horizons [ACM93]. Host [Chu03, OBSB23, Shu06]. Host-Based [Chu03]. Hosted [Coo98]. hosting [USR09]. hosts [SH13]. Hot [Roe02, Wis03]. Hot-Failover [Roe02]. Hotel [ACM93]. hours [BS98, Bal99b, Gay99, Hsi01, SBR01, Tay03, WT99, Wel99a]. House [Ste01f]. HP [Kuo99, Pon05, She06]. HP-UX [Kuo99]. HP/OSLO [She06]. HPCS’06 [IEE06]. HTM [HRW09]. HTML [Kro99, Pau95]. Hudson [Teo13, Teo13]. HULK [VTS^{+}23]. HULK-V [VTS^{+}23]. Hungarian [Lás05]. Hurd [Ano01a, Epp00a]. Hybrid [MRW99]. Hydrogenated [GZK^{+}01]. Hype [Gla99]. Hypertext [FES^{+}99]. hypervisor [MSZ09, SLQP07]. hypervisors [SPF^{+}07].

i-nodes [Len97]. I/O [BJL^{+}06, CKR08, GPK^{+}07, LSP07, PSK08, Rus08, dBB08]. i386 [Ano01a, Ano01b]. IA [Gig06]. IA-64 [Gig06]. IA32 [ST03]. ia64 [EM00]. IBM [AJ05, AV04, BBB^{+}20, G^{+}06a, P^{+}08, Sta05b, SvBC^{+}04, ZST^{+}07]. ICE [Ano06]. Iceberg [AHSM21]. IDE [Ano02d, Ano02c, Ano03d]. Identification [HLT^{+}21]. IDEs [Sor01]. idiot [Ric98]. IEC [ISO05]. IEEE [ACM04, ACM05a, Ale00, HKP09, PPSM96]. IEEE-1394 [HKP09]. IEEE-488 [PPSM96]. IEEE/ACM [ACM04]. IFT [DRL^{+}02]. Igniting [ACM03a]. Ihre [RB06a]. II [Ano01c, Eri97, uRH96, MS02, Neg06b, Sai01, Sul02]. III [BBB^{+}20]. Illinois [ACM05b]. illustration [LF95]. im [KGG00]. image [SKP96]. images [LHB24, Ran06, TD97]. Imaging [Kro99]. imake [Hus94]. Implantación [VD01]. Implantation [VD01]. implement [MS07]. implementación [RÓ01]. Implementation [GD18, JHT^{+}07, SB99, SGHL01, BGM99, BBS96, CP01, CP02, Daw94, Gab07, KAS^{+}06, LZ07, Lla06, PC01, RÓ01, Tor94, Wan02, WC22, Wer10]. implementations [BFN^{+}06]. Implemented [YKS^{+}02]. Implementing [CC97, Lam99, SMTW96, SW98, Wel95a, XZY^{+}10]. important [CK06b]. Improved [HRD^{+}07]. improvements [SvBC^{+}04]. Improving [GPK^{+}07, WSL^{+}99, SAF07]. In-Situ [RGK99]. IN-Tune [RGK99]. Increased [Vad00]. independent [Lin02b, Lin02c]. India [R.99]. individuals [Daw94]. Inducing [JJW^{+}24]. industrial [Fli03, Ran04]. industrial-strength [Fli03, Ran04]. Industry [Bar00b, Edw98]. Inel [SuS01]. Inel- [SuS01]. inferring [LPH^{+}07]. info [Moh97b]. Information [CGDB24, KD96, Kro00, KYB^{+}07, SS05b, BC08, DKK07, FOP06, GHRS05, RPB^{+}09, Wal01b]. Infrastructure [Gar99, MKSB22]. Infrastructures [SHB23]. Ingol [Hid04]. initial [Ano03b]. initialization [DIN05]. Innovation [ACM03a]. Inputs [CGP^{+}08]. ins [DF00]. insights [ETKF07]. Inslaw [Eri97]. instalacion [Bla97]. installable [BBS96]. InstallAnywhere [Ano02c]. Installation [Bor01, EDK00, G^{+}00, G^{+}02, GA04, GO99, KGG00, Kof97, SU94c, Str94a, SU95b, SE96, VRJ97, WF03, Ano96b, Jan01, Kof99, Kof00, Mag00, Mag01a, Pod97, Ron05b, SU94b, SUM95, SUM96, SMM97, VRJ95, VRJ96, VRFJ98, Wel94, Wel95b]. Installations [Kro00]. Installing [Hag00]. InstallShield [Ano02d]. Instructional [DW08]. instrument [OMCB07]. Instrumentation [PKFH02, DCH^{+}22, OMCB07]. Instrumenting [MZG14]. Insurance [Eri97]. Integrated [Ano02d, BSR06]. Integrating [McC01, TSC02]. Integration [Ano02c, CPJ^{+}98, Koh06, LBN^{+}18, LS01, Bru02, MB05, WG18]. Integrity [HHD^{+}16, Pon05, SLQP07]. Intel [Ano01b, Ano97b, Ano02b, Ano03d, Bar00b, Hua06, Red98]. intellectuelle [Cor00]. Intelligent [Ano02d]. Interactive [LD05, Kom02]. interactivity [TCM07]. interface [Ano96a, DOF96, Ker10, MHKT09, PPSM96, SNM22, ZS93]. Interfaces [KD96, BSR06, SH13]. Interfacing [Hol02]. Intermediate [Ros98]. InterMezzo [Bra00]. Internals [BBD^{+}96, Bar00a, B^{+}98b, Ben06]. International [ACM05b, ACM06, Ano94b, Ano94a, DW94, IEE93, IEE04, IEE06, ISO05, KD96, Pop94, SS05b]. Internet [CK06b, DF00, FST^{+}09, KGG00, CK06b, CS95, CP01, CP02, Don96b, HE95, HNH03, Kha02, Kro00, Le 96, PC01, SR96, UPS^{+}05, UPS^{+}07, USC^{+}08, USR09, VRFJ97b, Zic01, ZS93, ZCC00]. Internet-based [HNH03]. Internet-challenged [Kro00]. Internetkommunikation [CK06b, CK06d, CK06c]. Internetprogramme [CK06b]. Internetprogrammen [CK06d, CK06c]. Internetworking [CS01]. Interoperability [Bar00b, BM02, Wis03]. Interrupt [LP10, RLB08]. interrupt-driven [LP10]. Interview [Li94, Ray99b, Pub11]. Intervoice [Ano03c]. Introduces [Ano03b, Ano03e]. Introducing [MW08]. Introduction [BB99a, CK06a, CK06b, RB05b, Bok95b, Han97, Sho12, Wün13]. intrusion [WC22]. Invasion [ASW^{+}01]. invasive [RGK99]. Investigations [She06]. IO [PSK08]. iOS [KMP^{+}20]. IoT [DCH^{+}22]. IP [Str94a, BSR06, BB99b, CS01, CP01, CP02, FES^{+}99, PC01, SH13, WJS07, WW02, Wer10]. IP-forwarding [SH13]. IPsec [YW07]. IPSIS [But02b]. iptables [Pur04, Ras07]. IPv6 [AB04, Sta05b]. IRAF [TF96]. ISBN [Hid04, Jen05, Koh06, Sch04, Sta07b, Tho05]. ISDN [KGG00, Don97a]. isn’t [Daw94]. ISO/IEC [ISO05]. isolating [ALL13]. isolation [GNDB16, WRA05]. Issue [Eri97, FOP06, FOP06]. Issues [GZXZ23, TB05]. Italian [Cor05b]. Itanium [Ano06, Cha06, HKS^{+}04, Pon05]. Itanium-based [Ano06, Pon05]. ITI [KD96]. Itself [Bou95, Bou96a, Bou96b]. IV [BSSS96, BSS^{+}96].

J2EE [Ano02c]. J2EE-Compatible [Ano02c]. January [USE96b]. Java [Ano02d, Ano03a, Ano03c, Ano03e, DDJ98a, Gar98, HKS02, Jen02, Jon02, Kro99, Kro00, Oua97b, SKP^{+}02, Tim03, XSaJ08, YKS^{+}02]. JavaScript [Ric14, Gab07]. JDK [Gar98]. JDS [AH04]. jedermann [Str94a]. Jermy [Koh06]. jetzt [KGG00]. Jiao [RB08]. JIT [OMCB07]. Job [FES^{+}99]. Jobs [Bar01b]. John [Hid04, IEE06]. Journal [Eri95, Lin94]. Journaling [LBN14]. joy [HP01]. Jr [ACM99]. July [Ano94d]. June [ACM05b, ACM06, DW94, KD96, Tim99]. Just [TD01].

K42 [AAB^{+}05, KAR^{+}06]. Kali [Ano14]. Kanotix [CK06e, CK06e]. Karl [Jen05]. Kaufman [Ano95a]. Kaydara [Ano02c]. KDE [KGG00, SuS01, Böl01, Gri01a, MWM00, Mot02, Swe01, Wel99a]. KDE/QT [Gri01a]. Keep [Bar01b]. Keeping [Sta06, Jan06b]. Kernel [AR22, AMM^{+}18, Ano02a, Ano05a, BLH20, BBD^{+}96, BC02, BC03, CDM98, CMR^{+}21, EM00, Far01, GZXZ23, GD18, Geo94, HLT^{+}21, Hua06, JKS^{+}15, JMF14, JJW^{+}24, Kie99b, KLK^{+}02, LCHL21, MAP^{+}22, Max01, MFBWW20, MRW99, Nut01, PQM^{+}21, PKFH02, RMF19, RFS05, RFS06, SGWP02, SB99, SuS01, UJE^{+}22, Wel95a, XZY^{+}10, dOCC23, ALL13, Ano03d, AAB^{+}05, BF08, B^{+}98b, BYVF08, BC00, BC06, BV06, CIC13, CDD^{+}05, CDD^{+}07, DFC^{+}23, ETKF07, HNH03, HHAW19, JGS^{+}08, JYX^{+}16, KS19, Kor06, KH07, LSAF16, Lov05, Lov10, MMW96, Mau08, Mau96, MW08, NJ11, NH14, NAGL10, PFHLN22, RLPA18, SLQP07, SWF^{+}22, ST03, Tor94, WG08, WC22, YSCO04, dOdO16, BC01, HBB13]. Kernel-level [CMR^{+}21, Ano03d]. kernels [MR07, WYW^{+}13]. Kheops.96 [Ano96b]. Kheops.96-Ver [Ano96b]. Kickstart [O’K00a, McC07]. Kinder [CK06k]. King [ACM99]. Kit [Ano02d, Dul06, PFH97, SU95a, Gay99, SU94b, SU96a, SU97, Tau95, Tau96]. Kittyhawk [AUW08]. Klassische [RB06a]. KLogger [ETKF07]. Knoppix [CK06f, Ano01c, Neg06a, Neg07a, CK06f, Ran04, Ran05, Ran07]. Know [Wel99c, Lov07, War04, War15]. knowledge [RS14]. Knows [Coc01]. Kommunikation [Lin02a]. kompatibilni [WKD97]. kompromisslos [Bra04]. Konfiguration [Bor01, WF03]. Konvertieren [Gün02]. Kopper [Jen05]. Korn [May01, May01, Olc01]. Korn-Shell [May01]. Kubuntu [CK06g, CK06g]. kurz [BW04]. Kylix [BJ01, TMG01].

L4 [KEP07]. L4oprof [KEP07]. Lab [AB04, Coo09b, Coo09d, Tom11]. Laboratories [DW08]. laboratory [MB95, FES^{+}99]. Lake [ACM03b]. Laminar [RPB^{+}09]. LAMP [G^{+}06b, LW03]. LAN [RÓ01]. Landing [ACM03b]. Langages [Bla02]. language [Cas17, Dun00]. Languages [CPJ^{+}98]. laptop [HC15]. Lar [Ano95a]. Large [GVK^{+}03, MFPP03, DFC^{+}23]. Large-Scale [GVK^{+}03, MFPP03, DFC^{+}23]. latency [CC97]. Later [MFBWW20, PTS^{+}11]. LaTeX [Str94a]. Launches [Ano03f, Ano03d]. Law [CPJ^{+}98, Bar01b, MSSvK08]. Laws [CKM^{+}99]. layer [Ano03c]. LCA [Mag04]. LCP [Mag04]. LCP/LCA [Mag04]. LDP [Ray98]. Learn [Dos99b, LH01]. Learning [HLT^{+}21, Mac02, MAP^{+}22, McC99a, McC99b, McC02, McC03, McC04, Wel99b, Cha11, SWF^{+}22]. Learning-Based [HLT^{+}21]. Legal [Pom04]. Legit [Sta06]. Legitimizing [mH99a]. Lehrbuch [May01]. Lern [CK06k, Kuo99]. Lern- [CK06k, Kuo99]. Lernprogramme [CK06h, CK06i]. Lernprogrammen [CK06h, CK06i]. less [PLKH08]. Letters [ASW^{+}01, Ano94c, CPJ^{+}98, DM97, FES^{+}99, WSL^{+}99]. Level [Ros98, Ano03d, Cas17, CMR^{+}21, DCH^{+}22, LZ07, Mag00, Mag01a, Mag01b, Mag01c, MS02, Sai01, Sul02]. leveraging [ZBG^{+}05]. Liability [CKM^{+}99]. libC [Gar00]. libero [Cor05b]. libraries [PS07, PLKH08]. Library [But02a, Ano03d, Lov07]. libres [Cor00]. Life [Cas17]. Lifecycle [Kro99]. lifetime [SLQP07]. Lightweight [JKS^{+}15, OBSB23, JYX^{+}16]. Like [Sta04, Moh97b]. likely [Pub11]. limit [Pur95]. Limiting [CKM^{+}99]. Lindows.com [Ano03d]. line [Blu08, Blu11, Sho12, Ste99]. link [KS09b]. Linus [CPJ^{+}98, Flo94, Li94]. LINUX [Bra04, KGG00, Str94a, Ano99a, Ano99b, Ano00a, Ano00b, Ano01a, Ano01c, Ano01b, Ano02a, Lin02b, Lin02c, Ano05a, Ano05b, Ano06, Bar01c, BW02, BR95, Bon99, CV00, CK06a, CK06b, CK06d, CK06c, CK06f, CK06e, CK06g, CK06i, CK06j, Cor05b, DW06, Dan99b, DWV06, Den99b, Dos99b, Dun99, DF00, EEK^{+}04, G^{+}00, G^{+}02, GA04, Gan04, G^{+}06a, Gra99, Gri01b, Gün02, GGZ02, HS99, HS00, Hid04, HH05a, uRH95, uRH96, Jen05, Jor04, Kad95c, Kad95b, Kle00, Koh06, Kuo99, Lin02a, Mag04, May01, Neg01, Neg02, Ran99, Ric14, RÓ01, Ron05a, Sav11, Sch04, SC98, Sob03, Sta07b, SU96d, SuS01, Tho05, USE00, USE01, Hag00, WF03, Wei04, BH98, Han99, Kof00, McN99, Mea99, Lás05, RÓ01, SSSS97, SU94c, Str94a, SU94a]. LINUX [Tau96, TSEK01, VRFJ98, Wel94, ALC20, Abb03, ALP^{+}21, ABB19, AT05, AH04, AT07, ALL13, AB04, Ale00, AMM^{+}18, ASW^{+}01, AOAH14, AOAH15, AC11, AJ05, Ana99, And99, Ano93, Ano94c, Ano94e, Ano94g, Ano95b, Ano95c, Ano95d, Ano95f, Ano95g, Ano95e, Ano95h, Ano95i, Ano95l, Ano96b, Ano97a, Ano97b, Ano98a, Ano98b, Ano99c, Ano99d, Ano99f, Ano99e, Ano01b, Lin02b, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano14, AAB^{+}05, AR07, Asb00, Aul01, BGM99, BKB^{+}18, BLH20, Bai94, BS98, B^{+}98a, Bal99a, Bal99b, BP00, Bal00, BSSS96, BSS^{+}96, BBD^{+}00, BN01, Bar00a, Bar01a, Bar00c, Bar01b, Bar96a, Bar96b, Bar97, Bar98, Bar99a, Bar99b, BBK^{+}16, BSB03, BW04, Bar04, Bau02, Bau03, BsswLB05, BDP05, Bea04, BBD^{+}96, B^{+}98b]. Linux [Bee00, Bee16, BMF15, Bel00a, Bel00b, BYVF08, Ben96, Ben06, BFS07, Bes05, BBHL08, Bla99, Bla02, Bla04, BTS09, Bla97, Blu00, Blu08, Blu11, Bok95a, Bok95b, BR99, BJ01, BB99a, Bol99, Böl01, Bon01, Boo03, Bor01, BS05, BBB^{+}20, Bou95, Bou96a, Bou96b, BC00, BC01, BC02, BC03, BC06, BB99b, BM02, BGO02, Bra96, BV06, Bri05a, Bri05b, Bro00, BR^{+}00, Bro06, BDD^{+}06, Bru02, Bur97, Bur04, But02a, But02b, BF97, CF20, Cam04, CIC13, Car99, CDM97, CDM98, Car98, CDD00, CMMST20, Car96, Cas14, CMR^{+}21, CM06, CDD^{+}05, CDD^{+}07, Cha11, Cha99, Cha06, Chi02, CK06a, CK06b, C^{+}03, Chu03, Coc01, Coc03b, Coc03a, Col97, Col02, CS01, Com00, Com03, Coo98, Coo09a, Coo09b, Coo09c, Coo09d, CRKH05, Cor00, Cor05b, Cor97]. Linux [CKM^{+}99, CP01, CP02, DDJ98b, DDJ98a, DH02, DR01, Dan99a, Dan01, DJ01, DRP01, Daw94, Dea01, Dec04, Ded09, Den99a, DG00, Dik06, DIN05, DL08, DLP10, DRN^{+}23, Don04, Don22, Don96a, Don96b, Don97a, Don97b, Don98, DM19, DCH^{+}22, DFC^{+}23, Dul06, Dul14, Dum05, Dun00, Dup05, DRL^{+}02, Eck11, EEMS03, EDK00, Elb01, EJMT94, Emm11, Epp00b, EM00, Eri97, Fab97, Far94, Far01, Fel96a, Fel96b, Fel13, FFPC22, FWS^{+}03, FSW02, Fis02, FM02, Fli03, Flo99, FES^{+}99, FT95, Gab07, Gag00b, Gag01, Gag06, Gag07, GJJ05, Gan03, GA04, Gan08, Gar99, Gar98, Gar00, GZXZ23, Gat01, Gay99, Gay00, GD18, Geh98, GGG^{+}98, GGG^{+}99, G^{+}06a, GW94, Geo97, Geo94, GAH16, GVK^{+}03, G^{+}06b, GJ08, Gig06, GSN^{+}01, GSHL03, Gla99, GD96, GO99, Goe00, Gol95, GBC^{+}18]. Linux [Gon99, GZK^{+}01, GP00, Gra06a, Gra04, Gra05, Gra06b, Gre94, Gre19, GLS03, GNDB16, Gue19, GMTW08, Gün02, GGZ02, GHRS05, Hah09, HMK^{+}20, Hal05, mH99b, mH99a, HS99, HP01, HHJC99, Hal06, Hal07, Han00, HKS02, Han97, HE95, HBB13, Har99, HZ99, HST08, HLK01, Haw00, Haw02, Hea97, Hei99, Hek97, HNV98, HT97, HH04, Her99, HNH03, uRH95, uRH96, HLT^{+}21, HKS^{+}04, Hol02, Hol96, Hom00, Hou96, Hsi01, Hua06, HHD^{+}16, Hub02, HHAW19, HH05b, HH08, Hug97, Hug98, Hun01a, Hun99, Hun00a, Hun00c, Hun00b, Hun01b, Hun02, HP^{+}95, HP96a, HP96b, IBM05, ISO05, ILG19, Inf95, JKS^{+}15, JHT^{+}07, JGS^{+}08, Jan01, Jan03, Jan04, Jan06a, Jan06b, JMF14, JWKZ99, JW99, JPS00, JYX^{+}16, JJW^{+}24, JT98, JT05, JB99, Jon05, Kab99, Kad95d, Kad95a, KBPV99, Kal16, KIKT01, KEB^{+}99]. Linux [Kar16, Kas14, Kau96, Ker10, Kha02, Kie98, Kie99b, Kie99a, KLK^{+}02, Kir94, KC95, Kir95a, Kir95b, Kir95c, KD00, K^{+}06, Kla93, Kof97, Kof99, KS19, Kom96, KC98, Kom00, KC00, Kom02, KAS^{+}06, Kop05, Kor06, Kos00, KH07, Kro99, Kro00, Ksh04, Cor05a, LH01, Lac06, Lam99, LBN^{+}18, LHB24, Lat02, Le 96, LW03, gL02, LS01, Lei99, LSAF16, LR97, Lev99, Lev95, LCHL21, LH00, LW00, LZAH22, LZAH23, Lin02a, Lin00, Lip99, LWL06, Lom01, Lov05, Lov07, Lov10, Loz05, LADADL14, Luc00a, Luc00b, Luc05, Lum01, LCJH05, Mac02, Mac99, Mag00, Mag01a, Mag01b, Mag01c, MS02, Mag04, MAT08, MSSvK08, MR18, Man01, Man99, MM00, Mar04, Mar00, Mar01, Mar06, MAP^{+}22, MP98, Mat99a, MSB^{+}96, MS99, MS^{+}00, Mat99b, Mau08, Max00, Max01, May01, MMC06]. Linux [MKSB22, McC99a, McC99b, McC02, McC03, McC04, McC05, McC01, MM06, MW08, MFBWW20, MRW99, MWM00, MSW06, MZG14, MM01, Mil01, Mil96, MYH00, MY03, Min01, MOS01, Moe02, Moh97a, Moh97b, Moh98, MR99, MK04, Mol14, MKO^{+}17, Mon97, Moo01b, Moo01a, Mor94, Mor97, MB05, MLE^{+}00, Mus02, NJ11, NH14, NR03, NW03, Neg06a, Neg06c, Neg06b, Neg07a, Neg07b, NAGL10, Nel00, NSH02, NSH06, NSHW10, Neu06, NV06, Nie00, Nis01c, Nis03, Nor99, Nor02, Nut01, O’G01, O’K00b, OR96, Olc01, Oom13, OBSB23, Oua97a, Oua97b, PLM06, PLHM08, PBM^{+}22, PTS^{+}11, PTS^{+}14, PDG02, PS22, PS07, PKB01, PKB03, PNK02, Par96, P^{+}97, Par98b, Par98a, Par99, Par00, P^{+}08, PQM^{+}21, PFHLN22, Pau95, PKFH02, Pet96, Pet98b, Pet98a, Pet99, Pet00b, Pet00a, Pet01b, Pet01a]. Linux [Pet02a, Pet02b, PB05, Pet00c, PWN01, Pfa99, Pfa00, Pfa01, Phi00, PC01, Pit98a, Pit98b, PB98a, PB98b, Pod97, Pon05, Pri99, P^{+}06, PF96, PFH97, PF97, Pub11, Pur95, PR96a, PR96b, Pur97, PR97, Pur04, Qui00, QH01, R.99, RR15a, Ran97, Ran00, Ran06, Ras07, Rau01, Ray99a, Ray01a, RRR09, RR15b, Ray98, Ray99c, Ray99b, Ray01b, Red98, RMF19, RLB08, Reg11, Reh99, RP02, RGRG05, RLPA18, Ric98, RBHB99, Rob04, Rob95a, Rob95b, RBM13, RGK99, RFS05, RFS06, Roe02, Ron01a, Ron01b, Ron05b, Ron05a, Ros98, RS14, Rub98, RC01, RC98, SGP^{+}23, Sai01, SK10, SCB^{+}13, SH13, SBR01, San98, SR96, SS05a, Sch05, Sch07, SAL19, SJP^{+}09, SG04, Ser98, SH99a, Ser00, SK01, Sha00, Sha01, SGWP02, SK04, She02, SB99, She03]. Linux [She06, SWF^{+}22, ST03, SAN08, Sho12, Shu06, Sie00, SH99b, S^{+}00, SWL^{+}01, SFW03, Sie05, S^{+}05, SHB23, SMTW96, Slo05, Smi99, Smi00a, Smi00b, Smi01, Smi02, Smi03, SD05a, SD05b, Smi05a, Smi05b, S^{+}07, Smi12, Sob97, SC98, Sob06, SGL^{+}00, SGHL01, SY00, Sor01, Spe00b, Spr06, Spr07, SS01, SS02, Sta04, Sta05a, Sta05b, Sta06, Sta07a, SNF04, Ste01a, Ste02, Ste01c, Ste01f, Sto00, SU94b, Str94b, SU94d, SU95a, SUM95, SU95b, SU96a, SU96b, SE96, SUM96, SU96d, SU96c, SU97, SMM97, SE97, Stu01, Stu04, Sul02, Sun99, SKP^{+}02, SW98, Swa98, Swa00b, Tim99, TR99, TG95, TGB95a, TGB95b, TGBO96, TG96, TGW97, Tac97, TBD98, TB98, Tan00, Tau95, TW96, Tay99a, Tay99b, TD96, Tem96, TEGF98, TSC02, Thi04, Tho06, Thr01, Tim03, TCM07]. Linux [Tor94, Tor99, Tot99, Tox07, Toz16, Tra96, TD97, Tro96a, Tro96b, TECB03, TF04, TB05, Tur05, Tyl06, UJE^{+}22, VD01, Vad00, Vad02, VTS^{+}23, Ven08, VMKB05, VE99, VRJ95, VRJ96, VJR96, VRJ97, VRFJ97b, VRFJ97c, VRFJ97a, Vol97, VR00, VJ06, Vre03, WWW99, WSL^{+}99, Wal00, Wal01a, Wal95, WKDD03, Wal01c, Wan02, Wan11, WYW^{+}13, WC22, War00, War01, War04, War15, Way00, WDJ04, Weh01, Wel98, Wel99c, Wel99b, WT99, WY94, Wt95a, Wel95a, W^{+}95, Wt95b, Wel95b, WK95, WCC95, WK96, WKD97, WKC^{+}97, WDK99, WLS^{+}01, WKDD02, Wie93, Wie96, Wie97, WWS98, Wie99b, WPL^{+}99, Wie99a, WG18, WB06, Wil98, Wis03, WRA05, WWH97, WWH98, Woe03, Wol04a, WTKW08, Wor98, WBS01, WXX08, Wün13, WM00, Wyl95, XZY^{+}10, XCS^{+}22, Yag96, Yag03, YL08, YKS^{+}02, Yeo05, Ygg93]. Linux [Ygg94, Yor99, Lin94, YSCO04, Yu07, Zac01, Zad01, ZB19, ZST^{+}07, Zic01, Zie00, ZC02, dBB08, dOdO16, dOCC23, vH02, vH06, vH07, Hul05, Men12, Wer10, Ano95a]. Linux- [Her99]. Linux-Anwenderhandbuch [Ano01b, Ano01b, Ron01a, Ron01b]. Linux-based [ABB19, AOAH15, CMMST20, CM06, DCH^{+}22, MKSB22, PBM^{+}22]. Linux-basierte [Ric14]. Linux-Betriebssystem [SuS01].

Linux-Capable [OBSB23]. Linux-Cluster [GVK^{+}03]. Linux-compatible [AAB^{+}05]. Linux-Hosted [Coo98]. Linux-Ready [PS22, ZB19]. Linux-Systeme [Gün02]. Linux-Tasks [RB06a]. Linux-USB [Reg11]. Linux/GNU [Jan01]. Linux/GNU/ [Ygg93]. Linux/GNU/X [Ygg94]. Linux/ia64 [EM00]. Linux/Java [HKS02, YKS^{+}02]. Linux/Posix [CS01]. Linux/Unix [Gab07, Sor01, Ano03a]. LinuxONE [BBB^{+}20]. LinuxWorld [Bar00b]. Linuxzeitung [Lin02b, Lin02c]. LISA’96 [USE96a]. Lite [SC98]. Live [CCZ^{+}06, KS09a, Neg06c, Neg07b, Lum13]. liveness [BC08]. Living [Don99, Man99]. Load [Kro99, Mon00a]. Loadable [Wel95a]. local [TK07]. localities [DJC07]. lockdown [Chi02]. Locking [CCVLW23]. LOCKSMITH [PFH06]. log [Gat01, JHT^{+}07, KAS^{+}06]. log-structured [Gat01, JHT^{+}07, KAS^{+}06]. logging [ETKF07, PFHLN22]. Logic [LS01, BFN^{+}06, BDH^{+}09]. logical [HRS^{+}10]. logiciels [Cor00]. Look [Kro00, Neu06, Ben96, Kas14]. Looking [ASW^{+}01, Ano94f]. Lookup [DRL^{+}02]. loose [Sta05b]. Losing [Sta05a]. Lösungen [Bra04]. Lotus [RMZ01]. Low [Rob95b, Alm96, Ano03d, Cas17, CC97, GFRD96, MKO^{+}17, VTS^{+}23]. low-cost [Ano03d, GFRD96]. low-end [Alm96]. low-level [Cas17]. low-power [MKO^{+}17]. Loyal [Ded09, Gla99]. LPI [Dea01, P^{+}06]. LST [PF96, PF97]. LST-distribution [PF96, PF97]. Lucke [Tho05]. Luther [ACM99]. LVDS [Ano02d]. Lxrun [Sun99]. Lyngby [DW94].

M [Wer10]. Mac [AC11, Cha99, TECB03]. Mach [GD96]. Machine [Roe02, BDF^{+}03, CK06a]. Machines [CCO^{+}05, LD05, ABB19, WW01b, ZBG^{+}05]. Macintosh [Mor97]. Macmillan [TG95]. MacOS [May01, May01]. Macromedia [Ano02c]. made [Gra05, Mus02]. Mail [Blu00]. main [HHAW19]. Mainframe} [Hid04, EEMS03]. mainstream [BBHL08]. maintainability [YSCO04]. maintainer [BJM^{+}22]. Maintaining [PKB01, Boo03]. Maintenance [FES^{+}99]. Make [FES^{+}99, Cas17]. makes [Hol96]. Making [LLA07, O’S02]. malware [CMMST20, Rau01]. manage [BFS07, Tro96b]. Manageable [PKB01, PKB03]. Management [Ano02d, BMF15, Gag00a, Jan06b, Ano94b, CKR08, Cus04, DJC07, GR09, LZZ^{+}08, LZAH22, LZAH23, Max00, ON06, SAF07, VD01, ZYCZ10]. Manager [Kro00]. Managing [Cam04, Kop00, McC07, Spe00a, Vad02, RHP^{+}07, Sav11]. Manual [DF00, Olc01, Tom11]. manufacturing [VMKB05]. many [Hou96]. Map [Sta04]. March [SS05b]. Maria [Hid04]. Mark [Men12]. Martin [ACM99]. Masquerading [BB99b]. Mass [Coc01]. Master [Bel00b, Gri14]. Mastering [Dan99a, Dan99b, Dan01, DJ01, Gri14, Wan11, WB06]. Matching [BJ01, BDH^{+}09]. Mathematical [IBM05]. Mathematics [uRH95, uRH96]. Mathematik [uRH95, uRH96]. Matt [Ano95a]. matters [Yag96]. Matthaeu [Hid04]. Matthew [Cha13, Teo13]. Mature [Bra04]. maximize [Gri14]. Maximum [Bai97, Coc01, HRW09, Ray01a, TR99]. May [Bar01b, IEE06]. MCC [JMF14]. mean [SKP96]. mean-filtering [SKP96]. means [CDOY09]. measurement [KS09b, Lla06, PPSM96]. Mechanism [MSSvK08, GMTW08]. Mecum [Lip99]. Meets [Nis01c]. Melbourne [Ano94f]. Memory [GSN^{+}01, GSHL03, KLK^{+}02, LBN14, RRP^{+}08, WQR13, CDD^{+}07, GZXZ23, GCN^{+}10, GNDB16, GMTW08, HHAW19, IZG^{+}07, LZZ^{+}08, PLKH08, RRP^{+}07, RLPA18, RHP^{+}07, RRH^{+}08, SMXH01, WRA05]. MemTo [GSN^{+}01].

Merge [WG18]. Merge-Tree [WG18]. Merits [AHSM21]. mesh [KS09b]. Message [NJ11, AWW08, PS07]. Messages [DR01]. MetaCreation [Kro99]. MetaTM [RRP^{+}07, RRP^{+}08, RRH^{+}08]. MetaTM/TxLinux [RRP^{+}07, RRP^{+}08]. Method [HHD^{+}16, JYX^{+}16, WG08]. methodology [FFPC22, Wil98]. Methods [Geo94]. metric [MB08]. Metrowerks [Ano02d, Ano03b, Ano03c, Ano03d, Kro99]. Mexican [Bar01b]. mezzi [Zic01]. MICRO [IEE04, Mau96]. MICRO-37 [IEE04]. Microarchitecture [Bar00b, IEE04]. microelectronics [PPSM96]. Microkernel [Mor97, KEP07, Sto07]. microkernel-based [Sto07]. microprobe [JB95]. Microscope [Ano03e]. Middleware [Kro00, AJ05]. MIDI [Geo94]. Migration [BDD^{+}06, Cha06, CAW08, KS09a, LP10]. Minding [BYVF08]. Mine [DDJ98b]. Mini [Gra99]. Mini-Review [Gra99]. Minimal [WM00, HRW09]. mining [GAH16]. MiniRTL [WM00]. Minitel [DM19]. Minix [Wil98]. minutes [Ray99a, RBHB99]. MIPS [Swe07]. misses [ST03]. Misunderstood [DM97]. Mixed [Koh06, MB05]. Mixing [LD05]. mixture [HMK^{+}20]. mixture-based [HMK^{+}20]. mk [LHB24]. MkLinux [Mor97]. MLFQ [TCM07]. MLS [HRS^{+}10]. MMU [PLKH08]. MMU-less [PLKH08]. Mo [Coc01]. Mobile [AKL^{+}96, Ano03d, Jon02, SFE10, Sta06, SNF04, YKS^{+}02, Ano03c, Cha11, ISTE08]. mobile-platform [Ano03c]. Mode [AB04, Dik06, TW96]. Model [CPJ^{+}98, Coc01, NR03, Ste00a, ALC20, BDH^{+}09, GCN^{+}10, KS19, SSO^{+}10, UPS^{+}05, ZBG^{+}05]. Model-Based [CPJ^{+}98]. modeling [UPS^{+}07, WJY^{+}07]. Modellbahnsteuerungssoftware [Ano01b]. models [HMK^{+}20]. Modern [Cer03, Tan92, Tan01, Tan08, Gue19]. mods [Kra07]. Modularity [NR03, BTS09].

Module [Cam04, PNK02, SB99, ARJ08]. Modules [Wel95a, BTS09, GJJ05]. monadic [LZ07]. Monatsblatt [Lin02b, Lin02c]. mondo [Cor05b]. Mondrian [WRA05]. Mondrix [WRA05]. Monito [SGL^{+}00]. monitor [ALL06]. Monitoring [GSN^{+}01, Meh00, SGL^{+}00, WR03, HR95, KEP07, WJY^{+}07]. monitors [BDF^{+}03]. monthly [Lin02b, Lin02c]. Moore [Bar01b]. Moskowitz [Koh06]. Most [Kau96, CK06b, Pub11]. MotionBuilder [Ano02c]. Motivation [HNH03]. Motorola [Ano02d]. Mount [Ano02c]. Move [Bar00b, BB99a]. movement [Way00]. Moving [Gag06, Gag07, Ste01b]. Mowgli [AKL^{+}96]. mp3 [Zic01]. MPI [MK04, PSK08, SMTW96, Slo05]. MPI-2 [MK04]. MPI-IO [PSK08]. Multi [MKSB22, SGP^{+}23, XZY^{+}10, CAW08, FD10, GBC^{+}18, LPH^{+}07, RGK99, UPS^{+}05, USC^{+}08]. Multi-Core [SGP^{+}23, XZY^{+}10, CAW08, FD10]. Multi-tenant [MKSB22]. multi-threaded [RGK99]. multi-tier [UPS^{+}05, USC^{+}08]. multi-user [GBC^{+}18]. multi-variable [LPH^{+}07]. Multicast [MRW99]. Multicode [Mud97]. Multicore [PGW^{+}20, LZZ^{+}08]. multihoming [AMS^{+}08, AMSS08]. Multilingual [Mud97]. Multimedia [KLK^{+}02, PNK02, SuS01, Tra96, Ano03d, LL94b, Ran06, SMXH01, YY96]. multiple [Com00, SH13]. multiprocessing [RGK99]. Multiprocessor [WR03, GMTW08, ISTE08, LWV^{+}10, LBH09, MHKT09]. Multiprocessors [LBN^{+}18, SSO^{+}10, TAS07]. multistream [GB07]. multithreaded [XSaJ08]. multitier [UPS^{+}07]. MuPAD [uRH95, uRH96, uRH95, uRH96]. Music [Phi00]. Musings [Ray99c, Ray01b]. MUVI [LPH^{+}07]. MySQL [Gab07, G^{+}06b, LW03, TSC02, Wün13].

Nachschlagewerk [Kuo99]. named [Rob95a]. namespace [WDG^{+}06]. Namespaces [Min01]. Nan [RB08]. Nanotechnogy [Ano03e]. National [Ano02d]. Navigating [Bon01]. Navigator [SC98]. Near [Wol04b]. needs [Lov07]. nell’era [Zic01]. NeOS [BBS96]. Net [Kro99, GB96]. Netlink [NAGL10]. Netscape [SC98]. NetWare [WF03]. Network [FES^{+}99, Hom00, Jan03, JWKZ99, Kir95a, Kir95b, Kir95c, Kro00, LBN^{+}18, Meh00, Nel00, Ser98, Ser00, Ste00a, Ano95c, Ano96a, BDP05, Ben06, BP96, BSR06, BJL^{+}06, BH98, Chi02, DOF96, Hun99, Hun02, Kir94, KD00, LZ07, Man01, Max00, MSZ09, SH13, Smi05a, TR99, Tot99, TK07, WYW^{+}13]. Network-based [Kro00]. Networked [KLK^{+}02, BV06]. Networking [ACM04, CF20, SS05b, SuS01, Alm96, Col02, Mag01b, MB95, Pfa01, Sch07, SG04, Smi00b, Smi02]. Networks [cFJH^{+}03, AOAH14, AOAH15, KS09b, Ano03c]. Netzwerk [KGG00]. Netzwerkadministration [Kuo99]. Netzwerke [WF03]. Netzwerkkonfiguration [WF03]. Netzwerkspiele [Str94a]. neuem [Ano01c]. neural [Ano96a, DOF96]. Newfoundland [IEE06]. News [Bar00b, Bar00c, Bar01b, Coc01, Coc03b, Coc03a, DDJ98b, DDJ98a, Eri97, Gar99, Gla99, Sta04, Sta05a, Sta05b, Sta06, Sta07a, ZS93]. Next [HKS02, Thi04, ISTE08, Mat99a]. Next-Generation [HKS02]. NFS [Zad01]. NIC [MFPP03]. NIC-based [MFPP03]. NIDS [SK10]. NIST [Bar00b, Coc03a]. NixOS [DL08, DLP10]. nm [ZB19]. No [CPJ^{+}98, Jen05, Sta05a, HC15, Ran97, Ran99, Ran00]. NodeOS [SGWP02]. nodes [Len97]. Noise [dOCC23, ALL13]. Non [GSHL03, Gra04, SGHL01, Gra06b, MR07, RGK99]. Non-Dedicated [GSHL03, SGHL01]. Non-Geeks [Gra04, Gra06b]. non-invasive [RGK99]. non-quiescent [MR07]. Nondeterministic [MR18]. Nonvolatile [LBN14]. Noorda [Ano95b]. Noosphere [BV06]. normality [WG08]. Norton [Nor99]. noses [Kau96]. Notebook [Dum05]. Noted [CKM^{+}99]. Notes [DDJ98b]. Nothing [Sta04]. notification [BF08]. nouvelle [Bar99a]. novel [OMCB07]. Novell [WF03]. November [ACM93, ACM99, ACM03a, ACM04, ACM05a, MS05, USE01]. novice [Tho06]. noyau [CDM97, Car98]. NPACI [PKB01, PKB03]. NSA [McC05]. NT [Bur97, MYH00, VE99]. NT/2000 [MYH00]. NTT [PS22]. nuclear [JB95]. nuls [WWH98]. Number [AR22, Don22]. Numerical [DHH96, GZK^{+}01, KEB^{+}99, EP05, LWL06]. Nuts [Ano99f]. Nutshell [Hek97, HT97, Dea01, HNV98, KH07, P^{+}06, SH99b, S^{+}00, SFW03, Sie05, S^{+}05, Sch04]. Nvidia [Ano03e].

O [BJL^{+}06, CKR08, GPK^{+}07, LSP07, PSK08, Rus08, dBB08]. Oakland [USE01]. Oberfläche [Ste00a]. Oberon [WF03]. Object [JGS^{+}08, Ste00a, BJ01, Kro99]. Object-oriented [JGS^{+}08]. objektorientiert [Ste00a]. Observability [CF20, Gre19]. OCE [Sav11]. October [ACM03b, Ano06, IEE93, Pop94, USE96a, USE00]. off [DRL^{+}02, MS07, WSL^{+}99]. off-the-shelf [MS07]. Office [And99, Ano02c, Don98, Ano03c, NW03, Neg06b, Smi00b, Tot99]. Official [Ano97b, Red98, Cha13, Gra05, HBB^{+}06, H^{+}07a, H^{+}07b, SAN08]. offloading [MKO^{+}17]. Oh [Swa00a]. old [Ano94c, Wis03]. Omnicore [Ano02d]. on-the-fly [BGM99, SZD^{+}08a, SZD^{+}08b, SZD^{+}08c]. One [CS95, UJE^{+}22, Dul14]. Online [Sta06, LWV^{+}10]. Only [Bar00b, Daw94]. Ontario [ACM06]. Open [Ano94f, Lin02b, Lin02c, Cha11, CPJ^{+}98, DOS99a, Fra19, Gar09, Gla99, ILG19, JB99, LW03, Lew99a, Lew99b, Lin02a, MSSvK08, MLE^{+}00, PGW^{+}20, Ray99c, Ray99b, San98, SK04, SC98, Sor01, Vad00, ZST^{+}07, AC11, AAB^{+}05, BV06, CM06, Cus04, GHL^{+}04, HP01, Lla06, MM00, McC05, Moh97a, Moo01b, PPSM96, Ray01b, Tim99, Wis03, Wol02, Zic01, Ano95l, Lin02b, DOS99a, HNH03, Lin02a, Moo01a, NR03]. Open-Source [PGW^{+}20, Ray99b, AAB^{+}05, Lla06, Wol02]. OpenBSD [AT05, DIN05, SY00]. OpenLinux [Cal00, SRBB99]. openMosix [Slo05]. OpenMP [KKH03]. Openoffice [Ano01b]. OpenOffice.org [Ano03c]. Opens [Kau96]. OpenSSL [WN15]. OpenSUSE [CK06d, CK06c, CK06i, CK06j, CK06j]. OpenVMS [Wis03]. operacni [WKD97]. Operating [ACM03b, Ano98b, BSSS96, BSS^{+}96, Bok95a, Loz05, O’G01, RRP^{+}08, Sie00, TT01, Tan92, Tan08, Var94, WR03, dOCC23, Ano95c, Ano97a, Ano97b, AAB^{+}05, Ben96, BCP08, Bok95b, CCZ^{+}06, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CK06a, CK06b, Col97, Com00, CLDA07, Don04, FRRJ07, KAR^{+}06, LB96, LAK09, Mac99, MR07, NV06, OMCB07, PBR^{+}08, RRP^{+}07, Red98, RLB08, RÓ01, RHP^{+}07, RRH^{+}08, SPF^{+}07, SAF07, Tan95, Tan01, Van06, Ygg93, Ygg94]. operating-system [AAB^{+}05]. operativo [RÓ01]. Opposition [Ded09, Gla99]. Optimal [GB07, Ste00a]. Optimistic [TT01]. Optimization [Kro99]. Optimizations [HKS^{+}04, UJE^{+}22]. Optimizing [Ano02b, cFJH^{+}03, Fis02, FRRJ07]. Option [Sta04]. Optional [TEGF98]. Oracle [Ano02c, Emm11, Sav11]. Order [SGP^{+}23]. Oregon [ACM99, IEE04]. O’Reilly [Ano95a, Sch04, JW99]. Organization [Asb00]. organizing [LAK09]. Oriented [Kro99, CBC^{+}08a, CBC^{+}08b, CBC^{+}08c, JGS^{+}08, SMS08]. Orlando [Ano94b]. OS/2 [DF00, Ano94c]. OSCAR [Slo05, VS06]. OSDL [Ano02c]. OSes [Den99a, SLQP07]. OSLO [She06]. osobni [WKD97]. other [Neg06a, Neg07a]. Otorhinolaryngological [GVK^{+}03]. Ottawa [ACM06]. Our [HBB13, Sta05a]. Out-of-Order [SGP^{+}23]. overbooking [USR09]. Overdog [CPJ^{+}98]. Overhead [FST^{+}09, GD18]. overlap [PSK08]. Overshadow [CGL^{+}08a, CGL^{+}08b, CGL^{+}08c]. Overview [Bol99]. Own [Pau95, Tay99b, Gib17]. owned [Hou96].

P [AUW08]. P.R.O.S.E [Van06]. PA [ACM04, IEE93]. Pacific [ACM93, Ano03e]. Pack [G^{+}02]. Package [Oom13]. packages [BJM^{+}22, Bra04, Mil96, VD01]. Packaging [Kro99]. Packing [SKP96]. packt [Gün02]. page [ST03]. Pager [GGZ02]. pages [Ano95a]. Pagination [GGZ02]. pain [Gra06b]. pain-free [Gra06b]. Pajamas [Lew99a]. Pan [ACM93]. paquetes [VD01]. para [MR99, RÓ01, VD01, DW94]. ParADE [KKH03]. Paradigms [Swa00a]. Parallel [DW94, MAT08, TT01, CC97, GCN^{+}10, LWL06]. parallelism [XSaJ08]. Parametric [Cop00]. parity [AR07]. Part [BSSS96, BSS^{+}96, She03]. parted [SuS01]. partitioned [Van06]. partitioning [BJL^{+}06]. Partnership [Gar99]. Party [Loz05]. Pascal [BJ01]. passing [PS07]. Patch [HLT^{+}21, Jan06b]. Patching [CCVLW23]. PatchNet [HLT^{+}21]. Patenting [Coc01]. Patents [FES^{+}99]. path [DDA08]. path-sensitive [DDA08]. Patterns [GS00]. Paul [Teo13]. PBS [FMR07]. PC [SU94c, Str94a, SU94a, SU95b, SE96, SU96d, Ano03b, Ano03d, Cha99, EJMT94, Kad95c, Kad95a, KEB^{+}99, LR97, LWL06, PV02, SU94b, Str94b, SU94d, SU96a, SU96c, SE97, TD96, Tem96, TF96, TSEK01, Tro96b]. PC-Clusters [TSEK01]. PC-IRAF [TF96].

PC-Pentium [LR97]. PC/Linux [KEB^{+}99, TD96, Tem96]. PCIXCC [AV04]. pClock [GMV07]. PCs [SuS01, Hou96]. PDEs [MAT08]. PDF [DF00]. PDF-Dateien [DF00]. Penetration [Ano14]. Penguin [Nis01a]. Penguins [Ana99]. Pentium [LB96, LR97, RGK99]. People [ASW^{+}01, LH01, IEE93]. PEP [GB96]. perceived [ON06]. Perfect [FES^{+}99, Le 96]. Perforce [Ano03e]. Performance [ACM04, AT05, CF20, DRN^{+}23, DRL^{+}02, Gig06, Gre19, IEE06, KBPV99, LWL06, Lum01, MK04, MLE^{+}00, PKB01, SK10, SH13, SSSS97, WSL^{+}99, WR03, ZB19, ZYCZ10, AWW08, AMS^{+}08, AMSS08, AJ05, Ano94b, Ano03d, Bes05, CKE^{+}07, FSW02, FM02, FD10, FST^{+}09, KEP07, LB96, Lla06, NTW07, PBH^{+}07, RGK99, SBRM09, Slo05, SMXH01, SPF^{+}07, SKP^{+}02, TB05]. Performance-aware [ZYCZ10]. performance-monitoring-unit-based [KEP07]. performance-setting [FM02]. Performing [Ano03e, AC11]. Periodic [LD05]. Perl [DF00, Bla02, Bla04, DDJ98a, JW99, LW03, Pau95]. Persistent [GZXZ23]. Persistent-memory-related [GZXZ23]. Personal [SSSS97, YKS^{+}02, Tro96b]. Perspective [Nor02, WYW^{+}13]. Perspectives [Cop00]. Persuasive [CKM^{+}99]. Perturbing [EP05]. Peter [Nor99]. Petri [GB96]. Petroleum [Ano94d]. Philadelphia [IEE93]. philosophies [Cor05b]. Philosophy [Gan03]. Phoenix [ACM03a]. photography [AJD^{+}10]. PHP [G^{+}06b, LW03, WN15]. PHP5 [Gab07]. Phrasebook [Gra06a]. Physics [BSSS96]. Phyton [DF00]. PI [Gib17]. Picking [CPJ^{+}98]. Pining [FES^{+}99]. pipes [Rob95a]. PipesFS [dBB08]. Pittsburgh [ACM04]. PKI [Ahm08]. PL8 [GHL^{+}04]. placement [GJJ05, NLK09]. Plain [VRFJ97a]. Plan [Gan08]. Planning [CPJ^{+}98, FSW02].

Plate [Cha99]. Plate-forme [Cha99]. Platform [Coc03a, HKS02, KIKT01, Lac06, OBSB23, RP02, TSEK01, ARJ08, AJD^{+}10, AC11, Ano03b, Ano03c, AUW08, BJM^{+}22, BS05, Bro03, CC97, Don04, ISTE08, USR09, Wol02, YL08, Lac06, Ste01g]. Platforms [Ano06, IZG^{+}07, NV06, PBM^{+}22]. Play [Ano95h, LH01, Ano93, Ano94e, Ano94c]. PLD [Ano03b]. Plug [Ano93, Ano94e, Ano95h, DF00, Jen02, OBSB23, Ano94c]. Plug-and-play [Ano94e]. Plug-In [Jen02, OBSB23]. Plug-ins [DF00]. pocitace [WKD97]. Pocket [Fri03, Bar04, Bri05a, Pur04, Ran05]. point [Mat99a]. pointers [Moh97b]. points [LLA07, SLP^{+}09]. points-to [LLA07]. Policies [Oom13, CAW08, IZG^{+}07, RKSV20]. Policy [Mar00, Mar01, Mar06, HRS^{+}10, YW07]. portability [RS14]. Portal [But02b, Kro00]. Portals [McN99]. ported [Ano95e]. Porting [Mon00b, MSW06]. Portland [ACM99, IEE04]. ports [Gar98]. POSIX [Bla99, CS01]. PostgreSQL [Dup05, Wei04]. Potato [Ano01a, Kle00]. pour [Cor00, WWH98]. Pouzivame [WKD97]. Power [Daw94, G^{+}02, PFH97, Swa98, Gri14, HH04, K^{+}06, MKO^{+}17, PF96, PF97, Smi03, TMG01, VTS^{+}23, WJY^{+}07, Mor97]. Powered [SFE10, Sta04]. powerful [Hou96, Moh97a]. Powerpack [Gan04]. PowerPC [RFS05, RFS06]. powłoki [RB06b]. PPC [ASW^{+}01]. Practical [BC08, Gia99, Kal16, Kha02, Lat02, Loz05, Sob97, Sob03, Sta07b, TK07, Hal06, Hal07, LLA07, Men12, MHKT09, RPB^{+}09, Sob06, Wün13]. Practice [Bol99, SAL19]. practices [Sul02]. prakticky [Bra96]. praktischer [SU94c, SU95b, SE96]. pratique [LR97, Tau96]. Praxisführer [Bor01]. pre [Ano01c]. precision [PV02]. Predictive [SGHL01]. PREEMPT [dOdO16]. PREEMPT RT [RMF19]. prefetching [GB07, LSP07]. Prentice [Hid04, Sta07b, Tho05]. Preparation [Mag04]. Prepared [ASW^{+}01]. Prerelease [Lev95]. Press [Jen05]. Presto [FES^{+}99]. preview [Ano95c]. Primer [Sta07b, Hal06, Hal07, Man01, RFS05, RFS06]. primitives [Ano03d, LZ07, NLK09]. principle [LAK09]. Principles [ACM03b]. Print [TECB03, Jan03]. Printing [GGZ02]. priority [FMR07]. priority-based [FMR07]. privacy [Mag01c]. Private [Min01, FFPC22]. pro [Pod97, Pur95, WKD97]. Problem [Bar00b, War00, WB06]. Problems [YK15, FD10]. Proceedings [ACM93, ACM04, Ano94a, Ano94f, KD96, USE96b, USE00, USE01, ACM03b, ACM06, DW94, MS05, SS05b, ACM05a, ACM05b, Ano06, IEE04, IEE06]. Process [GS00, GGZ02, Kro99, MZG14, LHB24, LF95, Wil98]. Process-Oriented [Kro99]. processes [IEE93]. Processing [HKS^{+}04, ZB19, BJL^{+}06, GGZ02, Wün13]. Processor [SGP^{+}23, XZY^{+}10, BJL^{+}06]. Processors [Neu06]. Producing [Coc01]. Product [Ano94g, Ano03b]. Products [Ano02c, Ano02d, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Kro00, IEE93]. Professional [BBD^{+}00, G^{+}06b, Hid04, Mag04, MS^{+}00, Mau08, Sto00, SuS01, IEE93, Tho06, Ano95i, Ano02d, Gri01b]. Professional/Administrator [Mag04]. professionellen [DF00]. professionelles [SuS01]. Profile [UJE^{+}22]. Profile-Guided [UJE^{+}22]. profiling [KEP07, USR09]. Program [Coo09a, Coo09b, Yor99, Bra04, BDH^{+}09, Gar09, Mil01]. Programmation [CDM97, Car98, WWS98]. Programme [Str94a, SuS01]. Programmer [DDJ98b, DDJ98a, GS00, Gri01b, Tay99b, Woe02, Woe03, Lov07, Pet98a, Pet00a, SSO^{+}10]. Programmieren [Ste00a, Jor04]. Programmierung [Jor04, RB06a, May01]. Programming [ASW^{+}01, Ano03e, Böl01, CF20, CS01, DDJ98b, Eri95, Gay00, Goe00, Jor04, KKH03, LH01, LO97, MS99, Nel00, Pin02, QH01, Rob04, Ste00b, Ste01c, Ste01d, Ste01b, Ste01e, Ste01f, Ste01g, Swa00a, Tan00, Tem96, VRFJ97c, Wal00, Weh01, Woe02, Wri00, BFS07, CBC^{+}08a, CBC^{+}08b, CBC^{+}08c, Dun00, Gay99, Gri01a, Gue19, JPS00, JB99, Jon05, Ker10, Kha02, Lin00, Lov07, MSB^{+}96, MS^{+}00, May01, MOS01, Olc01, Pet01a, Sto00, TSC02, VJR96, WWW99, Wal01a, Wal01c, War01, Gri01b]. Programmpakete [Bra04]. Programowanie [RB06b]. Programs [MR18, CS06, CK06b, EP05, XSaJ08]. Project [EM00, Kro00, Gra06b, Tay99a, AUW08, But02b]. project-based [Gra06b]. projects [CIC13, Gri13, Gri14, HNH03, NW03, Neg06b, Nut01]. Projekte [Ric14]. propagation [RGGL^{+}09]. properties [Mau96]. propriété [Cor00]. Pros [Sie00]. Prosecutions [Coc03b]. Protect [LCHL21]. Protecting [BBK^{+}16, TR99]. protection [CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, WRA05]. protezione [Zic01]. Protocol [MRW99, BWH^{+}07, CP01, CP02, FST^{+}09, PC01]. protocols [BP96, SMS08]. ProtoVIew [Kro99]. Provably [GNDB16]. provide [SLQP07]. Provides [Sta04]. Providing [BS05]. provisioning [USC^{+}08]. Proxy [Reh99]. Prozess [GGZ02]. Prozess-Systems [GGZ02]. Prüfungsvorbereitung [Mag04]. pruvodce [Bra96]. psad [Ras07]. pseudo [WN15]. PTG [Kom02]. PTR [Sta07b, Tho05]. Public [Bar01b]. Pula [KD96]. purely [DL08, DLP10]. purging [BC08]. Push [Eri97, Gar98]. Put [Den99b]. PVM [SGL^{+}00, SSSS97, Zem94]. PXROS [Mau96]. PYMES [RÓ01]. Python [Lac06, Ric14, Bla02, Bla04, DHH96, GJ08, Lac06, TSC02]. Python-based [Lac06].

Q&A [Jen02]. Qemu [CK06a, CK06b, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i, CK06j, CK06k, MZG14, CK06a, CK06b, CK06d, CK06g, CK06h, CK06k]. QoS [GMV07, HKP09, IZG^{+}07]. QT [Gri01a, DH02, War01]. Qt/Embedded [DH02]. quality [KS09b]. Quantum [Coc01]. Quasistatic [PLKH08]. Quelltext [DF00]. quels [Cor00]. Quemu [CK06i]. Quest [Ano03c]. questionable [Lev95]. Questions [CPJ^{+}98, Pub11]. quick [Dea01, Hek97, Hug97, Hug98, SH99b, S^{+}00, S^{+}05]. quiescent [MR07].

R [Oom13, Wün13]. R&D [BYVF08]. r0a [Ano05b]. R5 [RMZ01]. rabbits [SJP^{+}09]. race [PFH06]. races [JYX^{+}16]. Rack [Ano02c]. Rack-Mount [Ano02c]. RAD [Ano03f]. Rage [Ste01e]. RAID [BV06, Vad00, Vad02]. Raiding [BV06]. Raksha [DKK07]. Random [AR22, Don22, WN15]. Ranges [Gar99]. Ransom [Sta05b]. Ransom-ware [Sta05b]. rapid [LP10]. RAppArmor [Oom13]. Rating [Nis01b]. Ray [Ano95b]. RCU [KS19, MFBWW20]. Re [WSL^{+}99]. read [GMTW08, KS19]. read-copy [KS19]. read-copy-update [GMTW08]. readahead [WXX08]. Reading [Eri97]. Ready [PS22, ZB19]. Real [Abb03, Dou02, LW00, LD05, Mar04, Nis01b, PNK02, RMF19, SAL19, She03, Sta05a, Sta07b, Swa98, TEGF98, WSL^{+}99, WM00, ABB19, BP96, GMTW08, Hal06, Hal07, HST08, HR95, LLA07, PBM^{+}22, Rau01, RLB08, SJP^{+}09, Epp00b, HST08, Nis03, She03]. Real-Time [Abb03, Dou02, LW00, Mar04, PNK02, RMF19, TEGF98, WM00, LD05, SAL19, She03, ABB19, BP96, GMTW08, HR95, PBM^{+}22, RLB08, SJP^{+}09, Epp00b, Nis03, She03, HST08]. Real-world [Sta07b, Hal06, Hal07]. Really [DDJ98a, FES^{+}99, KMP^{+}20]. realtime [Mau96]. Rebel [Moo01a, Moo01b]. Recipe [Moe02]. recompilation [EP05]. Reconfigurable [LS01, Don04]. reconfiguration [PBM^{+}22]. Record [Ano03e]. Record-Performing [Ano03e]. recoverable [LAK09]. Recovery [LAK09, Pre07]. recursos [MR99]. Red [Jan04, Men12, Pri99, LW00, Wan02, Ano03a, Bar98, Bar99b, Bar01c, BW02, Bel00b, Bon99, Col02, Dan99b, Dos99b, Dun99, G^{+}06a, Gra99, HS99, HS00, HH05b, HP96b, Kab99, Kom00, KC00, Kom02, Kro99, Max00, McC99b, McC02, McC03, McC04, Neg01, Neg02, Pet02b, Pit98a, Pit98b, PB98a, PB98b, Ran99, SBR01, SH99a, Ser00, Sob03, Sob06, Tom11, Hag00]. RED-Linux [LW00, Wan02]. Redhat [Ano97a, Ano97b, Red98]. reducing [ST03]. Reduction [KLK^{+}02, MFPP03, CDD^{+}07, PLKH08]. redundant [Gar09]. Reference [Ano02d, Fri03, GGZ02, Hid04, Pet98b, Pet00b, Dea01, Emm11, Gar00, Haw00, Haw02, Hek97, Hug97, Hug98, Kie98, MSW06, Pet96, Pet98a, Pet99, Pet00a, Pet01b, Pet02a, Pet02b, Pet00c, Pfa00, Pur97, Pur04, Ran05, SH99b, S^{+}00, S^{+}05, HNV98]. Referenz [GGZ02]. Reflections [BTS09, Cus04]. Regarding [CKM^{+}99]. Rehearsal [Ahm08]. Rejected [Eri97]. related [GZXZ23, LPH^{+}07]. Relay [Coo98]. Release [Ano02c, RR15a, Ano01a]. Released [Bar00b, Coc01, Coc03b, Coc03a, DDJ98a]. Releases [Ano03e, Den99a, Ano03c]. Reliability [Vad00, GPK^{+}07]. Reliable [MRW99, Van06]. Remote [BCP08, MKO^{+}17]. rendszerek [Lás05]. Repeats [Bou95, Bou96a, Bou96b]. Replay [MR18, LWV^{+}10, MHKT09]. replication [ALC20]. Report [Ano02a, Ano05a, Bla99, EM00, Far01, BFS07]. reporting [HRD^{+}07]. request [SZD^{+}08a, SZD^{+}08b, SZD^{+}08c]. required [Cas17]. Rescue [ASW^{+}01, SLP^{+}09]. Research [HHD^{+}16, KIKT01, Pop94]. réseau [KC95]. Reservoir [EJMT94]. Resource [BMF15, USR09, Bru02, CKR08, FT95, HH04, Inf95, MS02, Moh98, VFMM08]. Resources [Cas17, DM19, Gib17, FOP06, Gol95]. Respec [LWV^{+}10]. respecting [YL08]. response [ON06, Ras07]. Ressource [Ano01a, Ano01c, Ano01b]. ressources [Par99]. restart [BBHL08]. Restricted [SB99]. results [HR95]. Rethink [NVCF08]. retrofitting [CGL^{+}08a, CGL^{+}08b, CGL^{+}08c]. retuschieren [DF00]. Reuse [Hub02]. Reverse [HHD^{+}16]. Review [Ano95a, Cha13, Gra99, Hid04, Hul05, Jen05, Kad95c, Kad95b, Koh06, Men12, Sch04, Sta07b, Teo13, Tho05, Wer10]. Revolution [DOS99a, Moo01a, Moo01b]. Revolutionary [Ray99c, TD01, Ray01b]. RH302 [Jan04]. RHCE [Jan04, Tom11]. RHCSA [Tom11]. RHEL [P^{+}08]. RHIC [Thr01]. Rich [Lac06, Lac06]. richtig [DF00]. rigorous [SSO^{+}10]. Ring [MYH00]. RISC [DRN^{+}23, ILG19, OBSB23, PS22, PGW^{+}20, SGP^{+}23, SNM22, VTS^{+}23, ZB19]. RISC-V [DRN^{+}23, ILG19, OBSB23, PS22, PGW^{+}20, SGP^{+}23, SNM22, VTS^{+}23, ZB19]. RISCOS [WW01b]. RMC [MRW99]. Road [Ste01e]. roadmap [Jan03]. Robert [Tho05]. robin [LBH09]. Robot [Coc03a, Gri14]. robot-building [Gri14]. robotic [Gri13]. Robotics [Bar00b, Gri14]. Rocks [PKB01, PKB03, Slo05]. Rogue [Ano03d]. ROM [Ale00, Kom02, OR96]. ROMs [Ano01c, HS99]. Rootkit [LCHL21]. rootkits [WG08]. rootless [GBC^{+}18]. round [LBH09]. round-robin [LBH09]. route [AMS^{+}08, AMSS08]. Router [DRL^{+}02, Fel96b]. routers [Man01]. Routing [Mar00, Mar06, Ano94c, Mar01]. RPM [Bai97, McC07]. RPM-based [McC07]. RT [dOdO16, LS01, TEGF98]. RT-Linux [LS01, TEGF98]. Ruby [Bla02, Bla04, Tom02]. Ruling [Eri97]. Run [Sun99, Swe07, Com00]. Running [DW06, Le 96, OR96, SNF04, WKDD03, WK95, WK96, WDK99, WKDD02, Ano95a]. runs [Hou96, SAN08]. Runtime [BMF15, WJY^{+}07]. rute [She02].

s [WKD97]. S.u.S.E [KGG00]. safe [CLDA07, MSZ09]. safety [Gar09, RLPA18]. Sagamore [ACM03b]. Sair [Mag04, DRP01, Jan01, Mag00, Mag01a, Mag01b, Mag01c, Mag04, Sul02]. sale [Mat99a]. Salm [Hid04]. SALT [Ano03c]. SALT-based [Ano03c]. Samba [ECBK00, Hun00b, MS02, Smi01, TECB03, Wol04a]. Sameer [Wer10]. sampler [FT95]. Sams [BS98, Bal99b, Gay99, Hsi01, LH00, Ray99a, RBHB99, SBR01, Tay03, WT99, Wel99a]. San [ACM93, USE96b]. Sandboxing [Oom13]. Sarge [Ano05b]. Satan [RR95]. Satellites [Bar01b]. Savvy [Kro99]. Say [Sta05a]. Says [Coc03a]. SC2003 [ACM03a]. SC’99 [ACM99]. Scalability [She06]. Scalable [MFPP03, WR03, AAB^{+}05, DDA08, LZ07, LBH09, Mau96, SPF^{+}07, Bar00b]. Scale [GVK^{+}03, MFPP03, AUW08, DFC^{+}23]. Scaling [Neu06, Spe00b]. Scan [Bar01b, GFRD96]. ScanSoft [Ano03d]. Scattering [PQM^{+}21]. Scheduler [XZY^{+}10, FMR07, TCM07, WTKW08, XSaJ08]. Scheduling [GW94, LW00, LD05, ABB19, CAW08, CKE^{+}07, CKR08, LP10, LSAF16, LBH09, MB08, Sto07, TAS07, TK07, ZYCZ10]. Scheme [DF00, KLK^{+}02, DJC07, SKP96]. Schritt [Str94a]. SCI [LS01, TSEK01]. SCI-Based [TSEK01]. SCI-Cluster [LS01]. Science [Bar01b]. Scientific [TD96, Tem96, DW94, Spe00b, Wol02]. SCM [Ano03e]. SCMFS [WQR13]. SCO [DF00, Kuo99]. SCO-Unix [Kuo99]. SCOPE [Gar09]. scratch [Bee00, Bee16]. Script [RB08]. Scripting [Bur04, RB05a, Blu08, Blu11]. Scripts [Bla04, BJM^{+}22, Bla02, RB05b]. SD [AOAH15]. SDE [Ano02d]. SDK [Jon02, Kro99]. Seattle [ACM05a]. Sebastopol [Ano95a]. seccomp [XCS^{+}22]. Second [ACM06, Hul05]. seconds [Bri05b, PBR^{+}08]. Secret [Dat14, MYH00]. secretos [Bar97]. Secrets [Ray98, Bar96a, Bar96b, Bar98, Bar99a, Bar99b, Bar01c, HLK01, HC15]. Secure [BBK^{+}16, Bau02, BBB^{+}20, CLDA07, Lin02a, GNDB16, ISTE08, RKSV20]. Securing [Bea04, BR^{+}00, C^{+}03, Loz05, SFE10]. Security [AHSM21, Ano14, Bra04, C^{+}03, Chu03, DW08, Hsi01, Kal16, Lam99, Lin02a, Loz05, MS05, Oom13, SuS01, TB05, BSB03, BsswLB05, BTS09, DKK07, Flo99, FOP06, GJJ05, HLK01, Mag01c, MM00, PL06, Rau01, Ray01a, SCB^{+}13, Shu06, Smi05a, TR99, Tox07, GHRS05, Lin02a, MMC06, McC05]. Security-Enhanced [GHRS05]. SecVisor [SLQP07]. Sed [Bla02, Bla04]. See [Swe07]. SEED [DW08]. seinen [KGG00]. Selecting [BSSS96, BSS^{+}96, Bro00]. Self [WB06, Han97, SLP^{+}09]. self-contained [Han97]. self-healing [SLP^{+}09]. Self-service [WB06]. SELINUX [McC05, CGDB24, HRS^{+}10, MMC06, RKSV20, SFE10]. semantic [LPH^{+}07]. semantics [WDG^{+}06]. semi [MSZ09]. semi-automatic [MSZ09]. Semiconductor [Ano02d]. Seminar [Kom02, RGRG05]. sense [PFHLN22]. sensing [WJY^{+}07]. sensitive [DDA08, LLA07, PFH06]. sensors [ALL06]. September [Ano94f, USE96a]. sequences [WN15]. sequential [GB07, LSP07]. serious [Kra07]. Server [Ano02c, Blu00, Bra04, CS01, Don96b, Fel96a, Gag00b, Hun00a, Hun00b, Hun01b, Kro99, SWL^{+}01, TECB03, Vad00, VRFJ97b, Aul01, BsswLB05, FWS^{+}03, Fli03, HE95, Jan03, Kab99, NTW07, PBH^{+}07, SR96, SK01, Smi01, TR99, VJ06, Wel98, WLS^{+}01, AJ05, Rul07, SK10, SC98].

Server-Lösungen [Bra04]. Servers [Ano03e, Bau02, CS95, Nor02, BBK^{+}16, BBHL08, Hun99, Hun02, Sai01, TB05, Pon05]. Service [Coc01, KLK^{+}02, WB06]. Services [ACM93, Ano02c, Roe02, Ano03b, Ano03d, LZ07, UPS^{+}05]. Set [DDJ98b]. Seth [Wer10]. Setting [Mat99b, FM02]. Settled [Eri97]. SEU [Nis01c]. seven [Hun99, Tox07]. seventh [Cha13]. SGI [Ano03e, MLE^{+}00]. Shōkai [RB06c]. Shall [Sta04]. shape [CDOY09]. share [CAW08]. shared [GCN^{+}10, GB07, GMV07, GMTW08, PLKH08, USR09]. shared-memory [GMTW08]. shareware [Kau96]. Sharing [SW98, MS02, TAS07]. sharing-aware [TAS07]. Shelf [DRL^{+}02, MS07]. Shell [Bur04, May01, QH01, RB05a, RB06a, Tan00, Blu08, Blu11, Olc01, RB05b, Bla02, Bla04, May01, RB08]. Shell-Programmierung [RB06a, May01]. Shells [Qui02, Qui00]. shepherding [GPK^{+}07]. shieru [RB06c]. Ships [Ano02c, Ano03d]. shops [Kau96]. Short [DR01, AOAH15]. short-distance [AOAH15]. Shortage [DDJ98b, DDJ98a]. Shortchanging [Bar01b]. should [HBB13, War04, War15]. Showcase [Ano00a, USE00, USE01, Den99b, Ano99a, Ano99b]. sichere [Lin02a]. Sicherheit [Bra04]. Siever [Sch04]. Signal [She03, Gar09]. Signaling [FST^{+}09]. signatures [SZD^{+}08a, SZD^{+}08b, SZD^{+}08c]. Signs [Eri97]. SiLA [LHB24]. Silicon [Ano02d, GZK^{+}01]. Silk [SGWP02]. similarity [SWF^{+}22]. Simulation [GZK^{+}01, KEB^{+}99, TT01, LLT07, LWL06, MMW96, SvBC^{+}04]. simulator [Bro03]. Simulators [LBN^{+}18, Ano96a, DOF96]. Singapore [Ano06]. Single [CCO^{+}05]. Single-Computer [CCO^{+}05]. sins [Tox07]. SIP [AWW08, NTW07]. sistema [Bla97, RÓ01, VD01]. Site [Pau95, Le 96]. Situ [RGK99]. Size [MAP^{+}22]. skills [Gri14]. skryptów [RB06b]. sky [Pur95].

SL-A300 [YKS^{+}02]. SLACKWARE [Ano96b, Ano95i, LCJH05, P^{+}97, Vol97, Wal95]. SLACKWARE-Ver [Ano96b]. Sleep [BLH20]. Sleep-in-atomic-context [BLH20]. Slippery [SRTH15]. Slope [SRTH15]. Slow [CPJ^{+}98]. Small [AMM^{+}18, Kro00, Lat02, SH99a, SAN08, DH02]. Smaller [DDJ98a, DDJ98a]. SMP [KKH03, TAS07]. SMT [TAS07]. Snort [SK10]. Snow [SGWP02]. Sobell [Men12]. SoC [VTS^{+}23]. Society [CPJ^{+}98]. Socket [But02a, Gay00, Wal01c]. Sockets [CS01, NAGL10]. SoCs [PGW^{+}20]. Software [Ano95j, Lin02b, Lin02c, Ano02d, CPJ^{+}98, Coc01, Ded09, EJMT94, FES^{+}99, Gra99, Hol02, Jor04, Kro99, LZZ^{+}08, LO97, NR03, Pom04, San98, Ste01f, Sun99, Tro96a, YK15, ARJ08, Ano03c, BJM^{+}22, Cus04, DKK07, Don04, FRRJ07, HRD^{+}07, HNH03, KEP07, Kop05, LHB24, Mil96, MHKT09, SLP^{+}09, Way00, WJY^{+}07, YL08, Zic01, Lin02b, Ano02c, Ano03c, Ano03e, CK06k, Kro99]. software-hardware [MHKT09]. software-profiling [KEP07]. Softwaretools [Jor04]. SOHO [Blu00]. Solaris [DF00, WF03, BDD^{+}06, JW99, Kuo99, Loz05, MM06, Sun99, Wil98]. Solution [Hom00, LWL06, Tot99]. Solutions [Bra04, Chu03, Coo09b, Coo09d, Gig06, Koh06, Tro96a, YK15, HLK01, MAT08, MB05, TMG01, Ano02d]. Solved [Bar00b]. Solver [War00]. Some [Cas17]. Son [TD97]. Sophisticated [Kro00, Lum13]. SOSP [ACM03b]. Sound [DDA08, Phi00, Str94a]. Sounding [WSL^{+}99]. Source [Lin02c, Bon01, CPJ^{+}98, Gla99, ILG19, Lew99a, Lew99b, MSSvK08, MLE^{+}00, PGW^{+}20, Ray99c, Ray99b, San98, SK04, Sor01, Vad00, AC11, AAB^{+}05, CM06, Cha11, FRRJ07, Fra19, Gar09, HP01, JB99, LW03, Lla06, MM00, McC05, Moo01b, Ray01b, SNM22, Tim99, Wol02, Zic01, Lin02b, Lin02c, DOS99a, HNH03, Lin02a, Moo01a, NR03]. Sources [DOS99a, Ano94c]. sous [Bla02, Bla04, Dup05, Hom00, KC95, TD97]. Space [Coc03a, Nis01a, Nis01b, Nis01c, Nis03, SW98, NAGL10]. Spanish [RÓ01, VD01]. SPARC [Wil98]. Sparc64 [ST03]. spatial [DJC07]. Speaking [Bar00c]. Spec [Bar00b, DDJ98a]. Special [BN01, SRBB99, TGB95a, TG96, Mau96, FOP06]. specialization [BCP08, CDD^{+}05]. Specific [OBSB23, LHB24]. Specification [Coc03b, Coc03a, ISO05, BFN^{+}06, HRS^{+}10]. Specifying [CGDB24]. specimen [JB95]. Specs [Nis01a]. speculation [LWV^{+}10]. Speech [Eri97, Ram96]. speed [Alm96, AOAH14, AOAH15, HRX08]. Spell [Bar00c]. Spielesammlung [CK06k]. split [Ano94c]. spoofed [WJS07]. Spotlight [Ano95j]. Spotlighting [Ano06]. Spy [Bar01b]. Squeeze [Sta04]. St [IEE06]. Stabilität [Bra04]. Stability [Bra04]. Stabilization [RR15a]. Stable [HLT^{+}21]. stack [SWF^{+}22]. stage [JB95]. staging [Cou17]. Standard [Bla99, Coc01, ISO05, Yeo05, GHL^{+}04, KYB^{+}07, Rus08, DDJ98a]. Standardization [Bar01b, RS14]. Standards [Gar99]. Starch [Jen05]. StarOffice [Mea99, WT99]. Start [Han97]. Start-Up [Han97]. Started [Wt95b, Ric13, Wel94, Wel95b, WBS01]. Starter [Dul06, Gay99]. Startup [PS07]. State [Moh97b, Roe02]. Stateless [KS19, SMS08]. Statically [CCVLW23]. statistical [WN15]. Status [EM00]. stay [Oua97b]. Stefan [Kad95c, Kad95b]. stegomalware [CMR^{+}21]. Step [BR^{+}00, Dun00, Mil01]. step-by-step [Dun00]. Stephen [Sch04]. Steps [Sta05b]. Steuerung [GGZ02]. Stone [DDJ98b]. Storage [ACM04, GR09, Gig06, WQR13, GMV07, HKP09]. storia [Cor05b]. Stories [WSL^{+}99, Cor05b]. Story [TD01, Moo01b, Wal01b]. strap [Reg11]. strategic [Gon99]. strategies [PSK08]. strategy [Cus04]. strength [Fli03, Ran04]. StrikeIron [Ano03b]. Strobel [Kad95c, Kad95b]. structured [Gat01, JHT^{+}07, KAS^{+}06]. Study [cFJH^{+}03, Kie99b, LS01, LADADL14, PQM^{+}21, RB08, RÓ01, AOAH14, BTS09, CIC13, DIN05, FFPC22, FST^{+}09, GAH16, Jan04, Kor06, MW08, NH14, PSK08, PBR^{+}08, SG04, SMXH01, Yu07]. Studying [Coc03b]. Subsumes [LBN14]. Subsystem [IBM05]. subsystems [HKP09, MR07]. Success [Ray99b, WSL^{+}99]. Successful [Moe02]. Such [CPJ^{+}98]. Suchen [Gün02]. Suit [Moh97a]. Suite [DW08, Kro99, Ano03c, PL06, Ano02c]. sukuriputo [RB06c]. Summit [Far01, Ano02a, Ano05a]. Sun [Gar98, Kro99]. SunOS [Kuo99]. Super [Bar00c, EEK^{+}04, ST03, War04]. super-user [War04]. Supercomputer [Bar01b]. Supercomputing [ACM04, ACM05a, Mes00]. superuser [War15]. Support [Bra04, FES^{+}99, Gar99, MK04, Sta04, TT01, TEGF98, Ano03d, BS05, BV06, Don04, HHAW19, PBM^{+}22, RC98, SJP^{+}09]. Supporting [DRN^{+}23, LW00, MB95, AOAH15, GMTW08]. Survey [GD18, RMF19, HNH03]. Survival [C^{+}03, Par96, Par00]. SUSE [DWV06, Neg06a, Neg07a, Bal00, Bro06, EDK00, SK01]. SVGAlib [Lin00]. SVR [Kuo99]. SVR4 [Luc00a]. SVR4/Linux [Luc00a]. Swaine [Swa98]. Swan [Swa00b]. Swarm [Coc03a]. SWIFI [SB02]. Switches [DRL^{+}02, CS06]. Switching [Ron05a, SMM^{+}09]. sword [RR95]. Sybase [Ano03d, Ano03f]. Sybex [Koh06]. symbolic [BFN^{+}06]. symbolic-evaluation [BFN^{+}06]. symmetric [RGK99]. Symposium [ACM03b, Ano94a, IEE04, IEE06]. sync [NVCF08]. synchronization [NLK09]. Sysadmin [Tay99b]. System [Ano98b, Ano03e, BSSS96, BSS^{+}96, Bok95a, Bra00, BDD^{+}06, Cam04, CDD00, CDD^{+}05, Fri03, Gag01, G^{+}06a, GVK^{+}03, Gre19, GGZ02, uRH95, uRH96, Hun00c, ILG19, KC98, Kro99, Kuo99, Loz05, LADADL14, MLE^{+}00, O’K00b, P^{+}08, RRP^{+}08, Sie00, TT01, Var94, WQR13, ZST^{+}07, dOCC23, AT07, Ano95c, Ano97a, Ano97b, AAB^{+}05, AR07, BGM99, BKB^{+}18, Ben96, Bok95b, Car99, CK06b, Col97, Col02, CKR08, Don04, EEK^{+}04, FRRJ07, FMK^{+}07, Gat01, Gia99, GJ08, Gue19, GPK^{+}07, HR95, Hei99, JHT^{+}07, Kar16, Ker10, K^{+}06, Kom96, KAS^{+}06, Kra05, KAR^{+}06, LL94b, Lov07, Mac99, MR07, MM00, NH14, NSHW10, Par96, Par00, PPSM96, PBR^{+}08, Pon05, RRP^{+}07, Red98, RÓ01, RHP^{+}07, RRH^{+}08, SMXH01, SPF^{+}07, SS01, SS02, SAF07, Tay03, VD01, Van06, VR00, WC22, Wel99b]. system [WKD97, XCS^{+}22, Ygg93, Ygg94, YW07, GR09, O’K00b]. System- [Kuo99]. system-driven [FRRJ07]. System-wide [CDD^{+}05]. Systeme [Gün02, Jor04, WF03, WCC95, WKC^{+}97, CDM97, Car98, Cor00, Jor04]. Systemen [Kuo99]. Systemprogrammierung [Her99]. Systems [ACM03b, Cam04, Coc03a, DDJ98a, FES^{+}99, Gag00a, GGZ02, HKS^{+}04, Jor04, KKH03, Nor02, Pop94, SRTH15, Tan92, Tan08, TD96, Tem96, TEGF98, USE96a, WR03, Woe02, Yag03, ZST^{+}07, Tho05, ARJ08, ALC20, Alm96, AC11, Ano94b, Ano94f, Bar01a, BK13, BP96, BCP08, BBS96, CCZ^{+}06, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, CK06a, Com00, CBC^{+}08a, CBC^{+}08b, CBC^{+}08c, CLDA07, CP01, CP02, DH02, Don04, FD10, GCN^{+}10, GAH16, Gue19, GMV07, GMTW08, HMK^{+}20, Jan06b, Kha02, LB96, LLT07, LAK09, LZZ^{+}08, Lla06, Luc05, McC07, NV06, O’G01, OMCB07, PLKH08, PC01, PL06, Pon05, RLB08, RGGL^{+}09, Sto07, Sut02, Tan95, Tan01, Wis03, Kro99]. systems-on-chip [Don04]. tackle [Wol02]. Taipei [SS05b]. Taiwan [SS05b]. taming [Ran06]. Target [Bar00b, Ste01b, AC11]. Task [MB08, ZYCZ10]. Tasking [SB99]. Tasks [RB06a, ABB19, LP10]. tcl [DF00, Bla04, Bla02]. TCP [Str94a, ALP^{+}21, AOAH14, AOAH15, BSR06, BFN^{+}06, CS01, CP01, CP02, HRX08, PC01, WW02, Wer10]. TCP-friendly [HRX08]. TCP/IP [Str94a, BSR06, CS01, CP01, CP02, PC01, WW02, Wer10]. Teach [RBHB99, BS98, Bal99b, Gay99, Hsi01, LH00, Ray99a, SBR01, Tay03, WT99, Wel99a]. Teaching [Shu06, NV06]. Teams [Den99b, Ano03b]. Tech [Bar01b, RC98, Way00]. Technical [Hid04, IEE93, USE96b]. Techniques [NLK09, Stu01, Bes05, BC08, GFRD96, Kra05, Mol14, PKB03, Stu04, TSC02]. Technologies [CKM^{+}99, Ano03c, Ano03e]. technologique [Cor00]. Technology [Cus04, KD96, Kro99, ZB19, Cor05b, MW08, Wal01b]. tecnologia [Cor05b]. tecnologica [Zic01]. temperature [MB08, WJY^{+}07]. temperature-aware [MB08]. Temporal [LBN^{+}18, BDH^{+}09, DJC07]. ten [PTS^{+}11]. tenant [MKSB22]. terminal [DM19]. terminals [ISTE08]. Test [AT05, Ano02c, Bri05b, Coc03b, FES^{+}99, Kro99, Lew99b]. Testament [PR96a, WY94, W^{+}95]. tester [GFRD96]. Testing [Ano14, Geo94, VS06, BFN^{+}06, WN15]. Tests [MSSvK08]. Text [CPJ^{+}98, DF00, Mud97, GGZ02]. textbook [May01]. Textdateien [GGZ02]. Textverarbeitung [Str94a]. Textwerkzeuge [Gün02]. Thee [Sta04]. their [Kau96]. theory [SAL19]. There [CPJ^{+}98, Sta04]. Thermal [XZY^{+}10, LZZ^{+}08, WJY^{+}07, ZYCZ10]. Thermal-Aware [XZY^{+}10]. Thing [CPJ^{+}98]. Things [Wal99, Gra06b]. Think [LHB24]. Thinking [DDJ98b]. Third [Loz05]. Third-Party [Loz05]. Thomas [Kad95c, Kad95b, Koh06]. Thompson [FES^{+}99]. Thou [Dec04]. Thread [TAS07, SKP^{+}02]. threaded [RGK99]. threads [LZ07, SMM^{+}09, SKP^{+}02]. Three [Cas17]. Throughput [Vad00]. tier [UPS^{+}05, USC^{+}08]. Time [Abb03, Dou02, LW00, Mar04, Nis01b, PNK02, RMF19, SW98, TEGF98, WM00, ABB19, BP96, GMTW08, HST08, HR95, LD05, MMW96, ON06, PBM^{+}22, RLB08, SAL19, SJP^{+}09, She03, Epp00b, Nis03, She03, HST08]. timeliness [RLB08]. timer [PBR^{+}08]. Times [Lin02b, Lin02c]. TimeSys [Ano03f]. Timing [dOdO16, PV02]. Tiniest [Bar00c]. Tiny [Coc03b, Bar00c, SLQP07, SAN08, Kas14]. Tip [AHSM21]. Tips [Ste00a, Stu01, Bes05, Dos99b, Fli03, Ran04, Ran06, Stu04]. titans [Way00]. Tk [Bla02, Bla04]. TLB [ST03]. Tolerant [LW00]. Tom [Swa00b]. Tool [Ano02c, Ano02d, Ano03e, Ano03f, GSN^{+}01, Jen02, Kro99, MK04, PKFH02, SB02, SGL^{+}00, YKS^{+}02, Ano03b, Ano03c, Ano03d, GBC^{+}18, GB96, LHB24, MM06, RGK99]. Toolkit [JW99, Ser00, Wol04b, Wol04a, Bru02, Ser98, Wel98]. toolkits [Ano03d]. Tools [Cha06, DRN^{+}23, Dun99, Gre19, GS00, GGZ02, Jor04, Kro99, PKB01, SuS01, Wol04b, AC11, Ano03c, Fli03, MS07, MM00, Max00, Mol14, PKB03, Ran04, Ran06, Smi03, Tro96b]. Toolset [Ano02c, Ano03d]. top [RFS05, RFS06]. top-down [RFS05, RFS06]. TOP100 [EEK^{+}04]. Torvalds [Flo94, Li94]. tous [Cha99]. toutes [Wie99b]. toys [NW03, Neg06b]. TPC [Lla06]. TPC-C [Lla06]. TPCC [Lla06]. TPCC-UVa [Lla06]. Trace [MZG14, SWF^{+}22]. Tracers [GD18]. tracing [CMR^{+}21]. Tracking [She06]. Trademark [Eri97]. tradition [dBB08]. traditional [ALP^{+}21]. traffic [WJS07]. Training [Kom00, Kom02, QH01, Sta05b]. Transaction [HKS^{+}04]. Transactional [RRP^{+}08, RRP^{+}07, RHP^{+}07, RRH^{+}08]. Transfer [Jan03, MAP^{+}22, HH04]. transformations [FRRJ07]. transforms [Zem94]. Transistor [Bar00c]. Transmission [PNK02]. Transparent [Reh99, ST03]. Transport [FST^{+}09, BWH^{+}07]. trash [DM19]. Travelling [Rau01]. treasure [DM19]. Tree [RBM13, WG18]. Trenches [Ano99e]. Trends [Ahm08, Edw98, Var94, CMMST20, vD06]. Tricks [Ste00a, Wis03]. trio [CM06]. troubles [HBB13]. Troubleshooting [WF03, K^{+}06]. True [FES^{+}99]. Truly [Mud97]. Trusted [HHD^{+}16, ARJ08]. TSO [SSO^{+}10]. Tune [RGK99]. tuning [Bes05, FSW02, RGK99]. Turbulence [KEB^{+}99, LWL06]. Turing [Coc03b]. Turn [DM19, Gib17, Kau96]. turns [Toz16]. tutorial [She02]. TwinDrivers [MSZ09]. TWLinuX [TT01]. Two [Ahm08, Ano03f, Dou02, SJP^{+}09, SC98]. Two-Axis [Dou02]. TX [Ano94d]. TxLinux [RRP^{+}07, RRP^{+}08, RHP^{+}07, RRH^{+}08].

U.S. [Coc03a]. ubiquitous [Gar98].

Ubuntu

[Cha13, Gag06, Gag07, Gra06b, HHH11, HBB^{+}06, H^{+}07a, H^{+}07b, HH06, HH07, HH08, HH09, HHHT10, Kra07, Neg06a, Neg07a, ORC06, Tho06, vH06, vH07, Teo13]. udocker [GBC^{+}18]. Uhl [Kad95c, Kad95b]. Ultra [VTS^{+}23]. Ultra-low-power [VTS^{+}23]. Umbrella [GR09]. umfassende [Kuo99]. Umgang [DF00, Kuo99, Ste00a]. Umgebng [CK06j]. Umgebung [CK06a, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i, CK06k]. UML [Ano03e]. UMLinux [SB02]. Umsteiger [Ron05a]. unabhängige [Lin02b, Lin02c]. Uncompromising [Bra04]. Undercover [Ray98, Ste00b]. undercut [Way00]. Understanding [ALL13, Ben06, BC00, BC01, BC02, BC03, BC06, GZXZ23, JJW^{+}24, ON06, PLM06, Yu07]. Uneasy [Far94]. Unicode [DM97]. unification [WDG^{+}06]. Unified [HKS02, LBN14, WR03, FMR07]. unit [KEP07]. Universe [Kad95d, SU94b, SUM95, SU96a, SUM96, SMM97, Kad95b]. Unix [DF00, Her99, Kuo99, May01, RB06a, Str94a, AC11, Bau03, Bou96b, Bra96, Chi02, Chu03, Cor05b, DG00, FES^{+}99, Ker10, LB96, Len97, Loz05, Mon00a, Mon00b, Mus02, NSHW10, PDG02, Pom04, Qui02, Tan00, Ano95b, Ano95e, Ano03a, Cha99, Daw94, Gab07, Gan03, GJ08, Gue19, Hah09, Han00, Hou96, Kuo99, May01, MSW06, Moh97a, Moh97b, Olc01, RRR09, RR15b, Rob95b, RC98, Sor01, Tay03, TECB03, WDG^{+}06, dBB08]. Unix- [Her99]. Unix-Grundlagen [Str94a]. Unix-like [Moh97b]. Unix-Systemen [Kuo99]. Unix/Linux [RB06a, Chu03, Han00, Cha99]. Unix/Linux-Tasks [RB06a]. Unixem [WKD97]. UnixWare [Kuo99]. Unleashed [HH06, NR03, BP00, Bal00, CV00, HHH11, HH05b, HH07, HH08, HH09, HHHT10, HP^{+}95, HP96a, HP96b, P^{+}97, Par98a, PB98a, PB98b, WWW99, Wal01a, Teo13]. Unleashing [Kad95c, Kad95a, Str94b, SU94d, SU94b, SU96a, SU96d, SU96c, SE97]. unlocking [XSaJ08]. unmodified [Sun99]. unwired [WDJ04]. Update [JWKZ99, GMTW08, KS19, Moh97b]. updates [Ano03c, Ano03d, MR07]. updating [CCZ^{+}06]. Upgrades [Ano02d, Ano03e, Ano03f, Ano03b, Ano03c, HBB13]. Ups [Chu03, AMM^{+}18]. Upstream [LZAH22]. USA [ACM03b, ACM05b, MS05, USE96b, USE00]. usable [BBK^{+}16, SSO^{+}10]. Usage [GO99, MFBWW20, PBR^{+}08]. USB [Reg11]. Use [GA04, Jen02, Kof97, Stu01, TD96, Tem96, Bal99a, Kof99, Kof00, Ron05b, Stu04, VD01].

USENIX

[ACM05b, DDJ98b, Den99a, Den99b]. User [ACM93, AB04, Dik06, DF00, Gag01, GGZ02, Lat02, Moe02, Ron05a, Ano96a, BF08, DOF96, GBC^{+}18, Gre94, Moh98, NAGL10, Sto07, War04, Wil98]. user-controlled [Sto07]. User-Mode [AB04]. user-space [NAGL10]. Users [McC02, Ron05a, Gra05, HH04, K^{+}06, Kom96, Mil01, Mil96, She02]. Userspace [GD18]. Uses [Ano03f, GAH16, HRD^{+}07]. Using [AB04, ALL06, B^{+}98a, BMF15, CCO^{+}05, Don98, ECBK00, Fel96b, KBPV99, LW00, LD05, MS07, Mar04, Mar06, Nie00, Nor02, Oom13, SFE10, SRBB99, Smi99, TGB95a, TGB95b, Tac97, TB98, TECB03, VMKB05, AC11, AJ05, BN01, BFS07, BC08, BDH^{+}09, CCZ^{+}06, DG00, GVK^{+}03, Gue19, HR95, HKP09, LW03, LBH09, Lum13, Mar01, MMC06, NAGL10, NV06, PDG02, RÓ01, RHP^{+}07, SLP^{+}09, TG96, Tro96b, WJS07, WC22, WRA05]. uso [Bla97, VD01]. usuario [MR99]. usw [Ano01b]. utilisation [Fab97]. utilizando [RÓ01]. utilized [LHB24]. Utopia [Ano02d]. Utopia-LVDS [Ano02d]. UVa [Lla06]. UX [Kuo99]. üzemeltetése [Lás05].

V [DRN^{+}23, ILG19, OBSB23, PS22,

PGW^{+}20, SGP^{+}23, SNM22, VTS^{+}23, ZB19]. V1 [Ano95g]. Vade [Lip99]. validate [Reg11]. value [Lev95]. variability [NH14]. variable [LPH^{+}07]. variant [Ano95e, HRX08]. Variants [MAP^{+}22, AOAH14]. Varna [Pop94]. Vecotr [Bar00b]. vectors [MB08]. VEE [ACM05b, ACM06]. Venkatesulu [Wer10]. Ver [Ano96b, Ano96b]. Verification [HHD^{+}16]. Verifying [CGDB24, GHRS05]. Verilog [Kro99]. Versatile [SB02]. Versatility [WDG^{+}06]. Version [Ano01a, Ano02d, BR95, CS01, HS99, OST05, Daw94, GAH16, Hou96]. Versionen [DF00]. Versions [MAP^{+}22]. versus [ASW^{+}01]. Vertigo [FM02]. via [LWV^{+}10, LBN14, ZYCZ10]. Video [Sta06, Ran06, TK07]. Videotex [DM19]. viele [Ste00a]. View [Wal99]. Views [Bar00b, Bar00c, Bar01b, Coc01, Coc03b, Coc03a, DDJ98b, DDJ98a, Eri97]. Vintage [DDJ98b]. Virginia [MS05]. virtio [Rus08]. Virtual [ACM05b, ACM06, AB04, BDF^{+}03, BBHL08, CCO^{+}05, LD05, Nor02, TB05, ABB19, CK06a, CLDA07, NV06, Rus08, ZBG^{+}05, LCHL21]. Virtualisierung [Spr06, Spr07]. Virtualization [Spr07, BDF^{+}03, CCZ^{+}06, CGL^{+}08a, CGL^{+}08b, CGL^{+}08c, G^{+}06a, P^{+}08, SPF^{+}07, vD06]. virtualization-based [CGL^{+}08a, CGL^{+}08b, CGL^{+}08c]. virtuelle [WF03]. virtuellen [CK06a, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i, CK06j, CK06k]. virtues [BTS09]. Visualizing [WG18]. Visually [Bel00b]. VM [G^{+}06a, P^{+}08, TB05]. VMM [ALL06]. VMM-based [ALL06]. VMware [CK06c, KGG00, Bor01, CK06c, Com00, Com03, NL00, Rul07, WF03]. voice [Gib17]. VoiceGenie [Ano03c]. Voices [DOS99a]. VoiceXML [Ano03c]. VoIP [Ano03e]. Vol [CS01, VJ06]. volume [Ano95g]. volunteer [Lea12]. vom [SU94c, SU94a, SU95b, SE96, SU96d]. vorbereiten [DF00]. Vorstellung [CK06b, CK06d, CK06c, CK06h, CK06i, CK06k]. vos [Cha99]. VR [GVK^{+}03]. VR-Based [GVK^{+}03]. vs [AT05, SKP^{+}02]. VSched [LD05]. Vulnerability [JJW^{+}24, LZAH23].

W [Tho05]. W3C [Bar01b]. WA [ACM05a]. Wall [LCHL21]. ware [Sta05b]. warp [MMW96]. WARPED [MMW96]. was [Ste99]. Watch [KIKT01]. Wave [ALP^{+}21, Ano03d]. Wavelet [Zem94]. waves [Hol96]. Way [Bea04, CPJ^{+}98, Com00, EP05]. Ways [Nis01b]. Weapon [DDJ98b]. Wearable [KIKT01]. Web [Pau95, Ano02c, Ano03b, Ano03d, Aul01, But02b, Fel96a, FWS^{+}03, Gab07, G^{+}06b, Hun01b, JB99, Kro99, Kro00, LW03, PBH^{+}07, Sai01, SWL^{+}01, Sta04, TSC02, Wel98, WLS^{+}01]. Web-Savvy [Kro99]. WebCompiler [Kro99]. Weber [Sch04]. Webmin [Cam04]. WebSphere [AJ05]. Week [Ano94g]. weekend [BW02]. weighted [LBH09]. Weighting [Sie00]. Well [Kau96]. Welsh [Ano95a]. Werkzeugen [KGG00]. WGS [Pur95]. Where [Dec04]. White [Ste01f]. whitelist [XCS^{+}22]. Who [Cas14]. wichtigsten [CK06b, Gün02]. Wide [Pau95, CDD^{+}05]. Will [FES^{+}99, Tim99]. Win32 [Ano94c, Mon00a, Mon00b]. window [ALP^{+}21]. Windows [Bor01, DF00, Koh06, WF03, Wei04, AC11, Ano02b, Ano03d, Bar00b, BM02, Bri05b, Bru02, Bur97, Chi02, CK06a, CK06f, CK06e, CK06j, Ded09, HH04, Jan03, Kro00, McC01, Mil01, MYH00, MY03, MB05, SK10, SCB^{+}13, Smi05b, SSSS97, Sta07a, Tim03, TF04, VE99, War01, Wol04a]. Windows-to-Linux [Wol04a]. Windoze [CPJ^{+}98]. Wireless [SNF04, KS09b]. within [MKSB22]. without [EP05, Len97, Loz05, PV02]. wizards [Ano01c]. wonglediff [EP05]. woody [Ano01b]. WordPerfect [Smi99, HZ99]. Work [Asb00, Ste01d, Gon99]. Workloads [HKS^{+}04, BBB^{+}20]. Workplace [HHJC99]. Works [Bar01b, War04, War15]. Workshop [DW94]. Workstation [Ano03f, Bor01, Kad95c, Kad95a, SU94c, Str94a, Str94b, SU94d, SU95b, SE96, SU96d, WF03, SU94b, SU96a, SU96c, SE97, SU94a]. Workstations [cFJH^{+}03, TD96, Tem96, Ano95e, MAT08, RGK99]. World [Ano94f, BR99, CPJ^{+}98, Pau95, Cor05b, Daw94, Hal06, Hal07, LLA07, Smi05b, Sta07b]. WOS [Ano01c]. WOS-Doku [Ano01c]. would [Daw94]. wrappers [JGS^{+}08]. Write [LH01]. Writing [Coo09c, Coo09d, Don97b, Mot02, Cas14]. WSDL [Bar01b].

X [May01, Ste00a, Ano95l, May01, TECB03, Ygg93, Ygg94]. X/Open [Ano95l]. x86 [SSO^{+}10, RFS06, RFS05, SSO^{+}10]. x86-TSO [SSO^{+}10]. Xandros [Gra05]. Xen [BDF^{+}03, Spr06, Spr07, VS06]. XFree [Ano01b]. XFree86 [DF00, SuS01, Geo97]. XFree86/23.3.6 [DF00]. XFS [MLE^{+}00]. Xi [RB08]. Xilinx [Ano02d]. XIP [PLKH08]. XML [Gar99, Kro00]. XML-based [Kro00]. Xoc [CBC^{+}08a, CBC^{+}08b, CBC^{+}08c]. XP [Woe03]. XScale [Bar00b]. Xue [RB08]. XXI [ACM93].

Y2K [CKM^{+}99, DDJ98b, Eri97, Kie99a, Sta07a]. YAPP [Ste01g]. Year [Sta07a]. Years [Ahm08, MFBWW20, PTS^{+}11]. Yellow [WBS01]. Yggdrasil [Ygg94]. York [ACM03b]. you’ll [Pub11]. You’re [Coc01]. Yourself [RBHB99, BS98, Bal99b, Bur97, Gay99, Hsi01, LH00, Ray99a, Reg11, SBR01, Tay03, WT99, Wel99a]. Yum [McC07].

Z [May01, Cor97, May01, G^{+}06a, P^{+}08, TB05]. Z-Shell [May01]. z/VM [G^{+}06a, P^{+}08, TB05]. z15 [BBB^{+}20]. z9 [ZST^{+}07]. z990 [SvBC^{+}04]. Zaurus [HKS02]. zero [HMK^{+}20, Ano02c]. zero-day [HMK^{+}20]. Zertifizierung [Mag04]. Zhi [RB08]. Zipf [MSSvK08]. zSeries [BS05]. zum [Kuo99, Mag04]. zur [Jor04, KGG00, Kuo99, SU94c, SU94a, SU95b, SE96, SU96d]. zusätzlich [Ano01b].

References

Appavoo:2005:EKO

[AAB^{+}05] J. Appavoo, M. Auslander, M. Butrico, D. da Silva, O. Krieger, M. Mergen, M. Ostrowski, B. Rosenburg, R. W. Wisniewski, and J. Xenidis. Experiences with K42, an opensource, Linux-compatible, scalable operating-system kernel. IBM Systems Journal, 44(2): 427–??, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                                    8670.  URL http://www.
                                                    research.ibm.com/journal/
                                                    sj/442/appavoo.pdf.

Al-Buraiky:2004:BVI

[AB04] Salah M. S. Al-Buraiky. Building a virtual IPv6 lab using user-mode Linux. ;login: the USENIX Association newsletter, 29(1):??, February 2004. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                                    usenix.org/publications/
                                                    login/2004-02/pdfs/al-buraiky.
                                                    pdf.

Abbott:2003:LER

[Abb03]  Doug Abbott.  Linux for

Embedded and Real-Time Applications. Embedded technology series. Newnes Press, Amsterdam, The Netherlands, 2003. ISBN 0-7506-7546-2 (paperback). xii + 255 pp. LCCN QA76.76.O63 A24 2003. US$49.95. URL

                                                    ftp://uiarchive.cso.uiuc.
                                                    edu/pub/etext/gutenberg/;
                                                     http://www.loc.gov/catdir/
                                                    description/els041/2003277020.
                                                    html;   http://www.loc.gov/
                                                    catdir/toc/els041/2003277020.
                                                    html.

Abeni:2019:HSR

[ABB19] Luca Abeni, Alessandro Biondi, and Enrico Bini. Hierarchical scheduling of real-time tasks over Linux-based virtual machines. The Journal of Systems and Software, 149(??):234–249, March 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S016412121830270X.

Altheide:2011:DFO

[AC11] Cory Altheide and Harlan A.

Carvey. Digital forensics with open source tools: using open source platform tools for performing computer forensics on target systems: Windows, Mac, Linux, UNIX, etc. Syngress Publishing, Inc., Rockland, MA, USA, 2011. ISBN 1-59749-586-7 (paperback), 1-59749-587-5 (e-book). xvii + 264 pp. LCCN HV8079.C65 A474 2011. URL http:

//www.sciencedirect.com/
science/book/9781597495868.

ACM:1993:PUS

[ACM93] ACM, editor. Proceedings:

User Services Conference XXI, the Pan Pacific Hotel, San Diego, California, November 7– 10, 1993: toward new horizons. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-631-X. LCCN QA74.A26

ACM:1999:SPO

[ACM99] ACM, editor. SC’99: Oregon Convention Center 777 NE Martin Luther King Jr. Boulevard, Portland, Oregon, November 11–18, 1999. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. LCCN ????

ACM:2003:SII

[ACM03a] ACM, editor. SC2003: Igniting Innovation. Phoenix, AZ, November 15–21, 2003. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 1-58113-695-1. LCCN ????

ACM:2003:SPA

[ACM03b] ACM, editor. SOSP ’03: proceedings of the 19th ACM Symposium on Operating Systems Principles: the Sagamore, Bolton Landing, Lake George, New York, USA, October 19–22, 2003, volume 37(5) of Operating systems review. ACM Press, New York, NY 10036, USA, December 2003. ISBN 1-58113-757-5. ISSN 0163-5980 (print), 1943-586X (electronic). URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://uclibs.org/PID/34720.

ACM order number 534030.

ACM:2004:SHP

[ACM04] ACM, editor. SC 2004: High Performance Computing, Networking and Storage: Bridging communities: Proceedings of the IEEE/ACM Supercomputing 2004 Conference, Pittsburgh, PA, November 6–12, 2004. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2153-3. LCCN ????

ACM:2005:PAI

[ACM05a] ACM, editor. Proceedings of the 2005 ACM/IEEE conference on Supercomputing 2005, Seattle, WA, November 12–18 2005. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 1-59593-061-2. LCCN ????

ACM:2005:PFA

[ACM05b] ACM, editor. Proceedings of the First ACM/USENIX International Conference on Virtual Execution Environments: VEE ’05: June 11–12, 2005, Chicago, Illinois, USA. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-047-7. LCCN QA76.9.V5 I575 2005. URL http:

//www.loc.gov/catdir/toc/
fy0611/2006530661.html. ACM

order number 548059.

ACM:2006:VPS

[ACM06] ACM, editor. VEE 2006: proceedings of the Second International Conference on Virtual Execution Environments, June 14-16, 2006, Ottawa, Ontario, Canada. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-332-6 (??invalid ISBN??). LCCN QA76.9.V4.

Adelstein:2004:EJL

[AH04] Tom Adelstein and Sam Hiser. Exploring the JDS Linux Desktop. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00752-3 (paperback). xiv + 390 pp. LCCN QA76.76.O63 A353 2004. US$34.95, CAN$50.95, UK£24.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        jds/index.html.

Ahmad:2008:ATT

[Ahm08] David Ahmad. Attack trends: Two years of broken crypto: Debian’s dress rehearsal for a global PKI compromise. IEEE Security & Privacy, 6(5):70–73, September/October 2008. CO-DEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Alexopoulos:2021:TIM

[AHSM21] Nikolaos Alexopoulos, Sheikh Mahbub Habib, Steffen Schulz, and Max Mühlhäuser. The tip of the iceberg: On the merits of finding security bugs. ACM Transactions on Privacy and Security (TOPS), 24(1): 3:1–3:33, January 2021. CO-DEN ???? ISSN 2471-2566 (print), 2471-2574 (electronic).

URL https://dl.acm.org/
doi/10.1145/3406112.

Anand:2005:MPC

[AJ05] V. K. Anand and W. C. Jamison. A middleware performance characterization of Linux using IBM WebSphere Application Server. IBM Systems Journal, 44(2):353–??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/442/anand.pdf.

Adams:2010:FEP

[AJD^{+}10] Andrew Adams, David E. Jacobs, Jennifer Dolson, Marius Tico, Kari Pulli, Eino-Ville Talvala, Boris Ajdin, Daniel Vaquero, Hendrik P. A. Lensch, Mark Horowitz, Sung Hee Park, Natasha Gelfand, Jongmin Baek, Wojciech Matusik, and Marc Levoy. The Frankencamera: an experimental platform for computational photography. ACM Transactions on Graphics, 29(4):29:1–29:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alanko:1996:MCB

[AKL^{+}96] T. Alanko, M. Kojo, H. Laamanen, K. Raatikainen, and M. Tienari. Mobile computing based on GSM: the Mowgli approach. In Jose L. Encarnacao and Jan M. Rabey, editors, Mobile Communications. Technology, Tools, Applications, Authentication and Security. IFIP World Conference on Mobile Communications, pages ??–?? (of ix + 342). Chapman and Hall, Ltd., London, UK, 1996. ISBN 0-412-75580-7. LCCN TK6570.M6 I55 1996.

Abbes:2020:DRF

[ALC20] Heithem Abbes, Thouraya Louati, and Christophe Cérin. Dynamic replication factor model for Linux containersbased cloud systems. The Journal of Supercomputing, 76(9): 7219–7241, September 2020. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s11227-020-03158-5.

Alexander:2000:LIC

[Ale00] William F. Alexander. A Linux IEEE 1394 configuration ROM decoder. Dr. Dobb’s Journal of Software Tools, 25(8): 80, 82, 84, August 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_08/
linx1394.txt;   http://www.
ddj.com/ftp/2000/2000_08/
linx1394.zip.

Asrigo:2006:UVB

[ALL06] Kurniadi Asrigo, Lionel Litty, and David Lie. Using VMMbased sensors to monitor honeypots. In ACM [ACM06], pages 13–23. ISBN 1-59593-332-6 (??invalid ISBN??). LCCN QA76.9.V4.

Akkan:2013:UIN

[ALL13] Hakan Akkan, Michael Lang, and Lorie Liebrock. Understanding and isolating the noise in the Linux kernel. The International Journal of High Performance Computing Applications, 27(2):136–146, May 2013. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://hpc.

sagepub.com/content/27/2/
136.full.pdf+html.

Almesberger:1996:HAN

[Alm96] W. Almesberger. High-speed ATM networking on low-end computer systems. In Conference Proceedings of the 1996 IEEE Fifteenth Annual International Phoenix Conference on Computers and Communications, pages ??–?? (of xvii + 519). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3255-5. LCCN TK5105.5 .I332 1996. IEEE catalog number 96CH35917.

Abdelsalam:2021:TWL

[ALP^{+}21] A. Abdelsalam, M. Luglio, N. Patriciello, C. Roseti, and F. Zampognaro. TCP Wave over Linux: a disruptive alternative to the traditional TCP window approach. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):Article 107633, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1389128620312585.

Alglave:2018:FSC

[AMM^{+}18] Jade Alglave, Luc Maranget, Paul E. McKenney, Andrea Parri, and Alan Stern. Frightening small children and disconcerting grown-ups: Concurrency in the Linux kernel. ACM SIGPLAN Notices, 53(2): 405–418, February 2018. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Akella:2008:CPB

[AMS^{+}08] Aditya Akella, Bruce Maggs, Srinivasan Seshan, Anees Shaikh, and Ramesh K. Sitaraman. Corrections to ‘On the performance benefits of multihoming route control’. IEEE/ ACM Transactions on Networking, 16(2):496, April 2008. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See [AMSS08].

Akella:2008:PBM

[AMSS08] Aditya Akella, Bruce Maggs, Srinivasan Seshan, and Anees Shaikh. On the performance benefits of multihoming route control. IEEE/ACM Transactions on Networking, 16(1): 91–104, February 2008. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic). See corrections [AMS^{+}08].

Anandakrishnan:1999:PEG

[Ana99] Sridhar Anandakrishnan. Penguins everywhere: GNU/Linux in Antarctica. IEEE Software, 16(6):90–96, November/ December 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://computer.org/
software/so1999/s6090abs.
htm;   http://dlib.computer.
org/so/books/so1999/pdf/
s6090.pdf.

Andregg:1999:ALO

[And99] Bryan C. Andregg. The all Linux office. SysAdmin, 8(6): 33–38, June 1999. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Anonymous:1993:PPL

[Ano93] Anonymous. Plug and play Linux, 1993. ISBN 1-883601-11-8. 2 CD ROMs + 1 book.

Anonymous:1994:ICS

[Ano94a] Anonymous, editor. 1994 International Computer Symposium Conference Proceedings. Nat. Chiao Tung Univ, Hsinchu, Taiwan, 1994. Two volumes.

Anonymous:1994:CIC

[Ano94b] Anonymous, editor. CMG ’94: 20th International conference for the management and performance evaluation of enterprise computing systems — December 1994, Orlando, FL, CMG — CONFERENCE — 1994//V2. The Computer Measurement Group, Inc, Westmont, IL, USA, 1994.

Anonymous:1994:LPP

[Ano94c] Anonymous. Letters: Plug and Play new and old, split routing, Linux sources, OS/2 with Win32, and Ghost Busters. BYTE Magazine, 19(11):18–??, November 1994. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:PCC

[Ano94d] Anonymous, editor. Petroleum computer conference: — July 1994, Dallas, TX, Papers — Society of Petroleum Engineers of AIME 1994. Society of Petroleum Engineers, Richardson, TX, USA, 1994.

Anonymous:1994:PL

[Ano94e]  Anonymous.  Plug-and-play

Linux, 1994. 1 computer laser disc.

Anonymous:1994:PAC

[Ano94f] Anonymous, editor. Proceedings of the AUUG94 Conference: Open systems. Looking into the future: 6–9 September 1994, World Congress Centre, Melbourne, Australia. AUUG Inc, Kensington, NSW, Australia, 1994. ISBN 0-646-20108-5. LCCN TK5105.58.A87 1994.

Anonymous:1994:PWL

[Ano94g] Anonymous. Product of the week: Linux 1.0. PC Week, 11 (14):162–??, April 1994. ISSN 0740-1604.

Anonymous:1995:BRRb

[Ano95a] Anonymous. Book review:

Running linux: By Matt Welsh and Lar Kaufman. O’Reilly & Associates, Inc., Sebastopol, CA. (1995). 575 pages. $24.95. Computers and Mathematics with Applications, 29(11):110, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0898122195902708.

Anonymous:1995:CFR

[Ano95b] Anonymous. Caldera, funded by Ray Noorda, develops commercial applications based on the free Unix clone Linux. ComputerWorld, 29(16):32–??, April 1995. CODEN CMP-WAB. ISSN 0010-4841.

Anonymous:1995:CND

[Ano95c] Anonymous. Caldera network desktop Linux operating system: preview, 1995. 1 computer laser optical disk.

Anonymous:1995:DEW

[Ano95d] Anonymous. Desktop engineering: What about Linux? Electronic engineering times, ?? (865):96–??, September 1995. ISSN 0192-1541.

Anonymous:1995:LFV

[Ano95e] Anonymous. Linux, a free variant of Unix, is being ported to Digital workstations. ComputerWorld, 29(13):10–??, March 1995. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1995:LA

[Ano95f] Anonymous. Linux for astronomy, 1995. 1 computer laser optical disk.

Anonymous:1995:LAV

[Ano95g] Anonymous. Linux for astronomy, volume V1, 1995. 1 laser computer disk.

Anonymous:1995:PPL

[Ano95h] Anonymous. Plug and play Linux. InfoWorld, 17(22):75– ??, May 1995. CODEN IN-WODU. ISSN 0199-6649.

Anonymous:1995:SPL

[Ano95i] Anonymous. Slackware Professional Linux 2.3. InfoWorld, 17 (32):62–??, August 1995. CO-DEN INWODU. ISSN 0199-6649.

Anonymous:1995:SS

[Ano95j] Anonymous. Software in the spotlight. ACM Fortran Forum, 14(4):4–??, December 1, 1995. ISSN 1061-7264 (print), 1931-1311 (electronic).

Anonymous:1995:WN

[Ano95k]  Anonymous.  What’s new.

BYTE Magazine, 20(4):232–??, April 1, 1995. CODEN BYT-EDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1995:XBL

[Ano95l] Anonymous. X/Open branding for Linux. LAN: the network solutions magazine, 4(2): 8–??, July 1, 1995. CODEN LANNER. ISSN 1038-359X (or 1069-5621??).

Anonymous:1996:FGU

[Ano96a] Anonymous. Flexible graphical user interface for embedding heterogeneous neural network simulators. IEEE transactions on education, 39(3):367– 374, August 1996. CODEN IEEDAB. ISSN 0018-9359.

Anonymous:1996:LFI

[Ano96b] Anonymous. Linux facile!: installation: distributions Kheops.96-Ver 3.0.F, SLACKWARE-Ver 3.0. Logiciels du Soleil-JC les Pleyades, Cagnes sur mer (1 rue Pascalini 06802), France, 1996. ISBN 2-911067-13-4. 152 pp. LCCN ????

Anonymous:1997:CRL

[Ano97a] Anonymous. The complete Redhat Linux operating system. Macmillan Digital USA, Indianapolis, IN, USA, 1997. ISBN 0-672-31057-0, 1-888172-65-7 (users guide). 13 (booklet) + 223 (guide) pp. LCCN ???? Includes two CD-ROMs.

Anonymous:1997:ORL

[Ano97b] Anonymous. Official Redhat 5.0 Linux operating system: for Intel computers. Red Hat Software, Inc., Westport, CT, USA, 1997. ISBN 1-888172-94-0, 1-888172-97-5 (Installation guide). 304 pp. LCCN ???? Includes two CD-ROMs.

Anonymous:1998:L

[Ano98a] Anonymous. Linux. Simon & Schuster Macmillan (France), Paris, France, 1998. ISBN 2-7440-0539-8. xvi + 586 pp. LCCN ????

Anonymous:1998:LOS

[Ano98b] Anonymous. The Linux operating system. IEEE Software, 15(5):17, September/ October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1998/pdf/
s5017.pdf.

Anonymous:1999:AAA

[Ano99a] Anonymous. Announcement: 3rd Annual Atlanta Linux Showcase. ;login: the USENIX Association newsletter, 24 (3):??, June 1999. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
linuxshowcase.org.

Anonymous:1999:AAL

[Ano99b] Anonymous. Announcement: 3rd Annual Linux Showcase. ;login: the USENIX Association newsletter, 24(5s): ??, September 1999. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
linuxshowcase.org. Special

issue on intrusion detection.

Anonymous:1999:FL

[Ano99c] Anonymous. Focus: Linux.

IEEE Software, 16(1):??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:FLE

[Ano99d] Anonymous. Focus: Linux — essay. IEEE Software, 16(1):??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:FLT

[Ano99e] Anonymous. Focus: Linux — from the trenches. IEEE Software, 16(1):??, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:FLN

[Ano99f] Anonymous. Focus: Linux — nuts & bolts. IEEE Software, 16(1):??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:2000:AAL

[Ano00a] Anonymous. Announcement:

4th Annual Linux Showcase and Conference, Atlanta. ;login: the USENIX Association newsletter, 25(2):??, April 2000. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.linuxshowcase.org.

Anonymous:2000:CDG

[Ano00b] Anonymous. Citius Debian GNU/Linux 2.2. Investigación y Desarrolo Agora, Navarra, Spain, 2000. ISBN 84-607-1193-5. LCCN ???? One volume and ten CD-ROMs.

Anonymous:2001:DGLa

[Ano01a] Anonymous. Debian GNU, Linux - i386, GNU, Hurd - i386, Version 2.2 - “Potato”, release 2 [Elektronische Ressource]. Lehmann, Köln, Germany, 2001. ISBN 3-931253-74-0. LCCN ???? Includes six CD-ROMs.

Anonymous:2001:DGLc

[Ano01b] Anonymous. Debian GNU, Linux 3.0 (final) [Elektronische Ressource]: woody; i386 (Intel, AMD, Cyrix usw.); zusätzlich auf CD: Openoffice 1.0, XFree 86 4.2, Modellbahnsteuerungssoftware, Debian GNU/Linux-Anwenderhandbuch, Linux-Anwenderhandbuch. Lehmann, Berlin, Germany, 2001. ISBN 3-931253-89-9. LCCN ???? Includes eight CD-ROMs.

Anonymous:2001:DGLb

[Ano01c] Anonymous. Debian GNU, Linux 3.0 (pre) [Elektronische Ressource]: wizards of OS; 6 CD-ROMs mit WOS-Doku I und II und neuem KNOP-PIX. Lehmann, Berlin, Germany, 2001. ISBN 3-931253-81-3. LCCN ???? Includes six CD-ROMs.

Anonymous:2002:CRLb

[Ano02a] Anonymous. Conference report: 2002 Linux Kernel Developers Summit. ;login: the USENIX Association newsletter, 27(5):??, October 2002. CODEN LOGNEM. ISSN 1044-6397. URL http:

//www.usenix.org/events/
kernel02/2002Summit.pdf.

Anonymous:2002:OAI

[Ano02b] Anonymous. Optimizing applications with the Intel C++ and Fortran compilers for Windows and Linux. Technical report, Hewlett–Packard Corporation, Palo Alto, CA, USA, 2002. URL http:

//www.intel.com/software/
products/compilers/c60/techtopics/
Compiler_Optimization_6.
pdf.

Anonymous:2002:PNR

[Ano02c] Anonymous. Products: New rack-mount server from Apple; OSDL’s Linux database test tool; DigiTerra ships backoffice integration toolset; Oracle’s Developer Suite for business applications and Web services; Macromedia release J2EE-compatible application server and IDE; Motion-Builder 4.0 from Kaydara; Zero G Software’s InstallAnywhere 5. Computer, 35(7):96– 97, July 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2002/07/r7096.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/07/r7096.
                                        pdf.

Anonymous:2002:POU

[Ano02d] Anonymous. Products: Omnicore upgrades Java IDE CodeGuide emWare’s SDE for intelligent device management; Metrowerks’ CodeWarrior for Embedded Linux; integrated software environment form Xilinx; new version of InstallShield Professional; Motorola’s 32-bit CAN reference design; Utopia-LVDS bridge reference design kit from National Semiconductor; First Silicon Solutions’ analysis tool for flash-based FPGAs. Computer, 35(11):78–79, November 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/co/2002/11/ry078.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/co/2002/11/ry078.
                                        pdf.

Anonymous:2003:LUE

[Ano03a] Anonymous. Linux/Unix extend Red Hat 8’s functionality by adding Flash, extra fonts and Java. Personal computer world, 26(3):202–207, 2003. CODEN PCWODU. ISSN 0142-0232.

Anonymous:2003:PAI

[Ano03b] Anonymous. Products: Apple introduces 64-bit PC with 1-GHz bus; Altera upgrades PLD design tool; StrikeIron’s initial product analyzes Web services; Metrowerks teams up with AMD on Linux platform. Computer, 36(8):88, August 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/08/r8088.
htm;   http://csdl.computer.
org/dl/mags/co/2003/08/r8088.
pdf.

Anonymous:2003:PCN

[Ano03c] Anonymous. Products: Clear-Sight Networks releases applicationlayer analyzer; Intervoice announces first SALT-based components; VoiceGenie Technologies upgrades VoiceXML platform; AppForce enhances mobile-platform design software; Metrowerks upgrades tools for embedded Linux products; OpenOffice.org updates Linux office tool suite; Quest Software releases Java analysis tools. Computer, 36(11):86–87, November 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/11/ry086.
htm;   http://csdl.computer.
org/dl/mags/co/2003/11/ry086.
pdf.

Anonymous:2003:PIU

[Ano03d] Anonymous. Products: Intel updates multimedia performance primitives library; Lindows.com launches lowcost Linux PC; Metrowerks adds Linux kernel-level debugging; Rogue Wave announces C/C++ tool for Web services; ScanSoft announces toolkits to support Windows Mobile 2003; CoCreate launches collaboration toolset; Sybase ships IDE for Windows CE-based devices. Computer, 36(10):100–101, October 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/10/rx100.
htm;   http://csdl.computer.
org/dl/mags/co/2003/10/rx100.
pdf.

Anonymous:2003:PSA

[Ano03e] Anonymous. Products: SGI announces record-performing Linux servers; Pacific Nanotechnogy releases atomic force microscope; Nvidia extends Cg programming environment; Empirix introduces VoIP analyzer; Codagen Technologies upgrades UML tool; Perforce Software enhances SCM system; Compuware upgrades Java development tool. Computer, 36(2):80–81, February 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/02/r2080.
htm;   http://csdl.computer.
org/dl/mags/co/2003/02/r2080.
pdf.

Anonymous:2003:PSU

[Ano03f] Anonymous. Products: Sybase upgrades RAD tool; haptic workstation uses two hands; TimeSys launches Linux 4.0. Computer, 36(4):89, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/04/r4089.
htm;   http://csdl.computer.
org/dl/mags/co/2003/04/r4089.
pdf.

Anonymous:2005:CRLb

[Ano05a] Anonymous. Conference report: 2005 Linux Kernel Developers Summit. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/
login/2005-10/openpdfs/kernel05.
pdf.

Anonymous:2005:DGL

[Ano05b] Anonymous. Debian GNU/ Linux 3.1 r0a “Sarge”. Lehmann, Berlin, Germany, 2005. ISBN 3-86541-064-2. LCCN ???? EUR9.95. Two DVD-ROMs.

Anonymous:2006:PGI

[Ano06] Anonymous, editor. Proceedings of Gelato ICE: Itanium Conference and Expo: Spotlighting Linux on Itanium-based Platforms, October 1–4, 2006, Biopolis, Singapore. ????, ????, 2006. ISBN ???? LCCN

????  URL http://www.
ice.gelato.org/;   http:/
/www.ice.gelato.org/about/
oct06_presentations.php.

Anonymous:2014:KLA

[Ano14]  Anonymous.  Kali Linux

— assuring security by penetration testing. Network Security, 2014(8):4, August 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1353485814700777.

Alrshah:2014:CSH

[AOAH14] Mohamed A. Alrshah, Mohamed Othman, Borhanuddin Ali, and Zurina Mohd Hanapi. Comparative study of high-speed Linux TCP variants over high-BDP networks. Journal of Network and Computer Applications, 43(??):66– 75, August 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1084804514000903.

Alrshah:2015:ASL

[AOAH15] Mohamed A. Alrshah, Mohamed Othman, Borhanuddin Ali, and Zurina Mohd Hanapi. Agile-SD: a Linux-based TCP congestion control algorithm for supporting high-speed and short-distance networks. Journal of Network and Computer Applications, 55(??):181–190, September 2015. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1084804515001186.

Apte:2007:APL

[AR07] Himani Apte and Meenali Rungta. Adding parity to the Linux ext3 file system. Operating Systems Review, 41(1):56– 65, January 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Abhishek:2022:RNG

[AR22] Kunal Abhishek and E. George Dharma Prakash Raj. On random number generation for kernel applications. Fundamenta Informaticae, 185(4):285–311, ???? 2022. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aaraj:2008:ADH

[ARJ08] Najwa Aaraj, Anand Raghunathan, and Niraj K. Jha. Analysis and design of a hardware/software trusted platform module for embedded systems. ACM Transactions on Embedded Computing Systems, 8(1): 8:1–8:??, December 2008. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Asbury:2000:ELW

[Asb00] Steve Asbury. Enterprise Linux at Work: How to Build 10 Distributed Applications for Your Organization. Wiley, New York, NY, USA, 2000. ISBN 0-471-36349-9. 448 (est.) pp. LCCN QA76.76.O63 A78 2000. US$49.99; CDN$77.50. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/onix06/99059493.html.

Allison:2001:LLE

[ASW^{+}01] Dennis Allison, Randy Schrickel, Reid Womack, Jeremy C. Reed, Ashley Tate, and Paul Munsey. Letters: Looking for early PPC [People’s Computing Company] people; being prepared for invasion; BetterBASIC; Linux versus BSD; Diffie–Hellman to the rescue; the future of programming. Dr. Dobb’s Journal of Software Tools, 26(6):10, 12, June 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Adamo:2005:LVO

[AT05] Massimiliano Adamo and Mauro Tablò. Linux vs. OpenBSD: A firewall performance test. ;login: the USENIX Association newsletter, 30(6): ??, December 2005. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/
login/2005-12/pdfs/adamo.
pdf.

Adelstein:2007:LSA

[AT07] Tom Adelstein and Falko Timme. Linux system administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-00952-6 (paperback). xiii + 279 pp. LCCN QA76.76.O63; QA76.76.O63 A34 2007eb. URL http:/

/www.oreilly.com/catalog/
9780596009526.

Aulds:2001:LAW

[Aul01] Charles Aulds. Linux Apache Web server administration. The Craig Hunt Linux library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2734-7. xxiii + 615 pp. LCCN TK5105.8885.A63 A95 2001.

Appavoo:2008:PKB

[AUW08] Jonathan Appavoo, Volkmar Uhlig, and Amos Waterland. Project Kittyhawk: building a global-scale computer: Blue Gene/P as a generic computing platform. Operating Systems Review, 42(1):77–84, January 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Arnold:2004:IPN

[AV04] T. W. Arnold and L. P. Van Doorn. The IBM PCIXCC: A new cryptographic coprocessor for the IBM eServer. IBM Journal of Research and Development, 48(3/4):475–487, ???? 2004. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

                                        research.ibm.com/journal/
                                        rd/483/arnold.html;   http:
                                        //www.research.ibm.com/journal/
                                        rd/483/arnold.pdf.

Acharya:2008:SMC

[AWW08] Arup Acharya, Xiping Wang, and Charles Wright. SIP message classification: design and performance. Operating Systems Review, 42(1):100–101, January 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ball:1998:UL

[B^{+}98a] Bill Ball et al. Using Linux. Que Corporation, Indianapolis, IN, USA, 1998. ISBN 0-7897-1623-2. xix + 732 pp. LCCN QA76.76.O63U718 1998.

Beck:1998:LKI

[B^{+}98b] Michael Beck et al. Linux kernel internals. Addison-Wesley, Reading, MA, USA, second edition, 1998. ISBN 0-201-33143-

8. xvi + 480 pp. LCCN QA76.76.O63L54813 1998. Includes CD-ROM. Translation of the German edition Linux-Kernel-Programmierung.

Bailey:1994:L

[Bai94] David Bailey. Linux. UNIX review, 12(13):85–??, December 1994. CODEN UNRED5. ISSN 0742-3136.

Bailey:1997:MR

[Bai97] Edward C. Bailey. Maximum RPM. Red Hat Press, Indianapolis, IN, USA, 1997. ISBN 0-672-31105-4. xx + 450 pp. LCCN QA76.9.F5B35 1997.

Ball:1999:HUL

[Bal99a] Bill Ball. How to use Linux.

SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31545-9 (paperback). 323 pp. LCCN QA76.76.O63B357 1999. US$24.99.

Ball:1999:STY

[Bal99b] Bill Ball. Sams’ teach yourself Linux in 24 hours. SAMS Publishing, Indianapolis, IN, USA, second edition, 1999. ISBN 0-672-31526-2. xv + 574 pp. LCCN QA76.76.O63 B358 1999. Edited by Jeff Koch and Jane Brownlow.

Ball:2000:SLU

[Bal00] Bill Ball. SuSE Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31780-X. xxix + 1096 pp. LCCN QA76.76.O63 B355 2000.

Barkakati:1996:LSa

[Bar96a] Nabajyoti Barkakati. Linux secrets. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-798-X (paperback). xxxiv + 900 pp. LCCN QA76.76.O63B366 1996.

Barkakati:1996:LSb

[Bar96b] Nabajyoti Barkakati. Linux secrets. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 2-7361-2134-1. xxiiii + 1051 pp. LCCN ????

Barkakati:1997:SL

[Bar97] Naba Barkakati. Los secretos de Linux. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-415-0097-5. 788 pp. LCCN ????

Barkakati:1998:RHL

[Bar98]  Nabajyoti Barkakati.  Red

Hat Linux secrets. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 1998. ISBN 0-7645-3175-1. xxx + 895 pp. LCCN QA76.76.O63B36617 1998.

Barkakati:1999:LNE

[Bar99a] Nabajyoti Barkakati. Linux nouvelle édition secrets. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 2-7361-3231-9. xxx + 1010 pp. LCCN ????

Barkakati:1999:RLS

[Bar99b] Nabajyoti Barkakati. Red Hat Linux secrets. I D G Books Worldwide, Indianapolis, IN, USA, third edition, 1999. ISBN 0-7645-4639-2. xxviii + 976 pp. LCCN QA76.76.O63 B37 1999.

Bar:2000:LI

[Bar00a] Moshe Bar. Linux internals. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-212598-5. xv + 351 pp. LCCN QA76.76.O63 B362 2000.

Baran:2000:NVC

[Bar00b] Nicholas Baran. News and views: Computing problem solved; Scalable Vecotr Graphics spec released; Intel announces XScale microarchitecture; Windows not the only target at LinuxWorld; NIST, robotics industry move towards interoperability. Dr. Dobb’s Journal of Software Tools, 25 (11):18, November 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVE

[Bar00c] Nicholas Baran. News and views: ERP can spell disaster; Linux goes super; dot-com companies bet the farm; the tiniest transistor; and speaking of tiny. . .. Dr. Dobb’s Journal of Software Tools, 25 (2):18, February 2000. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ieee.
org/conference/iedm.

Bar:2001:LFS

[Bar01a] Moshe Bar. Linux file systems.

Network professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212955-7 (set), 0-07-212954-9 (book), 0-07-212953-0 (CD-ROM). xiv + 348 pp. LCCN QA76.76.O63 B3626 2001. Includes CD-ROM.

Baran:2001:NVW

[Bar01b] Nicholas Baran. News and views: WSDL goes to W3C for standardization; shortchanging science; EUVL may keep Moore’s Law going; spy satellites to generate high-tech jobs; Mexican government adopts Linux; supercomputer on a chip in the works; brain scan database goes public. Dr. Dobb’s Journal of Software Tools, 26(6):18, June 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Barkakati:2001:RLS

[Bar01c]  Nabajyoti Barkakati.  Red

Hat Linux 7.1 secrets. Hungry Minds, Indianapolis, IN, 2001. ISBN 0-7645-4771-2 (paperback). xxxiv + 1135 pp. LCCN QA76.76.O63 B366166 2001.

Barrett:2004:LPG

[Bar04]  Daniel J. Barrett.  Linux

pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00628-4. viii + 191 pp. LCCN QA76.76.O63 B3685 2004. URL http:/

/www.oreilly.com/catalog/
9780596006280.

Bauer:2002:BSS

[Bau02] Michael D. Bauer. Building Secure Servers with Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00217-3. xv + 430 pp. LCCN TK5105.59 .B38 2002. US$44.95. URL http:/

/www.oreilly.com/catalog/
bssrvrlnx.

Bauer:2003:AUL

[Bau03]  Kirk Bauer.  Automating

UNIX and Linux Administration. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-212-3. 600 (est.) pp. LCCN ????

Bollinger:1999:GEI

[BB99a] Terry Bollinger and Peter H.

Beckman. Guest Editors’ introduction: Linux on the move. IEEE Software, 16(1):30–35, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1999/pdf/
s1030.pdf.

Bowden:1999:LFM

[BB99b] Terrehon Bowden and Bodo Bauer. Linux firewall and masquerading: The IP chains concept in Linux 2.2. SysAdmin, 8(6):23–24, 26–30, 32, June 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Borntrager:2020:SYC

[BBB^{+}20] C. Bornträger, J. D. Bradbury, R. Bündgen, F. Busaba, L. C. Heller, and V. Mihajlovski. Secure your cloud workloads with IBM Secure Execution for Linux on IBM z15 and LinuxONE III. IBM Journal of Research and Development, 64(5/6):2:1–2:11, 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Beck:1996:LKI

[BBD^{+}96] Michael Beck, Harold Bohme, Mirko Dzladzka, Ulrich Kunitz, Robert Magnus, and Dirk Verworner. Linux Kernel Internals. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-87741-4, 0-201-87742-2 (with CD-ROM). xvii + 438 pp. LCCN QA76.76.O63 L55 1996. US$38.68. Forward by Linus Torvalds, creator of LINUX.

Banahan:2000:PLD

[BBD^{+}00] Mike Banahan, Michael Boerner, Ian Dickson, Jonathon Kelly, Nikhilesh Kumar Mandalay, Richard Ollerenshaw, Jonathan Pinnock, Ganesh Prasad, Joel Rowbottom, Geoff Sherlock, and Mark Wilcox. Professional Linux Deployment. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-287-4. xvii + 653 pp. LCCN QA76.76.O63 P76124 2000. US$49.99.

Bhattiprolu:2008:VSC

[BBHL08] Sukadev Bhattiprolu, Eric W. Biederman, Serge Hallyn, and Daniel Lezcano. Virtual servers and checkpoint/restart in mainstream Linux. Operating Systems Review, 42(5):104– 113, July 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Barlev:2016:SYU

[BBK^{+}16] S. Barlev, Z. Basil, S. Kohanim, R. Peleg, S. Regev, and A. Shulman-Peleg. Secure yet usable: Protecting servers and Linux containers. IBM Journal of Research and Development, 60(4):12:1–12:10, July/August 2016. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

URL  http://ieeexplore.
ieee.org/document/7523363/

.

Bujanovic:1996:NIF

[BBS96] I. Bujanovic, R. Benis, and H. Stipetic. NeOS installable file systems: design and implementation. In D. Kalpic and V. H. Dobric, editors, ITI ’96. Proceedings of the 18th International Conference on Information Technology Interfaces, page ?? ????, ????, 1996. ISBN ???? LCCN ????

Bovet:2000:ULK

[BC00] D. (Daniele) Bovet and Marco Cesati. Understanding the Linux kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 0-596-00002-2. xvi + 684 pp. LCCN QA76.76.O63 B665 2001.

Bovet:2001:ULK

[BC01] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux Kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00002-2. xvi + 684 pp. LCCN QA76.76.O63 B674 2001.

Bovet:2002:ULK

[BC02] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux Kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00213-0. xv + 765 pp. LCCN QA76.76.O63 B683 2003. US$49.95. URL http:/

/www.oreilly.com/catalog/
linuxkernel2.

Bovet:2003:ULK

[BC03] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux Kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00213-0. xv + 765 pp. LCCN QA76.76.O63 B683 2003. US$49.95. URL http:/

/www.oreilly.com/catalog/
9780596002138;   http://
www.oreilly.com/catalog/
linuxkernel2.

Bovet:2006:ULK

[BC06] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux kernel. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-00565-2 (paperback). xvi + 923 pp. LCCN QA76.76.O63 B683 2006.

Boutcher:2008:PTP

[BC08] David Boutcher and Abhishek Chandra. Practical techniques for purging deleted data using liveness information. Operating Systems Review, 42(5):85– 94, July 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bhatia:2008:RSE

[BCP08] Sapan Bhatia, Charles Consel, and Calton Pu. Remote specialization for efficient embedded operating systems. ACM Transactions on Programming Languages and Systems, 30(4): 22:1–22:32, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Brown:2006:SLM

[BDD^{+}06] Mark Brown, Chuck Davis, William Dy, Paul Ionescu, Jeff Richardson, Kurt Taylor, and Robbie Williamson. Solaris to Linux Migration: A Guide for System Administrators. IBM redbooks. IBM Corporation, San Jose, CA, USA, 2006. ISBN 0-7384-9608-1. xx + 424 pp. LCCN ???? URL

                                        http://www.redbooks.ibm.
                                        com/redbooks/pdfs/sg247186.

pdf. IBM order number SG24-7186-00.

Barham:2003:VMM

[BDF^{+}03] Paul Barham, Boris Dragovic, Keir Fraser, Steven Hand, Tim Harris, Alex Ho, Rolf Neugebauer, Ian Pratt, and Andrew Warfield. Virtual machine monitors: Xen and the art of virtualization. In ACM [ACM03b], pages 164–177. ISBN 1-58113-757-5. ISSN 0163-5980 (print), 1943-586X (electronic). URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://uclibs.org/PID/34720.

ACM order number 534030.

Brunel:2009:FFB

[BDH^{+}09] Julien Brunel, Damien Doligez, René Rydhof Hansen, Julia L. Lawall, and Gilles Muller. A foundation for flow-based program matching: using temporal logic and model checking. ACM SIGPLAN Notices, 44(1): 114–126, January 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bautts:2005:LNA

[BDP05] Tony Bautts, Terry Dawson, and Gregor N. Purdy. Linux network administrator’s guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2005. ISBN 0-596-00548-2. xxii + 338 pp. LCCN QA76.76.O63; QA76.76.O63 K566 2005; QA76.76.O63 K566 2005eb. URL http:/

/www.oreilly.com/catalog/
9780596005481.

Beale:2004:SLB

[Bea04] Jay Beale. Securing Linux the Bastille Way. Addison-Wesley, Reading, MA, USA, 2004. ISBN ????

????  pp.  LCCN ????

URL http://www.bastille-
linux.org/;   http://www.
linuxsecurity.com/feature_
stories/feature_story-59.

html. To appear.

Beekmans:2000:LS

[Bee00]  Gerard Beekmans.  Linux

from scratch. IUniverse.com, Inc., San Jose, CA, USA, 2000. ISBN 0-595-13765-2. xvi + 95 pp. LCCN QA76.76.O63. URL http://

www.linuxfromscratch.org/.

Beekmans:2016:LS

[Bee16]  Gerard Beekmans.  Linux

from scratch. Web site.,

                                        2016.  URL http://www.
                                        linuxfromscratch.org//lfs/
                                        faq.html.

Bellomo:2000:DGL

[Bel00a]  Michael Bellomo.  Debian

GNU/Linux for dummies. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-0713-3. xxii + 324 pp. LCCN QA76.76.O63 B44978 2000.

Bellomo:2000:MRL

[Bel00b]  Michael Bellomo.  Mas-

ter Red Hat Linux Visually. IDG Books, San Mateo, CA, USA, 2000. ISBN 0-7645-3436-X. 56 pp. LCCN ?QA76.76.O63 B44982 2000??? US$34.99. URL

                                        http://catalog.idgbooks.
                                        com/product.asp?isbn=076453436X.

Bentson:1996:ILL

[Ben96]  Randolph Bentson.  Inside

Linux: a look at operating system development. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1996. ISBN 0-916151-89-1. xvii + 291 pp. LCCN QA76.76.O63 B464 1996.

Benvenuti:2006:ULN

[Ben06] Christian Benvenuti. Understanding Linux network internals. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00255-

6. xxiv + 1035 pp. LCCN QA76.76.O63; QA76.76.O63 B46 2006eb; QA76.76.O63 B46 2006. URL http:/

/www.oreilly.com/catalog/
9780596002558.

Best:2005:LDP

[Bes05] Steve (Steve Francis) Best.

Linux debugging and performance tuning: tips and techniques. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2005. ISBN 0-13-149247-0. ???? pp. LCCN QA76.76.O63 B4756 2005.

URL ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/
;   http://www.loc.gov/catdir/
toc/ecip0514/2005017192.
html;   http://www.phptr.
com/title/0131492470.

Butzen:1997:LD

[BF97] Fred Butzen and Dorothy Forbes. The Linux Database. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-491-5. xiv + 561 pp. LCCN QA76.9.D3B893 1997. US$39.95, CDN$55.95. URL

http://www.mispress.com/.

Bahmann:2008:EFK

[BF08] Helge Bahmann and Konrad Froitzheim. Extending futex for kernel to user notification. Operating Systems Review, 42(5): 18–26, July 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bishop:2006:ELH

[BFN^{+}06] Steve Bishop, Matthew Fairbairn, Michael Norrish, Peter Sewell, Michael Smith, and Keith Wansbrough. Engineering with logic: HOL specification and symbolic-evaluation testing for TCP implementations. ACM SIGPLAN Notices, 41(1):55–66, January 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beshers:2007:ERU

[BFS07] Clifford Beshers, David Fox, and Jeremy Shaw. Experience report: using functional programming to manage a Linux distribution. ACM SIGPLAN Notices, 42(9):213– 218, September 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

B:1999:DIF

[BGM99] Praveen B., Deepak Gupta, and Rajat Moona. Design and implementation of a file system with on-the-fly data compression for GNU/ Linux. Software—Practice and Experience, 29(10):863–874, August 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
63001367;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=63001367&
PLACEBO=IE.pdf.

Bramer:2002:DGL

[BGO02] Michael Bramer, John Goerzen, and Ossama Othman. Debian GNU/Linux guide. Linuxland, München, Germany, 2002. ISBN 3-936759-00-6. 404 (est.) pp. LCCN ????

Butzen:1998:LN

[BH98] Fred Butzen and Christopher Hilton. The LINUX network. M and T Books slackware series. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 1-55828-589-X. xix + 526 pp. LCCN TK5105.5.B889 1998.

Boling:2001:EHK

[BJ01] Eli Boling and Chuck Jazdzewski.

Exception handling in Kylix: Matching up Object Pascal, Linux, and exception handling. Dr. Dobb’s Journal of Software Tools, 26(11):66, 68–71, November 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_11/
kylix.txt.

Brecht:2006:ENP

[BJL^{+}06] Tim Brecht, G. (John) Janakiraman, Brian Lynn, Vikram Saletore, and Yoshio Turner. Evaluating network processing efficiency with processor partitioning and asynchronous I/O. Operating Systems Review, 40(4):265–278, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Becker:2022:CPA

[BJM^{+}22] Benedikt Becker, Nicolas Jeannerod, Claude Marché, Yann Régis-Gianas, Mihaela Sighireanu, and Ralf Treinen. The CoLiS platform for the analysis of maintainer scripts in Debian software packages. International Journal on Software Tools for Technology Transfer (STTT), 24(5):717– 733, October 2022. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10009-022-00671-1.

Barrett:2013:BBC

[BK13] Steven F. Barrett and Jason Kridner. Bad to the Bone: crafting electronic systems with BeagleBone and BeagleBone Black, volume 41 of Synthesis lectures on digital circuits and systems. Morgan and Claypool, San Rafael, CA, USA, 2013. ISBN 1-62705-137-6 (paperback), 1-62705-138-4. ISSN 1932-3174. xxv + 397 pp. LCCN QA76.5 .B37 2013; TJ223.P76.

Bagherzadeh:2018:ADL

[BKB^{+}18] Mojtaba Bagherzadeh, Nafiseh Kahani, Cor-Paul Bezemer, Ahmed E. Hassan, Juergen Dingel, and James R. Cordy. Analyzing a decade of Linux system calls. Empirical Software Engineering, 23(3):1519– 1551, June 2018. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10664-017-9551-z.

Blanco:1997:LIA

[Bla97] Vicente J. Blanco. Linux: instalacion, administracion y uso del sistema. Alfaomega, Mexico, DF, Mexico, 1997. ISBN 970-15-0326-0. xiv + 320 pp. LCCN ????

Blackwood:1999:SRL

[Bla99] David Blackwood. Standard report: Is Linux the future of POSIX? ;login: the USENIX Association newsletter, 24(2):??, April 1999. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/
login/standards/26.linux.
html.

Blaess:2002:LSS

[Bla02] Christophe Blaess. Langages de scripts sous Linux: Shell Bash, Sed, Awk, Perl, Tcl, Tk, Python, Ruby. Eyrolles, Paris, France, 2002. ISBN 2-212-11028-6. xx + 733 pp. LCCN QA76.7 B4 2002.

Blaess:2004:SSL

[Bla04] Christophe Blaess. Scripts sous Linux: Shell Bash, Sed, Awk, Perl, TCL, Tk, Python, Ruby. Eyrolles, Paris, France, second edition, 2004. ISBN 2-212-11405-2. xxi + 761 pp. LCCN QA76.76O63; QA76.7.

Bai:2020:EDS

[BLH20] Jia-Ju Bai, Julia Lawall, and Shi-Min Hu. Effective detection of sleep-in-atomic-context bugs in the Linux kernel. ACM Transactions on Computer Systems, 36(4):10:1–10:30, June 2020. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3381990.

Blum:2000:SLM

[Blu00] Richard Blum. SOHO Linux Mail Server. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31834-2. x + 526 pp. LCCN TK5105.73 .B58 2000. US$34.95. Edited by Angie Wethington, Neil Rowe, and Tony Amico.

Blum:2008:LCL

[Blu08] Richard Blum. Linux command line and shell scripting bible. Wiley, New York, NY, USA, 2008. ISBN 0-470-25128-X (paperback). xxx + 809 pp. LCCN QA76.76.O63 B598

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2008012238-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0827/2008012238-
t.html;   http://www.loc.
gov/catdir/enhancements/
fy0828/2008012238-b.html.

Blum:2011:LCL

[Blu11] Richard S. Blum. Linux command line and shell scripting bible. Wiley, New York, NY, USA, second edition, 2011. ISBN 1-118-00442-6 (paperback). ???? pp. LCCN ????

Bradford:2002:LWI

[BM02] Ed Bradford and Lou Mauget.

Linux and Windows Interoperability Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-032477-9. xv + 600 pp. LCCN QA76.76.O63 B7168 2002 Bar. US$44.99.

URL http://www.phptr.com/
ptrbooks/ptr_0130324779.

html. Configuring, programming, and administering mixed Linux/Windows environments.

Bellasi:2015:ERR

[BMF15] Patrick Bellasi, Giuseppe Massari, and William Fornaciari. Effective runtime resource management using Linux control groups with the BarbequeRTRM framework. ACM Transactions on Embedded Computing Systems, 14(2): 39:1–39:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Bandel:2001:SEU

[BN01] David A. (David Allan) Bandel and Robert Napier. Special edition using Linux. Que Corporation, Indianapolis, IN, USA, sixth edition, 2001. ISBN 0-7897-2543-6. xv + 817 pp. LCCN QA76.76.O63 B3625 2001.

Bokhari:1995:LOSa

[Bok95a] Shahid H. Bokhari. The Linux operating system. Computer, 28(8):74–79, August 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bokhari:1995:LOSb

[Bok95b] Shahid H. Bokhari. The Linux operating system: an introduction. ICASE report 95-49 and NASA contractor report; NASA CR-199178, Institute for Computer Applications in Science and Engineering NASA Langley Research Center, Hampton, VA, USA, 1995.

Bollinger:1999:LPO

[Bol99] Terry Bollinger. Linux in practice: An overview of applications. IEEE Software, 16(1):72–79, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

                                        URL http://computer.org/
                                        software/so1999/s1072abs.
                                        htm;   http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s1072.pdf.

Boloni:2001:PKC

[Böl01] Lotzi Bölöni. Programming KDE 2.0: Creating Linux Desktop Applications. CMP Books, Lawrence, KS, 2001.

ISBN 1-929629-13-3. ix + 265 pp. LCCN QA76.9.U83B65 2001. US$49.95. Includes CD-ROM.

Bonnetain:1999:RL

[Bon99] Pierre-Yves Bonnetain, trad.

Red Hat Linux 5.2. Le starter. Campus Press, Paris, France, 1999. ISBN 2-7440-0613-0. 598 pp. LCCN ????

Bonang:2001:NLS

[Bon01] James Bonang. Navigating Linux source code. C/C++ Users Journal, 19(3):18–??, March 2001. CODEN CCU-JEX. ISSN 1075-2838.

Bookman:2003:LCB

[Boo03] Charles Bookman. Linux clustering: building and maintaining Linux clusters. New Riders Publishing, Carmel, IN, USA, 2003. ISBN 1-57870-274-7. xv + 265 pp. LCCN QA76.76.O63 B427 2003.

Born:2001:VWP

[Bor01] Günter Born. VMware Workstation Praxisführer: Installation, Konfiguration, Anwendung unter Windows und Linux. SuSE-PRESS, Nürnberg, Germany, 2001. ISBN 3-934678-81-5. xxii + 372 pp. LCCN ???? DM 79.00, EUR 40.00.

Bourne:1995:HRI

[Bou95] P. E. Bourne. History repeats itself with Linux. Digital systems journal, 17(6):11–??, ????

1995. CODEN DSJOEE. ISSN 1067-7224.

Bourne:1996:HRI

[Bou96a] Philip E. Bourne. History repeats itself with Linux. Digital Age, 15(1):34–??, January 1, 1996. ISSN 1084-9076.

Bourne:1996:UNH

[Bou96b] Philip E. Bourne. UNIX now — history repeats itself with Linux. The DEC Professional, 15(1):34–??, ???? 1996. CO-DEN DECPDJ. ISSN 0744-9216.

Bertolissi:1996:ARC

[BP96] E. Bertolissi and C. Preece. Assessment of real-time communications capabilities of network protocols in distributed systems. In Kalpic and Dobric [KD96], pages 155–161. ISBN ???? LCCN T58.5.I56 1996.

Ball:2000:LU

[BP00] Bill Ball and David Pitts. Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, fourth edition, 2000. ISBN 0-672-31688-9. xl + 1468 pp. LCCN QA76.76.O63 B35 2000.

Barnes:1995:RCL

[BR95] Donnie Barnes and Red Hat Software, Inc. Red Hat Commercial Linux: Version 2.1, 1995. 4 computer laser optical disks.

Bokhari:1999:LDW

[BR99] Shahid H. Bokhari and Rafeequr Rehman. Linux and the developing world. IEEE Software, 16(1):58–64, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://computer.org/
software/so1999/s1058abs.
htm;   http://dlib.computer.
org/so/books/so1999/pdf/
s1058.pdf.

Brotzman:2000:SLS

[BR^{+}00] Lee E. Brotzman, David A.

Ranch, et al. Securing Linux: Step by Step. The SANS Institute, Bethesda, MD, USA, 2000. ISBN 0-9672992-0-9. ???? pp. LCCN ???? US$49.00.

Brandejs:1996:ULP

[Bra96] Michal Brandejs. UNIX — Linux: prakticky pruvodce. Grada, Praha, Czech Republic, 1996. ISBN 80-7169-170-4 (broz.). 340 pp. LCCN ????

Braam:2000:IHA

[Bra00] Peter J. Braam. The Inter-Mezzo high-availability file system. Journal of Linux Technology, 1(2):8–10, 12–15, 2000. ISSN 1527-2761.

Bramer:2004:DGL

[Bra04] Michael Bramer. Debian GNU LINUX 2.2: kompromisslos in Sicherheit & Stabilität: über 4000 Programmpakete: ausgereifte Desktop- & Server-Lösungen: mit Support & Handbuch. (German) [Debian GNU LINUX 2.2: Uncompromising in Security and Stability: Over 4000 program packages: Mature Desktop and Server Solutions: With Support and Handbook]. LinuxLand Internationale, Müenchen, Germany, 2004. ISBN 3-00-005785-4. ???? pp. LCCN ????

Brickner:2005:LDP

[Bri05a] David Brickner. Linux desktop pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10104-X. xvi + 181 pp. LCCN QA76.76.O63 B7423 2005.

Brickner:2005:TDL

[Bri05b] David Brickner. Test driving Linux: from Windows to Linux in 60 seconds. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00754-X. xvii + 341 pp. LCCN QA76.76.O63 B744 2005.

Brockmeier:2000:SGL

[Bro00] Joe “Zonker” Brockmeier. Selecting a GUI for Linux. SysAdmin, 9(1):59–61, January 2000. CODEN SYADE7. ISSN 1061-

                                        2688.  URL http://www.
                                        samag.com/.

Brooke:2003:DDC

[Bro03] Thomas C. Brooke. Development of a distributed, crossplatform simulator. ACM SIGADA Ada Letters, 23(1): 12–21, March 2003. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Brown:2006:SL

[Bro06] Chris Brown. SuSE Linux.

O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10183-X. 446 (est.) pp. LCCN ???? US$39.99.

Brunson:2002:LWI

[Bru02] Ross Brunson. Linux and Windows 2000 integration toolkit: a complete resource. Wiley, New York, NY, USA, 2002. ISBN 0-471-41746-7 (paperback). xii + 371 pp. LCCN QA76.6 .B7778 2002.

Ball:1998:STY

[BS98] Bill Ball and Stephen Smoogen.

Sams’ teach yourself Linux in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 1998. ISBN 0-672-31162-3. xviii + 380 pp. LCCN QA76.76.O63 B358 1998. Includes CD-ROM.

Borntrager:2005:PLS

[BS05] C. Bornträger and M. Schwidefsky. Providing Linux 2.6 support for the zSeries platform.

IBM Systems Journal, 44(2): 331–??, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/borntraeger.pdf.

Barrett:2003:LSC

[BSB03] Daniel J. Barrett, Richard E. Silverman, and Robert G. Byrnes. Linux security cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00391-9. xvii + 311 pp. LCCN TK5105.59 .B36 2003.

URL http://www.oreilly.
com/catalog/9780596003913.

Binkert:2006:INI

[BSR06] Nathan L. Binkert, Ali G. Saidi, and Steven K. Reinhardt. Integrated network interfaces for high-bandwidth TCP/IP. ACM SIGPLAN Notices, 41 (11):315–324, November 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balluder:1996:SOS

[BSS^{+}96] Karsten Ballüder, John A. Scales, Claus Schroeter, Martin L. Smith, and James R. Matey. Selecting an operating system, Part IV: Linux. Computers in Physics, 10(1): 17–??, January 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822351.

Balluder:1996:EPS

[BSSS96] Karsten Balluder, John A.

Scales, Claus Schroeter, and Martin L. Smith. Experimental physics: Selecting an operating system, part IV: Linux. Computers in Physics, 10(1):17–??, ???? 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Bauer:2005:LSS

[BsswLB05] Michael D. Bauer and Michael D. Building secure servers with Linux Bauer. Linux server security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00670-5. xvii + 522 pp. LCCN TK5105.59 .B38 2005; TK5105.59 .B38 2005eb; TK5105.59. URL http:/

/www.oreilly.com/catalog/
9780596006709.

Blaich:2009:RVM

[BTS09] Andrew Blaich, Douglas Thain, and Aaron Striegel. Reflections on the virtues of modularity: a case study in Linux security modules. Software—Practice and Experience, 39(15):1235– 1251, October ??, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Burgess:1997:LWH

[Bur97] Reginald P. Burgess. Linux, DOS and Windows: a how to build yourself a 95/NT clone. American Group Publishing, Wilmington, DE, USA, 1997. ISBN 1-891950-02-9. 155 guide pp. LCCN ????

Burtch:2004:LSS

[Bur04] Ken O. Burtch. Linux Shell Scripting with Bash. SAMS Publishing, Indianapolis, IN, USA, 2004. ISBN 0-672-32642-6. xvi + 412 pp. LCCN QA76.76.O63 B8725 2004. US$34.99.

But:2002:CSL

[But02a] Jason But. A C++ socket library for Linux. Dr. Dobb’s Journal of Software Tools, 27 (6):19–22, 24, June 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_06/
                                        socketcc.txt;   http://www.
                                        ddj.com/ftp/2002/2002_06/
                                        socketcc.zip.

Butorac:2002:PIW

[But02b] Danko Butorac. Project IP-SIS — Web portal and Linux for the blind. Lecture Notes in Computer Science, 2398:585– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/

0558/bibs/2398/23980585.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2398/23980585.
pdf.

Breuer:2006:RNO

[BV06] Peter T. Breuer and Marisol García Valls. Raiding the Noosphere: the open development of networked RAID support for the Linux kernel. Software—Practice and Experience, 36(4):365– 395, ???? 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barkakati:2002:RLW

[BW02] Naba Barkakati and Kurt Wall.

Red Hat Linux 7.2 weekend crash course. Hungry Minds, 909 Third Avenue, New York, NY 10022, USA, 2002. ISBN 0-7645-3642-7. xix + 347 pp. LCCN QA76.76.O63 B366167 2002. URL http:

//www.loc.gov/catdir/toc/
wiley021/2001092935.html.

Barrett:2004:LKG

[BW04] Daniel J. Barrett and Torsten Wilhelm. Linux — kurz and gut. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 3-89721-501-2 (paperback). 204 pp. LCCN ???? EUR 9.90. URL http:

                                        //www.gbv.de/dms/ilmenau/
                                        toc/388679999.PDF.

Bridges:2007:CET

[BWH^{+}07] Patrick G. Bridges, Gary T. Wong, Matti Hiltunen, Richard D. Schlichting, and Matthew J. Barrick. A configurable and extensible transport protocol. IEEE/ACM Transactions on Networking, 15(6):1254–1265, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Ben-Yehuda:2008:MGR

[BYVF08] Muli Ben-Yehuda, Eric Van Hensbergen, and Marc Fiuczynski. Minding the gap: R&D in the Linux kernel. Operating Systems Review, 42(5): 1–3, July 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chuvakin:2003:SLS

[C^{+}03] Anton Chuvakin et al. Securing Linux: A Survival Guide for Linux Security. The SANS Institute, Bethesda, MD, USA, 2003. ISBN ???? 104 pp. LCCN ???? US$39.00.

CS:2000:LSA

[Cal00]  Caldera Systems.  Caldera

OpenLinux eDesktop 2.4. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 2000. ISBN 0-13-088247-X. ???? pp. LCCN ???? US$59.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_013088247X.html.

Cameron:2004:MLS

[Cam04] Jamie Cameron. Managing Linux Systems with Webmin: System Administration and Module Development. Bruce Perens’ Open source series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2004. ISBN 0-13-140882-8. xvi + 792 pp. LCCN QA76.76.O63 2004. US$44.99.

Carter:1996:GHL

[Car96] S. Carter. Getting to the hardware from Linux. Forth Dimensions, 17(6):30–??, ???? 1996. CODEN FODMD5. ISSN 0884-0822.

Card:1998:PLA

[Car98] Remy Card. Programmation Linux 2.0: API système et fonctionnement du noyau. Eyrolles, Paris, France, 1998. ISBN 2-212-09018-8. xviii + 520 pp. LCCN ????

Carasik:1999:LSA

[Car99] Anne H. Carasik. Linux system administration. The M and T books slackware series. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1999. ISBN 0-7645-7008-0. xxv + 451 pp. LCCN QA76.76.O63C3729 1999.

Cass:2014:WWL

[Cas14] Stephen Cass. Who’s writing Linux? [data flow]. IEEE Spectrum, 51(2):72, February 2014.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2017:SAL

[Cas17] Stephen Cass. Some assembly (language) required — three games that make low-level coding fun [resources geek life]. IEEE Spectrum, 54(5):19–20, May 2017. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Choffnes:2008:MPM

[CAW08] David Choffnes, Mark Astley, and Michael J. Ward. Migration policies for multi-core fair-share scheduling. Operating Systems Review, 42(1):92– 93, January 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cox:2008:XEOa

[CBC^{+}08a] Russ Cox, Tom Bergan, Austin T. Clements, Frans Kaashoek, and Eddie Kohler. Xoc, an extension-oriented compiler for systems programming. ACM SIGARCH Computer Architecture News, 36(1): 244–254, March 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Cox:2008:XEOb

[CBC^{+}08b] Russ Cox, Tom Bergan, Austin T. Clements, Frans Kaashoek, and Eddie Kohler. Xoc, an extension-oriented compiler for systems programming. Operating Systems Review, 42(2):244–254, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cox:2008:XEOc

[CBC^{+}08c] Russ Cox, Tom Bergan, Austin T. Clements, Frans Kaashoek, and Eddie Kohler. Xoc, an extension-oriented compiler for systems programming. ACM SIGPLAN Notices, 43(3):244–254, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chiola:1997:ILC

[CC97] G. Chiola and G. Ciaccio. Implementing a low cost, low latency parallel platform. Parallel Computing, 22(13):1703– 1717, ???? 28, 1997. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Childs:2005:SCG

[CCO^{+}05] Stephen Childs, Brian Coghlan, David O’Callaghan, Geoff Quigley, and John Walsh. A single-computer Grid gateway using virtual machines. In Shih and Shibata [SS05b], pages 310–315. ISBN 0-7695-2249-1 (paperback). ISSN 1550-445X. LCCN TK5105.5 .I5616 2005.

URL  http://ieeexplore.
ieee.org/servlet/opac?punumber=

9746. IEEE Computer Society Order Number P2249.

Cruz-Carlon:2023:PLB

[CCVLW23] Juan Cruz-Carlon, Mahsa Varshosaz, Claire Le Goues, and Andrzej Wasowski. Patching locking bugs statically with crayons. ACM Transactions on Software Engineering and Methodology, 32(3):56:1–56:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3548684.

Chen:2006:LUO

[CCZ^{+}06] Haibo Chen, Rong Chen, Fengzhe Zhang, Binyu Zang, and Pen-Chung Yew. Live updating operating systems using virtualization. In ACM [ACM06], pages 35–44. ISBN 1-59593-332-6 (??invalid ISBN??). LCCN QA76.9.V4.

Carling:2000:LSA

[CDD00] M. Carling, Stephen Degler, and James T. Dennis. Linux System Administration. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 0-562-05934-3. x + 337 pp. LCCN QA76.76.O63 C3745 2000. US$29.99.

Chanet:2005:SWC

[CDD^{+}05] Dominique Chanet, Bjorn De Sutter, Bruno De Bus, Ludo Van Put, and Koen De Bosschere. System-wide compaction and specialization of the Linux kernel. ACM SIGPLAN Notices, 40(7):95–104, July 2005.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chanet:2007:ARM

[CDD^{+}07] Dominique Chanet, Bjorn De Sutter, Bruno De Bus, Ludo Van Put, and Koen De Bosschere. Automated reduction of the memory footprint of the Linux kernel. ACM Transactions on Embedded Computing Systems, 6(4):23:1–23:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Card:1997:PLA

[CDM97] Remy Card, Eric Dumas, and Franck Mevel. Programmation Linux 2.0: API système et fonctionnement du noyau. Eyrolles, Paris, France, 1997. ISBN 2-212-08932-5. xviii + 519 pp. LCCN ????

Card:1998:LKB

[CDM98] Remy Card, Eric Dumas, and Franck Mevel. The Linux Kernel book. Wiley, New York, NY, USA, 1998. ISBN 0-471-98141-9. xxx + 518 pp. LCCN QA76.76.O63 C374 1998. US$44.99. Translated by Chris Skrimshire.

Calcagno:2009:CSA

[CDOY09] Cristiano Calcagno, Dino Distefano, Peter O’Hearn, and Hongseok Yang. Compositional shape analysis by means of biabduction. ACM SIGPLAN Notices, 44(1):289–300, January 2009. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ceruzzi:2003:HMC

[Cer03] Paul E. Ceruzzi. A History of Modern Computing. History of computing. MIT Press, Cambridge, MA, USA, second edition, 2003. ISBN 0-262-53203-4 (paperback). xi + 445 pp. LCCN QA76.17.C47 2003.

Calavera:2020:LOB

[CF20] David Calavera and Lorenzo Fontana. Linux Observability with BPF: Advanced Programming for Performance Analysis and Networking. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2020. ISBN 1-4920-5020-2. xiii + 162 pp. LCCN QA76.774.L46. URL http://

proquest.safaribooksonline.
com/?fpi=9781492050193. Fore-

word by Jessie Frazelle.

Feng:2003:OGE

[cFJH^{+}03] Wu chun Feng, Justin, Hurwitz, Harvey B. Newman, Sylvain Ravot, Roger Les Cottrell, Olivier Martin, Fabrizio Coccetti, Cheng Jin, David Wei, and Steven Low. Optimizing 10-gigabit Ethernet in networks of workstations, clusters, and grids: A case study. In ACM [ACM03a], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/

inter_cal/inter_cal_detail.
php?eventid=10702#1; http:
//www.sc-conference.org/
sc2003/paperpdfs/pap293.
pdf.

Ceragioli:2024:SVI

[CGDB24] Lorenzo Ceragioli, Letterio Galletta, Pierpaolo Degano, and David Basin. Specifying and verifying information flow control in SELinux configurations. ACM Transactions on Privacy and Security (TOPS), 27(4):31:1–31:??, November 2024. CODEN ???? ISSN 2471-2566 (print), 2471-2574 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3690636.

Chen:2008:OVBa

[CGL^{+}08a] Xiaoxin Chen, Tal Garfinkel, E. Christopher Lewis, Pratap Subrahmanyam, Carl A. Waldspurger, Dan Boneh, Jeffrey Dwoskin, and Dan R. K. Ports. Overshadow: a virtualizationbased approach to retrofitting protection in commodity operating systems. ACM SIGARCH Computer Architecture News, 36(1):2–13, March 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Chen:2008:OVBb

[CGL^{+}08b] Xiaoxin Chen, Tal Garfinkel, E. Christopher Lewis, Pratap Subrahmanyam, Carl A. Waldspurger, Dan Boneh, Jeffrey Dwoskin, and Dan R. K. Ports.

Overshadow: a virtualizationbased approach to retrofitting protection in commodity operating systems. Operating Systems Review, 42(2):2–13, March 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chen:2008:OVBc

[CGL^{+}08c] Xiaoxin Chen, Tal Garfinkel, E. Christopher Lewis, Pratap Subrahmanyam, Carl A. Waldspurger, Dan Boneh, Jeffrey Dwoskin, and Dan R. K. Ports. Overshadow: a virtualizationbased approach to retrofitting protection in commodity operating systems. ACM SIGPLAN Notices, 43(3):2–13, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cadar:2008:EAG

[CGP^{+}08] Cristian Cadar, Vijay Ganesh, Peter M. Pawlowski, David L. Dill, and Dawson R. Engler. EXE: Automatically generating inputs of death. ACM Transactions on Information and System Security, 12(2): 10:1–10:??, December 2008. CODEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Chappuy:1999:GTV

[Cha99] Gilles Chappuy. Graver tous vos CD: Mac, PC, Plate-forme, Unix/Linux. Pro-micro. Osman Eyrolles Multimedia, Paris, France, 1999. ISBN 2-212-25009-6. 417 pp. LCCN ????

Chatterjee:2006:BML

[Cha06] Soumitra Chatterjee. 64-bit migration to Linux on Itanium: Challenges, advantages, and tools. In Anonymous [Ano06], page ?? ISBN ???? LCCN

????  URL http://www.
ice.gelato.org/;   http:/
/www.ice.gelato.org/about/
oct06_presentations.php.

Chao:2011:OSM

[Cha11] Lee Chao, editor. Open source mobile learning: mobile Linux applications. Information Science Reference, Hershey, PA, USA, 2011. ISBN 1-60960-613-2. ???? pp. LCCN LB1044.87 .O64 2011.

Chapelle:2013:BRO

[Cha13] Gregory Chapelle. Book review: The official Ubuntu book, seventh edition by Matthew Helmke and Amber Graner. ACM SIGSOFT Software Engineering Notes, 38(1): 54–55, January 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Chirillo:2002:HAD

[Chi02] John Chirillo. Hack attacks denied: a complete guide to network lockdown for UNIX, Windows, and Linux. Wiley, New York, NY, USA, second edition, 2002. ISBN 0-471-23283-1. xiii + 689 pp. LCCN QA76.9.A25 C47 2002.

Chuvakin:2003:UDU

[Chu03] Anton Chuvakin. Ups and downs of UNIX/Linux hostbased security solutions. ;login: the USENIX Association newsletter, 28(2):??, April 2003. CODEN LOGNEM. ISSN 1044-6397. URL http://

www.usenix.org/publications/
login/2003-04/pdfs/chuvakin.
pdf.

Capiluppi:2013:EEF

[CIC13] Andrea Capiluppi and Daniel Izquierdo-Cortázar. Effort estimation of FLOSS projects: a study of the Linux kernel. Empirical Software Engineering, 18(1):60–88, February 2013. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s10664-011-
9191-7.

Chryselius:2006:DQE

[CK06a] Toralf Chryselius and Andrea Kuntz. Debian unter Qemu Einführung in das Betriebssystem Debian Linux in der virtuellen Umgebung Qemu unter Windows. (German) [Debian under Qemu: Introduction in the Debian Linux operating systems in the Qemu virtual machine under Windows], volume 17 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-116-3 (book), 3-86768-716-1 (DVD). 159 pp. LCCN ????

Chryselius:2006:IDQ

[CK06b] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in Debian unter Qemu Einführung in das Betriebssystem Debian Linux in Qemu und Vorstellung der wichtigsten Internetprogramme. (German) [Internet Communication in Debian under Qemu: Introduction in the Debian Linux operating system in Qemu and creation of the most important Internet programs], volume 18 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-117-1 (book), 3-86768-717-X (DVD). 109 pp. LCCN ????

Chryselius:2006:IOV

[CK06c] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in OpenSUSE unter VMware [Qemu] Einführung in das Betriebssystem Open-SUSE Linux und Vorstellung von Internetprogrammen in der virtuellen Umgebung VMware, volume 66 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-165-1 (book). 117 pp. LCCN ????

Chryselius:2006:IOQ

[CK06d] Toralf Chryselius and Andrea Kuntz. Internetkommunikation mit OpenSUSE unter Qemu: Einführung in das Betriebssystem OpenSUSE Linux und Vorstellung von Internetprogrammen in der virtuellen Umgebung Qemu, volume 66 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-165-1. 104 pp. LCCN ????

Chryselius:2006:KLQb

[CK06e] Toralf Chryselius and Andrea Kuntz. Kanotix Linux unter Qemu Einführung in das Betriebssystem Kanotix Linux in der virtuellen Umgebung Qemu unter Windows, volume 33 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-132-5 (book), 3-86768-732-3 (DVD). 156 pp. LCCN ????

Chryselius:2006:KLQa

[CK06f] Toralf Chryselius and Andrea Kuntz. Knoppix Linux unter Qemu Einführung in das Betriebssystem Knoppix Linux in der virtuellen Umgebung Qemu unter Windows, volume 49 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-148-1 (book), 3-86768-748-X (DVD). 142 pp. LCCN ????

Chryselius:2006:KQE

[CK06g] Toralf Chryselius and Andrea Kuntz. Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu Linux in der virtuellen Umgebung Qemu, volume 5 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-104-X (book), 3-86768-704-8 (DVD). 158 pp. LCCN ????

Chryselius:2006:LDQ

[CK06h] Toralf Chryselius and Andrea Kuntz. Lernprogramme in Debian unter Qemu Einführung in das Betriebssystem Debian und Vorstellung von Lernprogrammen in der virtuellen Umgebung Qemu, volume 19 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-118-X (book), 3-86768-718-8 (DVD). 141 pp. LCCN ????

Chryselius:2006:LOL

[CK06i] Toralf Chryselius and Andrea Kuntz. Lernprogramme mit OpenSUSE Linux unter Qemu: Einführung in das Betriebssystem, OpenSUSE Linux und Vorstellung von Lernprogrammen in der virtuellen Umgebung Quemu, volume 63 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-166-X, 3-86768-766-8. 147 pp. LCCN ????

Chryselius:2006:OLQ

[CK06j] Toralf Chryselius and Andrea Kuntz. OpenSuSE Linux unter Qemu Einführung in das Betriebssystem Open-SUSE Linux in der virtuellen Umgebng Qemu unter Windows, volume 65 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-164-3 (book). 168 pp. LCCN ????

Chryselius:2006:SKD

[CK06k] Toralf Chryselius and Andrea Kuntz. Software für Kinder in Debian unter Qemu Einführung in das Betriebssystem Debian und Vorstellung der Lernund Spielesammlung Gcompris in der virtuellen Umgebung Qemu, volume 20 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-119-8 (book), 3-86768-719-6 (DVD). 113 pp. LCCN ????

Choi:2007:CPE

[CKE^{+}07] Gyu Sang Choi, Jin-Ha Kim, Deniz Ersoz, Andy B. Yoo, and Chita R. Das. A comprehensive performance and energy consumption analysis of scheduling alternatives in clusters. The Journal of Supercomputing, 40 (2):159–184, May 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=40&issue=2&spage=
159.

Crawford:1999:FFL

[CKM^{+}99] Diane Crawford, Leon Kappelman, Ken Mandefrot, Chuck Crawford, Dennis E. Hamilton, Orville E. Wheeler, Albert L. LeDuc, Max Hailperin, Michael E. Whitman, Anthony M. Townsend, Robert J. Aalberts, Charles Hixson, Daniel Berdichevsky, Erik Neunschwander, Bength Klenbergs, and Linux Torvalds. Forum: The folly of laws limiting Y2K liability; regarding Glass; duly noted; persuasive technologies; Linux goal. Communications of the ACM, 42(8): 11–15, August 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
cacm/1999-42-8/p11-crawford/

.

Craciunas:2008:RMT

[CKR08] Silviu S. Craciunas, Christoph M. Kirsch, and Harald Röck. I/O resource management through system call scheduling. Operating Systems Review, 42(5):44– 54, July 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Criswell:2007:SVA

[CLDA07] John Criswell, Andrew Lenharth, Dinakar Dhurjati, and Vikram Adve. Secure virtual architecture: a safe execution environment for commodity operating systems. Operating Systems Review, 41(6):351–366, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chae:2006:ATA

[CM06] Bongsug (Kevin) Chae and Roger McHaney. Asian trio’s adoption of Linux-based open source development. Communications of the ACM, 49(9): 95–99, September 2006. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Carrillo-Mondejar:2020:CLB

[CMMST20] J. Carrillo-Mondéjar, J. L. Martínez, and G. Suarez-Tangil. Characterizing Linuxbased malware: Findings and recent trends. Future Generation Computer Systems, 110(??):267–281, September 2020. CODEN FGSEVI.

ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X19325002.

Caviglione:2021:KLT

[CMR^{+}21] Luca Caviglione, Wojciech Mazurczyk, Matteo Repetto, Andreas Schaffhauser, and Marco Zuppelli. Kernel-level tracing for detecting stegomalware and covert channels in Linux environments. Computer Networks (Amsterdam, Netherlands: 1999), 191(??): ??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1389128621001249.

Cochran:2001:NVL

[Coc01] Shannon Cochran. News and views: Linux Standard Base 1.0 released; now everyone knows you’re a dog; maximum fiberoptic capacity calculated; mass producing quantum chips; mo’ better batteries; patenting the software service model. Dr. Dobb’s Journal of Software Tools, 26(10):18, October 2001. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Cochran:2003:NVe

[Coc03a] Shannon Cochran. News and views: MIT honors builder of robot swarm; developing space; Embedded Linux platform specification released; biometric systems at U.S. borders, says NIST. Dr. Dobb’s Journal of Software Tools, 28(5):14, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7927/ddj0305o/.

Cochran:2003:NVT

[Coc03b] Shannon Cochran. News and views: Tiny executable contest concludes; DeCSS, DMCA prosecutions fail; studying for the Turing test; Embedded Linux specification released. Dr. Dobb’s Journal of Software Tools, 28(3):14, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/documents/s=7826/
                                        ddj0303n/.

Colburn:1997:ADB

[Col97]  Scot Colburn.  An ATM

driver for the BT8233 and the Linux operating system. Thesis (m.s.), University of Colorado, Boulder, CO, USA, 1997. ix + 69 pp.

Collings:2002:RLN

[Col02] Terry Collings. Red Hat Linux networking and system administration. Transworld, London, UK, 2002. ISBN 0-7645-3632-X. xxix + 843 pp. LCCN QA76.76.O63 C625 2002.

Compton:2000:VLB

[Com00] Jason Compton. VMware 2 for Linux: [a better way to run multiple operating systems on Linux]. Prima Tech, Rocklin, CA, USA, 2000. ISBN 0-7615-2764-8. xxii + 406 pp. LCCN QA76.76.O63 C656 2000.

Compton:2003:VL

[Com03] Jason Compton. VMware 2 for Linux. Premier Press, Rocklin, CA, USA, 2003. ISBN 0-7615-2764-8. xxii + 406 pp. LCCN QA76.76.O63 C656 2000. US$39.99. Includes CD-ROM.

Cook:1998:LHF

[Coo98] Anthony L. Cook. Linuxhosted frame relay. SysAdmin, 7(1):8, 10, 14, 16, 18–19, January 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Cooperstein:2009:LPDa

[Coo09a] Jerry Cooperstein. Linux Program Development: a guide with exercises. CreateSpace, Scotts Valley, CA, USA, 2009. ISBN 1-4499-0602-8. 290 (est.) pp. LCCN ????

Cooperstein:2009:LPDb

[Coo09b] Jerry Cooperstein. Linux Program Development: Lab Solutions: a guide with exercises. CreateSpace, Scotts Valley, CA, USA, 2009. ISBN 1-4499-0604-4. 204 (est.) pp. LCCN ????

Cooperstein:2009:WLDa

[Coo09c] Jerry Cooperstein. Writing Linux Device Drivers: a guide with exercises, volume 3. CreateSpace, Scotts Valley, CA, USA, 2009. ISBN 1-4486-7238-4. 394 (est.) pp. LCCN ????

Cooperstein:2009:WLDb

[Coo09d] Jerry Cooperstein. Writing Linux Device Drivers: Lab Solutions: a guide with exercises. CreateSpace, Scotts Valley, CA, USA, 2009. ISBN 1-4495-3124-5. 270 (est.) pp. LCCN ????

Copp:2000:PP

[Cop00] Karlo Copp. Parametric perspectives. Journal of Linux Technology, 1(1):25–??, 2000. ISSN 1527-2761.

Cornes:1997:LZ

[Cor97] Phil Cornes. The Linux A-Z. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-234709-1 (paperback), 0-13-742867-7. xv + 531 pp. LCCN QA76.76.O63 C675 1997.

Coris:2000:CTP

[Cor00] Marie Coris. Compétition technologique et propriété intellectuelle: quels enjeux pour les logiciels libres ? le cas du système d’exploitation GNU/Linux. Master’s thesis, Mém. D.E.A.: écon. de l’environnement innovation et aménagement, Bordeaux 4, France, 2000.

LSBT:2005:BAL

[Cor05a] Core Members of the Linux Standard Base Team. Building applications with the Linux Standard Base. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2005. ISBN 0-13-145695-4. xxvi + 246 pp. LCCN QA76.76.O63 B8375 2004. URL http:

//www.freestandards.org/;
http://www.lanana.org/;
http://www.linuxbase.org/
;   http://www.linuxbase.
org/test/registered.html;
http://www.phptr.com/title/
0131456954;   https://www.
linux-foundation.org/en/

Book. Foreword by Theodore Ts’o. Includes CD-ROM.

Cornelio:2005:MLG

[Cor05b] Pietro Cornelio. Il mondo libero di GNU/Linux e UNIX BSD: storia, filosofia, tecnologia. (Italian) [The free world of GNU/Linux and BSD UNIX: stories, philosophies, technology]. Duke Italia, Milano, Italy, 2005. ISBN 88-86460-10-4. 104 pp. LCCN ???? Suppement to Linux Journal (Italian edition), number 59, October 2005.

Courtes:2017:CSG

[Cou17] Ludovic Courtès. Code staging in GNU Guix. ACM SIGPLAN Notices, 52(12):41–48, December 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crowcroft:2001:TIL

[CP01] Jon Crowcroft and Iain Phillips.

TCP/IP and the Linux protocol implementation: systems code for the Linux Internet. Wiley Networking Council series. Wiley, New York, NY, USA, 2001. ISBN 0-471-40882-4 (cloth). xlix + 925 pp. LCCN TK5105.585 .T34 2001.

Crowcroft:2002:TIL

[CP02] Jon Crowcroft and Iain Phillips. TCP/IP and Linux protocol implementation: systems code for the Linux Internet. Wiley Networking Council series. Wiley, New York, NY, USA, 2002. ISBN 0-471-40882-4 (cloth). xlix + 925 pp. LCCN TK5105.585 .T34 2002.

Clinton:1998:LBM

[CPJ^{+}98] Terry Clinton, Tom Parsons, Capers Jones, William Adams, Garth Klatt, Eric Haines, Ted Lewis, Philip Machanik, Stig Nilsson, Karl Reed, Howard R. Stearns, Neville Holmes, and John Brownie. Letters: The benefits of model-based integration; documentation is not green; picking on the overdog; buggy, slow windoze; there’s no such thing as free software; Linus’ law of open source development; bug-free development? no way; governmental IT planning and the Computer Society; text encoding questions; encoding the world’s languages. Computer, 31(11):4, 5– 7, 11, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1998/pdf/

ry004.pdf. Two letters discuss Unicode and Multicode [Mud97].

Corbet:2005:LDD

[CRKH05] Jonathan Corbet, Alessandro Rubini, and Greg Kroah-Hartman. Linux device drivers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2005. ISBN 0-596-00590-3. xviii + 615 pp. LCCN QA76.76.D49 R92 2005; QA76.76.D49 R92 2005eb; QA76.76.D49. URL http:/

/www.oreilly.com/catalog/
9780596005900.

Corcoran:1995:ISU

[CS95] Cate T. Corcoran and Mark Schlack. Internet servers for under $100? here’s one! Datamation, 41(12):41, 43, July 1, 1995. CODEN DTMNAT. ISSN 0011-6963.

Comer:2001:ITI

[CS01] Douglas E. Comer and David L.

Stevens. Internetworking with TCP/IP, Vol. 3: Client-Server Programming and Applications, Linux/Posix Sockets Version. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-032071-4. ???? pp. LCCN TK5105.585 .C66 2001. URL http:

//www.phptr.com/ptrbooks/
esm_0130320714.html.

Chivers:2006:CSD

[CS06] Ian D. Chivers and Jane Sleightholme. Compiler switches for developing Fortran programs. ACM Fortran Forum, 25(3):5–8, December 2006. CO-DEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Cusumano:2004:TSMc

[Cus04] Michael A. Cusumano. Technology strategy and management: Reflections on free and open software. Communications of the ACM, 47(10):25– 27, October 2004. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Camou:2000:DGL

[CV00] Mario Camou and Aaron Von Cowenberghe. Debian GNU/Linux 2.1 unleashed. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31700-1. xxvii + 1119 pp. LCCN QA76.76.O63 C355 2000.

Danesh:1999:ML

[Dan99a] Arman Danesh. Mastering Linux. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2341-4 (paperback). xxviii + 928 pp. LCCN QA76.76.O63D3428 1999.

Danesh:1999:MRL

[Dan99b] Arman Danesh. Mastering Red Hat Linux 6. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2613-

8. xxxvi + 905 pp. LCCN QA76.76.O63 D354 1999.

Danesh:2001:MCL

[Dan01] Arman Danesh. Mastering Corel Linux. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2852-1. xxix + 734 pp. LCCN QA76.76.O63 D3429 2001.

Datko:2014:BSA

[Dat14] Josh Datko. BeagleBone for Secret Agents. Community Experience Distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78398-604-2, 1-78398-605-0 (e-book), 1-322-15302-7 (e-book). LCCN QA76.774.L46. URL http:

//lib.myilibrary.com?id=
646557;   http://public.
eblib.com/choice/PublicFullRecord.
aspx?p=1797268.

Dawson:1994:PCL

[Daw94] J. Bruce Dawson. Power of cooperation: Linux is a freely available version of Unix, developed by individuals from all over the world. but it isn’t a bare-bones Unix clone. it is full of the features you would expect to find only in a commercial implementation of Unix. BYTE Magazine, 19(9):167– ??, September 1994. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

deBruijn:2008:PFL

[dBB08] Willem de Bruijn and Herbert Bos. PipesFS: fast Linux I/O in the Unix tradition. Operating Systems Review, 42(5):55– 63, July 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Du:2022:FIF

[DCH^{+}22] Xuechao Du, Andong Chen, Boyuan He, Hao Chen, Fan Zhang, and Yan Chen. AflIot: Fuzzing on Linux-based IoT device with binary-level instrumentation. Computers & Security, 122(??):Article 102889, November 2022. CO-DEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0167404822002838.

Dillig:2008:SCS

[DDA08] Isil Dillig, Thomas Dillig, and Alex Aiken. Sound, complete and scalable path-sensitive analysis. ACM SIGPLAN Notices, 43(6):270–280, June 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DDJStaff:1998:NVS

[DDJ98a] DDJ Staff. News and views: A Standard Linux? cryptography contest; drives get smaller and Smaller; Perl conference; really embedded systems; programmer shortage?; Beowulf: Linux clustering; Java SPEC released.

Dr. Dobb’s Journal of Software Tools, 23(11):16, November 1998. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

DDJ:1998:NVP

[DDJ98b] DDJ staff. News and views:

Programming contests announced; Y2K set in stone; USENIX notes; thinking about clusters; “it’s mine, all mine!”; vintage computers; fighting the programmer shortage; Linux: A dangerous weapon. Dr. Dobb’s Journal of Software Tools, 23(9):16, September 1998. CODEN DDJOEB. ISSN 1044-789X.

Dean:2001:LLC

[Dea01] Jeffrey Dean. LPI Linux certification in a nutshell: a desktop quick reference. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 1-56592-748-6. xvi + 551 pp. LCCN QA76.76.O63 D43 2001. Covers exams 101 and 102 for LPI level 1.

Decrem:2004:DLW

[Dec04] Bart Decrem. Desktop Linux:

Where art thou? ACM Queue: Tomorrow’s Computing Today, 2(3):48–56, May 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Dedeke:2009:LOL

[Ded09] Adenekan (Nick) Dedeke. Loyal opposition: Is Linux better than Windows software? IEEE Software, 26(3):104, 103, May/ June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Deno:1999:NRB

[Den99a] Cynthia Deno. New releases of *BSD and Debian Linux OSes given away at USENIX Annual Conference. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL http://www.
                                        usenix.org/publications/
                                        login/1999-6/newreleases.
                                        html.

Deno:1999:UTP

[Den99b] Cynthia Deno. USENIX teams up to put on the 1999 Atlanta Linux Showcase. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/
                                        login/1999-6/als.html.

Dutt:2000:GBG

[DF00] Christoph Dutt and Joachim Freiburg. GiMP: Bilder gestalten, Fotos retuschieren; [Grundlagen der professionellen Bildbearbeitung, der Umgang mit Fotos, Grafiken und Text, Bilder für das Internet richtig vorbereiten; auf der CD: GIMP für Windows, SCO Unix, Debian GNU Linux, Solaris, OS/2 und BSD, Quelltext aller GIMP- und GTK-Versionen, über 300 Plug-ins in C, Perl, tcl, Phyton und Scheme, XFree86/23.3.6, GIMP User Manual als PDF-Dateien]. C und L, Böblingen, Germany, 2000. ISBN 3-932311-64-7. 522 + 98 pp. LCCN ???? Includes CD-ROM.

Duan:2023:DLS

[DFC^{+}23] Guoyun Duan, Yuanzhi Fu, Minjie Cai, Hao Chen, and Jianhua Sun. DongTing: a large-scale dataset for anomaly detection of the Linux kernel. The Journal of Systems and Software, 203(??):??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0164121223001401.

Dent:2000:GUU

[DG00] Jack Dent and Tony Gaddis. Guide to UNIX: using Linux. Course Technology, Cambridge, MA, USA, 2000. ISBN 0-7600-1096-X. xiv + 552 pp. LCCN QA76.76.O63 D4 2000. Includes CD-ROM.

Dalheimer:2002:EDQ

[DH02] Matthias Kalle Dalheimer and Steffen Hansen. Embedded development with Qt/ Embedded: Linux for small systems. Dr. Dobb’s Journal of Software Tools, 27(3):48, 50, 52–54, March 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_03/qtembed.
                                        txt;   http://www.ddj.com/
                                        ftp/2002/2002_03/qtembed.
                                        zip.

Dubois:1996:NP

[DHH96] P. F. Dubois, K. Hinsen, and J. Hugunin. Numerical Python. Computers in Physics, 10(3): 262–267, May–June 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Dike:2006:UML

[Dik06]  Jeff Dike.  User Mode

Linux. Bruce Perens Open Source series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-186505-6 (paperback). ???? pp. LCCN QA76.76.O63 D545 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip068/2006004225.html.

Dodge:2005:SIL

[DIN05] Catherine Dodge, Cynthia Irvine, and Thuy Nguyen. A study of initialization in Linux and OpenBSD. Operating Systems Review, 39(2):79–93, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Danesh:2001:ML

[DJ01] Arman Danesh and Michael H. Jang. Mastering Linux. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 2001. ISBN 0-7821-2915-3. xxxv + 955 pp. LCCN QA76.76.O63 D3428 2001.

Ding:2007:BCM

[DJC07] Xiaoning Ding, Song Jiang, and Feng Chen. A buffer cache management scheme exploiting both temporal and spatial localities. ACM Transactions on Storage, 3(2):5:1–5:??, June 2007. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Dalton:2007:RFI

[DKK07] Michael Dalton, Hari Kannan, and Christos Kozyrakis. Raksha: a flexible information flow architecture for software security. ACM SIGARCH Computer Architecture News, 35(2): 482–493, May 2007. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Dolstra:2008:NPF

[DL08] Eelco Dolstra and Andres Löh.

NixOS: a purely functional Linux distribution. ACM SIG-PLAN Notices, 43(9):367–378, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dolstra:2010:NPF

[DLP10] Eelco Dolstra, Andres Löh, and Nicolas Pierron. NixOS: A purely functional Linux distribution. Journal of Functional Programming, 20(5– 6):577–615, November 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        product/C1ACBA2A51D2E5466820F5B5086EA2CE.

Do:1997:LEU

[DM97] James Do and Muhammed Mudawwar. Letters to the editor: Unicode misunderstood. Computer, 30(6):6, 9, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response and rebuttal to [Mud97].

Driscoll:2019:TTT

[DM19] K. Driscoll and J. Mailland. From trash to treasure: Turn a Minitel terminal into a Linux terminal or Videotex display — [resources hands on]. IEEE Spectrum, 56(3):16–17, March 2019. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

deOliveira:2023:OSN

[dOCC23] Daniel Bristot de Oliveira, Daniel Casini, and Tommaso Cucinotta. Operating system noise in the Linux kernel. IEEE Transactions on Computers, 72(1):196–207, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deOliveira:2016:TAP

[dOdO16] Daniel Bristot de Oliveira and Romulo Silva de Oliveira. Timing analysis of the PREEMPT RT Linux kernel. Software —Practice and Experience, 46 (6):789–819, June 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Drossu:1996:FGU

[DOF96] R. Drossu, Z. Obradovic, and J. Fletcher. A flexible graphical user interface for embedding heterogeneous neural network simulators. IEEE transactions on education, 39(3):367– 374, August 1996. CODEN IEEDAB. ISSN 0018-9359.

Donkers:1996:BLF

[Don96a] Arthur Donkers. Building a Linux firewall. SysAdmin, 5 (6):??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:LIS

[Don96b] Arthur Donkers. Linux as an Internet server. SysAdmin, 5 (1):??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1997:IL

[Don97a] Arthur Donkers. ISDN and Linux. SysAdmin, 6(10):51–53, 55, 57–60, October 1997. CO-DEN SYADE7. ISSN 1061-2688.

Donkers:1997:WLD

[Don97b] Arthur Donkers. Writing Linux device drivers. SysAdmin, 6(1): 8–??, January 1997. CODEN SYADE7. ISSN 1061-2688.

Donkers:1998:ULO

[Don98] Arthur Donkers. Using Linux in an office environment. SysAdmin, 7(1):45–46, 48, 50, 51, January 1998. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Donkers:1999:LE

[Don99] Arthur Donkers. Living on the edge. SysAdmin, 8(1):29–30, 33–35, January 1999. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Donato:2004:SPS

[Don04] Alberto Donato. A software platform to support dynamically reconfigurable systems-on-chip under the GNU/Linux operating system. Politecnico, Milano, Italy, 2004. ISBN ???? xvi + 95 pp. LCCN ????

Donenfeld:2022:RNG

[Don22] Jason A. Donenfeld. Random number generator enhancements for Linux 5.17 and 5.18. Web document, March

18, 2022. URL https://www.
zx2c4.com/projects/linux-
rng-5.17-5.18/.

DiBona:1999:OSV

[DOS99a] Chris DiBona, Sam Ockman, and Mark Stone, editors. Open Sources: Voices from the Open Source Revolution. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-582-3. viii + 272 pp. LCCN QA76.754 .O63 1999; QA76.76.S46 O64 1999. US$24.95. URL http:/

/www.oreilly.com/catalog/
9781565925823;   http://
www.oreilly.com/catalog/
opensources/.  Includes a

chapter The GNU Operating System and the Free Software Movement, by Richard Stallman, and a chapter Future of Cygnus Solutions: An Entrepreneur’s Account, by Michael Tiemann.

Doss:1999:LRH

[Dos99b] George M. Doss. Learn Red Hat Linux OS 5.2 tips. Wordware, Plano, TX, USA, 1999. ISBN 1-55622-715-9 (paperback). ???? pp. LCCN QA76.76.O63D683 1999.

Dougan:2002:TAR

[Dou02] Cort Dougan. Two-axis, realtime camera control. Dr. Dobb’s Journal of Software Tools, 27(10):60, 62, 64– 67, October 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_10/
rtlcam.txt;   http://www.
ddj.com/ftp/2002/2002_10/
rtlcam.zip.

Danelutto:2001:FSM

[DR01] M. Danelutto and A. Rampini.

Fast “short” messages on a Linux cluster. Lecture Notes in Computer Science, 2110:393– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2110/21100393.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2110/21100393.
pdf.

Duret:2002:HRF

[DRL^{+}02] Christian Duret, Francis Rischette, Joël Lattmann, Valéry Laspreses, Pim Van Heuven, Steven Van den Berghe, and Piet Demeester. High router flexibility and performance by combining dedicated lookup hardware (IFT), off the shelf switches and Linux. Lecture Notes in Computer Science, 2345:1117– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2345/23451117.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2345/23451117.
pdf.

Domingos:2023:SRV

[DRN^{+}23] Joao Mario Domingos, Tiago Rocha, Nuno Neves, Nuno Roma, Pedro Tomás, and Leonel Sousa. Supporting RISC-V performance counters through Linux performance analysis tools. In IEEE, editor, 2023 IEEE 34th International Conference on Applicationspecific Systems, Architectures and Processors (ASAP): ASAP 2023, 19–21 July 2023, Porto, Portugal. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, July 2023.

Danesh:2001:SLG

[DRP01] Arman Danesh, James Russell, and Richard Petersen. SAIR Linux & GNU certified administrator: exam guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213205-1. xxv + 1008 pp. LCCN QA76.76.O63 D342855 2001. URL ftp:

//uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/description/
mh022/2002277480.html;
http://www.loc.gov/catdir/
toc/mh021/2002277480.html.

Dulaney:2006:LSK

[Dul06]  Emmett Dulaney.  Linux

Starter Kit. SAMS Publishing, Indianapolis, IN, USA, 2006. ISBN 0-672-32887-9. ???? pp. LCCN ???? US$31.99.

Dulaney:2014:LAO

[Dul14] Emmett A. Dulaney. Linux allin-one for dummies. For dummies. Wiley, New York, NY, USA, fifth edition, 2014. ISBN 1-118-84435-1. xviii + 550 pp. LCCN QA76.76.O63 D8557

2014. URL http://www.loc.
gov/catdir/enhancements/
fy1506/2013958402-b.html;

                                         http://www.loc.gov/catdir/
                                        enhancements/fy1506/2013958402-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1506/2013958402-t.html.

Dumbill:2005:DGL

[Dum05] Edd Dumbill. Debain GNU/ Linux: An Explorer‘s Notebook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00883-X. 256 (est.) pp. LCCN ???? EUR 24.00.

Dunne:1999:RLA

[Dun99] Paul E. Dunne. Red Hat Linux 6.0 Application Development Tools. McGraw-Hill, New York, NY, USA, 1999. ISBN 0-07-134147-1. ca. 512 pp. LCCN ???? US$39.99.

Duntemann:2000:ALS

[Dun00] Jeff Duntemann. Assembly language step-by-step: programming with DOS and Linux. Wiley, New York, NY, USA, second edition, 2000. ISBN 0-471-37523-3 (paper/CD-ROM). xxv + 613 pp. LCCN QA76.73.A8 D88 2000. Revised edition of Assembly language, 1992.

Dupuy:2005:CBD

[Dup05] Stéphane Dupuy. Conception d’une base de données PostgreSQL sous Linux. Mémoire de diplôme d’ingénieur, Centre d’enseignement, Université de soutenance, Bordeaux, France, 2005. 99 pp.

Dongarra:1994:PSC

[DW94] Jack Dongarra and Jerzy Wasniewski, editors. Parallel scientific computing: First International Workshop, PARA ’94, Lyngby, Denmark, June 20–23, 1994: proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). LCCN QA76.58 .P35 1994. DM104.00.

Dalheimer:2006:RL

[DW06] Matthias Kalle Dalheimer and Matt Welsh. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2006. ISBN 0-596-00760-4 (paperback). xviii + 951 pp. LCCN QA76.76.O63 R855

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0715/2007270110-d.html.

Du:2008:SSI

[DW08] Wenliang Du and Ronghua Wang. SEED: A suite of instructional laboratories for computer security education. ACM Journal on Educational Resources in Computing (JERIC), 8(1):3:1–3:??, March 2008. CODEN ???? ISSN 1531-4278.

Davies:2006:SLB

[DWV06] Justin Davies, Roger Whittaker, and William Von Hagen. SUSE Linux 10 bible. Wiley, New York, NY, USA, 2006. ISBN 0-471-75488-9 (paperback/DVD). xlvii + 852 pp. LCCN QA76.76.O63 D34992 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0519/2005027885.html.

Eckstein:2000:US

[ECBK00] Robert Eckstein, David Collier-Brown, and Peter Kelly. Using Samba. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-449-5. xi + 398 pp. LCCN QA76.9.C55 E267 2000; QA76.9.C55 E267 2000; TK5105.5 .E35 2000. US$34.95.

                                        URL http://www.oreilly.
                                        com/catalog/9781565924499.

Includes CD-ROM with complete mirror of Samba site with Samba 2.0.5a, including sources, documentation, binaries, and supplementary tools.

Eckert:2011:LGL

[Eck11] Jason W. Eckert. Linux+ guide to Linux certification. Cengage Learning - Delmar, Clifton Park, NY, USA, third edition, 2011. ISBN 1-4188-3721-0, 1-111-54153-1. ???? pp. LCCN ????

El-Dirghami:2000:SLI

[EDK00] Nazeeh Amin El-Dirghami and Youssef A. Abu Kwaik. SuSE Linux Installation and Configuration Handbook. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2355-7. xx + 756 pp. LCCN QA76.76.O63 E4 2000. US$39.99. Includes CD-ROM.

Edwards:1998:ITC

[Edw98]  John Edwards.  Industry

trends: The changing face of freeware. Computer, 31(10): 11–13, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1998/pdf/
rx011.pdf.

Edmundsson:2004:DET

[EEK^{+}04] Niklas Edmundsson, Erik Elmroth, Bo Kågström, Markus Mårtensson, Mats Nylén, Åke Sandgren, and Mattias Wadenstein. Design and evaluation of a TOP100 Linux Super Cluster system. Concurrency and Computation: Practice and Experience, 16(8):735–750, July 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eilert:2003:LM

[EEMS03] John Eilert, Maria Eisenhaendler, Dorothea Matthaeu, and Ingol Salm, editors. Linux on the mainframe. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2003. ISBN 0-13-101415-3 (paperback). xxviii + 430 pp. LCCN QA76.76.O63 L54553 2003.

Emanuel:1994:ARE

[EJMT94] A. S. Emanuel, R. C. Jones, R. S. McWilliams, and J. C. Tsaio. Adaptation of reservoir engineering software to a DOS/ Linux PC. In Anonymous [Ano94d], pages 373–380.

Elboth:2001:LB

[Elb01] David Elboth. The Linux Book. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-032765-4. xxii + 501 pp. LCCN QA76.76.O63 E42 2001. US$34.99.

Eranian:2000:LIP

[EM00] Stéphane Eranian and David Mosberger. The Linux/ ia64 project: Kernel design and status report. Technical Report HPL-2000-85, Hewlett–Packard Corporation, Palo Alto, CA, USA, 2000.

                                        URL  http://www.hpl.hp.
                                        com/techreports/.

Emmons:2011:LOD

[Emm11] Jon Emmons. Linux for the Oracle DBA: the definitive reference, volume 40 of Oracle infocus series. Rampant Tech-Press, Kittrell, NC, USA, 2011. ISBN 0-9823061-9-9. ???? pp. LCCN ????

Eggert:2005:PEN

[EP05] P. R. Eggert and D. S. Parker.

Perturbing and evaluating numerical programs without recompilation — the wonglediff way. Software—Practice and Experience, 35(4):313–322, April 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Epplin:2000:IDH

[Epp00a] Jerry Epplin. Inside Debian Hurd. Dr. Dobb’s Journal of Software Tools, 25(12):21–22, 24, 26, December 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Epplin:2000:IRT

[Epp00b] Jerry Epplin. Inside Real-Time Linux. Dr. Dobb’s Journal of Software Tools, 25(3):72, 74, 76, 78, March 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_03/rtlinux.
txt;   http://www.ddj.com/
ftp/2000/2000_03/rtlinux.
zip.

Erickson:1995:DDJ

[Eri95] Jonathan Erickson. Dr. Dobb’s Journal Excellence In Programming Awards. Dr. Dobb’s Journal of Software Tools, 20(3): 18–??, March 1995. CODEN DDJOEB. ISSN 1044-789X.

Erickson:1997:DDN

[Eri97] Jonathan Erickson. Dr. Dobb’s news and views: Linux trademark issue settled; Inslaw ruling rejected; push over?; reading signs for the blind; free speech I; free speech II; Y2K insurance. Dr. Dobb’s Journal of Software Tools, 22(11): 16, November 1997. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
iplawyers.com/text/linux.

htm. A patent and trademark claim dispute on the name Linux has been resolved in favor of Linus Torvalds and the Linux community. A U.S. District Court Judge has rulled that source code is protected speech under the First Amendment to the U.S. Constitution, and that the U.S. Commerce Department acted illegally in requiring academics to obtains a government license before discussing cryptographic research with scholars on the Internet.

Etsion:2007:FGK

[ETKF07] Yoav Etsion, Dan Tsafrir, Scott Kirkpatrick, and Dror G. Feitelson. Fine grained kernel logging with KLogger: experience and insights. Operating Systems Review, 41(3):259– 272, June 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Fabrot:1997:LDU

[Fab97]  Bernard Fabrot.  Linux:

découverte et utilisation. Reference informatique. Marabout, Alleur, Belgique, 1997. ISBN 2-501-02867-8. 490 pp. LCCN ????

Farrow:1994:CUB

[Far94] Rik Farrow. Cheap and Uneasy: BSD/ 386 and Linux. UnixWorld’s Open Computing, 11(9):83–??, September 1994. CODEN OPCOEB. ISSN 1072-4044.

Farrow:2001:RLK

[Far01] Rik Farrow. A report on the Linux 2.5 Kernel Developers Summit. ;login: the USENIX Association newsletter, 26(3):??, June 2001. CODEN LOGNEM. ISSN 1044-6397. URL http:

//www.usenix.org/events/
kernel01/summit.pdf.

Fournier:2010:ABD

[FD10] Pierre-Marc Fournier and Michel R. Dagenais. Analyzing blocking to debug performance problems on multi-core systems. Operating Systems Review, 44(2):77–87, April 2010. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Feldman:1996:BLW

[Fel96a] Jonathan Feldman. Building a Linux Web server. SysAdmin, 5 (10):41–??, October 1996. CO-DEN SYADE7. ISSN 1061-2688.

Feldman:1996:ULR

[Fel96b]  Jonathan Feldman.  Using

Linux as a router. SysAdmin, 5 (1):??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Felten:2013:LBA

[Fel13] Edward Felten. The Linux backdoor attempt of 2003. Web site., 2013. URL https:

//freedom-to-tinker.com/
blog/felten/the-linux-backdoor-
attempt-of-2003/.

Frankston:1999:LHN

[FES^{+}99] Bob Frankston, Alec Effrat, Fred L. Schneider, Ed Viguerie, William D. Jackson, Jim Geringer, Kenneth Nichols, Alexander Fronk, Pearl Brereton, David Budgen, and Geoff Hamilton. Letters: The home network will be an IP network; pining for the perfect job; what the dickens make Thompson fault Linux?; UNIX Systems Laboratory; presto! A true test of AI; are software patents really different?; support for hypertext maintenance. Computer, 32(6): 6–9, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/
r6006.pdf.

Fernandez-Fuentes:2022:DFA

[FFPC22] Xosé Fernández-Fuentes, Tomás F. Pena, and José C. Cabaleiro. Digital forensic analysis methodology for private browsing: Firefox and Chrome on Linux as a case study. Computers & Security, 115(??):Article 102626, April 2022. CODEN CPSEDU.

ISSN 0167-4048 (print), 1872-6208 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167404822000256.

Fischer:2002:OEL

[Fis02] Todd Fischer. Optimizing Embedded Linux. Dr. Dobb’s Journal of Software Tools, 27 (5):51–57, May 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_05/emlinux.
txt.

Flickenger:2003:LSH

[Fli03] Rob Flickenger. Linux server hacks: 100 industrial-strength tips and tools. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00461-3. xvii + 221 pp. LCCN QA76.9.C55 F58 2003.

URL http://www.oreilly.
com/catalog/9780596004613.

Floyd:1994:CLT

[Flo94] Michael A. Floyd. A conversation with Linus Torvalds. Dr. Dobb’s Journal of Software Tools, 19(5):27–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Flowers:1999:LS

[Flo99] John S. Flowers. Linux security. New Riders professional series. New Riders Publishing, Carmel, IN, USA, 1999. ISBN 0-7357-0035-4. 400 pp. LCCN ???? Edited by Laurie Petrycki and others.

Flautner:2002:VAP

[FM02] Krisztián Flautner and Trevor Mudge. Vertigo: automatic performance-setting for Linux. Operating Systems Review, 36 (5S):105–116, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Frost:2007:GFS

[FMK^{+}07] Christopher Frost, Mike Mammarella, Eddie Kohler, Andrew de los Reyes, Shant Hovsepian, Andrew Matsuoka, and Lei Zhang. Generalized file system dependencies. Operating Systems Review, 41(6): 307–320, December 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Feng:2007:PUP

[FMR07] Hanhua Feng, Vishal Misra, and Dan Rubenstein. PBS: a unified priority-based scheduler. ACM SIGMETRICS Performance Evaluation Review, 35(1):203–214, June 2007. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Frincke:2006:ESI

[FOP06] D. Frincke, S. Oudekirk, and B. Popovsky. Editorial: Special issue on resources for the computer security and information assurance curriculum: Issue 1. ACM Journal on Educational Resources in Computing (JERIC), 6(3):1:1–1:??, September 2006. CODEN ???? ISSN 1531-4278.

Frazelle:2019:OSF

[Fra19] Jesse Frazelle. Open source firmware. Communications of the ACM, 62(10):34–38, October 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:

//cacm.acm.org/magazines/
2019/10/239673/fulltext.

Frisch:2003:ESAb

[Fri03] Æleen Frisch. Essential System Administration Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00449-4. vi + 137 pp. LCCN QA76.76.O63 F782 2003. US$14.95. URL http:/

/www.oreilly.com/catalog/
9780596004491;   http://
www.oreilly.com/catalog/
esapr.

Fei:2007:EOS

[FRRJ07] Yunsi Fei, Srivaths Ravi, Anand Raghunathan, and Niraj K. Jha. Energy-optimizing source code transformations for operating system-driven embedded software. ACM Transactions on Embedded Computing Systems, 7(1):2:1–2:26, December 2007. CODEN ????

ISSN 1539-9087 (print), 1558-3465 (electronic).

Fu:2009:OPS

[FST^{+}09] Xiaoming Fu, Henning Schulzrinne, Hannes Tschofenig, Christian Dickmann, and Dieter Hogrefe. Overhead and performance study of the General Internet Signaling Transport (GIST) protocol. IEEE/ACM Transactions on Networking, 17(1): 158–171, February 2009. CO-DEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Fink:2002:LPT

[FSW02] Jason Fink, Matt Sherer, and Kurt Wall. Linux performance tuning and capacity planning. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32081-9. xi + 317 pp. LCCN QA76.76.O63 F557 2002.

Frazier:1995:LSL

[FT95] Belinda Frazier and Laurie Tucker. The Linux sampler: a Linux resource guide. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1995. ISBN 0-916151-74-3 (paperback). viii + 240 pp. LCCN QA76.76.O63 L46 1995.

Figgin:2003:LWS

[FWS^{+}03] Stephen Figgin, Aaron Weber, Ellen Siever, Matt Welsh, George Reese, Ben Laurie, Peter Laurie, Rasmus Lerdorf, Stas Bekman, and Eric Cholet, editors. The Linux Web server CD bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, version 2.0. edition, 2003. ISBN 0-596-00529-6. LCCN QA76.76.O63.

Ganten:2000:DGL

[G^{+}00] Peter H. Ganten et al. Debian GNU, Linux: Grundlagen, Installation, Administration und Anwendung. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-65841-6, 3-540-66384-3. xvi + 792 pp. LCCN ????

Ganten:2002:DGL

[G^{+}02] Peter H. Ganten et al. Debian GNU/Linux Power Pack: Grundlagen, Installation, Administration und Anwendung. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2002. ISBN 3-540-43250-7, 3-540-43267-1. 800 (est.) pp. LCCN ???? Includes four CD-ROMs.

Geiselhart:2006:IZV

[G^{+}06a] Gregory Geiselhart et al., editors. IBM z/VM and Linux on IBM System z: virtualization cookbook for Red Hat Enterprise Linux 4. Number SG24-7272-00 in IBM redbooks. IBM Corporation, San Jose, CA, USA, 2006. ISBN 0-7384-9495-X (paperback). xiv + 218 pp. LCCN QA76.76.O63 I28 2006.

Gerner:2006:PLL

[G^{+}06b] Jason Gerner et al., editors. Professional LAMP: Linux, Apache, MySQL, and PHP Web development. Wiley, New York, NY, USA, 2006. ISBN 0-7645-9723-X. xxii + 379 pp. LCCN TK5105.888 .P677

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0623/2005026487-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0654/2005026487-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0654/2005026487-t.html.

Ganten:2004:DGLa

[GA04] Peter H. Ganten and Wulf Alex. Debian GNU/Linux: Grundlagen, Installation, Administration und Anwendung. (German) [Debian GNU/Linux: Basis, Installation, Administration, and Use]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-43267-1. xxii + 946 pp. LCCN ????

Gabarro:2007:WAD

[Gab07]  Steven A. Gabarró.  Web

application design and implementation: Apache 2, PHP5, MySQL, JavaScript, and Linux/Unix. Quantitative software engineering series. Wiley, New York, NY, USA, 2007. ISBN 0-471-77391-3 (cloth). xv + 295 pp. LCCN TK5105.8883 .G33 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0613/2006014999.html.

Gagne:2000:EDS

[Gag00a] Marcel Gagné. EnlightenDSM — distributed systems management. SysAdmin, 9(1): 41–42, 44–46, 48, January 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Gagne:2000:LES

[Gag00b] Marcel Gagné. A Linux email server. SysAdmin, 9(2):53– 58, February 2000. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Gagne:2001:LSA

[Gag01] Marcel Gagné. Linux System Administration — A User’s Guide. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-71934-7. xxi + 532 pp. LCCN QA76.76.O63 G34 2002.

Gagne:2006:MUL

[Gag06] Marcel Gagné. Moving to Ubuntu Linux. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-42722-X (paperback). xxii + 463 pp. LCCN QA76.76.O63 .G3455 2006. US$34.99.

URL http://www.aw-bc.com/
catalog/academic/product/
0,1144,032142722X,00.html.

Includes CD-ROM.

Gagne:2007:MUL

[Gag07] Marcel Gagné. Moving to Ubuntu Linux. Addison-Wesley, Reading, MA, USA, 2007. ISBN 0-321-42722-X (paperback). xxii + 463 pp. LCCN QA76.76.O63 G3455 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0616/2006021595.html.

German:2016:CMD

[GAH16] Daniel M. German, Bram Adams, and Ahmed E. Hassan. Continuously mining distributed version control systems: an empirical study of how Linux uses Git. Empirical Software Engineering, 21(1): 260–299, February 2016. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s10664-014-
9356-2.

Gancarz:2003:LUP

[Gan03] Mike Gancarz. Linux and the Unix Philosophy. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2003. ISBN 1-55558-273-7. xxvii + 220 pp. LCCN QA76.76.O63G364 2003. US$34.99. URL http://

www.loc.gov/catdir/description/
els031/2003051482.html;
http://www.loc.gov/catdir/
toc/els031/2003051482.html.

Ganten:2004:DGLb

[Gan04] Peter H. Ganten. Debian-GNU-Linux-Powerpack. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-66384-3. ???? pp. LCCN ????

Ganti:2008:PAL

[Gan08] Ashwin Ganti. Plan 9 authentication in Linux. Operating Systems Review, 42(5):27–33, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gardner:1998:SPJ

[Gar98] Dana Gardner. Sun ports JDK 1.2 to Linux in push for ubiquitous Java. JavaWorld: IDG’s magazine for the Java community, 3(11):??, 1998. CO-DEN ???? ISSN 1091-8906.

URL http://www.javaworld.
com/javaworld/jw-11-1998/
jw-11-iw-jdklinux.htm.

Garber:1999:NBL

[Gar99]  Lee Garber.  News briefs:

Linux support ranges from GUI to Big Blue; new XML standards; partnership for advanced computational infrastructure. Computer, 32(5): 20–22, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/
r5020.pdf.

Garzik:2000:GCR

[Gar00] Jeff Garzik. Glibc: a comprehensive reference to GNU/ Linux libC. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 1-57870-202-X. ???? pp. LCCN ???? Edited by Laurie Petrycki and others.

Gardinier:2009:OSD

[Gar09] Mark Gardinier. Open source development of a safety critical dual redundant (ada95/ C++) signal control program environment (SCOPE). ACM SIGADA Ada Letters, 29(3): 23–30, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gatwood:2001:CCL

[Gat01] David A. Gatwood. Cleaning the cleaner for the Linux logstructured file system. Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2001.

Gay:1999:STY

[Gay99] Warren W. Gay. Sams’ teach yourself Linux programming in 24 hours: complete starter kit. Macmillan Computer Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31582-3. xviii + 526 pp. LCCN QA76.76.O63 G398 1999. Edited by Brian Gill, Ron Gallagher, and Gus Miklos.

Gay:2000:LSP

[Gay00] Warren W. Gay. Linux Socket Programming by Example. Que Corporation, Indianapolis, IN, USA, 2000. ISBN ???? ???? pp. LCCN ????

Grahlmann:1996:PTP

[GB96] B. Grahlmann and E. Best. PEP — more than a Petri net tool. In T. Margaria and B. Steffen, editors, Tools and Algorithms for the Construction and Analysis of Systems. Second International Workshop, TACAS ’96. Proceedings, volume 1055 of Lecture notes in computer science, pages ??– ?? (of xi + 434). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61042-1. LCCN QA76.9.S88 T33 1996.

Gill:2007:OMS

[GB07] Binny S. Gill and Luis Angel D. Bathen. Optimal multistream sequential prefetching in a shared cache. ACM Transactions on Storage, 3(3):10:1– 10:??, October 2007. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Gomes:2018:ERL

[GBC^{+}18] Jorge Gomes, Emanuele Bagnaschi, Isabel Campos, Mario David, Luís Alves, João Martins, João Pina, Alvaro López-García, and Pablo Orviz. Enabling rootless Linux Containers in multi-user environments: the udocker tool. Computer Physics Communications, 232(??):84–97, November 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0010465518302042.

Gelado:2010:ADS

[GCN^{+}10] Isaac Gelado, Javier Cabezas, Nacho Navarro, John E. Stone, Sanjay Patel, and Wen mei W. Hwu. An asymmetric distributed shared memory model for heterogeneous parallel systems. ACM SIGPLAN Notices, 45(3):347–358, March 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goel:1996:LDD

[GD96] S. Goel and D. Duchamp. Linux device driver emulation in Mach. In USENIX Association [USE96b], pages 65–73. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

Gebai:2018:SAK

[GD18] Mohamad Gebai and Michel R. Dagenais. Survey and analysis of kernel and userspace tracers on Linux: Design, implementation, and overhead. ACM Computing Surveys, 51 (2):26:1–26:??, June 2018. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Gehrke:1998:La

[Geh98] Frank Gehrke. Linux. Mega livre. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, nouvelle edition, 1998. ISBN 2-7361-3120-7. 668 pp. LCCN ????

Georgiev:1994:MDL

[Geo94] L. Georgiev. A MIDI driver for the Linux kernel — methods of beta-testing. In Popov [Pop94], pages 154–158. ISBN 954-8329-06-9 (??invalid checksum??). LCCN TJ212.2 .I572 1994.

Georgescu:1997:LX

[Geo97] Emil Georgescu. Linux et XFree86. Eyrolles, Paris, France, 1997. ISBN 2-212-08953-8. x + 404 pp. LCCN ????

Guettaf:1996:LDT

[GFRD96] A. Guettaf, O. Florent, E. Rejouan, and A. Derieux. A low-cost digital tester based on boundary-scan techniques. In G. Kamarinos, N. Guillemot, and B. Courtois, editors, Proceedings of the European Workshop Microelectronics Education: Grenoble, France, 5–6 Feb 1996, pages ??–?? (of xv + 302). World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2653-5. LCCN TK7874 .E8 1996.

Gehrke:1998:Lb

[GGG^{+}98] Frank Gehrke, Peter Glinsky, Forbjon Gripp, Michel Joinville, Trad., and Jean-Louis Greco, Trad. Linux. Mega livre. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1998. ISBN 2-7361-2795-1. xii + 630 pp. LCCN ????

Gehrke:1999:L

[GGG^{+}99] Frank Gehrke, Peter Glinsky, Forbjon Gripp, Michele Joinville, and Jean-Louis Greco. Linux. Mega livre. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, nouvelle edition, 1999. ISBN 2-7361-3218-1. xii + 704 pp. LCCN ????

Gunther:2002:LUR

[GGZ02] Karsten Günther, Kester Grelck, and Thorsten Zilm, editors. Linux User Referenz: [Linux Drucksysteme: BSD und CUPS: Steuerung des Prozess-Systems: CDs erstellen und Textdateien bearbeiten: Pager, Formatierung und die glimpse-Tools]. (German) [Linux User Reference: Linux Printing Systems: BSD and CUPS: System Process Control: CD creation and text file processing; Pagination, Formatting and glimpse tools]. mitp Verlag, Bonn, Germany, 2002. ISBN 3-8266-0709-0. 1055 (est.) pp. LCCN ????

Gellerich:2004:GBP

[GHL^{+}04] W. Gellerich, T. Hendel, R. Land, H. Lehmann, M. Mueller, P. H. Oden, and H. Penner. The GNU 64-bit PL8 compiler: Toward an open standard environment for firmware development. IBM Journal of Research and Development, 48 (3/4):543–556, ???? 2004. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
rd/483/gellerich.html;
http://www.research.ibm.
com/journal/rd/483/gellerich.
pdf.

Guttman:2005:VIF

[GHRS05] Joshua D. Guttman, Amy L.

Herzog, John D. Ramsdell, and Clement W. Skorupka. Verifying information flow goals in Security-Enhanced Linux. Journal of Computer Security, 13(1):115–134, ???? 2005. CO-DEN JCSIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Giampaolo:1999:PFS

[Gia99] Dominic Giampaolo. Practical file system design with the BE file system. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-497-9. x + 237 pp. LCCN QA76.9.F5 G49 1999. US$34.95. Includes comparison with Apple Macintosh, Linux, and Microsoft Windows file systems.

Gibbs:2017:BYO

[Gib17] W. W. Gibbs. Build your own Amazon Echo — turn a PI into a voice controlled gadget [resources hands on]. IEEE Spectrum, 54(5):20–21, May 2017. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gigante:2006:HPS

[Gig06] Mike Gigante. High-performance storage solutions on IA-64 Linux. In Anonymous [Ano06], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        ice.gelato.org/;   http:/
                                        /www.ice.gelato.org/about/
                                        oct06_presentations.php.

Gift:2008:PUL

[GJ08] Noah Gift and Jeremy M. Jones. Python for Unix and Linux system administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51582-0. xix + 433 pp. LCCN QA76.73.P98 G54 2008.

Ganapathy:2005:APA

[GJJ05] Vinod Ganapathy, Trent Jaeger, and Somesh Jha. Automatic placement of authorization hooks in the Linux security modules framework. In Meadows and Syverson [MS05], pages 330–339. ISBN 1-59593-226-7. LCCN QA76.9.A25. ACM order number 459050.

Lehey:2002:WBB

[gL02] Greg “groggy” Lehey. Why BSD is better than Linux. IBM Linux Technology Center, Ozlabs, Brisbane, Queensland, Australia, February 7, 2002.

                                        URL http://www.lemis.com/

linux.conf.au.pdf. 197-slide

conference talk.

Glass:1999:NBL

[Gla99] Robert L. Glass. News briefs:

Loyal opposition: Of open source, Linux . . . . . . and hype. IEEE Software, 16 (1):128, 126–127, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1999/pdf/
s1128.pdf.

Gropp:2003:BCC

[GLS03] William Gropp, Ewing Lusk, and Thomas Lawrence Sterling, editors. Beowulf cluster computing with Linux. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, second edition, 2003. ISBN 0-262-69292-9 (paperback). xxxix + 618 pp. LCCN QA76.58.B46 2003; QA76.58.B46.

Guniguntala:2008:RCU

[GMTW08] D. Guniguntala, P. E. McKenney, J. Triplett, and J. Walpole. The read-copy-update mechanism for supporting real-time applications on shared-memory multiprocessor systems with Linux. IBM Systems Journal, 47(2):221–??, April/June 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/472/guniguntala.html.

Gulati:2007:PAC

[GMV07] Ajay Gulati, Arif Merchant, and Peter J. Varman. pClock: an arrival curve based approach for QoS guarantees in shared storage systems. ACM SIG-METRICS Performance Evaluation Review, 35(1):13–24, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Guanciale:2016:PSM

[GNDB16] Roberto Guanciale, Hamed Nemati, Mads Dam, and Christoph Baumann. Provably secure memory isolation for Linux on ARM. Journal of Computer Security, 24(6):793– 837, ???? 2016. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Goerzen:1999:DGL

[GO99] John Goerzen and Ossama Othman. Debian GNU/Linux: Guide to Installation and Usage. New Riders Publishing, Carmel, IN, USA, 1999. ISBN 0-7357-0914-9. 158 pp. LCCN QA76.76.O63 .G634 1999. US$24.99. URL http:/

                                        /www.newriders.com/books/
                                        title.cfm?isbn=0735709149.

Goerzen:2000:LPB

[Goe00] John Goerzen. Linux Programming Bible. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-4657-0. 837 pp. LCCN QA76.76.O63 G6343 2000. US$39.99.

Goldberg:1995:LCR

[Gol95] Howard L. Goldberg. The Linux commands and resources directory. Cruising the Cutting Edge Pub., Houston, TX, USA, 1995. ISBN 0-00-017192-1. 360 pp. LCCN QA76.76.O63 G6373 1995. Version 1.2.4.

Goncalves:1999:LAW

[Gon99] Marcus Goncalves. Linux at work: building strategic applications for business. Wiley, New York, NY, USA, 1999. ISBN 0-471-33349-2 (paperback). xx + 363 pp. LCCN HF5548.4.L5G66 1999. Includes CD-ROM.

Grace:2000:LG

[GP00] Rich Grace and Tim Parker.

Linux to go. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-999269-3. 300 pp. LCCN QA76.76.O63G415

2000.  URL http://www.
phptr.com/ptrbooks/ptr_0139992693.
html.

Gunawi:2007:IFS

[GPK^{+}07] Haryadi S. Gunawi, Vijayan Prabhakaran, Swetha Krishnan, Andrea C. Arpaci-Dusseau, and Remzi H. Arpaci-Dusseau. Improving file system reliability with I/O shepherding. Operating Systems Review, 41(6):293–306, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Garrison:2009:UFS

[GR09] John A. Garrison and A. L. Narasimha Reddy. Umbrella File System: Storage management across heterogeneous devices. ACM Transactions on Storage, 5(1): 3:1–3:??, March 2009. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Gray:1999:SMR

[Gra99] Bob Gray. Software minireview: Red Hat Linux 6.0. ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL http://www.
                                        usenix.org/publications/
                                        login/1999-8/features/redhat.
                                        html.

Grant:2004:LNG

[Gra04] Rickford Grant. Linux for Non-Geeks. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-034-8. xx + 308 pp. LCCN QA76.76.O63 G723 2004. URL http:/

                                        /www.oreilly.com/catalog/

1593270348/. Includes two CD-ROMs with Fedora.

Grant:2005:LME

[Gra05] Rickford Grant. Linux made easy: the official guide to Xandros 3 for everyday users. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-057-7. xxvi + 463 pp. LCCN QA76.76.O63 G7245 2005.

Granneman:2006:LP

[Gra06a]  Scott Granneman.  Linux

Phrasebook. SAMS Publishing, Indianapolis, IN, USA, 2006. ISBN 0-672-32909-3; 0-672-32838-0 (paperback). xvii + 382 pp. LCCN QA76.76.O63 G72 2006. EUR 13.90.

Grant:2006:ULN

[Gra06b]  Rickford Grant.  Ubuntu

Linux for non-geeks: a painfree, project-based, get-things-done guidebook. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-118-2. xx + 334 pp. LCCN QA76.76.O63 G7246 2006eb.

URL http://site.ebrary.
com/lib/ucmerced/Doc?id=
10137836;   http://site.
ebrary.com/lib/ucsc/Doc?
id=10137836;   http://www.
loc.gov/catdir/enhancements/
fy0715/2006015576-d.html;
http://www.loc.gov/catdir/
toc/ecip0613/2006015576.
html.

Greenfield:1994:LUG

[Gre94] Larry Greenfield. The Linux user’s guide. ????, ????, 1994. ix + 150 pp.

Gregg:2019:BPT

[Gre19] Brendan Gregg. BPF Performance Tools: Linux System and Application Observability. Addison-Wesley Professional, ????, 2019. ISBN 0-13-655482-2, 0-13-658887-5. LCCN ???? Foreword by Alexei Starovoitov, creator of the new BPF.

Griffith:2001:KQP

[Gri01a] Arthur Griffith. KDE/QT programming bible. I D G Books Worldwide, Indianapolis, IN, USA, 2001. ISBN 0-7645-4682-1. ???? pp. LCCN QA76.73.C153 .G7426 2001.

Grinzo:2001:PBP

[Gri01b] Lou Grinzo. Programmer’s bookshelf: Professional Linux Programming. Dr. Dobb’s Journal of Software Tools, 26 (4):151–152, April 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Grimmett:2013:BRP

[Gri13] Richard Grimmett. Beagle-Bone robotic projects: create complex and exciting robotic projects with the BeagleBone Black. Community experience distilled. Packt Publishing, Birmingham, UK, 2013. ISBN 1-78355-932-2, 1-78355-933-0 (e-book), 1-306-28021-4 (e-book). iii + 228 pp. LCCN TJ211.45 .G75 2013.

URL http://proquest.tech.
safaribooksonline.de/9781783559329.

Grimmett:2014:MBR

[Gri14] Richard Grimmett. Mastering BeagleBone robotics: master the power of the Beagle-Bone Black to maximize your robot-building skills and create awesome projects. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78398-890-8 (paperback), 1-78398-891-6 (ebook), 1-322-55407-2 (e-book). iii + 217 pp. LCCN TJ211.45 .G7566 2014.

Grinzo:2000:PBG

[GS00] Lou Grinzo and Jacques Surveyer. Programmer’s bookshelf: GNU tools and process patterns. Dr. Dobb’s Journal of Software Tools, 25(12):159– 160, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Gine:2003:DMC

[GSHL03] Francesc Giné, Francesc Solsona, Porfidio Hernández, and Emilio Luque. Dealing with memory constraints in a nondedicated Linux cluster. The International Journal of High Performance Computing Applications, 17(1):39–48, Spring 2003. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Gine:2001:MMM

[GSN^{+}01] Francesc Giné, Francesc Solsona, Xavi Navarro, Porfidio Hernández, and Emilio Luque. MemTo: A memory monitoring tool for a Linux cluster. Lecture Notes in Computer Science, 2131:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2131/21310225.
htm;   http://link.springer-

                                        ny.com/link/service/series/
                                        0558/papers/2131/21310225.
                                        pdf.

Guerrieri:2019:HSP

[Gue19] Alex Guerrieri. Hands-on system programming with Go: build modern and concurrent applications for Unix and Linux systems using Golang. Packt Publishing, Birmingham, 2019. ISBN 1-78980-336-5 (e-book), 1-78980-407-8. xii + 437 pp. LCCN QA76.73.G63. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781789804072.

Gunther:2002:LGP

[Gün02] Karsten Günther. Linux gepackt: [die wichtigsten Befehle für GNU/Linux-Systeme; Textwerkzeuge: Drucken, Konvertieren, Suchen und Anpassen; CDs erstellen unter Linux]. MITP Verlag, Bonn, Germany, 2002. ISBN 3-8266-0946-8. 560 (est.) pp. LCCN ???? EUR 14.95 (DE), EUR 15.40 (AT).

Gerndt:2003:LSC

[GVK^{+}03] Andreas Gerndt, Thomas Van Reimersdahl, Torsten Kuhlen, Christian Bischof, Ingolf Hörschler, Matthias Meinke, and Wolfgang Schröder. Large-scale CFD data handling in a VR-based otorhinolaryngological CAS-system using a Linux-cluster. The Journal of Supercomputing, 25(2):143–154, June 2003.

CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//ipsapp009.kluweronline.
com/content/getfile/5189/
44/5/abstract.htm;   http:
//ipsapp009.kluweronline.
com/content/getfile/5189/
44/5/fulltext.pdf.

Geist:1994:DSL

[GW94] R. M. Geist and J. M. Westall. Disk scheduling in Linux. In Anonymous [Ano94b], pages 739–746.

Gorbachev:2001:DSA

[GZK^{+}01] Y. E. Gorbachev, M. A. Zatevakhin, V. V. Krzhizhanovskaya, A. A. Ignatiev, V. K. Protopopov, N. V. Sokolova, and A. B. Witenberg. Distributed simulation of amorphous hydrogenated silicon films: Numerical experiments on a Linux based computing environment. Lecture Notes in Computer Science, 2073:483– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2073/20730483.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2073/20730483.
pdf.

Gatla:2023:UPM

[GZXZ23] Om Rameshwar Gatla, Duo Zhang, Wei Xu, and Mai Zheng. Understanding persistentmemory-related issues in the Linux kernel. ACM Transactions on Storage, 19(4):36:1– 36:??, November 2023. CO-DEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3605946.

Hill:2007:OUBa

[H^{+}07a] Benjamin Mako Hill et al., editors. The official Ubuntu book. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2007. ISBN 0-13-243594-2 (paperback). xxxiv + 412 pp. LCCN QA76.76.O63 O34348 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0613/2006016172.html.

Hill:2007:OUBb

[H^{+}07b] Benjamin Mako Hill et al., editors. The official Ubuntu book. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2007. ISBN 0-13-235413-6 (paperback). xlii + 463 pp. LCCN QA76.76.O63 O34348 2007a. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0717/2007018467.html.

VonHagen:2000:IRL

[Hag00] Bill Von Hagen. Installing Red Hat Linux 7. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31826-1. 198 pp. LCCN QA76.76.O63 V66 2000.

Hahn:2009:HHG

[Hah09] Harley Hahn. Harley Hahn’s guide to Unix and Linux.

McGraw-Hill Higher Education, Boston, MA, USA, 2009. ISBN 0-07-313361-2 (paperback). xxxiv + 926 pp. LCCN QA76.76.O63 H3378

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0808/2007052525-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0808/2007052525-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0810/2007052525-t.html.

Haletky:2005:DLD

[Hal05] Edward Haletky. Deploying Linux on the desktop. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2005. ISBN 1-55558-328-8. ???? pp. LCCN QA76.76.O63 H34345 2005.

Hallinan:2006:ELP

[Hal06] Christopher Hallinan. Embedded Linux primer: a practical, real-world approach. Prentice Hall open source software development series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-167984-8 (paperback). xxix + 537 pp. LCCN QA76.76.O63 H34462 2006. US$44.99. URL http:

//www.loc.gov/catdir/toc/
ecip0612/2006012886.html.

Hallinan:2007:ELP

[Hal07] Christopher Hallinan. Embedded Linux primer: a practical, real-world approach. Prentice Hall open source software development series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2007. ISBN 0-13-167984-8 (paperback). xxix + 537 pp. LCCN QA76.76.O63 H34462 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0612/2006012886.html.

Hantelmann:1997:LSG

[Han97] Fred Hantelmann. Linux Start-Up Guide: a self-contained introduction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-62676-X. xvi + 344 pp. LCCN QA76.76.O63H3475 1997. US$28.00.

Hantelmann:1999:LD

[Han99] F. Hantelmann. LINUX fur Durchstarter. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2., vollst. ubera. edition, 1999. ISBN 3-540-65215-9. ca. 480 pp. LCCN ????

Hancock:2000:NCU

[Han00] Bill Hancock. New classes of Unix/Linux attacks. Computers & Security, 19(7): 570–571, November 1, 2000. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167404800070024.

Harlow:1999:DLA

[Har99] Eric Harlow. Developing Linux applications with GTD+ and GDK. New Riders Publishing, Carmel, IN, USA, 1999.

ISBN 0-7357-0021-4. xii + 488 pp. LCCN QA76.76.O63H3487 1999. Edited by Laurie Petrycki and others.

Hawkins:2000:LDR

[Haw00] Scott Hawkins. Linux desk reference. Open source technology series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-016391-0. xxvii + 545 pp. LCCN QA76.76.O63 H38 2000 Sci-Eng.

Hawkins:2002:LDR

[Haw02] Scott Hawkins. Linux desk reference. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2002. ISBN 0-13-061989-2. xxi + 583 pp. LCCN QA76.76.O63 H386 2002.

Hill:2006:OUB

[HBB^{+}06] Benjamin Mako Hill, Jono Bacon, Corey Burger, Jonathan Jesse, and Ivan Krstić, editors. The official Ubuntu book. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-243594-2 (paperback). 320 (est.) pp. LCCN QA76.76.O63 O34348 2006. US$34.99. URL http://

vig.prenhall.com/catalog/
academic/product/0,1144,
0132435942,00.html; http:
//www.loc.gov/catdir/toc/
ecip0613/2006016172.html.

Harji:2013:OTL

[HBB13] Ashif S. Harji, Peter A. Buhr, and Tim Brecht. Our troubles with Linux Kernel upgrades and why you should care. Operating Systems Review, 47(2): 66–72, July 2013. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Huang:2015:LNS

[HC15] A. Huang and S. Cross. A laptop with no secrets. IEEE Spectrum, 52(11):40–56, November 2015. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hare:1995:BLI

[HE95] R. Christopher E. Hare and George Eckel. Building a Linux Internet server. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-525-9. xiv + 340 pp. LCCN QA76.9.C55 H36 1995.

Heath:1997:EL

[Hea97] Steve Heath. Essential Linux. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1997. ISBN 1-55558-177-3. vii + 257 pp. LCCN QA76.76.O63H42 1997.

Hein:1999:LCS

[Hei99] Jochen Hein. Linux companion for system administrators. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-36044-6. xvii + 501 pp. LCCN QA76.76.O63H4523 1999.

Hekman:1997:LND

[Hek97] Jessica Perry Hekman. Linux in a Nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-167-4. xi + 424 pp. LCCN QA76.76.O63 H453 1997. US$19.95. URL http:/

/www.oreilly.com/catalog/
9781565921672;   http://
www.oreilly.com/catalog/
linuxnut.

Herold:1999:LUS

[Her99] Helmut Herold. Linux- Unix-Systemprogrammierung. Addison-Wesley, München, Germany, 1999. ISBN 3-8273-1512-

3. xi + 1179 pp. LCCN ???? DM 99,90; EUR 51,08.

Hertzog:2004:DHT

[Her04] Raphaël Hertzog. Debian. Eyrolles, Paris, France, 2004. ISBN 2-212-11398-6. xi + 246 pp. LCCN ????

Hentzen:2004:LTW

[HH04] Whil Hentzen and Chris Herborth. Linux transfer for Windows power users [electronic resource]. Hentzenwerke Publishing, Whitefish Bay, WI, USA, 2004. ISBN 1-930919-42-5. LCCN QA76.76.O63 H468 2004b. URL ftp:

//uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//link.library.utoronto.
ca/eir/EIRdetail.cfm?Resources\
_\_ID=77297&T=resource.

Hill:2005:DGL

[HH05a] Benjamin Mako Hill and David B. Harris. Debian GNU/Linux 3.X bible. Wiley, New York, NY, USA, 2005. ISBN 0-7645-7644-5 (paper + CD-ROM). xxvii + 640 pp. LCCN QA76.76.O63 H57135 2005. URL ftp:

                                        //uiarchive.cso.uiuc.edu/
                                        pub/etext/gutenberg/;   http:
                                        //www.loc.gov/catdir/toc/
                                        ecip054/2004027963.html. In-

cludes two CD-ROMs.

Hudson:2005:RFL

[HH05b] Paul Hudson and Andrew Hudson. Red Hat Fedora 5 Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, 2005. ISBN 0-672-32847-X (paperback). ???? pp. LCCN ???? Includes DVD.

Hudson:2006:UU

[HH06] Andrew Hudson and Paul Hudson. Ubuntu Unleashed. SAMS Publishing, Indianapolis, IN, USA, 2006. ISBN 0-672-32909-3 (hardcover). 800 (est.) pp. LCCN ???? US$66.99.

Hudson:2007:UU

[HH07] Andrew Hudson and Paul Hudson. Ubuntu unleashed. SAMS Publishing, Indianapolis, IN, USA, 2007. ISBN 0-672-32909-3 (paperback: CD-ROM). xxvi + 879 pp. LCCN QA76.76.O63 H8167 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        fy0707/2006286996.html.

Hudson:2008:ULU

[HH08] Andrew Hudson and Paul Hudson. Ubuntu 7.10 Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, third edition, 2008. ISBN 0-672-32969-7 (paperback). xxiv + 811 pp. LCCN QA76.76.O63 H81665 2008.

Hudson:2009:UU

[HH09] Andrew Hudson and Paul Hudson. Ubuntu unleashed. SAMS Publishing, Indianapolis, IN, USA, fourth edition, 2009. ISBN 0-672-32993-X (paperback). xxiv + 826 pp. LCCN QA76.76.O63 H81666 2009. URL http:

//www.loc.gov/catdir/toc/
ecip0818/2008021150.html.

Includes DVD.

Huber:2019:FCL

[HHAW19] Manuel Huber, Julian Horsch, Junaid Ali, and Sascha Wessel. Freeze and crypt: Linux kernel support for main memory encryption. Computers & Security, 86(??):420–436, September 2019. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0167404818310435.

Huang:2016:RLT

[HHD^{+}16] Chenlin Huang, Chuanwang Hou, Huadong Dai, Yan Ding, Songling Fu, and Mengluo Ji. Research on Linux trusted boot method based on reverse integrity verification. Scientific Programming, 2016(??): 4516596:1–4516596:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https://

www.hindawi.com/journals/
sp/2016/4516596/.

Helmke:2011:UUC

[HHH11] Matthew Helmke, Andrew Hudson, and Paul Hudson. Ubuntu unleashed: covering 10.10 and 11.04. SAMS Publishing, Indianapolis, IN, USA, sixth edition, 2011. ISBN 0-672-33344-9. ???? pp. LCCN QA76.76.O63 U36 2010.

Hudson:2010:UUH

[HHHT10] Andrew Hudson, Paul Hudson, Matthew Helmke, and Ryan Troy, editors. Ubuntu unleashed: covering 9.10 and 10.4. SAMS Publishing, Indianapolis, IN, USA, fifth edition, 2010. ISBN 0-672-33109-8. 864 (est.) pp. LCCN QA76.76.O63 U385 2010.

Hallen:1999:LW

[HHJC99] Jacob Hallén, Anders Hammarqvist, Fredrik Juhlin, and Anders Chrigström. Linux in the workplace. IEEE Software, 16(1):52–57, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://computer.org/
software/so1999/s1052abs.
htm;   http://dlib.computer.
org/so/books/so1999/pdf/
s1052.pdf.

Hidgson:2004:BRL

[Hid04] J. P. E. Hidgson. Book review: Linux on the Mainframe by John Eilert, Maria Eisenhaendler, Dorothea Matthaeu, and Ingol Salm, Prentice Hall Professional Technical Reference, 2003, $49.99, ISBN: 0-13-101415-3. ACM Queue: Tomorrow’s Computing Today, 2 (1):92, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Huang:2009:QSS

[HKP09] Chih-Yuan Huang, Tei-Wei Kuo, and Ai-Chun Pang. QoS for storage subsystems using IEEE-1394. ACM Transactions on Storage, 4(4):12:1– 12:??, January 2009. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Haneda:2002:LJU

[HKS02] I. Haneda, T. Kiyohara, and N. Shiraishi. Linux/Java — unified platform for the nextgeneration of Zaurus. Sharp Technical Journal, 84:18–21, 2002. CODEN ???? ISSN 0285-0362.

Hoflehner:2004:COT

[HKS^{+}04] Gerolf Hoflehner, Knud Kirkegaard, Rod Skinner, Daniel Lavery, Yong fong Lee, and Wei Li. Compiler optimizations for transaction processing workloads on Itanium Linux systems. In IEEE [IEE04], pages 294–303. ISBN 0-7695-2126-6. ISSN 1072-4451. LCCN QA76.9.A73. IEEE Computer Society Order Number P2126. ACM Order Number 520040.

Hatch:2001:HLE

[HLK01] Brian Hatch, James Lee, and George Kurtz. Hacking Linux exposed: Linux security secrets and solutions. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212773-2. xxx + 566 pp. LCCN QA76.76.O63 H377 2001.

Hoang:2021:PHD

[HLT^{+}21] Thong Hoang, Julia Lawall, Yuan Tian, Richard J. Oentaryo, and David Lo. PatchNet: Hierarchical deep learningbased stable patch identification for the Linux kernel. IEEE Transactions on Software Engineering, 47(11):2471– 2486, November 2021. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Haider:2020:FHF

[HMK^{+}20] Waqas Haider, Nour Moustafa, Marwa Keshk, Amanda Fernandez, Kim-Kwang Raymond Choo, and Abdul Wahab. FGMC-HADS: Fuzzy Gaussian mixture-based correntropy models for detecting zeroday attacks from Linux systems. Computers & Security, 96(??):Article 101906, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167404820301826.

Hertel:2003:MSD

[HNH03] Guido Hertel, Sven Niedner, and Stefanie Herrmann. Motivation of software developers in Open Source projects: an Internet-based survey of contributors to the Linux kernel. Research Policy, 32(7): 1159–1177, 2003. CODEN ???? ISSN 0048-7333 (print), 1873-7625 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0048733303000477.

Hekman:1998:LNG

[HNV98] Jessica Perry Hekman, Alain Nadeau, and Jean-Michel Vansteene. Linux in a nutshell: guide de référence. O’Reilly France, Paris, France, 1998. ISBN 2-84177-031-1. xi + 424 pp. LCCN ????

Holmes:1996:LMW

[Hol96]  Lewis M. Holmes.  Linux

makes waves. Computers in Physics, 10(1):3–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.
1063/1.4822352.

Hollabaugh:2002:ELH

[Hol02] Craig Hollabaugh. Embedded Linux: Hardware, Software, and Interfacing. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-672-32226-9 (paperback). viii + 419 pp. LCCN QA76.76.O63 H6475

Hombourger:2000:SNC

[Hom00] Cédric Hombourger. Solution de network computing sous GNU/Linux. Mém. d.e.s.s.: Techniques nouvelles en informatique [t.n.i.], Université Montpellier II Sciences et Techniques du Languedoc, Montpellier 2, France, 2000. 59 pp.

Houston:1996:LLI

[Hou96] M. Houston. Linux: Linux is a powerful version of Unix that runs on PCs and is collectively owned by its many authors. Personal computer world, 19 (5):120–??, ???? 1996. CODEN PCWODU. ISSN 0142-0232.

Husain:1995:LU

[HP^{+}95] Kamran Husain, Tim Parker, et al. Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, 1995. ISBN 0-672-30705-7. xxxv + 1058 pp. LCCN QA76.76.O63 L5547 1995. US$49.99. Includes CD ROM.

Husain:1996:LU

[HP96a] Kamran Husain and Tim Parker. Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, second edition, 1996. ISBN 0-672-30908-4. xlii + 1176 pp. LCCN QA76.76.O63L5547 1996.

Husain:1996:RLU

[HP96b] Kamran Husain and Tim Parker. Red Hat Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, 1996.

ISBN 0-672-30962-9. xxxv + 1128 pp. LCCN QA76.76.O63 H87 1996.

Hall:2001:JLG

[HP01] Michael Hall and Brian Proffitt. The joy of Linux: a gourmet guide to open source. Prima Publishing, Roseville, CA, USA, 2001. ISBN 0-7615-3151-3. 340 pp. LCCN QA76.76.O63 H34354 2001.

Hein:1995:FRU

[HR95] Gunter Hein and Bernhard Riedl. First results using the new DGPS real-time deformation monitoring system ‘DREAMS’. Proceedings of ION GPS, 2:1647–1475, 1995. CODEN PIGPEU.

Ha:2007:IER

[HRD^{+}07] Jungwoo Ha, Christopher J.

Rossbach, Jason V. Davis, Indrajit Roy, Hany E. Ramadan, Donald E. Porter, David L. Chen, and Emmett Witchel. Improved error reporting for software that uses black-box components. ACM SIGPLAN Notices, 42(6):101–111, June 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hicks:2010:LSA

[HRS^{+}10] Boniface Hicks, Sandra Rueda, Luke St. Clair, Trent Jaeger, and Patrick McDaniel. A logical specification and analysis for SELinux MLS policy. ACM Transactions on Information and System Security, 13 (3):26:1–26:??, July 2010. CO-DEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Hofmann:2009:MBM

[HRW09] Owen S. Hofmann, Christopher J. Rossbach, and Emmett Witchel. Maximum benefit from a minimal HTM. ACM SIGPLAN Notices, 44(3):145– 156, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ha:2008:CNT

[HRX08] Sangtae Ha, Injong Rhee, and Lisong Xu. CUBIC: a new TCP-friendly high-speed TCP variant. Operating Systems Review, 42(5):64–74, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Hall:1999:RLD

[HS99] Jon Hall and Paul G. Sery. Red Hat Linux for Dummies with CDROM: The Complete Version of Red Hat Linux 6.1 on 2 CD-ROMs. I D G Books Worldwide, Indianapolis, IN, USA, 1999. ISBN 0-7645-0663-3. xxiv + 374 pp. LCCN QA76.76.O63 H34349 2000.

Hall:2000:RLD

[HS00] Jon Hall and Paul G. Sery. Red Hat Linux 7 for Dummies with CDROM. IDG Books, San Mateo, CA, USA, 2000. ISBN 0-7645-0795-8. xxii + 393 pp. LCCN QA76.76.O63 H34348 2001. US$29.99. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/wiley038/00108205.
html;   http://www.loc.gov/
catdir/toc/wiley021/00108205.
html.

Hsiao:2001:STY

[Hsi01] Aron Hsiao. Sams teach yourself Linux Security basics in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-32091-6 (paperback). xv + 412 pp. LCCN QA76.76.O63 H755 2001.

Hart:2008:RTL

[HST08] D. Hart, J. Stultz, and T. Ts’o.

Real-time Linux in real time. IBM Systems Journal, 47(2): 207–??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/472/hart.html.

Hekmann:1997:LN

[HT97] Jessica Perry Hekmann and The Staff of O’Reilly and Associates. Linux in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-167-4. xi + 424 pp. LCCN QA76.76.O63 H453 1997. US$19.95.

Huang:2006:CLK

[Hua06] Feilong Huang. Compiling the Linux kernel with the Intel compiler. In Anonymous [Ano06], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        ice.gelato.org/;   http:/
                                        /www.ice.gelato.org/about/
                                        oct06_presentations.php.

Hubbard:2002:BCR

[Hub02] John F. Hubbard. Binary code reuse in a Linux environment. C/C++ Users Journal, 20(3): ??, March 2002. CODEN CCU-JEX. ISSN 1075-2838.

Hughes:1997:LDQ

[Hug97] Phil Hughes. Linux for dummies quick reference. I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-0302-2. xiv + 209 pp. LCCN QA76.76.O63 H844 1997. US$17.99.

Hughes:1998:LDQ

[Hug98] Phil Hughes. Linux for dummies quick reference. For dummies. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 1998. ISBN 0-7645-0422-3 (paperback). xvi + 219 pp. LCCN QA76.76.O63 H844 1998.

Hulse:2005:RBC

[Hul05] Paul Hulse. Review: Beowulf Cluster Computing with Linux, second edition. The Computer Journal, 48(3):379–380, May 2005. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/48/3/379.

Hunt:1999:LNS

[Hun99] Craig Hunt. Linux network servers 24 × seven. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2506-9. xxvi + 626 pp. LCCN QA76.9.C55 H86 1999.

Hunt:2000:LDS

[Hun00a] Craig Hunt. Linux: DNS Server Administration. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2736-3. xxii + 423 pp. LCCN QA76.76.O63 H863 2000. US$39.99. URL http://

www.sybex.com/cgi-bin/rd_
bookpg.pl?2736back.html.

Hunt:2000:LSS

[Hun00b] Craig Hunt. Linux: Samba Server Administration. The Craig Hunt Linux library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2740-1. xxiii + 629 pp. LCCN QA76.76.O63 S58847 2001.

Hunt:2000:LSA

[Hun00c] Craig Hunt. Linux: System Administration. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2735-5. xxviii + 657 pp. LCCN A76.76.O63 S7346 2001.

Hunger:2001:DGL

[Hun01a] Steven Hunger. Debian GNU/ Linux bible. I D G Books Worldwide, Indianapolis, IN, USA, 2001. ISBN 0-7645-4710-0. ???? pp. LCCN QA76.76.O63 H847 2001.

Hunt:2001:LAW

[Hun01b] Craig Hunt. Linux: Apache Web Server Administration. The Craig Hunt Linux library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2734-7. xxiii + 615 pp. LCCN TK5105.8885.A63 A95 2001.

Hunt:2002:LNS

[Hun02] Craig Hunt. Linux network servers. Craig Hunt Linux library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2002. ISBN 0-7821-4123-4. LCCN QA76.76.O63 H86 2002b; **See. URL http:

                                        //library.books24x7.com/
                                        library.asp?B&isbn=0782141234.

Husain:1994:EI

[Hus94] K. Husain. Extending imake. Dr. Dobb’s Journal of Software Tools, 19(6):70, 72, 74–76, June 1994. CODEN DDJSDM. ISSN 1044-789X.

Harris:1999:WLB

[HZ99] Stephen E. Harris and Erwin Zijleman. WordPerfect for Linux bible. IDG Books, San Mateo, CA, USA, 1999. ISBN 0-7645-3374-6. ???? pp. LCCN Z52.5.W65H336 1999.

IBM:2005:MAS

[IBM05] IBM Corporation. Mathematical Acceleration Subsystem for Linux. World Wide Web document, 2005. URL http://

www-306.ibm.com/software/
awdtools/mass/linux/mass-
linux.html.

IEEE:1993:NFT

[IEE93] IEEE, editor. The new face of technical communication: people, processes, products: International professional communication conference — October 1993, Philadelphia, PA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1466-2, 0-7803-1465-4, 0-7803-1467-0. LCCN T 10.5 I54 1993.

IEEE:2004:PIS

[IEE04] IEEE, editor. Proceedings:

37th International Symposium on Microarchitecture, MICRO-37: 4–8 December 2004, Portland, Oregon. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2126-6. ISSN 1072-4451. LCCN QA76.9.A73. IEEE Computer Society Order Number P2126. ACM Order Number 520040.

IEEE:2006:PIS

[IEE06]  IEEE, editor.  Proceed-

ings of the 20th International Symposium on High-Performance Computing in an Advanced Collaborative Environment (HPCS’06), St. John’s, Newfoundland, May 14–17, 2006. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2582-2. ISSN 1550-5243. LCCN QA76.88. ACM product number E2582.

Ince:2019:BOS

[ILG19] M. Numan Ince, Joseph Ledet, and Melih Gunay. Building an open source Linux computing system on RISC-V. In IEEE, editor, 2019 1st International Informatics and Software Engineering Conference (UBMYK), pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2019.

InfoMagic:1995:LDR

[Inf95] InfoMagic, Inc. Linux: developer’s resource, 1995. 4 CD ROMs and booklet.

ISO:2005:IID

[ISO05] ISO. ISO/IEC DIS 23360: Draft International Standard: Linux Standard Base Core Specification 2.0.1. International Organization for Standardization, Geneva, Switzerland, 2005. ISBN ???? 1214 pp. LCCN ????

URL http://www.iso.org/
iso/en/CatalogueDetailPage.
CatalogueDetail?CSNUMBER=
41481&scopelist=PROGRAMME;
http://www.linuxbase.org/
LSBWiki/IsoBallot.

Inoue:2008:FAC

[ISTE08] Hiroaki Inoue, Junji Sakai, Sunao Torii, and Masato Edahiro. FIDES: an advanced chip multiprocessor platform for secure next generation mobile terminals. ACM Transactions on Embedded Computing Systems, 8(1):1:1–1:??, December 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Iyer:2007:QPA

[IZG^{+}07] Ravi Iyer, Li Zhao, Fei Guo, Ramesh Illikkal, Srihari Makineni, Don Newell, Yan Solihin, Lisa Hsu, and Steve Reinhardt. QoS policies and architecture for cache/memory in CMP platforms. ACM SIG-METRICS Performance Evaluation Review, 35(1):25–36, June 2007. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Jang:2001:SLG

[Jan01] Michael H. Jang. Sair Linux/ GNU installation and configuration. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-57610-953-4. xxx + 351 pp. LCCN QA76.3 .J345 2001.

Jang:2003:LTW

[Jan03] Michael Jang. Linux Transfer for Windows Network Admins: A roadmap for building a Linux file and print server. Hentzenwerke Publishing, Milwaukee, WI 53217-5361, USA, 2003. ISBN 1-930919-46-8. 310 (est.) pp. LCCN ???? US$49.95. Edited by Elizabeth Zinkann.

Jang:2004:RRC

[Jan04] Michael Jang. RHCE Red Hat certified engineer: Linux study guide (exam RH302). Osborne/McGraw-Hill, Berkeley, CA, USA, fourth edition, 2004. ISBN 0-07-225365-7. 755 pp. LCCN QA76.3 .R46 2002.

Jang:2006:LAG

[Jan06a] Michael H. Jang. Linux annoyances for geeks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-00801-5. xv + 484 pp. LCCN QA76.76.O63; QA76.76.O63 J36 2006eb; QA76.76.O63 J36 2006. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596008017.

Jang:2006:LPM

[Jan06b]  Michael H. Jang.  Linux

Patch Management: keeping Linux systems up to date. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-236675-4. ???? pp. LCCN QA76.76.O63 J368 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0519/2005028070.html.

Jamieson:1995:CCS

[JB95] D. N. Jamieson and W. A.

Belcher. Computer controlled specimen stage goniometer for a nuclear microprobe. Nuclear Instruments and Methods in Physics Research, Section B (Beam Interactions with Materials and Atoms), B104(1-4): 124–129, September 1995. CO-DEN NIMBEU. ISSN 0168-583X.

Jones:1999:OSL

[JB99] Christopher A. Jones and Drew Batchelor. Open source Linux Web programming. IDG Books, San Mateo, CA, USA, 1999. ISBN 0-7645-4619-8. xxi + 476 pp. LCCN QA76.76.O63 J662 1999. US$39.99.

Jennings:2002:JQ

[Jen02] Mike Jennings. Java Q&A:

How do I use the Java plugin tool? Dr. Dobb’s Journal of Software Tools, 27(1): 93–94, January 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_01/jqa0102.
zip.

Jenkins:2005:RLE

[Jen05] G. K. Jenkins. Review of “The Linux Enterprise Cluster”, by Karl Kopper, No Starch Press, 2005, $49.95 ISBN: 1-59327-036-4. ACM Queue: Tomorrow’s Computing Today, 3(9): 58, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Janakiram:2008:OOW

[JGS^{+}08] D. Janakiram, Ashok Gunnam, N. Suneetha, Vineet Rajani, and K. Vinay Kumar Reddy. Object-oriented wrappers for the Linux kernel. Software —Practice and Experience, 38 (13):1411–1427, November 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jambor:2007:ILL

[JHT^{+}07] Martin Jambor, Tomas Hruby, Jan Taus, Kuba Krchak, and Viliam Holub. Implementation of a Linux log-structured file system with a garbage collector. Operating Systems Review, 41(1):24–32, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jiang:2024:UVI

[JJW^{+}24] Muhui Jiang, Jinan Jiang, Tao Wu, Zuchao Ma, Xiapu Luo, and Yajin Zhou. Understanding vulnerability inducing commits of the Linux kernel. ACM Transactions on Software Engineering and Methodology, 33(7):170:1–170:??, September 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672452.

Jacobsen:2015:LCD

[JKS^{+}15] Charles Jacobsen, Muktesh Khole, Sarah Spall, Scotty Bauer, and Anton Burtsev. Lightweight capability domains: Towards decomposing the Linux kernel. Operating Systems Review, 49(2):44–50, December 2015. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jbara:2014:HMF

[JMF14] Ahmad Jbara, Adam Matan, and Dror G. Feitelson. High-MCC functions in the Linux kernel. Empirical Software Engineering, 19(5):1261– 1298, October 2014. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s10664-013-
9275-7.

Jones:2002:JMA

[Jon02] M. Tim Jones. Java mobile agents and the Aglets SDK. Dr. Dobb’s Journal of Software Tools, 27(1):42, 44, 46–48, January 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_01/
aglet.txt.

Jones:2005:GLA

[Jon05]  M. Tim Jones.  GNU/

Linux application programming. Charles River Media programming series. Charles River Media, Hingham, MA, USA, 2005. ISBN 1-58450-371-8. xxv + 486 pp. LCCN QA76.76.O63 J665 2005. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        toc/ecip052/2004024882.html.

Jordan:2004:ESL

[Jor04] Edmund Jordan. Embedded Systeme mit Linux programmieren: GNU-Softwaretools zur Programmierung ARMbasierender Systeme. (German) [Embedded Systems with Linux Programming: GNU Software Tools for Programming ARMbased Systems]. Franzis, Feldkirchen, Germany, 2004. ISBN 3-7723-5599-4. 384 pp. LCCN ????

Jepson:2000:DAP

[JPS00] Brian Jepson, Joan Peckham, and Ram Sadasiv. Database application programming with Linux. Wiley, New York, NY, USA, 2000. ISBN 0-471-35549-6 (paperback). xii + 516 pp. LCCN QA76.9.D26 J48 2000.

Johnson:1998:LAD

[JT98] Michael K. Johnson and Erik W. Troan. Linux Application Development. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-30821-5. xxiii + 538 pp. LCCN QA76.76.O63J635 1998. US$45.95.

Johnson:2005:LAD

[JT05] Michael K. Johnson and Erik W. Troan. Linux application development. Addison-Wesley, Reading, MA, USA, second edition, 2005. ISBN 0-321-21914-7. xxxi + 702 pp. LCCN QA76.76.O63 J635 2005. URL http:

//www.loc.gov/catdir/toc/
ecip0421/2004017882.html.

Jepson:1999:OPT

[JW99] Brian Jepson and Larry Wall.

O’Reilly Perl Toolkit for Linux and Solaris. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-604-8. 120 pp. LCCN ???? US$49.95. URL http:/

/www.oreilly.com/catalog/
perltkit/.

Jepsen:1999:LUE

[JWKZ99] Thomas C. Jepsen, Stephen A.

Wright, Richard L. Klevans, and Ze Zhang. Linux update: An experimental ATM network. IEEE Software, 16(5):32–39, September/October 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://computer.org/
software/so1999/s5032abs.
htm;   http://dlib.computer.
org/so/books/so1999/pdf/
s5032.pdf.

Jiang:2016:DLM

[JYX^{+}16] Yunyun Jiang, Yi Yang, Tian Xiao, Tianwei Sheng, and Wenguang Chen. DRDDR: a lightweight method to detect data races in Linux kernel. The Journal of Supercomputing, 72 (4):1645–1659, April 2016. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s11227-016-1691-1.

Kirkland:2006:LTS

[K^{+}06] James Kirkland et al., editors. Linux troubleshooting for system administrators and power users. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-185515-8 (paperback). ???? pp. LCCN QA76.76.O63 L54875 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip066/2006000036.html.

Kabir:1999:RLS

[Kab99] Mohammed J. Kabir. The Red Hat Linux server. IDG Books, San Mateo, CA, USA, 1999. ISBN 0-7645-3337-1. ???? pp. LCCN QA76.76.O63K314 1999.

Kadhim:1995:LUW

[Kad95a] B. Kadhim. Linux: Unleashing the workstation in your PC. Operating Systems Review, 29 (2):2–??, ???? 1995. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kadhim:1995:BRLb

[Kad95b] Basim Kadhim. Book review:

Linux Universe, Stefan Strobel and Thomas Uhl. Operating Systems Review, 29(4):3, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kadhim:1995:BRLa

[Kad95c] Basim Kadhim. Book review:

Linux: Unleashing the Workstation in Your PC, Stefan Strobel and Thomas Uhl. Operating Systems Review, 29(2): 2–3, April 1995. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kadhim:1995:LU

[Kad95d] Basim Kadhim. Linux universe. Operating Systems Review, 29(4):2–??, ???? 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kalsi:2016:PLS

[Kal16]  Tajinder Kalsi.  Practi-

cal Linux Security Cookbook. Packt Publishing, Birmingham, 2016. ISBN 1-78528-642-0 (paperback), 1-78528-530-0 (e-book). LCCN T55.4-

60.8. URL http://proquest.
safaribooksonline.com/9781785286421.

Krieger:2006:KBC

[KAR^{+}06] Orran Krieger, Marc Auslander, Bryan Rosenburg, Robert W. Wisniewski, Jimi Xenidis, Dilma Da Silva, Michal Ostrowski, Jonathan Appavoo, Maria Butrico, Mark Mergen, Amos Waterland, and Volkmar Uhlig. K42: building a complete operating system. Operating Systems Review, 40(4):133–145, October 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Karpowicz:2016:EEC

[Kar16] Michał P. Karpowicz. Energyefficient CPU frequency control for the Linux system. Concurrency and Computation: Practice and Experience, 28(2):420– 437, February 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Konishi:2006:LIL

[KAS^{+}06] Ryusuke Konishi, Yoshiji Amagai, Koji Sato, Hisashi Hifumi, Seiji Kihara, and Satoshi Moriai. The Linux implementation of a log-structured file system. Operating Systems Review, 40(3):102–107, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kasanen:2014:CLT

[Kas14]  Lauri Kasanen.  Into the

core: A look at Tiny Core Linux. ????, ????, 2014. ISBN 952-93-3391-9. LCCN ????

URL http://tinycorelinux.
net/book.html.

Kaufman:1996:LOC

[Kau96] Lar Kaufman. Linux opens corporate IS doors — most corporate IS shops turn up their noses at shareware, but Linux is good code for free. well, almost. Datamation, 42(7):72–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Kale:1999:APL

[KBPV99] L. Kale, R. Brunner, J. Phillips, and K. Varadarajan. Application performance of a Linux cluster using Converse. Lecture Notes in Computer Science, 1586:483–??, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirch:1995:ARS

[KC95] Olaf Kirch and Rene Cougnenc. Administration réseau sous Linux. Editions O’Reilly International Thomson, Paris, France, 1995. ISBN 2-84177-007-9. xviii + 350 pp. LCCN ????

Komarinski:1998:LSA

[KC98] Mark F. Komarinski and Cary Collett. Linux System Administration Handbook. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-680596-5. 416 pp. LCCN QA76.76.O63 K6483 1998. US$39.95. URL http://

www.phptr.com/. Includes CD-

ROM.

Komarinski:2000:RLA

[KC00] Mark (Mark F.) Komarinski and Cary Collett. Red Hat Linux administration handbook. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-025395-2. ix + 405 pp. LCCN QA76.76.O63 K6485 2000.

Kalpic:1996:IPI

[KD96] Damir Kalpic and Vesna Hljuz Dobric, editors. ITI ’96: Proceedings of the 18th International Conference on Information Technology Interfaces, Pula, Croatia, June 18–21, 1996. University of Zagreb, University Computing Centre, Zagreb, Croatia, 1996. ISBN ???? LCCN T58.5.I56 1996.

Kirch:2000:LNA

[KD00] Olaf Kirch and Terry Dawson. Linux network administrator’s guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2000. ISBN 1-56592-400-2. xxix + 474 pp. LCCN QA76.76.O63 K566 2000. Revised edition of Linux system administration, 1999.

Karamanos:1999:DNS

[KEB^{+}99] George-Sosei Karamanos, Constantinos Evangelinos, Richard C. Boes, Robert M. Kirby, and George E. Karniadakis. Direct numerical simulation of turbulence with a PC/Linux cluster: Fact or fiction? In ACM [ACM99], page ?? LCCN ????

Kim:2007:LPM

[KEP07] Dohun Kim, Jugwan Eom, and Chanik Park. L4oprof: a performance-monitoring-unit-based software-profiling framework for the L4 microkernel. Operating Systems Review, 41 (4):69–76, July 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kerrisk:2010:LPI

[Ker10] Michael Kerrisk. The Linux programming interface: a Linux and UNIX system programming handbook. No Starch Press, San Francisco, CA, USA, 2010. ISBN 1-59327-220-0 (hardcover). ???? pp. LCCN QA76.76.O63 K496 2010.

Klappheck:2000:BLE

[KGG00] Günther Klappheck, Peter Glinsky, and Frank Gehrke. Das Buch — LINUX Edition 2000: [jetzt zur S.u.S.E. 6.4, Installation und Bedienung von LINUX und seinen Werkzeugen, LINUX im Netzwerk und Internet, LINUX und ISDN, der Desktop KDE 1.x, VMware unter LINUX]. Sybex, Düsseldorf, Germany, 2000. ISBN 3-8155-0175-X. xxviii + 906 pp. LCCN ???? DM 69.95.

Kroah-Hartman:2007:LKN

[KH07] Greg Kroah-Hartman. Linux kernel in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10079-5. xiii + 182 pp. LCCN QA76.76.O63 K786 2007. URL http://www.

                                        kroah.com/lkn/;   http://
                                        www.loc.gov/catdir/enhancements/
                                        fy0728/2007274361-d.html;
                                         http://www.loc.gov/catdir/
                                        toc/fy0713/2007274361.html.

Khan:2002:PLP

[Kha02]  Ashfaq A. Khan.  Practi-

cal Linux programming: device drivers, embedded systems, and the Internet. Charles River Media, Inc., Hingham, MA, USA, 2002. ISBN 1-58450-096-4 (paperback). xv + 420 pp. LCCN QA76.76.O63 K497 2002.

Kiesling:1998:LCR

[Kie98] Robert Kiesling, editor. Linux, the complete reference. Walnut Creek CDROM and Linux System Laboratory, Walnut Creek, CA, USA and Chesterfield, MI, USA, sixth edition, 1998. ISBN 1-57176-199-3. xxvi + 1631 pp. LCCN ????

Kiesling:1999:LYB

[Kie99a] Robert Kiesling. Linux and the Y2K bug. SysAdmin, 8(1):16, 18–20, January 1999. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Kiesling:1999:LKC

[Kie99b] Robert Kiesling. The Linux kernel: A case study for CVS 39. SysAdmin, 8(6):39, 41–43, June 1999. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Kamijoh:2001:LWH

[KIKT01] Noboru Kamijoh, Tadanobu Inoue, Kohichiroh Kishimoto, and Ken Tamagawa. Linux watch: Hardware platform for wearable computing research. Lecture Notes in Computer Science, 2195:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2195/21950001.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2195/21950001.
pdf.

Kirch:1994:LNA

[Kir94] Olaf Kirch. Linux network administrators’ guide. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1994. ISBN 0-916151-75-1 (paperback). 289 (or 348??) pp. LCCN ????

Kirch:1995:LNA

[Kir95a]  Olaf Kirch.  Linux Net-

work Administrators’ Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 1995. ISBN 1-56592-087-2. xxxi + 335 pp. LCCN QA76.76.O63 K566 1995. US$24.95. URL http:/

/www.oreilly.com/catalog/
9781565920873;   http://
www.oreilly.com/catalog/
linag.

Kirch:1995:LNAa

[Kir95b] Olaf Kirch. Linux Network Administrator’s Guide. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1995. ISBN 0-916151-75-1. 289 pp. US$18.95.

Kirch:1995:LNAb

[Kir95c] Olaf Kirch. Linux Network Administrators’ Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1995. ISBN 1-56592-087-2. xxxi + 335 pp. LCCN QA76.76.O63 K566 1995. US$24.95. URL http:/

/www.oreilly.com/catalog/
9781565920873;   http://
www.oreilly.com/catalog/
linag.

Kee:2003:POP

[KKH03] Yang-Suk Kee, Jin-Soo Kim, and Soonhoi Ha. ParADE: An OpenMP programming environment for SMP cluster systems. In ACM [ACM03a], page ?? ISBN 1-58113-695-1.

LCCN ???? URL http://www.
sc-conference.org/sc2003/
inter_cal/inter_cal_detail.
php?eventid=10708#0; http:
//www.sc-conference.org/
sc2003/paperpdfs/pap130.
pdf.

Klaus:1993:CL

[Kla93] Todd C. Klaus. Checking Out Linux. UNIX/world, 10(3):66– ??, March 1993. ISSN 0739-5922.

Klecker:2000:DGL

[Kle00] Joel Espy Klecker. Debian GNU-Linux 2.2 Potato. Alcove, Madrid, Spain, 2000. ISBN ???? LCCN ???? Includes four CD-ROMs.

Kim:2002:MCR

[KLK^{+}02] JeongWon Kim, YoungUhg Lho, YoungJu Kim, Kwang-Baek Kim, and SeungWon Lee. A memory copy reduction scheme for networked multimedia service in Linux kernel. Lecture Notes in Computer Science, 2510:188– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2510/25100188.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2510/25100188.pdf.

Knox:2020:WRH

[KMP^{+}20] Shawn Knox, Steven Moghadam, Kenny Patrick, Anh Phan, and Kim-Kwang Raymond Choo. What’s really ‘ Happning’? A forensic analysis of Android and iOS Happn dating apps. Computers & Security, 94(??):Article 101833, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167404820301073.

Kofler:1997:LIC

[Kof97] Michael Kofler. Linux: Installation, Configuration, and Use. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-17809-5. xxiii + 677 pp. LCCN QA76.76.O63K64413 1997. US$34.95. Includes CD-ROM.

Kofler:1999:LIC

[Kof99] Michael Kofler. Linux: installation, configuration and use. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-59628-8. xxv + 772 pp. LCCN QA76.76.O63 K64413 2000.

Kofler:2000:LIC

[Kof00] Michael Kofler. LINUX: installation, configuration, and use. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-59628-8. xxv + 772 pp. LCCN QA76.76.O63 K64413 2000. Includes CD-ROM.

Kohlhepp:2006:RWL

[Koh06] Bayard Kohlhepp. Review of “Windows and Linux Integration: Hands-on Solutions for a Mixed Environment by Jermy Moskowitz and Thomas Boutell,” Sybex Inc., 2005, $59.99, ISBN 0-7821-4428-4. ACM Queue: Tomorrow’s Computing Today, 4(6): 58, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [MB05].

Komarinski:1996:LCE

[Kom96] Mark F. Komarinski. Linux companion: the essential guide for users and system administrators. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-231838-5. xii + 191 pp. LCCN QA76.76.O63 K648 1996.

Komarinski:2000:CRL

[Kom00] Mark F. Komarinski. The Complete Red Hat Linux Training Course. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-088223-2. ???? pp. LCCN ???? US$99.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130882232.html.

Komarinski:2002:PIT

[Kom02] Mark Komarinski. PTG Interactive’s Training Course for Red Hat Linux: A Digital Seminar on CD-ROM. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-034799-X. ???? pp. LCCN ???? US$69.99.

Kopp:2000:MCC

[Kop00] Carlo Kopp. Managing cluster computers. Dr. Dobb’s Journal of Software Tools, 25(7): 21–26, July 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_07/cluster.
                                        txt.

Kopper:2005:LEC

[Kop05] Karl Kopper. The Linux Enterprise Cluster: build a highly available cluster with commodity hardware and free software. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-036-4. xxiv + 430 pp. LCCN QA76.58 .K67 2005.

Koren:2006:SLK

[Kor06] Oded Koren. A study of the Linux kernel evolution. Operating Systems Review, 40(2):110– 112, April 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Koskelin:2000:LG

[Kos00] Al Koskelin. Linux games. No Starch Press, San Francisco, CA, USA, 2000. ISBN 1-886411-33-6. ???? pp. LCCN QA76.76.O63K676 2000.

Krafft:2005:DSC

[Kra05] Martin F. Krafft. The Debian system: concepts and techniques. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-069-

0. ???? pp. LCCN QA76.76.O63 K68 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip0515/2005019963.
html.

Krawetz:2007:HUS

[Kra07] Neal Krawetz. Hacking Ubuntu:

serious hacks, mods, and customizations. Wiley Technology Publishing, Indianapolis, IN, USA, 2007. ISBN 0-470-10872-X. xviii + 388 pp. LCCN QA76.76.O63 K742

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0741/2007003316-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0741/2007003316-
d.html;   http://www.loc.
gov/catdir/toc/ecip079/2007003316.
html.

Kroeker:1999:NTN

[Kro99] Kirk L. Kroeker. New tools:

Net development: Sun’s Java embedded server; MetaCreation’s Web-savvy graphics tool; WebCompiler’s HTML packaging tool. component technology: ProtoVIew Development’s ActiveX tools; Avilon Software’s load balancing component system. software development: Verilog’s test checker; Red Hat and Metrowerks’ development tools for Linux; the Object Factory’s optimization tool; Acumen Systems’s SDK for imaging; Aonix’s process-oriented lifecycle environment; Baan’s embedded software development suite. Computer, 32(5): 103–107, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.
                                        org/co/books/co1999/pdf/
                                        r5103.pdf.

Kroeker:2000:PCL

[Kro00]  Kirk L. Kroeker.  Prod-

ucts: Controlling the look and feel of Windows; sophisticated project estimation for $99; Web applications for the Internet-challenged; networkbased digital asset manager; Linux middleware framework; XML-based information portal; Java development for small computing devices; build installations for Windows 2000. Computer, 33(3):105– 107, March 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.
                                        org/co/books/co2000/pdf/
                                        r3105.pdf.

Kadav:2009:LMD

[KS09a] Asim Kadav and Michael M. Swift. Live migration of direct-access devices. Operating Systems Review, 43(3):95–104, July 2009. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kim:2009:AAA

[KS09b] Kyu-Han Kim and Kang G.

Shin. On accurate and asymmetry-aware measurement of link quality in wireless mesh networks. IEEE/ACM Transactions on Networking, 17(4):1172–1185, August 2009. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Kokologiannakis:2019:SMC

[KS19] Michalis Kokologiannakis and Konstantinos Sagonas. Stateless model checking of the Linux kernel’s read-copy update (RCU). International Journal on Software Tools for Technology Transfer (STTT), 21(3):287–306, June 2019. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10009-019-00514-6;
https://link.springer.com/
content/pdf/10.1007/s10009-
019-00514-6.pdf.

Kshetri:2004:ELA

[Ksh04] Nir Kshetri. Economics of Linux adoption in developing countries. IEEE Software, 21(1):74–81, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Kuo:1999:UUL

[Kuo99] Peter Kuo. Unix: das umfassende Lern- und Nachschlagewerk; zum effektiven Umgang mit Unix-Systemen; zur System- und Netzwerkadministration unter AIX, BSD, HP-UX, Linux, SCO-Unix, SVR 4, SVR 5, UnixWare und SunOS Solaris. Markt & Technik, Buch- und Software-Verlag, Munich, Germany, 1999. ISBN 3-8272-5532-5. 981 (est.) pp. LCCN ???? Includes CD-ROM.

Krohn:2007:IFC

[KYB^{+}07] Maxwell Krohn, Alexander Yip, Micah Brodsky, Natan Cliffer, M. Frans Kaashoek, Eddie Kohler, and Robert Morris. Information flow control for standard OS abstractions. Operating Systems Review, 41(6): 321–334, December 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lacheiner:2006:EPB

[Lac06] Hermann Lacheiner. Entwicklung einer auf Python basierenden Rich Client Platform für Linux. (German) [Development of a Python-based Rich Client Platform for Linux]. Diplome-Arbeit, Universität Linz, Linz, Austria, 2006. v + 86 pp.

Lu:2014:SLF

[LADADL14] Lanyue Lu, Andrea C. Arpaci-Dusseau, Remzi H. Arpaci-Dusseau, and Shan Lu. A study of Linux file system evolution. ACM Transactions on Storage, 10(1):3:1–3:??, January 2014. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Lenharth:2009:RDO

[LAK09] Andrew Lenharth, Vikram S.

Adve, and Samuel T. King. Recovery domains: an organizing principle for recoverable operating systems. ACM SIGPLAN Notices, 44(3):49–60, March 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lambert:1999:ISL

[Lam99] Patrick Lambert. Implementing security on Linux. SysAdmin, 8(10):67–70, October 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Pere:2005:GLR

[Lás05] Pere László. GNU/LINUX rendszerek üzemeltetése. (Hungarian) []. Kiskapu, Pécs, Hungary, 2005. ISBN ???? ???? pp. LCCN ????

Lathrop:2002:LSB

[Lat02] John P. Lathrop. Linux in Small Business: A Practical User’s Guide. Apress, Berkeley, CA, USA, 2002. ISBN 1-893115-46-1. 336 (est.) pp. LCCN ???? US$36.95.

Lai:1996:PCU

[LB96] K. Lai and M. Baker. A performance comparison of UNIX operating systems on the Pentium. In USENIX Association [USE96b], pages 265–277. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

Li:2009:ESM

[LBH09] Tong Li, Dan Baumberger, and Scott Hahn. Efficient and scalable multiprocessor fair scheduling using distributed weighted round-robin. ACM SIGPLAN Notices, 44(4):65– 74, April 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:2014:UBC

[LBN14] Eunji Lee, Hyokyung Bahn, and Sam H. Noh. A unified buffer cache architecture that subsumes journaling functionality via nonvolatile memory. ACM Transactions on Storage, 10(1):1:1–1:??, January 2014. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Lamps:2018:TIE

[LBN^{+}18] Jereme Lamps, Vignesh Babu, David M. Nicol, Vladimir Adam, and Rakesh Kumar. Temporal integration of emulation and network simulators on Linux multiprocessors. ACM Transactions on Modeling and Computer Simulation, 28(1): 1:1–1:??, January 2018. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Li:2021:VWF

[LCHL21] Yong-Gang Li, Yeh-Ching Chung, Kai Hwang, and Yue-Jin Li. Virtual Wall: Filtering rootkit attacks to protect Linux kernel functions. IEEE Transactions on Computers, 70(10): 1640–1653, October 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lumens:2005:SLE

[LCJH05] Chris Lumens, David Cantrell, Logan Johnson, and Alan Hicks. Slackware Linux Essentials. Slackware Linux, Inc., 1164 Claremont Drive, Brentwood, CA 94513, USA, second edition, 2005. ISBN 1-57176-338-4. xxii + 262 pp. LCCN ???? URL ftp:

//ftp.slackbook.org/pub/
slackbook/slackbook-2.0.
pdf;   http://www.slackbook.
org/.

Lin:2005:VMB

[LD05] Bin Lin and Peter A. Dinda.

VSched: Mixing batch and interactive virtual machines using periodic real-time scheduling. In ACM [ACM05a], pages 8–?? ISBN 1-59593-061-2. LCCN ????

LeBlanc:1996:RPI

[Le 96] Dee-Ann Le Blanc. Running a perfect Internet site with Linux. Que Corporation, Indianapolis, IN, USA, 1996. ISBN 0-7897-0514-1. xvii + 388 pp. LCCN TK5105.875.I57L45 1996.

Leavengood:2012:DHH

[Lea12] Ryan Leavengood. The dawn of Haiku — how a volunteer crew brought a crack OS back. IEEE Spectrum, 49(5):40–54, May 2012. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Leibovitch:1999:BCL

[Lei99] Evan Leibovitch. The business case for Linux. IEEE Software, 16(1):40–44, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

                                        URL http://computer.org/
                                        software/so1999/s1040abs.
                                        htm;   http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s1040.pdf.

Lendecke:1997:UFI

[Len97] V. Lendecke. UNIX filesystems without i-nodes. Dr. Dobb’s Journal of Software Tools, 22 (2):60, 62, 64, 66, February 1997. CODEN DDJSDM. ISSN 1044-789X.

Levitt:1995:PCL

[Lev95] J. Levitt. Prerelease of commercial Linux of questionable value. Information Week, August 1995. CODEN INFWE4. ISSN 8750-6874.

Levine:1999:L

[Lev99] John Levine. Linux. Visual quickpro guide. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-35437-3. 400 pp. LCCN QA76.76.O63 D3584 1999.

Lewis:1999:BCA

[Lew99a]  Ted Lewis.  Binary critic:

Asbestos pajamas: An open source dialogue. Computer, 32(4):112, 108–111, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/r4112.

pdf. Continues the debate about open source software development, notably in GNU/ Linux [Lew99b].

Lewis:1999:BCO

[Lew99b]  Ted Lewis.  Binary critic:

The open source acid test. Computer, 32(2):128, 125– 127, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/

r2128.pdf. See continuation in [Lew99a].

Leach:1995:IDA

[LF95] Ronald J. Leach and Terrence L. Fuller. An illustration of the domain analysis process. ACM SIGSOFT Software Engineering Notes, 20(5): 78–82, December 1995. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Liberty:2000:STY

[LH00] Jesse Liberty and David B.

Horvath. Sams teach yourself C++ for Linux in 21 days. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31895-4. xxix + 1110 pp. LCCN QA76.73.C153 L526 2000.

LSI:2001:PLG

[LH01] Loki Software, Inc. and John Hall. Programming Linux Games: Learn to Write the Games Linux People Play. Linux Journal Press, San Francisco, CA, USA, 2001. ISBN 1-886411-49-2. xviii + 415 pp. LCCN QA76.76.C672 L65 2001. US$39.95.

Lange:2024:TIM

[LHB24] Ferdinand Lange, Tessa Habich, and Sascha Beutel. Think inside the (mk)box — a tool for creating and deploying application-specific Linux images utilized in a SiLA software development process. SoftwareX, 26(??): ??, May 2024. CODEN ???? ISSN 2352-7110. URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S2352711024000761.

Li:1994:ILT

[Li94] Sing Li. An interview with Linus Torvalds. Dr. Dobb’s Journal of Software Tools, 19 (5):26–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Young:1994:LJ

[Lin94] Robert F. Young, editor. Linux journal, 1994. ISSN 1075-3583 (print), 1938-3827 (electronic). Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155.

Link:2000:LGP

[Lin00] Jay Link. Linux graphics programming with SVGAlib. Coriolis Group Books, Scottsdale, AZ, USA, 2000. ISBN 1-57610-524-5. xxii + 513 pp. LCCN T385 .L564 2000.

Lingmann:2002:DSK

[Lin02a] Thomas Lingmann. Datenverschlüsselung: sichere Kommunikation mit Linux und BSD: Security mit Open Source. (German) [Data encoding: Secure communication with Linux and BSD: Security with Open Source]. C & L, Böblingen, Germany, 2002. ISBN 3-932311-87-8 (??invalid checksum??). 476 (est.) pp. LCCN ????

Anonymous:2002:LSU

[Lin02b] Anonymous, editor. Die Linuxzeitung: das unabhängige Monatsblatt für GNU/Linux, freie Software und Open Source. (German) [The Linux Times: the independent monthly for GNU/Linux, Free Software, and Open Source], 2002. ISSN ????

Anonymous:2002:LUM

[Lin02c] Anonymous, editor. Die Linuxzeitung: das unabhängige Monatsblatt für GNU/Linux, freie Software und Open Source. (German) [The Linux Times: the independent monthly for GNU/Linux, Free Software, and Open Source], 2002. URL

http://www.die.linuxzeitung.

de/. ????, Berlin, Germany. On-line journal.

Lipkin:1999:LLV

[Lip99] Bernice Sacks Lipkin. LaTEX for Linux: a Vade Mecum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98708-8 (paperback). xxxi + 568 pp. LCCN Z253.4 L38 L56 1999. US$49.95.

Lischner:2002:BCC

[Lis02]  Ray Lischner.  Borland’s

CLX component framework. Dr. Dobb’s Journal of Software Tools, 27(2):77, 79–81, February 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_02/clx.txt.

LDP:1994:DL

[LL94a] Linux Documentation Project and Linux System Labs. Dr. Linux. Linux System Laboratory, Chesterfield, MI, USA, 1994. ISBN 1-885329-02-4. iv + 1176 pp. LCCN ????

Liou:1994:MGF

[LL94b] Der-Ming Liou and In-Chen Lin. A multimedia groupware framework system. In Anonymous [Ano94a], pages 368–375 (vol. 1). Two volumes.

Llanos:2006:TUO

[Lla06]  Diego R. Llanos.  TPCC-

UVa: an open-source TPC-C implementation for global performance measurement of computer systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 35(4):6–15, December 2006. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Lattner:2007:MCS

[LLA07] Chris Lattner, Andrew Lenharth, and Vikram Adve. Making context-sensitive pointsto analysis with heap cloning practical for the real world. ACM SIGPLAN Notices, 42 (6):278–289, June 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lees:2007:DSA

[LLT07] Michael Lees, Brian Logan, and Georgios Theodoropoulos. Distributed simulation of agentbased systems with HLA. ACM Transactions on Modeling and Computer Simulation, 17(3): 11:1–11:??, July 2007. CO-DEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic).

Loukides:1997:PGS

[LO97] Mike Loukides and Andy Oram. Programming with GNU Software. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-112-7. xiv + 244 pp. LCCN QA76.76.O63L65 1997.

US$39.95. URL http://www.

                                        ora.com/catalog/prognu/;
                                         http://www.oreilly.com/
                                        catalog/prognu. Includes CD-

ROM.

Lombardo:2001:EL

[Lom01] John Lombardo. Embedded Linux. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-0998-X. xxvii + 192 pp. LCCN QA76.76.O63 L646 2001. US$39.99.

Love:2005:LKD

[Lov05] Robert Love. Linux kernel development. Novell Press, Indianapolis, IN, USA, second edition, 2005. ISBN 0-672-32720-1 (paperback). xvi + 401 pp. LCCN QA76.76.O63 L673 2005.

Love:2007:LSP

[Lov07] Robert Love. Linux system programming: system and library calls every programmer needs to know. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-00958-5 (paperback). xvii + 368 pp. LCCN QA76.76.O63 L69 2007.

Love:2010:LKD

[Lov10]  Robert Love.  Linux ker-

nel development. Developer’s library: essential references for programming professionals. Addison-Wesley, Reading, MA, USA, third edition, 2010. ISBN 0-672-32946-8 (paperback). xx + 440 pp. LCCN NLS PB8.210.508/14; QA76.76.O63 L674 2010.

Loza:2005:USL

[Loz05] Boris Loza. UNIX, Solaris and Linux: A Practical Security Cookbook: Securing UNIX Operating System without Third-Party Applications. Author-House, ????, 2005. ISBN 1-4208-4824-0. 368 (est.) pp. LCCN ????

Lee:2010:IHM

[LP10] Jupyung Lee and Kyu Ho Park. Interrupt handler migration and direct interrupt scheduling for rapid scheduling of interrupt-driven tasks. ACM Transactions on Embedded Computing Systems, 9(4): 42:1–42:??, March 2010. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lu:2007:MAI

[LPH^{+}07] Shan Lu, Soyeon Park, Chongfeng Hu, Xiao Ma, Weihang Jiang, Zhenmin Li, Raluca A. Popa, and Yuanyuan Zhou. MUVI: automatically inferring multivariable access correlations and detecting related semantic and concurrency bugs. Operating Systems Review, 41(6):103– 116, December 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Leon:1997:LPG

[LR97] Louis Leon and Patrick Rougeau. Linux sur PC-Pentium: guide pratique d’installation. Le Micro Bulletin Thématique. no 2. Centre national de la recherche scientifique, Paris, France, 1997. ISBN 2-9510137-1-X. 141 pp. LCCN ????

Lehmann:2001:CSI

[LS01] Thomas Lehmann and Andreas Schreckenberg. Case study of integration of reconfigurable logic as a coprocessor into a SCI-cluster under RT-Linux. Lecture Notes in Computer Science, 2147:633– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2147/21470633.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2147/21470633.
                                        pdf.

Lelli:2016:DSL

[LSAF16] Juri Lelli, Claudio Scordino, Luca Abeni, and Dario Faggioli. Deadline scheduling in the Linux kernel. Software —Practice and Experience, 46 (6):821–839, June 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Li:2007:CPC

[LSP07] Chuanpeng Li, Kai Shen, and Athanasios E. Papathanasiou.

Competitive prefetching for concurrent sequential I/O. Operating Systems Review, 41(3): 189–202, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lucas:2000:FSL

[Luc00a] Michael Lucas. FreeBSD for the SVR4/Linux administrator. SysAdmin, 9(3):8, 10, 12, 14, 16, March 2000. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Lucas:2000:LUF

[Luc00b] Michael Lucas. Linux under FreeBSD. SysAdmin, 9 (1):26, 29–30, 32, January 2000. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Lucke:2005:BCL

[Luc05]  Robert W. Lucke.  Build-

ing clustered Linux systems. Hewlett–Packard professional books. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2005. ISBN 0-13-144853-6. xxxviii + 606 pp. LCCN QA76.76.O63 L838 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip0418/2004014016.
html.

Lumb:2001:LCH

[Lum01] Ian Lumb. Linux clustering for high-performance computing. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CO-DEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/
login/2001-08/pdfs/lumb.
pdf.

Lumme:2013:BHA

[Lum13]  Juha Lumme.  BeagleBone

home automation live your sophisticated dream with home automation using BeagleBone. Community experience distilled. Packt Publishing, Birmingham, UK, 2013. ISBN 1-78328-573-7, 1-306-27005-7 (ebook), 1-78328-574-5 (e-book). LCCN TJ223.P76 L85 2013.

URL http://proquest.tech.
safaribooksonline.de/9781783285730.

Lin:2000:SFT

[LW00] Kwei-Jay Lin and Yu-Chung Wang. Supporting faulttolerant real-time applications using the RED-Linux general scheduling framework. Lecture Notes in Computer Science, 1800:692–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/1800/18000692.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1800/18000692.
pdf.

Lee:2003:OSW

[LW03] James Lee and Brent Ware. Open source Web development with LAMP: using Linux, Apache, MySQL, Perl, and PHP. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-77061-X (paperback). xxxiv + 460 pp. LCCN QA76.76.D47 L435 2003.

Liu:2006:PAP

[LWL06] Chun-Ho Liu, Chat-Ming Woo, and Dennis Y. C. Leung. Performance analysis of a parallel finite element solution to the direct numerical simulation of fluid turbulence on Linux PC clusters. Applied Mathematics and Computation, 172(2):731– 743, January 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Lee:2010:REO

[LWV^{+}10] Dongyoon Lee, Benjamin Wester, Kaushik Veeraraghavan, Satish Narayanasamy, Peter M. Chen, and Jason Flinn. Respec: efficient online multiprocessor replay via speculation and external determinism. ACM SIGPLAN Notices, 45(3):77–90, March 2010. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:2007:CET

[LZ07] Peng Li and Steve Zdancewic.

Combining events and threads for scalable network services implementation and evaluation of monadic, application-level concurrency primitives. ACM SIGPLAN Notices, 42(6):189– 199, June 2007. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lin:2022:UBM

[LZAH22] Jiahuei Lin, Haoxiang Zhang, Bram Adams, and Ahmed E. Hassan. Upstream bug management in Linux distributions. Empirical Software Engineering, 27(6):??, November 2022. CODEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10664-022-10173-y.

Lin:2023:VML

[LZAH23] Jiahuei Lin, Haoxiang Zhang, Bram Adams, and Ahmed E. Hassan. Vulnerability management in Linux distributions. Empirical Software Engineering, 28(2):??, March 2023. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10664-022-10267-7.

Lin:2008:STM

[LZZ^{+}08] Jiang Lin, Hongzhong Zheng, Zhichun Zhu, Eugene Gorbatov, Howard David, and Zhao Zhang. Software thermal management of DRAM memory for multicore systems. ACM SIG-METRICS Performance Evaluation Review, 36(1):337–348, June 2008. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

MacKinnon:1999:LOS

[Mac99] James G. MacKinnon. The Linux operating system: Debian GNU/Linux. Journal of Applied Econometrics, 14 (4):443–452, July–August 1999. CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

MacCarty:2002:LDG

[Mac02] Bill MacCarty. Learning Debian GNU Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-705-2. ???? pp. LCCN QA76.76.O63 M372 1999. US$34.95. Includes CD-ROM.

Maginnis:2000:SLG

[Mag00] Tobin Maginnis. Sair Linux and GNU certification level I: installation and configuration. Wiley, New York, NY, USA, 2000. ISBN 0-471-36978-0 (paperback). xix + 284 pp. LCCN QA76.3 .M3235 2000.

Maginnis:2001:SLGa

[Mag01a] Tobin Maginnis. Sair Linux and GNU certification level I: installation and configuration. Wiley, New York, NY, USA, 2001. ISBN 0-471-41797-1. xix + 508 pp. LCCN QA76.3 .M32337 2001.

Maginnis:2001:SLGb

[Mag01b] Tobin Maginnis. Sair Linux and GNU certification level I: networking. Wiley, New York, NY, USA, 2001. ISBN 0-471-36977-2. xx + 476 pp. LCCN QA76.3 .M32337 2000.

Maginnis:2001:SLGc

[Mag01c] Tobin Maginnis. Sair Linux and GNU certification level I: security, ethics, and privacy. Wiley, New York, NY, USA, 2001. ISBN 0-471-36975-6. xix + 395 pp. LCCN QA76.3 .M32339 2001.

Maginnis:2004:GLZ

[Mag04] Tobin Maginnis. GNU-Linux-Zertifizierung: Prüfungsvorbereitung zum Sair Linux and GNU Certified Professional/Administrator (LCP/LCA). (German) [GNU/ Linux Certification: Examination Preparation for Sair Linux and GNU Certified Professional/Administrator (LCP/ LCA)]. dpunkt-Verlag, Heidelberg, Germany, 2004. ISBN ???? ???? pp. LCCN ????

Mann:1999:CLL

[Man99] Charles C. Mann. Computers: Living with Linux. The Atlantic, 284(2):80–??, August 1999. CODEN ATLAEO. ISSN 0276-9077.

Mancill:2001:LRP

[Man01] Tony Mancill. Linux routers: a primer for network administrators. Prentice Hall series in computer networking and distributed systems. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-086113-8. xx + 345 pp. LCCN QA76.76.O63 M348 2001.

Martin:2022:TLA

[MAP^{+}22] Hugo Martin, Mathieu Acher, Juliana Alves Pereira, Luc Lesoil, Jean-Marc Jézéquel, and Djamel Eddine Khelladi. Transfer learning across variants and versions: The case of Linux kernel size. IEEE Transactions on Software Engineering, 48(11):4274–4290, November 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Marsh:2000:PRL

[Mar00] Matthew Marsh. Policy routing in Linux. SysAdmin, 9(1):49– 50, 52–58, January 2000. CO-DEN SYADE7. ISSN 1061-

2688.  URL http://www.
samag.com/.

Marsh:2001:PRU

[Mar01] Matthew G. Marsh. Policy routing using Linux. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-32052-5. vii + 205 pp. LCCN TK5105.543 .M37 2001.

Marchesin:2004:ULR

[Mar04] Armand Marchesin. Using Linux for real-time applications. IEEE Software, 21 (5):18–20, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Marsh:2006:PRU

[Mar06] Matthew G. Marsh. Policy Routing Using Linux. SAMS Publishing, Indianapolis, IN, USA, 2006. ISBN 0-672-32052-5. 205 (est.) pp. LCCN ???? EUR 35.90.

Matai:1999:LNG

[Mat99a] Seema C. Matai. Linux: the next generation in point sale. Thesis (m.b.a.), University of California, Riverside, Riverside, CA, USA, 1999. vi + 27 pp.

Matus:1999:SLF

[Mat99b] Juan Matus. Setting up a Linux firewall. ;login: the USENIX Association newsletter, 24(6s):??, November 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

                                        usenix.org/publications/
                                        login/1999-11/features/linux.

html. Special issue on security.

Mahmoud:2008:PAS

[MAT08] Feras A. Mahmoud and Mohammad H. Al-Towaiq. Parallel algorithm for the solutions of PDEs in Linux clustered workstations. Applied Mathematics and Computation, 200(1): 178–188, June 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Maurer:1996:PSM

[Mau96] D. Maurer. PXROS — a scalable micro kernel with special realtime properties. Informationstechnik und technische Informatik: IT + TI, 38 (2):26–32, April 1996. CODEN ITINEY. ISSN 0944-2774.

Mauerer:2008:PLK

[Mau08] Wolfgang Mauerer. Professional Linux kernel architecture. Wrox professional guides. Wiley, New York, NY, USA, 2008. ISBN 0-470-34343-5 (paperback). xxx + 1337 pp. LCCN QA76.9.A73 M38

2008eb. URL http://site.
ebrary.com/lib/princeton/
Doc?id=10257612.

Maxwell:2000:RLN

[Max00] Steven Maxwell. Red Hat Linux network management tools. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-212260-9, 0-07-212261-7 (CD-ROM), 0-07-212602-7 (CD-ROM), 0-07-212262-5 (set). xxviii + 683 pp. LCCN QA76.76.O63 M373339 2000.

Maxwell:2001:LCK

[Max01] Scott Andrew Maxwell. Linux Core Kernel commentary. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-58880-149-7. xviii + 717 pp. LCCN QA76.76.O63 M373337 2001. US$49.99.

Mayer:2001:SPU

[May01]  Alexander Mayer.  Shell-

Programmierung in Unix: [das Lehrbuch zu Bourne-Shell, Korn-Shell, Z-Shell, Bash für Linux, Unix, BSD und MacOS X]. (German) [Shell programming in Unix: The textbook for the Bourne Shell, Korn Shell, Z Shell, Bash for Linux, Unix, BSD and MacOS X]. Computer-&-Literatur-Verlag, Böblingen, Germany, 2001. ISBN 3-932311-78-7. 767 pp. LCCN ????

Mengel:1995:SNC

[MB95] S. A. Mengel and C. D. Bowling. Supporting networking courses with a hands-on laboratory. In D. Budny and B. Herrick, editors, Proceedings. Frontiers in Education 1995 25th Annual Conference. Engineering Education for the 21st Century, volume 2, pages 2–?? ????, ????, 1995. ISBN 0-7803-3022-6. LCCN ???? IEEE catalog number 95CH35867.

Moskowitz:2005:WLI

[MB05] Jeremy Moskowitz and Thomas Boutell. Windows and Linux integration: hands-on solutions for a mixed environment. Wiley, New York, NY, USA, 2005. ISBN 0-7821-4428-4. xix + 539 pp.

Merkel:2008:TAV

[MB08] Andreas Merkel and Frank Bellosa. Task activity vectors: a new metric for temperatureaware scheduling. Operating Systems Review, 42(4):1– 12, May 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McCarty:1999:LDG

[McC99a] Bill McCarty. Learning Debian GNU/Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-705-2. xiii + 343 pp. LCCN QA76.76.O63 M372 1999. US$34.95. URL http:/

/www.oreilly.com/catalog/
debian.

McCarty:1999:LRL

[McC99b] Bill McCarty. Learning Red Hat Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-627-7. xiii + 378 pp. LCCN QA76.76.O63 M379 1999. US$34.95. URL http:/

/www.oreilly.com/catalog/
redhat.

McCune:2001:ILW

[McC01]  Mike McCune.  Integrating

Linux and Windows. Prentice Hall PTR open source technology series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-030670-3. xxii + 344 pp. LCCN QA76.9.U83 M46 2000. US$39.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130306703.html.

McCarty:2002:LRL

[McC02] Bill McCarty. Learning Red Hat Linux: A Guide to Red Hat Linux for New Users. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00071-5. xx + 346 pp. LCCN QA76.76.O63 M376 2002. US$34.95. URL http://

                                        safari.oreilly.com/0596000715;
                                         http://www.oreilly.com/
                                        catalog/redhat2. Includes

CD-ROM with Red Hat Linux 7.2.

McCarty:2003:LRL

[McC03] Bill McCarty. Learning Red Hat Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2003. ISBN 0-596-00469-9. xvi + 319 pp. LCCN QA76.76.O63 M376 2003. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596004699.

McCarty:2004:LRE

[McC04] Bill McCarty. Learning Red Hat Enterprise Linux and Fedora. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2004. ISBN 0-596-00589-X. xvii + 326 pp. LCCN QA76.76.O63 M376 2004. URL http:/

/www.oreilly.com/catalog/
9780596005894.

McCarty:2005:SNO

[McC05]  Bill McCarty.  SELINUX:

NSA’s open source Security Enhanced Linux. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00716-7. xiii + 238 pp. LCCN TK5105.59 .M37 2005.

URL http://www.oreilly.
com/catalog/9780596007164.

McCallum:2007:MRB

[McC07] Ethan McCallum. Managing RPM-based systems with Kickstart and Yum. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51382-8. LCCN QA76.76.O63; QA76.76.O63 M34 2007eb. URL http:/

/www.oreilly.com/catalog/
9780596513825.

McNab:1999:BPL

[McN99] A. D. McNab. BSD Portals for LINUX 2.0. Technical report, ????, ????, 1999. 12 pp. NAS2-14303 ; RTOP 509-10-61.

Meadhra:1999:SLD

[Mea99] Michael Meadhra. StarOffice for LINUX for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1999. ISBN 0-7645-0576-9. xxiv + 404 pp. LCCN A76.76.I57 M427 1999. US$19.99. URL http://www.

                                        idgbooks.com/.

Mehat:2000:NMC

[Meh00] San Mehat. Network monitoring for clusters. Journal of Linux Technology, 1(1):8–??, 2000. ISSN 1527-2761.

Mendell:2012:BRP

[Men12] Matt Mendell. Book review: A practical guide to Fedora and Red Hat enterprise Linux by Mark G. Sobell. ACM SIGSOFT Software Engineering Notes, 37(1):36, January 2012. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Messina:2000:DSG

[Mes00]  Paul Messina.  Distributed

supercomputing for the grid. Journal of Linux Technology, 1 (1):30–??, 2000. ISSN 1527-2761.

McKenney:2020:RUL

[MFBWW20] Paul E. McKenney, Joel Fernandes, Silas Boyd-Wickizer, and Jonathan Walpole. RCU usage in the Linux kernel: Eighteen years later. Operating Systems Review, 54(1):47–63, August 2020. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3421473.3421481.

Moody:2003:SNB

[MFPP03] Adam Moody, Juan Fernandez, Fabrizio Petrini, and Dhabaleswar K. Panda. Scalable NIC-based reduction on large-scale clusters. In ACM [ACM03a], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.
sc-conference.org/sc2003/
inter_cal/inter_cal_detail.
php?eventid=10716#2; http:
//www.sc-conference.org/
sc2003/paperpdfs/pap316.
pdf.

Hall:1999:LL

[mH99a] Jon “maddog” Hall. Legitimizing Linux. SysAdmin, 8(6): 21–22, June 1999. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Hall:1999:LD

[mH99b] Jon “maddog” Hall. Linux for dummies. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 1999. ISBN 0-7645-0421-5. xxii + 360 pp. LCCN QA76.76.O63 W59 1999. Includes CD-ROM.

Montesinos:2009:CSH

[MHKT09] Pablo Montesinos, Matthew Hicks, Samuel T. King, and Josep Torrellas. Capo: a software-hardware interface for practical deterministic multiprocessor replay. ACM SIG-PLAN Notices, 44(3):73–84, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mills:1996:LUA

[Mil96] D. Mills. Linux users [astronomical software packages]. Astron. Soc. Pac. Conf. Ser. (USA), Astronomical Society of the Pacific Conference Series, 101(????):233–235, ???? 1996. ISSN 1050-3390.

Miller:2001:LWA

[Mil01] Michael Joseph Miller. Linux for Windows addicts: a 12step program for habitual Windows users. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213081-4. xvii + 414 pp. LCCN QA76.76.O63 M57 2001.

Minnich:2001:PNL

[Min01]  Ronald G. Minnich.  Pri-

vate namespaces for Linux. Dr. Dobb’s Journal of Software Tools, 26(12):23–24, 26, 28, 30, December 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/.

Mohror:2004:PTS

[MK04] Kathryn Mohror and Karen L. Karavanic. Performance tool support for MPI-2 on Linux. In ACM [ACM04], page 28. ISBN 0-7695-2153-3. LCCN ????

Montella:2017:ALA

[MKO^{+}17] Raffaele Montella, Sokol Kosta, David Oro, Javier Vera, Carles Fernández, Carlo Palmieri, Diana Di Luccio, Giulio Giunta, Marco Lapegna, and Giuliano Laccetti. Accelerating Linux and Android applications on low-power devices through remote GPGPU offloading. Concurrency and Computation: Practice and Experience, 29 (24):??, December 25, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Mbongue:2022:DMT

[MKSB22] Joel Mandebi Mbongue, Danielle Tchuinkou Kwadjo, Alex Shuping, and Christophe Bobda. Deploying multi-tenant FPGAs within Linux-based cloud infrastructure. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):19:1– 19:31, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3474058.

Mostek:2000:OSX

[MLE^{+}00] Jim Mostek, Steven Levine, William Earl, et al. Open source XFS: SGI’s highperformance file system for Linux. Journal of Linux Technology, 1(2):16–24, 2000. ISSN 1527-2761.

Mann:2000:LSS

[MM00] Scott Mann and Ellen L.

Mitchell. Linux system security: an administrator’s guide to open source security tools. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-015807-0. xxxvii + 564 pp. LCCN QA76.76.O63 M3515 2000. US$49.00.

Miles:2001:EL

[MM01] Kathy Miles and Ethan Metsger. Everyday Linux. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-091762-1. xxxiv + 369 pp. LCCN QA76.76.O63 M534 2001. US$29.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130917621.html.

McIlwain:2006:TCL

[MM06] Sean McIlwain and Barton P. Miller. A tool for converting Linux device drivers into Solaris compatible binaries. Software—Practice and Experience, 36(7):689–710, June 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mayer:2006:SEU

[MMC06] Frank Mayer, Karl MacMillan, and David Caplan. SELinux by example: using Security Enhanced Linux. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-196369-4 (paperback). 460 (est.) pp. LCCN QA76.76.O63 M3738 2006. US$44.99. URL http://

vig.prenhall.com/catalog/

academic/product/0,1144,
0131963694,00.html; http:
//www.loc.gov/catdir/toc/
ecip0612/2006012657.html.

Martin:1996:WTW

[MMW96] D. E. Martin, T. J. McBrayer, and P. A. Wilsey. WARPED: a time warp simulation kernel for analysis and application development. In H. El-Rewini and B. D. Shriver, editors, Proceedings of the Twenty-Ninth Hawaii International Conference on System Sciences, volume 1, pages 5–?? ????, ????, 1996. ISBN 0-8186-7324-9. LCCN ????

Moen:2002:RSL

[Moe02] Rick Moen. A recipe for a successful Linux user group. ;login: the USENIX Association newsletter, 27(2): ??, April 2002. CODEN LOGNEM. ISSN 1044-

6397.  URL http://www.
usenix.org/publications/
login/2002-04/pdfs/moen.
pdf.

Mohr:1997:LGF

[Moh97a] Jim Mohr. Linux in a gray flannel suit: It’s powerful. it’s open. it’s free. that’s why this Unix is entering corporate IS. BYTE Magazine, 22(3):96NA3, 96NA6, 96NA8, March 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mohr:1997:SLU

[Moh97b] Jim Mohr. The state of Linux — an update on this 32-bit Unix-like OS, with pointers to even more Linux info. BYTE Magazine, 22(1):49–50, January 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Mohr:1998:LUR

[Moh98] James Mohr. Linux user’s resource: developer’s resource. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-842378-4. xxxv + 795 pp. LCCN QA76.76.O63 M7453 1998.

                                        URL http://www.prenhall.
                                        com/allbooks/ptr_0138423784.
                                        htm.

Molloy:2014:EBT

[Mol14] Derek Molloy. Exploring BeagleBone: tools and techniques for building with embedded Linux. Wiley, New York, NY, USA, 2014. ISBN 1-118-93512-8 (paperback), 1-118-93513-6 (e-book). xxx + 564 pp. LCCN ????

MonteroRivero:1997:L

[Mon97] Raul Montero Rivero. Linux. Manuales avanzados. Anaya Multimedia, Madrid, Spain, 1997. ISBN 84-415-0353-2. xvii + 462 pp. LCCN ????

Mondal:2000:LBU

[Mon00a] Abdul Sakib Mondal. Load balancing for UNIX and Win32.

Dr. Dobb’s Journal of Software Tools, 25(7):32, 34, 36, 38–40, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
load.txt;   http://www.ddj.
com/ftp/2000/2000_07/load.
zip;   http://www.ddj.com/
ftp/2000/2000_07/xyalb.zip.

Mondal:2000:PAU

[Mon00b] Abdul Sakib Mondal. Porting across UNIX and Win32. Dr. Dobb’s Journal of Software Tools, 25(12):30, 32, 34, 36, December 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_12/
portuw.txt;   http://www.
ddj.com/ftp/2000/2000_12/
portuw.zip.

Moody:2001:RCL

[Moo01a] Glyn Moody. Rebel code: Linux and the Open Source Revolution. Allen Lane, London, UK and New York, NY, USA, 2001. ISBN 0-7139-9520-3. viii + 334 pp. LCCN QA76.76.O63 M645 2001.

Moody:2001:RCI

[Moo01b] Glyn Moody. The rebel code:

the inside story of Linux and the open source revolution. Perseus Publishers, Cambridge, MA, USA, 2001. ISBN 0-7382-0333-5. viii + 334 pp. LCCN QA76.76.O63 M663 2001. US$27.50.

Morin:1994:L

[Mor94] Richard Morin. Linux. UNIX review, 12(9):83–??, September 1994. CODEN UNRED5. ISSN 0742-3136.

Morin:1997:MML

[Mor97] Rich Morin. MkLinux: Microkernel Linux for the Power Macintosh. Prime Time Freeware, 370 Altair Way, Suite 150, Sunnyvale, CA, USA, 1997. ISBN 1-881957-24-1. 322 pp. LCCN ???? US$50.00. Includes CD-ROM.

Mitchell:2001:ALP

[MOS01] Mark Mitchell, Jeffrey Oldham, and Alex Samuel. Advanced Linux programming. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1043-0. xxiii + 340 pp. LCCN QA76.76.O63 M58

                                        2001.  URL http://www.
                                        advancedlinuxprogramming.
                                        com/.

Mott:2002:WKA

[Mot02] Jason Mott. Writing KDE applications. C/C++ Users Journal, 20(3):??, March 2002. CO-DEN CCUJEX. ISSN 1075-2838.

MartinPerez:1998:L

[MP98] Cesar Martin Perez and Ismael Perez Crespo. Linux. Guia Practica para Usuarios. Anaya Multimedia, Madrid, Spain, 1998. ISBN 84-415-0475-X, 84-415-0406-7. 352 pp. LCCN ????

Mohr:1999:LRP

[MR99] James Mohr and Sergio Ruiz, tr. Linux: recursos para el usuario. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 970-17-0192-5. xxxv + 789 pp. LCCN ????

Makris:2007:DAU

[MR07] Kristis Makris and Kyung Dong Ryu. Dynamic and adaptive updates of non-quiescent subsystems in commodity operating system kernels. Operating Systems Review, 41(3):327– 340, June 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Majeed:2018:DNF

[MR18] Shakaiba Majeed and Minsoo Ryu. Debugging nondeterministic failures in Linux programs through replay analysis. Scientific Programming, 2018(1):8939027:1–8939027:??, 2018. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL https:/

/onlinelibrary.wiley.com/
doi/epdf/10.1155/2018/8939027.

McKinley:1999:HRH

[MRW99] Philip K. McKinley, Ravi T.

Rao, and Robin F. Wright. H-RMC: A hybrid reliable multicast protocol for the Linux kernel. In ACM [ACM99], page ?? LCCN ????

Matthew:1999:BLP

[MS99] Neil Matthew and Richard Stones. Beginning Linux Programming. Wrox Press, Chicago, IL, USA, second edition, 1999. ISBN 1-86100-297-1. xxvii + 945 pp. LCCN QA76.76.O63 M37157 1999. US$39.99.

Matthew:2000:PLP

[MS^{+}00] Neil Matthew, Richard Stones, et al. Professional Linux programming. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-301-3. xviii + 1155 pp. LCCN QA76.76.O63 P754 2000.

Maginnis:2002:SLG

[MS02] Tobin Maginnis and Sair Development Team. Sair Linux and GNU certification: level II, Samba and resource sharing. Wiley, New York, NY, USA, 2002. ISBN 0-471-40535-3. ???? pp. LCCN QA76.3 .M3237 2002.

Meadows:2005:CHE

[MS05] Catherine Meadows and Paul Syverson, editors. CCS ’05: proceedings of the 12th ACM Conference on Computer and Communications Security: November 7–11, 2005, Alexandria, Virginia, USA. ACM Press, New York, NY 10036, USA, 2005. ISBN 1-59593-226-7. LCCN QA76.9.A25. ACM order number 459050.

Mahoney:2007:UCS

[MS07] William R. Mahoney and William L. Sousan. Using common off-the-shelf tools to implement dynamic aspects.

ACM SIGPLAN Notices, 42(2): 34–41, February 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matthew:1996:BLP

[MSB^{+}96] Neil Matthew, Rick Stones, Christopher Browne, Brad Clements, Andrew Froggatt, David Goodger, Ivan Griffin, Jeff Licquia, Udaya Ranawake, Harish Rawat, Marius Sundbakken, Deepak Thomas, Stephen Turnbull, Ronald van Loon, and David Woodhouse. Beginning Linux programming. Wrox Press, Chicago, IL, USA, 1996. ISBN 1-874416-68-0. xxi + 710 pp. LCCN QA76.76.O63 M37157 1996.

Maillart:2008:ETZ

[MSSvK08] T. Maillart, D. Sornette, S. Spaeth, and G. von Krogh. Empirical tests of Zipf’s Law mechanism in open source Linux distribution. Physical Review Letters, 101(21): 218701, November 19, 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL

http://link.aps.org/doi/
10.1103/PhysRevLett.101.
218701.

Mendoza:2006:ULP

[MSW06] Alfredo Mendoza, Chakarat Skawratananond, and Artis Walker. Unix to Linux porting: a comprehensive reference. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-187109-9 (paperback). ???? pp. LCCN QA76.76.O63 M437 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip066/2006000234.html.

Menon:2009:TSA

[MSZ09] Aravind Menon, Simon Schubert, and Willy Zwaenepoel. TwinDrivers: semi-automatic derivation of fast and safe hypervisor network drivers from guest OS drivers. ACM SIG-PLAN Notices, 44(3):301–312, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mudawwar:1997:MTM

[Mud97] Muhammed F. Mudawwar. Multicode: A truly multilingual approach to text encoding. Computer, 30(4):37–43, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also response and rebuttal [DM97], and letters [CPJ^{+}98].

Muster:2002:UME

[Mus02] John Muster. UNIX made easy: Unix and Linux basics and beyond. Osborne/McGraw-Hill, Berkeley, CA, USA, third edition, 2002. ISBN 0-07-219314-X. xviii + 1011 pp. LCCN QA76.76.O63 M877 2002.

McKenney:2008:ITL

[MW08] Paul E. McKenney and Jonathan Walpole. Introducing technology into the Linux kernel: a case study. Operating Systems Review, 42(5):4–17, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Meadhra:2000:KLD

[MWM00] Michael Meadhra, Kate Wrightson, and Joe Merlino. KDE for Linux for dummies. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-0658-7. xxii, 344 pp. LCCN QA76.9.U83 M46 2000.

Minasi:2003:LWA

[MY03] Mark Minasi and Dan York.

Linux for Windows administrators. Mark Minasi Windows administrator library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2003. ISBN 0-7821-4119-6. xxxiii + 522 pp. LCCN QA76.76.O63 M57385

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0613/2002113840-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0613/2002113840-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0613/2002113840-t.html.

Minasi:2000:LWN

[MYH00] Mark Minasi, Dan York, and Craig Hunt. Linux for Windows NT/2000 Administrators: The Secret Decoder Ring. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN ???? ???? pp. LCCN ???? URL http://

                                        www.sybex.com/cgi-bin/rd_
                                        bookpg.pl?2730back.html.

Mihajlovic:2014:DIQ

[MZG14] Bojan Mihajlović, Zeljko Zilić, and Warren J. Gross. Dynamically instrumenting the QEMU emulator for Linux process trace generation with the GDB debugger. ACM Transactions on Embedded Computing Systems, 13(5s):167:1– 167:??, November 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Neira-Ayuso:2010:CBK

[NAGL10] Pablo Neira-Ayuso, Rafael M. Gasca, and Laurent Lefevre. Communicating between the kernel and user-space in Linux using Netlink sockets. Software —Practice and Experience, 40 (9):797–810, August 2010. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Negus:2001:RLB

[Neg01] Christopher Negus. Red Hat Linux 7 Bible. IDG Books, San Mateo, CA, USA, 2001. ISBN ???? ???? pp. LCCN ????

Negus:2002:RLB

[Neg02] Chris Negus. Red Hat Linux 8 bible. Wiley, New York, NY, USA, 2002. ISBN 0-7645-4968-5. xviii + 1063 pp. LCCN QA76.73.O63 N44 2002.

Negus:2006:LBB

[Neg06a] Chris Negus. Linux bible: boot up to Fedora, KNOPPIX, Debian, SUSE, Ubuntu, and 7 other distributions. Wiley, New York, NY, USA, 2006 edition, 2006. ISBN 0-471-75489-7 (paper/DVD). xxix + 870 pp. LCCN QA76.76.O63 N42143 2006.

Negus:2006:LTI

[Neg06b] Chris Negus. Linux toys II: 9 cool new projects for home, office, and entertainment. ExtremeTech. Wiley, New York, NY, USA, 2006. ISBN 0-7645-7995-9 (paper/CD-ROM + e-book). xxii + 397 pp. LCCN QA76.76.O63 N4233

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0623/2005025375-d.html;
http://www.loc.gov/catdir/
enhancements/fy0662/2005025375-
b.html;   http://www.loc.
gov/catdir/toc/ecip0518/
2005025375.html.

Negus:2006:LLC

[Neg06c]  Chris Negus.  Live Linux

CDs: building and customizing bootables. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-243274-9 (paperback). 448 (est.) pp. LCCN QA76.76.O63 N4245 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0619/2006027573.html.

Negus:2007:LBB

[Neg07a] Chris Negus. Linux bible: boot up Ubuntu, Fedora, KNOP-PIX, Debian, SUSE, and 11 other distributions. Wiley, New York, NY, USA, 2007 edition, 2007. ISBN 0-470-08279-8 (paper/DVD). xxxi + 848 pp. LCCN QA76.76.O63 N422 2007. URL http:

//www.loc.gov/catdir/toc/
ecip078/2007000458.html.

Negus:2007:LLC

[Neg07b]  Chris Negus.  Live Linux

CDs: building and customizing bootables. Negus live Linux series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2007. ISBN 0-13-243274-9 (paperback). xix + 430 pp. LCCN QA76.76.O63 N4245 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0619/2006027573.html.

Nelson:2000:NPL

[Nel00] Erik Nelson. Network programming with Linux. C/C++ Users Journal, 18(9):26–??, September 2000. CODEN CCUJEX. ISSN 1075-2838.

Neuner:2006:ILS

[Neu06]  Steve Neuner.  An inside

look at scaling Linux to 1024 processors. In Anonymous [Ano06], page ?? ISBN ???? LCCN ???? URL http://

www.ice.gelato.org/oct06/
pres_pdf/gelato_ICE06oct_
scaling1024_neuner_sgi.pdf.

Nadi:2014:LKC

[NH14] Sarah Nadi and Ric Holt. The Linux kernel: a case study of build system variability. Journal of Software: Evolution and Process, 26(8):730–746, August 2014. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Nielsen:2000:ULD

[Nie00] Mark Nielsen. Using Linux to duplicate computers. SysAdmin, 9(4):16, 18, 20, 22, 25, April 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Nisley:2001:ESP

[Nis01a] Ed Nisley. Embedded space:

Penguin specs. Dr. Dobb’s Journal of Software Tools, 26 (6):139–141, June 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Nisley:2001:ESRc

[Nis01b] Ed Nisley. Embedded space:

Rating real time: Count the ways. Dr. Dobb’s Journal of Software Tools, 26(9):113– 115, September 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Nisley:2001:ESS

[Nis01c] Ed Nisley. Embedded space:

SEU meets Embedded Linux. Dr. Dobb’s Journal of Software Tools, 26(3):129–131, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/.

Nisley:2003:ESRb

[Nis03] Ed Nisley. Embedded space: Real-Time Linux. Dr. Dobb’s Journal of Software Tools, 28 (4):79–81, April 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7827/
                                        ddj0304j/.

Nadella:2011:MFH

[NJ11] Suneetha Nadella and Dharanipragada Janakiram. Message filters for hardening the Linux kernel. Software—Practice and Experience, 41(1):51–62, January 2011. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Nieh:2000:EV

[NL00] Jason Nieh and Ozgur Can Leonard. Examining VMware. Dr. Dobb’s Journal of Software Tools, 25(8):70, 72–74, 76, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Nicolau:2009:TEP

[NLK09] Alexandru Nicolau, Guangqiang Li, and Arun Kejariwal. Techniques for efficient placement of synchronization primitives. ACM SIGPLAN Notices, 44 (4):199–208, April 2009. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Norton:1999:CGL

[Nor99] Peter Norton. Peter Norton’s Complete guide to Linux. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-672-31573-4. xv + 581 pp. LCCN QA76.76.O63 N67792 2000. Edited by Bryan Gambrel.

Norton:2002:ISP

[Nor02]  Roger Norton.  IT sys-

tems perspective: Using virtual Linux servers. Computer, 35(11):106–107, November 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2002/11/ry106.
htm;   http://csdl.computer.
org/dl/mags/co/2002/11/ry106.
pdf.

Narduzzo:2003:MAG

[NR03] A. Narduzzo and A. Rossi.

Modularity in action: GNU/ Linux and Free/Open Source software development model unleashed. World-Wide Web document., May 12, 2003.

URL  http://opensource.
mit.edu/papers/narduzzorossi.
pdf.

Nemeth:2002:LAH

[NSH02] Evi Nemeth, Garth Snyder, and Trent Hein. Linux Administration Handbook. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-008466-2. xxxi + 890 pp. LCCN QA76.76.O63 N448 2002. US$49.99.

Nemeth:2006:LAH

[NSH06] Evi Nemeth, Garth Snyder, and Trent R. Hein. Linux administration handbook. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2006. ISBN 0-13-148004-9 (paperback). xxxvii + 1001 pp. LCCN QA76.76.O63 N448 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0620/2006030150.html.

Nemeth:2010:ULS

[NSHW10] Evi Nemeth, Garth Snyder, Trent R. Hein, and Ben Whaley. UNIX and Linux system administration handbook. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 2010. ISBN 0-13-148005-7. xlvii + 1279 pp. LCCN QA76.76.O63 N45 2010.

Nahum:2007:ESS

[NTW07] Erich M. Nahum, John Tracey, and Charles P. Wright. Evaluating SIP server performance. ACM SIGMETRICS Performance Evaluation Review, 35 (1):349–350, June 2007. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Nutt:2001:KPL

[Nut01] Gary J. Nutt. Kernel projects for Linux. Addison-Wesley Longman, Harlow, Essex CM20 2JE, England, 2001. ISBN 0-201-61243-7. xvi + 239 pp. LCCN QA76.76.O63 N885 2001.

Nieh:2006:ETO

[NV06] Jason Nieh and Chris Vaill. Experiences teaching operating systems using virtual platforms and Linux. Operating Systems Review, 40(2):100–104, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Nightingale:2008:RS

[NVCF08] Edmund B. Nightingale, Kaushik Veeraraghavan, Peter M. Chen, and Jason Flinn. Rethink the sync. ACM Transactions on Computer Systems, 26(3):6:1– 6:26, September 2008. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Negus:2003:LTC

[NW03] Chris Negus and Chuck Wolber. Linux toys: 13 cool projects for home, office, and entertainment. Extreme-Tech. Wiley, New York, NY, USA, 2003. ISBN 0-7645-2508-5. xxv + 330 pp. LCCN QA76.76.O63 N423

2003. URL http://www.loc.
gov/catdir/bios/wiley046/
2003101901.html;   http:/
/www.loc.gov/catdir/description/
wiley039/2003101901.html;
 http://www.loc.gov/catdir/
toc/wiley041/2003101901.
html.

Ottaviano:2023:CLL

[OBSB23] Alessandro Ottaviano, Thomas Benz, Paul Scheffler, and Luca Benini. Cheshire: a lightweight, Linux-capable RISC-V host platform for domain-specific accelerator plug-in. IEEE Transactions on Circuits and Systems II: Express Briefs, 70(10): 3777–3781, 2023. ISSN 1549-7747 (print), 1558-3791 (electronic).

OGorman:2001:OSL

[O’G01] John O’Gorman. Operating systems with Linux. Cornerstones of computing. Palgrave, Basingstoke, UK, 2001. ISBN 0-333-94745-2. xv + 462 pp. LCCN QA76.76.O63 O336 2001.

OKane:2000:K

[O’K00a]  James O’Kane.  Kickstart.

SysAdmin, 9(1):33–34, 36, January 2000. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

OKeefe:2000:GFS

[O’K00b] Matthew O’Keefe. The Global File System: A cluster file system for Linux. Journal of Linux Technology, 1(2):33–39, 2000. ISSN 1527-2761.

Olczak:2001:KSU

[Olc01] Anatole Olczak. The Korn shell: Unix and Linux programming manual. Addison-Wesley, Reading, MA, USA, third edition, 2001. ISBN 0-201-67523-4. xxi + 446 pp. LCCN QA76.73.K67 O38 2001.

Olszewski:2007:JIN

[OMCB07] Marek Olszewski, Keir Mierle, Adam Czajkowski, and Angela Demke Brown. JIT instrumentation: a novel approach to dynamically instrument operating systems. Operating Systems Review, 41(3):3–16, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Olshefski:2006:UMC

[ON06] David Olshefski and Jason Nieh. Understanding the management of client perceived response time. ACM SIGMET-RICS Performance Evaluation Review, 34(1):240–251, June 2006. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ooms:2013:RPE

[Oom13] Jeroen Ooms. The RAppArmor package: Enforcing security policies in R using dynamic sandboxing on Linux. Journal of Statistical Software, 55(7): ??, November 2013. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v55/i07.

OReilly:1996:RLC

[OR96] O’Reilly and Associates and Red Hat Software. Running Linux Companion CD-ROM. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-171-2. 100 pp. US$24.95. URL http:/

/www.ora.com/gnn/bus/ora/
item/runuxcd.html.

Oxer:2006:UH

[ORC06] Jonathan Oxer, Kyle Rankin, and Bill Childers. Ubuntu Hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52720-9. xix + 426 pp. LCCN QA76.76.O63 2006. EUR 29.00.

OSullivan:2002:MCA

[O’S02] Maureen O’Sullivan. Making copyright ambidextrous: An expose of copyleft. Journal of Information, Law & Technology (JILT), 2002(3), 2002. ISSN 1361-4169. URL

                                        http://elj.warwick.ac.uk/
                                        jilt/02-3/osullivan.html;
                                         http://www2.warwick.ac.
                                        uk/fac/soc/law/elj/jilt/
                                        2002_3/osullivan/.

Osvik:2005:CAC

[OST05] Dag Arne Osvik, Adi Shamir, and Eran Tromer. Cache attacks and countermeasures: the case of AES: (extended version). Technical report, Department of Computer Science and Applied Mathematics, Weizmann Institute of Science, Rehovot 76100, Israel, October 8, 2005. URL http:

                                        //www.wisdom.weizmann.ac.
                                        il/~tromer/papers/cache.
                                        pdf.

Oualline:1997:DL

[Oua97a] Steve Oualline. Discover Linux.

I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-3105-0. xxvi + 438 pp. LCCN QA76.76.O63O83 1997. US$24.99.

Oualline:1997:LGH

[Oua97b] Steve Oualline. Linux and the gorilla are here to stay; not so Java. HP Chronicle, 14(2):15, January 1997. ISSN 0892-2829.

Parker:1997:SLU

[P^{+}97] Timothy Parker et al. Slackware Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, third edition, 1997. ISBN 0-672-31012-0. xlvi + 1331 pp. LCCN QA76.76.O63S585 1997.

Pritchard:2006:LLC

[P^{+}06] Steven Pritchard et al., editors. LPI Linux certification in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-00528-8. xviii + 961 pp. LCCN QA76.76.O63; QA76.76.O63 L65 2006eb. URL http:/

/www.oreilly.com/catalog/
9780596005283.

Parziale:2008:ZVL

[P^{+}08] Lydia Parziale et al., editors.

z/VM and Linux on IBM System z: the virtualization cookbook for RHEL 5.2. Number SG24-7492-00 in IBM redbooks. IBM Corporation, San Jose, CA, USA, 2008. ISBN 0-7384-3181-8. xvi + 250 pp. LCCN QA76.76.O63 Z867

                                        2008. URL http://proquest.
                                        safaribooksonline.com/?fpi=
                                        0738431818.

Parker:1996:LSA

[Par96] Tim Parker. Linux system administrator’s survival guide. SAMS Publishing, Indianapolis, IN, USA, 1996. ISBN 0-672-30850-9. xxiii + 765 pp. LCCN QA76.76.O63 P36 1996.

Parker:1998:LU

[Par98a] Tim Parker. Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, third edition, 1998. ISBN 0-672-31372-3. xxxvi + 1114 pp. LCCN QA76.76.O63P36493 1998.

Parker:1998:L

[Par98b] Timothy Parker. Linux. Le tout en poche. Simon & Schuster Macmillan (France), Paris, France, 1998. ISBN 2-7440-0424-3. xi + 407 pp. LCCN ????

Parker:1999:LRD

[Par99] Tim Parker. Linux: ressources d’experts. CampusPress France, Paris, France, 1999. ISBN 2-7440-0581-9. xviii + 1034 pp. LCCN ????

Parker:2000:LSA

[Par00] Tim Parker. Linux system administrator’s survival guide. SAMS Publishing, Indianapolis, IN, USA, second edition, 2000. ISBN 0-672-31793-1. xx + 740 pp. LCCN QA76.76.O63 P36 2000.

Paulson:1995:BYO

[Pau95] Linus Paulson. Build Your Own Web Site for Almost Free: Create a World Wide Web Site with HTML, Perl, and Linux. Waite Group Press, Corte Madera, CA, USA, September 1995. ISBN 1-57169-053-0. 700 pp. LCCN ???? US$36.95.

Pitts:1998:RHLb

[PB98a] David Pitts and Bill Ball. Red Hat Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, third edition, 1998. ISBN 0-672-31410-X. xxxiv + 1020 pp. LCCN QA76.76.O63P5615 1998.

Pitts:1998:RLU

[PB98b] David Pitts and Bill Ball. Red Hat Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, third edition, 1998. ISBN 0-672-31410-X. xxxiv + 1020 pp. LCCN QA76.76.O63P5615 1998.

Petreley:2005:LDH

[PB05] Nick Petreley and Jono Bacon. Linux desktop hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00911-9. xxi + 318 pp. LCCN QA76.76.O63 P528645 2005. URL http:/

/www.oreilly.com/catalog/
9780596009113.

Pariag:2007:CPW

[PBH^{+}07] David Pariag, Tim Brecht, Ashif Harji, Peter Buhr, Amol Shukla, and David R. Cheriton. Comparing the performance of Web server architectures. Operating Systems Review, 41(3):231–243, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pagani:2022:LBS

[PBM^{+}22] Marco Pagani, Alessandro Biondi, Mauro Marinoni, Lorenzo Molinari, Giuseppe Lipari, and Giorgio Buttazzo. A Linuxbased support for developing real-time applications on heterogeneous platforms with dynamic FPGA reconfiguration. Future Generation Computer Systems, 129(??):125– 140, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X21004362.

Peter:2008:SES

[PBR^{+}08] Simon Peter, Andrew Baumann, Timothy Roscoe, Paul Barham, and Rebecca Isaacs. 30 seconds is not enough!: a study of operating system timer usage. Operating Systems Review, 42(4):205–218, May 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Phillips:2001:TIL

[PC01] Iain Phillips and Jon Crowcroft.

TCP/IP and the Linux protocol implementation: systems code for the Linux Internet. Wiley Networking Council series. Wiley, New York, NY, USA, 2001. ISBN 0-471-40882-4 (cloth). 925 pp. LCCN TK5105.585 .T34 2002.

Palmer:2002:GUU

[PDG02] Michael Palmer, Jack Dent, and Tony Gaddis. Guide to UNIX using Linux. Networking. Thomson/Course Technology, Boston, MA, USA, second edition, 2002. ISBN 0-619-12147-5. xx + 540 pp. LCCN QA76.76.O63 P35 2002. Previous edition written by Jack Dent and Tony Gaddis. Includes a copy of Red Hat Linux 7.2 publisher’s edition.

Pennington:1999:GGA

[Pen99] Havoc Pennington. GTK+/ Gnome application development. New Riders Publishing, Carmel, IN, USA, 1999. ISBN 0-7357-0078-8. xviii, 492 pp. LCCN QA76.76.D47 P46 1999.

Petersen:1996:LCR

[Pet96]  Richard Petersen.  Linux,

the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882189-4. xxiii + 840 pp. LCCN QA76.76.O63P523 1996. US$39.40.

Petersen:1998:LPR

[Pet98a] Richard Petersen. Linux programmer’s reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 1998. ISBN 0-07-882587-3. xiv + 303 pp. LCCN QA76.76.O63 P525 1998. US$16.99.

Petersen:1998:LCR

[Pet98b] Richard Petersen. Linux: The Complete Reference. McGraw-Hill, New York, NY, USA, second edition, 1998. ISBN 0-07-882461-3. xxvi + 1059 pp. LCCN QA76.76.O63P523 1998. US$49.99. Includes CD-ROM.

Petersen:1999:LCR

[Pet99] Richard Petersen. Linux: the complete reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, third edition, 1999. ISBN 0-07-212164-5. xxvii + 929 pp. LCCN QA76.76.O63 P523 1999.

Petersen:2000:LPR

[Pet00a] Richard Petersen. Linux programmer’s reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, second edition, 2000. ISBN 0-07-212355-9. xix + 443 pp. LCCN QA76.8.U65 P484 2000. US$19.99.

Petersen:2000:LCR

[Pet00b] Richard Petersen. Linux: The Complete Reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, fourth edition, 2000. ISBN 0-07-212940-9. xxxv + 1257 pp. LCCN QA76.76.O63 P523 2001. US$19.99.

Petron:2000:LER

[Pet00c] Ed Petron. Linux essential reference. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 0-7357-0852-5. xxiii + 332 pp. LCCN QA76.76.O63 P5286 2000.

Petersen:2001:LPB

[Pet01a] Richard Petersen. Linux programming: a beginner’s guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212743-0. xxiii + 422 pp. LCCN QA76.76.O63 P5228 2001.

Petersen:2001:LCR

[Pet01b] Richard Petersen. Linux: the complete reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, fourth edition, 2001. ISBN 0-07-212940-9. xxxv + 1257 pp. LCCN QA76.76.O63 P523 2001. Includes CD-ROM.

Petersen:2002:LCR

[Pet02a] Richard Petersen. Linux: the complete reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, fifth edition, 2002. ISBN 0-07-222505-X, 0-07-222506-8 (book), 0-07-222507-6 (DVD). xxvii + 911 pp. LCCN QA76.76.O63 P523 2002. DVD-ROM includes complete 5 CD-ROM set for the entire Red Hat distributions and comprehensive set of Linux software applications, including the GNU software packages (graphics, communications, publishing, editing, programming, games), as well as development tools, and Internet servers (FTP, Web, mail, news, and DNS).

Petersen:2002:RLC

[Pet02b] Richard Petersen. Red Hat Linux: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 2002. ISBN 0-07-219178-3. xxx + 1117 pp. LCCN QA76.76.O63 P5237 2002.

Probst:1996:PLL

[PF96] Stefan Probst and Ralf Flaxa. The power Linux: Linux 1.2, LST-distribution 2.1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-14556-7. vi + 196 pp. LCCN QA76.76.O63P758 1996.

Probst:1997:PLL

[PF97] Stefan Probst and Ralf Flaxa. The power Linux: Linux 2.0, LST-distribution 2.2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., international edition, 1997. ISBN 3-540-14556-7. vi + 196 pp. LCCN ????

Pfaffenberger:1999:LCE

[Pfa99] Bryan Pfaffenberger. Linux clearly explained. Academic Press, New York, NY, USA, 1999. ISBN 0-12-553169-9, 0-12-553170-2 (CD-ROM: Red Hat Linux 6.0), 0-12-553171-0 (CD-ROM: Corel WordPerfect 8.0). 350 pp. LCCN QA76.76.O63 P55 1999.

Pfaffenberger:2000:LCR

[Pfa00] Bryan Pfaffenberger. Linux command reference. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2748-7. 640 pp. LCCN QA76.76 .P435 2000.

Pfaffenberger:2001:LNC

[Pfa01] Bryan Pfaffenberger. Linux networking clearly explained. Academic Press, New York, NY, USA, 2001. ISBN 0-12-533171-1. ix + 390 pp. LCCN QA76.76.O63 P533 2001.

Probst:1997:PLK

[PFH97] S. Probst, R. Flaxa, and F. Hantelmann. The Power Linux Kit. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-98263-9. ???? pp. LCCN ???? US$65.95. Two volumes. Includes two CD-ROMs.

Pratikakis:2006:LCS

[PFH06] Polyvios Pratikakis, Jeffrey S.

Foster, and Michael Hicks. LOCKSMITH: context-sensitive correlation analysis for race detection. ACM SIGPLAN Notices, 41(6):320–331, June 2006. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Patel:2022:SLL

[PFHLN22] Keyur Patel, João Faccin, Abdelwahab Hamou-Lhadj, and Ingrid Nunes. The sense of logging in the Linux kernel. Empirical Software Engineering, 27(6):??, November 2022. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10664-022-10136-3.

Petrisko:2020:BAO

[PGW^{+}20] Daniel Petrisko, Farzam Gilani, Mark Wyse, Dai Cheol Jung, Scott Davidson, Paul Gao, Chun Zhao, Zahra Azad, Sadullah Canakci, Bandhav Veluri, Tavio Guarino, Ajay Joshi, Mark Oskin, and Michael Bedford Taylor. BlackParrot: An agile open-source RISC-V multicore for accelerator SoCs. IEEE Micro, 40(4):93– 102, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Phillips:2000:LMS

[Phi00] Dave Phillips. Linux Music and Sound. No Starch Press, San Francisco, CA, USA, 2000. ISBN 1-886411-34-4. xx + 408 pp. LCCN MT723 .P53 2000. US$39.95.

Pinto:2002:PGC

[Pin02] Paulo Pinto. Programming in GNOME with C++. C/C++ Users Journal, 20(5):28–??, May 2002. CODEN CCUJEX. ISSN 1075-2838.

Pitts:1998:BRL

[Pit98a] David Pitts. La biblia de Red Hat Linux. Informatica Personal-Profesional. Anaya Multimedia, Madrid, Spain, 1998. ISBN 84-415-0524-1. 790 pp. LCCN ????

Pitts:1998:RHLa

[Pit98b] David Pitts. Red Hat Linux.

SAMS Publishing, Indianapolis, IN, USA, second edition, 1998. ISBN 0-672-31173-9. xxv + 684 pp. LCCN QA76.76.O63R4325 1998.

Papadopoulos:2001:NRC

[PKB01] Philip M. Papadopoulos, Mason J. Katz, and Greg Bruno. NPACI rocks clusters: Tools for easily deploying and maintaining manageable high-performance Linux clusters. Lecture Notes in Computer Science, 2131:10– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2131/21310010.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2131/21310010.
pdf.

Papadopoulos:2003:NRT

[PKB03] Philip M. Papadopoulos, Mason J. Katz, and Greg Bruno. NPACI Rocks: tools and techniques for easily deploying manageable Linux clusters.

Concurrency and Computation: Practice and Experience, 15 (7–8):707–725, June/July 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pearce:2002:GDI

[PKFH02] David J. Pearce, Paul H. J. Kelly, Tony Field, and Uli Harder. GILK: A dynamic instrumentation tool for the Linux kernel. Lecture Notes in Computer Science, 2324:220– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2324/23240220.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2324/23240220.
pdf.

Poe:2006:BBS

[PL06] James Poe and Tao Li. BASS: a benchmark suite for evaluating architectural security systems. ACM SIGARCH Computer Architecture News, 34(4): 26–33, September 2006. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Padioleau:2008:DAC

[PLHM08] Yoann Padioleau, Julia Lawall, René Rydhof Hansen, and Gilles Muller. Documenting and automating collateral evolutions in Linux device drivers. Operating Systems Review, 42 (4):247–260, May 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Park:2008:QSL

[PLKH08] Jiyong Park, Jaesoo Lee, Saehwa Kim, and Seongsoo Hong. Quasistatic shared libraries and XIP for memory footprint reduction in MMU-less embedded systems. ACM Transactions on Embedded Computing Systems, 8(1):6:1–6:??, December 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Padioleau:2006:UCE

[PLM06] Yoann Padioleau, Julia L.

Lawall, and Gilles Muller. Understanding collateral evolution in Linux device drivers. Operating Systems Review, 40(4):59– 71, October 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Park:2002:RTM

[PNK02] Nam-Sup Park, Sang-Jun Nam, and Tai-Yun Kim. Realtime multimedia data transmission module based on Linux. Lecture Notes in Computer Science, 2343:504–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2343/23430504.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2343/23430504.pdf.

Poduska:1997:LPI

[Pod97] Paul R. Poduska. Linux pro: installation and more. Work-Group Solutions, Aurora, CO, USA, 1997. ISBN 1-888894-39-3 (paperback). 134 pp. LCCN ????

Pomerantz:2004:LHU

[Pom04] Gregory M. Pomerantz. The legal history of UNIX and free software. Web document., June 19, 2004. URL http:

                                        //wiki.tuhs.org/lib/exe/
                                        fetch.php?media=publications:
                                        theses:gmp_thesis.pdf.

Poniatowski:2005:LHI

[Pon05] Marty Poniatowski. Linux on HP Integrity Servers: system administration for Itaniumbased systems. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2005. ISBN 0-13-140000-2. xxvi + 332 pp. LCCN QA76.76.O63 P652 2005.

Popov:1994:SAE

[Pop94] Angel Popov, editor. Systems for automation of engineering and research: 8th International conference — October 1994, Varna, Bulgaria. SAER Forum Group, ????, 1994. ISBN 954-8329-06-9 (??invalid checksum??). LCCN TJ212.2 .I572 1994.

Perello:1996:OSI

[PPSM96] C. Perello, N. Poch, C. Schroeter, and J. Millan. An open system to interface IEEE-488 measurement devices designed in a microelectronics environment. In Quality Measurements: The Indispensable Bridge between Theory and Reality (No Measurements? No Science!) Joint Conference - 1996: IEEE Instrumentation and Measurement Technology Conference and IMEKO Technical Committee 7. Conference Proceedings: Sheraton Brussells Hotel and Tower, Brussels, Belgium, June 4–6, 1996, volume 1, pages 2–?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3312-8. LCCN TK7878 .I3295 1996. IEEE catalog number 96CH35936.

Passos:2021:SFS

[PQM^{+}21] L. Passos, R. Queiroz, M. Mukelabai, T. Berger, S. Apel, K. Czarnecki, and J. A. Padilla. A study of feature scattering in the Linux kernel. IEEE Transactions on Software Engineering, 47(1):146–164, 2021. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Purcell:1996:LBG

[PR96a] John Purcell and Amanda Robinson, editors. The Linux Bible: The GNU Testament. Yggdrasil Computing, Inc., Berkeley, CA, USA, fourth edition, 1996. ISBN 1-883601-10-X (??), 1-883601-20-7. v + 1886 pp. LCCN QA76.76.O63 L5458 1996. US$39.95. URL

http://www.yggdrasil.com/

                                        Products/bible4.html. In-

cludes CD ROM.

Purcell:1996:LE

[PR96b] John Purcell and Amanda Robinson. Linux encyclopedia. WorkGroup Solutions, Aurora, CO, USA, fifth edition, 1996. ISBN 0-9644309-2-4 (paperback). v + 1886 pp. LCCN ????

Purcell:1997:LE

[PR97] John Purcell and Amanda Robinson. Linux encyclopedia. WorkGroup Solutions, Aurora, CO, USA, fifth edition, 1997. ISBN 0-9644309-2-4. 1886 pp. LCCN ????

Preston:2007:BR

[Pre07] W. Curtis Preston. Backup and recovery. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10246-1 (paperback). xxviii + 729 pp. LCCN QA76.9.B32 P74

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0715/2007273944-d.html;
                                         http://www.oreilly.com/
                                        catalog/9780596102463/index.
                                        html.

Pritchard:1999:LRC

[Pri99] Kara J. Pritchard. Linux Red Hat Certified Engineer Exam Cram. Coriolis Group Books, Scottsdale, AZ, USA, 1999. ISBN 1-57610-487-7. xxvii + 380 pp. LCCN QA76.3 .P75 1999. US$29.99.

Pandey:2007:SCM

[PS07] Nirved Pandey and G. K.

Sharma. Startup comparison for message passing libraries with DTM on Linux clusters. The Journal of Supercomputing, 39(1):59–72, January 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=39&issue=1&spage=
59.

Paludo:2022:NAL

[PS22] Rogério Paludo and Leonel Sousa. NTT architecture for a Linux-ready RISC-V fullyhomomorphic encryption accelerator. IEEE Transactions on Circuits and Systems I: Regular Papers, 69(7):2669–2682, 2022. ISSN 1549-8328 (print), 1558-0806 (electronic).

Patrick:2008:CEO

[PSK08] Christina M. Patrick, Seung-Woo Son, and Mahmut Kandemir. Comparative evaluation of overlap strategies with study of I/O overlap in MPI-IO. Operating Systems Review, 42(6): 43–49, October 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Palix:2011:FLT

[PTS^{+}11] Nicolas Palix, Gaël Thomas, Suman Saha, Christophe Calvès, Julia Lawall, and Gilles Muller. Faults in Linux: ten years later. ACM SIGARCH Computer Architecture News, 39(1): 305–318, March 2011. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Palix:2014:FL

[PTS^{+}14] Nicolas Palix, Gael Thomas, Suman Saha, Christophe Calvès, Gilles Muller, and Julia Lawall. Faults in Linux 2.6. ACM Transactions on Computer Systems, 32(2):4:1–4:??, June 2014. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Publishers:2011:LIQ

[Pub11] Vibrant Publishers. Linux interview questions you’ll most likely be asked. Interview questions you’ll most likely be asked. Vibrant Publishers, Erie, CO, USA, 2011. ISBN 1-4564-7381-6. ???? pp. LCCN ????

Purcell:1995:WLP

[Pur95] John Purcell. WGS Linux pro “the sky’s the limit.”, 1995. ISBN 0-9644309-1-6. 4 computer laser optical discs.

Purcell:1997:LCC

[Pur97] John Purcell. Linux complete command reference. Red Hat Press, Indianapolis, IN, USA, 1997. ISBN 0-672-31104-6. xxxii + 1495 pp. LCCN QA76.76.O63P87 1997.

Purdy:2004:LIP

[Pur04] Gregor N. Purdy. Linux iptables: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2004. ISBN 0-596-00569-5. iii + 91 pp. LCCN QA76.76.O63 P873 2004. URL http:/

/www.oreilly.com/catalog/
9780596005696.

Pasztor:2002:PBP

[PV02] Attila Pásztor and Darryl Veitch. PC based precision timing without GPS. ACM SIG-METRICS Performance Evaluation Review, 30(1):1–10, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Petrovsky:2001:LDB

[PWN01] Michele Petrovsky, Stephen Wysham, and Mojo Nichols. Linux database bible. Hungry Minds, 909 Third Avenue, New York, NY 10022, USA, 2001. ISBN 0-7645-4641-4. xxi + 715 pp. LCCN QA76.76.O63 P5288 2001. URL http:

//www.loc.gov/catdir/toc/
wiley024/2001092731.html.

Quigley:2001:CLS

[QH01] Ellie Quigley and Scott Hawkins.

The Complete Linux Shell Programming Training Course. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-040676-7. ???? pp. LCCN ???? US$109.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130406767.html.

Quigley:2000:LSE

[Qui00] Ellie Quigley. Linux shells by example. Open source technology series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-014711-7. xviii + 761 pp. LCCN QA76.76.O63 Q538 2000.

Quigley:2002:USE

[Qui02] Ellie Quigley. UNIX Shells by Example. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2002. ISBN 0-13-066538-X (paperback). xix + 1015 pp. LCCN QA76.76.O63 Q54 2002.

R:1999:LI

[R.99] Shailaja V. R. Linux in India. Performance Computing, 17(11):15–17, October 1999. CODEN UNRED5. ISSN 0742-3136.

Raman:1996:ESA

[Ram96] T. V. Raman. Emacspeak — direct speech access. Annual ACM Conference on Assistive Technologies, Proceedings, pages 32–36, 1996.

Rankin:1997:NBG

[Ran97] Bob Rankin. The no B.S. guide to Linux. No Starch Press, San Francisco, CA, USA, 1997. ISBN 1-886411-04-2 (paperback). xxi + 335 pp. LCCN QA76.76.O63R365 1997.

Rankin:1999:NBG

[Ran99] Bob Rankin. The no B.S. guide to Red Hat Linux 6. No Starch Press, San Francisco, CA, USA, 1999. ISBN 1-886411-30-1 (paperback). ???? pp. LCCN QA76.76.O63R366 1999.

Rankin:2000:NBG

[Ran00] Bob Rankin. The no B.S. guide to Linux. No Starch Press, San Francisco, CA, USA, second edition, 2000. ISBN 1-886411-30-1. xxxi + 422 pp. LCCN QA76.76.O63 R366 2000. Includes CD-ROM.

Rankin:2004:KHI

[Ran04] Kyle Rankin. Knoppix hacks:

100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00787-6 (paperback). xx + 314 pp. LCCN QA76.76.O63 R368 2004.

Rankin:2005:KPR

[Ran05] Kyle Rankin. Knoppix pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10075-2. vii + 84 pp. LCCN QA76.76.O63 R36 2005.

Rankin:2006:LMH

[Ran06] Kyle Rankin. Linux multimedia hacks: tips and tools for taming images, audio and video. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10076-0. xviii + 310 pp. LCCN QA76.575; QA76.575 .R345 2006eb. URL http:/

/www.oreilly.com/catalog/
9780596100766.

Rankin:2007:KH

[Ran07] Kyle Rankin. Knoppix hacks. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51493-X (paperback). xxv + 391 pp. LCCN QA76.76.O63; QA76.76.O63 R36 2007eb. URL http:/

/www.oreilly.com/catalog/
9780596514938.

Rash:2007:LFA

[Ras07] Michael Rash. Linux firewalls: attack detection and re-

sponse with iptables, psad,

and fwsnort. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-141-7. xix + 308 pp. LCCN QA76.9.A25 R36 2007. URL http:/

/www.oreilly.com/catalog/
9781593271411.

Rautiainen:2001:TLM

[Rau01] Sami Rautiainen. Travelling with Linux malware: Is Linux security for real? Information Security Technical Report, 6(4):58–64, December 1, 2001. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S136341270100406X.

Raymond:1998:LU

[Ray98] Eric Raymond, editor. Linux Undercover: Linux Secrets from the LDP. Red Hat Software, Inc., Westport, CT, USA, 1998. ISBN 1-888172-05-3. 2018 pp. LCCN ???? UK£34.99, US$39.99. URL http://

genesis.ukpost.com/undercover.
html.

Ray:1999:STY

[Ray99a] John Ray. Sams teach yourself Linux in 10 minutes. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31524-

6. vii + 212 pp. LCCN QA76.76.O63 R3918 1999.

Raymond:1999:ILO

[Ray99b]  Eric Raymond.  Interview:

Linux and open-source success. IEEE Software, 16(1):85–89, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1999/pdf/
s1085.pdf.

Raymond:1999:CBM

[Ray99c] Eric S. Raymond. The Cathedral & the Bazaar: Musings on Linux and Open Source by an Accidental Revolutionary. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-724-9. 288 pp. LCCN QA76.76.O63 R396 1999. US$19.95. URL http:/

/www.oreilly.com/catalog/
cb.

Ray:2001:MLS

[Ray01a] John Ray. Maximum Linux security. SAMS Publishing, Indianapolis, IN, USA, second edition, 2001. ISBN 0-672-32134-3. 800 pp. LCCN QA76.9.A25 M387 2001.

Raymond:2001:CBM

[Ray01b] Eric S. Raymond. The cathedral and the bazaar: musings on Linux and open source by an accidental revolutionary. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, revised edition, 2001. ISBN 0-596-00131-2, 0-596-00108-8 (paperback). xiv + 241 pp. LCCN QA76.76.O63 R397 2001.

Robbins:2005:CSS

[RB05a] Arnold Robbins and Nelson H. F. Beebe. Classic Shell Scripting. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00595-4. xxii + 534 pp. LCCN QA76.76.O63 R633 2005. URL http:/

/www.oreilly.com/catalog/

shellsrptg/. Also available in Chinese [RB08], French [RB05b], German [RB06a], Japanese [RB06c], and Polish [RB06b] translations.

Robbins:2005:ISS

[RB05b] Arnold Robbins and Nelson H. F. Beebe. Introduction aux scripts shell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 2-84177-375-2. xxii + 558 pp. LCCN ????

URL http://www.silicon.
fr/getarticle.asp?id=14015.

French translation of [RB05a] by Eric Jacoboni.

Robbins:2006:KSP

[RB06a] Arnold Robbins and Nelson H. F. Beebe. Klassische Shell-Programmierung: [automatisieren Sie Ihre Unix/Linux-Tasks]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 3-89721-441-

5. xxiii + 572 pp. LCCN QA76.76.O63 R563 2005. EUR 44.00; EUR 45.65 (AT).

                                        URL  http://www.gbv.de/
                                        dms/hebis-darmstadt/toc/
                                        17645067X.pdf.  German

translation of [RB05a] by

Kathrin Lichtenberg.

Robbins:2006:PSP

[RB06b] Arnold Robbins and Nelson H. F. Beebe. Programowanie skryptów powłoki. Helion, Gliwice, Poland, 2006. ISBN 83-246-0131-7. 557 + 2 pp. LCCN ???? URL http:/

                                        /www.empik.com/b/o/19/f1/
                                        19f16b85e0d75ae1d3a1e7062569fbb0.
                                        jpg;   http://www.empik.
                                        com/programowanie-skryptow-
                                        powloki-ksiazka,360529,p.

Polish translation of [RB05a] by Przemysław Szeremiota.

Robbins:2006:SSS

[RB06c] Arnold Robbins and Nelson H. F. Beebe. Shōkai shieru sukuriputo. Orairī Japan, Tōkyō, Japan, 2006. ISBN 4-87311-267-2. 345 pp. LCCN ???? Japanese translation of

[RB05a] by Aoi Hyūga.

Robbins:2008:SJB

[RB08] Arnold Robbins and Nelson H. F. Beebe. Shell Jiao Ben Xue Xi Zhi Nan = Shell Script Study Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 7-111-25504-6. vi + 494 pp. LCCN ???? 79.00 renminbi. Simplified Chinese translation of [RB05a].

Ricart:1999:STY

[RBHB99] Manuel Alberto Ricart, Grace Buechlein, Gregory Harris, and Laura Bulcher. Sams Teach Yourself Linux in 10 minutes. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31524-6. vii + 212 pp. LCCN QA76.76.O63R3918 1999. US$12.99.

Rodeh:2013:BLB

[RBM13] Ohad Rodeh, Josef Bacik, and Chris Mason. BTRFS: The Linux B-tree filesystem. ACM Transactions on Storage, 9(3): 9:1–9:32, August 2013. CO-DEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Russel:1998:ULA

[RC98] Charlie Russel and Sharon Crawford. Unix and Linux answers: certified tech support. Osborne/McGraw-Hill, Berkeley, CA, USA, 1998. ISBN 0-07-882446-X (paperback). xxvii + 308 pp. LCCN QA76.76.O63R886 1998. US$24.99.

Rubini:2001:LDD

[RC01] Alessandro Rubini and Jonathan Corbet. Linux Device Drivers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2001. ISBN 0-596-00008-1. 564 pp. LCCN QA76.76.D49 R92 2001. US$39.95. URL http://

                                        safari.oreilly.com/0596000081;
                                         http://www.oreilly.com/
                                        catalog/linuxdrive2.

RedHatSoftware:1998:ORL

[Red98] Red Hat Software, Inc. Official Redhat 5.1 Linux operating system for Intel computers. Red Hat Software, Inc., Westport, CT, USA, 1998. ISBN 1-888172-07-X, 1-888172-09-6 (Installation guide). 306 pp. LCCN ???? Includes three CD-ROMs.

Regupathy:2011:BYL

[Reg11] Rajaram Regupathy. Bootstrap yourself with Linux-USB strap: design, develop, debug, and validate embedded USB. Cengage Learning, Boston, MA, USA, 2011. ISBN 1-4354-5786-2. ???? pp. LCCN ????

Rehman:1999:LTP

[Reh99] Rafeeq Ur Rehman. Linux transparent proxy. SysAdmin, 8(5):29–30, 32–34, May 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/.

Rodriguez:2005:LKP

[RFS05] Claudia Salzberg Rodriguez, Gordon Fischer, and Steven Smolski. The Linux Kernel primer: a top-down approach for x86 and PowerPC architectures. Prentice Hall open source software development series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2005. ISBN 0-13-118163-7 (paperback). ???? pp. LCCN QA76.76.O63 R633 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip0514/2005016702.
html;   http://www.phptr.
com/title/0131181637.

Rodriguez:2006:LKP

[RFS06] Claudia Salzberg Rodriguez, Gordon Fischer, and Steven Smolski. The Linux Kernel primer: a top-down approach for X86 and PowerPC architectures. Prentice Hall open source software development series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-118163-7 (paperback). xxvii + 616 pp. LCCN QA76.76.O63 R633 2006.

Rubio-Gonzalez:2009:EPA

[RGGL^{+}09] Cindy Rubio-González, Haryadi S.

Gunawi, Ben Liblit, Remzi H. Arpaci-Dusseau, and Andrea C. Arpaci-Dusseau. Error propagation analysis for file systems. ACM SIGPLAN Notices, 44(6):270–280, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rodgers:1999:TSN

[RGK99] Jeremy B. Rodgers, Rhonda Kay Gaede, and Jeffrey H. Kulick. IN-Tune: an In-Situ noninvasive performance tuning tool for multi-threaded Linux on symmetric multiprocessing Pentium workstations. Software—Practice and Experience, 29(9):775–792, July 25, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=
                                        62501865;   http://www3.
                                        interscience.wiley.com/cgi-
                                        bin/fulltext?ID=62501865&
                                        PLACEBO=IE.pdf.

Rehn-Gostenmeier:2005:LE

[RGRG05] Gudrun Rehn-Göstenmeier and Ralph Rehn-Göstenmeier. Linux — Das Einsteigerseminar. (German) [Linux — The Beginner Seminar]. bhv, Bonn, Germany, 2005. ISBN 3-8266-7375-1. 448 pp. LCCN ????

Rossbach:2007:TUM

[RHP^{+}07] Christopher J. Rossbach, Owen S. Hofmann, Donald E. Porter, Hany E. Ramadan, Bhandari Aditya, and Emmett Witchel. TxLinux: using and managing hardware transactional memory in an operating system. Operating Systems Review, 41(6): 87–102, December 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ricart:1998:CIG

[Ric98] Manuel Alberto Ricart. The complete idiot’s guide to Linux. Que Corporation, Indianapolis, IN, USA, 1998. ISBN 0-7897-1826-X. vi + 336 pp. LCCN QA76.76.O63R49 1998.

Richardson:2013:GSB

[Ric13]  Matt Richardson.  Getting

started with BeagleBone. Maker Media, Sebastopol, CA, 2013. ISBN 1-4493-4537-9 (paperback), 1-4493-4536-0, 1-4493-4535-2 (e-book), 1-4493-4533-6 (e-book). xiii + 126 pp. LCCN TK7895.E42 R43 2013.

Richardson:2014:BEL

[Ric14] Matt Richardson. BeagleBone für Einsteiger: [Linux-basierte Elektronik-Projekte mit Python und JavaScript]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2014. ISBN 3-95561-409-3. xii + 134 pp. LCCN ????

Radhika:2020:CAF

[RKSV20] B. S. Radhika, N. V. Narendra Kumar, R. K. Shyamasundar, and Parjanya Vyas. Consistency analysis and flow secure enforcement of SELinux policies. Computers & Security, 94(??):Article 101816, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167404820300948.

Regnier:2008:EIH

[RLB08] Paul Regnier, George Lima, and Luciano Barreto. Evaluation of interrupt handling timeliness in real-time Linux operating systems. Operating Systems Review, 42(6):52–63, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reshetova:2018:TLK

[RLPA18] Elena Reshetova, Hans Liljestrand, Andrew Paverd, and N. Asokan. Toward Linux kernel memory safety. Software—Practice and Experience, 48(12):2237–2256, December 2018. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Reghenzani:2019:RTL

[RMF19] Federico Reghenzani, Giuseppe Massari, and William Fornaciari. The real-time Linux kernel: a survey on PRE-EMPT RT. ACM Computing Surveys, 52(1):18:1–18:??, February 2019. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3297714.

Russell:2001:LDR

[RMZ01] Steve Russell, Kathleen Mc-Givney, and Zeljka Zoranovic. Lotus Domino R5 Clustering.

P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2001. ISBN 0-13-060836-X. xii + 255 pp. LCCN HF5548.4.L673 R87 2001. US$39.99. URL http:

//www.phptr.com/ptrbooks/
ptr_013060836X.html.

Rodriguez:2001:EIL

[RÓ01] Ameneiros Rodríguez and Ibán Óscar. Estudio e implementación de una LAN para PYMES utilizando GNU/ LINUX como sistema operativo. (Spanish) [Study and implementation of a LAN with PYMES using GNU/ Linux as operating system]. E.U. Politécnica, Ferrol, Spain, 2001. ISBN ???? LCCN ???? Includes one CD-ROM.

Roberts:1995:ENP

[Rob95a] Bruce A. Roberts. External named pipes for Linux. Thesis (m.s.), Azusa Pacific University, Azusa, CA, USA, 1995.

Roberts:1995:LLC

[Rob95b] Rick Roberts. Linux — the low cost Unix. C/C++ Users Journal, 13(1):43–??, January 1, 1995. CODEN CCUJEX. ISSN 1075-2838.

Robbins:2004:LPE

[Rob04] Arnold Robbins. Linux Programming by Example: The Fundamentals. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2004. ISBN 0-13-142964-7. xxxii + 687 pp. LCCN QA76.76.O63 R568 2004. US$39.99, CAN$57.99.

Roelle:2002:HFS

[Roe02] Harald Roelle. A hot-failover state machine for gateway services and its application to a Linux firewall. Lecture Notes in Computer Science, 2506:181– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2506/25060181.htm; http:
//link.springer.de/link/
service/series/0558/papers/
2506/25060181.pdf.

Ronneburg:2001:DGLa

[Ron01a]  Frank Ronneburg.  Debian

GNU, Linux-Anwenderhandbuch. Lehmann, Berlin, Germany, 2001. ISBN 3-931253-73-2. 600 (est.) pp. LCCN ???? Includes CD-ROM.

Ronneburg:2001:DGLb

[Ron01b]  Frank Ronneburg.  Debian

GNU, Linux-Anwenderhandbuch. Addison-Wesley, Reading, MA, USA, 2001. ISBN 3-8273-1748-7. 600 (est.) pp. LCCN ???? Includes CD-ROM.

Ronneburg:2005:DGLb

[Ron05a]  Frank Ronneburg.  Debian

GNU/Linux Anwenderhandbuch: für Einsteiger, Umsteiger und Fortgeschritten. (German) [Debian GNU/Linux User Handbook: for Beginning, Switching, and Advanced Users]. Addison Wesley in Pearson Education Deutschland, München, Germany, 2005. ISBN 3-8273-2148-4. 744 pp. LCCN ???? EUR49.95. Includes one DVD.

Ronneburg:2005:DGLa

[Ron05b]  Frank Ronneburg.  De-

bian GNU/Linux: installation, administration, exploitation. (French) [Debian GNU/ Linux: installation, administration, use]. Campus Press, Paris, France, 2005. ISBN 2-7440-1941-0. 624 pp. LCCN ???? Includes one CD-ROM.

Rossmeyer:1998:BIL

[Ros98]  Brian Rossmeyer.  Book-

shelf: Intermediate level Linux guide. IEEE Software, 15(4): 90, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1998/pdf/

s4087.pdf. Critical review of [SE97].

Ross:2010:D

[Ros10] P. E. Ross. Discharged. IEEE Spectrum, 47(1):36–38, January 2010. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rehman:2002:LDP

[RP02] Rafeeq Rehman and Christopher Paul. The Linux Development Platform. Bruce Perens’ Open source series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002.

ISBN 0-13-009115-4. xxii + 294 pp. LCCN QA76.76.O63 R435 2003. US$49.99.

Roy:2009:LPF

[RPB^{+}09] Indrajit Roy, Donald E. Porter, Michael D. Bond, Kathryn S. McKinley, and Emmett Witchel. Laminar: practical fine-grained decentralized information flow control. ACM SIGPLAN Notices, 44(6):63– 74, June 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ram:1995:SDS

[RR95] P. Ram and D. K. Rand. Satan: double-edged sword. Computer, 28(6):82–83, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rahman:2015:RSL

[RR15a] Md Tajmilur Rahman and Peter C. Rigby. Release stabilization on Linux and Chrome. IEEE Software, 32(2):81–88, March/April 2015. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

                                        URL http://www.computer.
                                        org/csdl/mags/so/2015/02/
                                        mso2015020081-abs.html.

Ray:2015:UL

[RR15b] Deborah S. Ray and Eric J. Ray. Unix and Linux. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, fifth edition, 2015. ISBN 0-321-99754-9 (paperback). xii + 396 pp. LCCN QA76.76.O63 R3915

2015. URL http://proquest.
safaribooksonline.com/9780133793871.

Rossbach:2008:TMT

[RRH^{+}08] Christopher J. Rossbach, Hany E.

Ramadan, Owen S. Hofmann, Donald E. Porter, Aditya Bhandari, and Emmett Witchel. TxLinux and MetaTM: transactional memory and the operating system. Communications of the ACM, 51(9):83–91, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ramadan:2007:MTT

[RRP^{+}07] Hany E. Ramadan, Christopher J. Rossbach, Donald E. Porter, Owen S. Hofmann, Aditya Bhandari, and Emmett Witchel. MetaTM/ TxLinux: transactional memory for an operating system. ACM SIGARCH Computer Architecture News, 35(2):92–103, May 2007. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Ramadan:2008:MTT

[RRP^{+}08] Hany E. Ramadan, Christopher J. Rossbach, Donald E. Porter, Owen S. Hofmann, Aditya Bhandari, and Emmett Witchel. MetaTM/TxLinux: Transactional memory for an operating system. IEEE Micro, 28(1):42–51, January/February 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ray:2009:UL

[RRR09] Deborah S. Ray, Eric J. Ray, and Deborah S. Unix Ray. Unix and Linux. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, fourth edition, 2009. ISBN 0-321-63678-3 (paperback). xiv + 393 pp. LCCN QA76.76.O63 R3915 2009.

Rubanov:2014:EPL

[RS14] Vladimir Rubanov and Denis Silakov. Ensuring portability of Linux applications through standardization and knowledge base driven analysis. Science of Computer Programming, 91 (part B)(??):234–248, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167642314000240.

Rubini:1998:LDD

[Rub98] Alessandro Rubini. Linux Device Drivers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-292-1. xviii + 421 pp. LCCN QA87.76.O63 R82 1998. US$29.95. URL http://www.

                                        ora.com/catalog/linuxdrive/

;   http://www.oreilly.com/
catalog/linuxdrive.

Rule:2007:HCC

[Rul07] David Rule. How to Cheat at Configuring VMware ESX Server. Syngress Publishing, Inc., Rockland, MA, USA, 2007. ISBN 1-59749-194-2. xvii + 372 pp. LCCN QA76.9.V5 H69 2007. URL http:

//www.sciencedirect.com/
science/book/9781597491945.

Russell:2008:VTF

[Rus08] Rusty Russell. virtio: towards a de-facto standard for virtual I/O devices. Operating Systems Review, 42(5):95–103, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Siever:2000:LND

[S^{+}00] Ellen Siever et al. Linux in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2000. ISBN 0-596-00025-1. xv + 797 pp. LCCN QA76.76.O63 L5459 2000.

Siever:2005:LND

[S^{+}05] Ellen Siever et al., editors.

Linux in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2005. ISBN 0-596-00930-5. xiv + 925 pp. LCCN QA76.76.O63 L56 2005; QA76.76.O63 L5459 2005.

URL http://www.oreilly.
com/catalog/9780596009304.

Smith:2007:LAD

[S^{+}07] Bob Smith et al., editors. Linux appliance design: a hands-on guide to building Linux appliances. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-140-9. xxi + 356 pp. LCCN QA76.76.O63 L545115 2007. URL http:/

/www.oreilly.com/catalog/
9781593271404.

Su:2007:AIC

[SAF07] Ya-Yunn Su, Mona Attariyan, and Jason Flinn. AutoBash: improving configuration management with operating system causality analysis. Operating Systems Review, 41(6): 237–250, December 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

SDT:2001:SLG

[Sai01] Sair Development Team. Sair Linux and GNU certification level II: Apache and Web servers. Wiley, New York, NY, USA, 2001. ISBN 0-471-40537-X. xix + 300 pp. LCCN A76.3 .M3236 2001.

Scordino:2019:RTE

[SAL19] Claudio Scordino, Luca Abeni, and Juri Lelli. Real-time and energy efficiency in Linux: theory and practice. ACM SIGAPP Applied Computing Review, 18(4):18–30, January 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3307624.3307627.

Sanders:1998:ELO

[San98] James Sanders. Edgewatch:

Linux, open source, and software’s future. IEEE Software, 15(5):88–91, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1998/pdf/
s5088.pdf.

Shingledecker:2008:ODS

[SAN08] Robert Shingledecker, John Andrews, and Chris Negus. The official Damn Small Linux book: the tiny adaptable Linux that runs on anything. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2008. ISBN 0-13-233869-6 (paperback). xx + 426 pp. LCCN QA76.76.O63 S555554 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0718/2007020589.html.

Savio:2011:OCL

[Sav11] Hubert Savio. Oracle Certified Linux Expert exam cram: OCE exam: 1Z0-046: managing Oracle on Linux Certified Expert, volume 38 of Oracle in-focus series. Rampant TechPress, Kittrell, NC, USA, 2011. ISBN 0-9844282-1-6. ???? pp. LCCN ????

Shen:1999:LKM

[SB99] Hongfeng Shen and Theodore P. Baker. A Linux kernel module implementation of restricted Ada tasking. ACM SIGADA Ada Letters, 19(2):96–103, June 1999. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Sieh:2002:UVS

[SB02] Volkmar Sieh and Kerstin Buchacker. UMLinux — A versatile SWIFI tool. Lecture Notes in Computer Science, 2485:159–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.de/link/service/
series/0558/bibs/2485/24850159.
htm; http://link.springer.
de/link/service/series/0558/
papers/2485/24850159.pdf.

Samson:2001:STY

[SBR01] Judith Samson, Jason Byars, and Dallas Releford. Sams teach yourself Red Hat Linux in 24 hours. Sams teach yourself in 24 hours series. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-31845-8 (paperback). xv + 429 pp. LCCN QA76.76.O63 S354 2001.

Saidi:2009:EEP

[SBRM09] Ali G. Saidi, Nathan L.

Binkert, Steven K. Reinhardt, and Trevor Mudge. End-to-end performance forecasting: finding bottlenecks before they happen. ACM SIGARCH Computer Architecture News, 37 (3):361–370, June 2009. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Sobell:1998:HLF

[SC98] Mark G. Sobell and Caldera, Inc. Hands-on Linux: featuring Caldera Open Linux Lite, Netscape Navigator Gold, and Netscape FastTrack Server on two CDs. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-32569-1. 1216 pp. LCCN QA76.76.O63S5939 1998. US$51.95. Includes CD-ROM.

Salah:2013:ASW

[SCB^{+}13] Khaled Salah, Jose M. Alcaraz Calero, Jorge Bernal Bernabé, Juan M. Marín Perez, and Sherali Zeadally. Analyzing the security of Windows 7 and Linux for cloud computing. Computers & Security, 34(??):113–122, May 2013. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0167404812001800.

Schweitzer:2004:BRL

[Sch04] Martin Schweitzer. Book review: Linux in a Nutshell, 4th ed., by Ellen Siever, Stephen Figgins, and Aaron Weber, O’Reilly Books, 2003, $39.95, ISBN: 0-596-00482-6. ACM Queue: Tomorrow’s Computing Today, 1(10):88, February 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Schroder:2005:LC

[Sch05] Carla Schroder. Linux cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00640-3 (paperback). xxiii + 553 pp. LCCN QA76.76.O63; QA76.76.O63 S377 2004eb; QA76.76.O63 S377 2004. URL http:/

/www.oreilly.com/catalog/
9780596006402.

Schroder:2007:LNC

[Sch07] Carla Schroder. Linux networking cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-10248-8 (paperback). xxiii + 612 pp. LCCN TK5105.5 .S384 2007.

Smith:2005:DL

[SD05a] Roderick Smith and Jeff Duntemann. Degunking Linux.

Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-933097-04-3. xx + 332 pp. LCCN QA76.76.O63 S58 2005.

URL http://www.oreilly.
com/catalog/9781933097046.

Smith:2005:DLH

[SD05b] Roderick Smith and Jeff Duntemann. Degunking Linux. Paraglyph Press, Scottsdale, AZ, USA, 2005. ISBN ???? xx + 332 pp. LCCN QA76.76.O63 S65 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://site.ebrary.com/
lib/ucsc/Doc?id=10080003.

Strobel:1996:LPW

[SE96] Stefan Strobel and Volker Elling. Linux — vom PC zur Workstation: Grundlagen, Installation und praktischer Einsatz. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60557-6, 3-540-58097-2. xiv + 521 pp. LCCN ???? Also available in English [SE97] and Hungarian [SU96b] translations.

Strobel:1997:LUW

[SE97] Stefan Strobel and Volker Elling. Linux, unleashing the workstation in your PC. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94880-5. xv + 587 pp. LCCN QA76.5 .S78513 1997. US$30. Foreword by Jurgen Gulbins. Translated by Robert Bach and Aileen Darling from the German edition [SE96]. See critical review in [Ros98].

Sery:1998:LNT

[Ser98] Paul G. Sery. Linux Network toolkit. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-3146-8. xi + 596 pp. LCCN QA76.76.O63S468 1998.

Sery:2000:RLN

[Ser00] Paul G. Sery. Red Hat Linux Network Toolkit with CD. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 2000. ISBN 0-7645-4656-2. xxviii + 439 pp. LCCN QA76.76.O63 S4693 2000.

Shabtai:2010:SAP

[SFE10] Asaf Shabtai, Yuval Fledel, and Yuval Elovici. Securing Android-powered mobile devices using SELinux. IEEE Security & Privacy, 8(3):36–44, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Siever:2003:LN

[SFW03] Ellen Siever, Stephen Figgins, and Aaron Weber. Linux in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2003. ISBN 0-596-00482-6. xiv + 928 pp. LCCN QA76.76.O63 S5582 2003.

Serral:2004:LNS

[SG04] René Serral and Marisa Gil. A Linux networking study. Operating Systems Review, 38(3): 1–11, July 2004. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Solsona:2001:PCI

[SGHL01] Francesc Solsona, Francesc Giné, Porfidio Hernández, and Emilio Luque. Predictive coscheduling implementation in a non-dedicated Linux cluster. Lecture Notes in Computer Science, 2150:732–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2150/21500732.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2150/21500732.
pdf.

Solsona:2000:MCM

[SGL^{+}00] Francesc Solsona, Francesc Giné, Josep Lérida, Porfidio Hernández, and Emilio Luque. Monito: A communication monitoring tool for a PVM–Linux environment. Lecture Notes in Computer Science, 1908:233–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.

                                        com/link/service/series/
                                        0558/bibs/1908/19080233.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/1908/19080233.
                                        pdf.

S:2023:DMC

[SGP^{+}23] Sajin S, Shubham Sunil Garag, Anuj Phegade, Deepshikha Gusain, and Kuruvilla Varghese. Design of a multi-core compatible Linux bootable 64bit out-of-order RISC-V processor core. In IEEE, editor, 2023 36th International Conference on VLSI Design and 2023 22nd International Conference on Embedded Systems (VLSID), pages 42–47. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2023.

Shalaby:2002:SSN

[SGWP02] Nadia Shalaby, Yitzchak Gottlieb, Mike Wawrzoniak, and Larry Peterson. Snow on silk: A NodeOS in the Linux kernel. Lecture Notes in Computer Science, 2546:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/
                                        service/series/0558/bibs/
                                        2546/25460001.htm; http:
                                        //link.springer.de/link/
                                        service/series/0558/papers/
                                        2546/25460001.pdf.

Sery:1999:RLS

[SH99a] Paul G. Sery and Eric Harper.

Red Hat Linux in Small Business. I D G Books Worldwide, Indianapolis, IN, USA, 1999. ISBN 0-7645-3335-

5. xxv + 392 pp. LCCN QA76.76.O63 S469 1999. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
bios/wiley044/99038069.html;
 http://www.loc.gov/catdir/
description/wiley038/99038069.
html;   http://www.loc.gov/
catdir/toc/wiley031/99038069.
html.

Siever:1999:LND

[SH99b] Ellen Siever and Jessica Perry Hekman. Linux in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1999. ISBN 1-56592-585-8. xiv + 612 pp. LCCN QA76.76.O63 .S4 1999. Revised edition of: Linux in a nutshell, by Jessica Perry Hekman. 1997.

Salah:2013:PIF

[SH13] K. Salah and M. Hamawi. Performance of IP-forwarding of Linux hosts with multiple network interfaces. Journal of Network and Computer Applications, 36(1):452–465, January 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S1084804512001063.

Shah:2000:LAB

[Sha00] Steve Shah. Linux administration: a beginner’s guide. Network professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212229-3. xxx + 542 pp. LCCN QA76.76.O63 S523 2000.

Shah:2001:LAB

[Sha01] Steve Shah. Linux administration: a beginner’s guide. Network professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 2001. ISBN 0-07-213136-5. xxvi + 843 pp. LCCN QA76.76.O63 S524 2001. Includes CD-ROM.

Simon:2023:CAH

[SHB23] Marek Simon, Ladislav Huraj, and Nicolas Búcik. A comparative analysis of high availability for Linux container infrastructures. Future Internet, 15 (8):253, July 28, 2023. CO-DEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.com/
                                        1999-5903/15/8/253.

Sheer:2002:LRU

[She02] Paul Sheer. Linux: rute users tutorial and exposition. Prentice Hall PTR open source technology series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-033351-4. xxxi + 630 pp. LCCN QA76.76.O63 S5527 2002. URL http://

                                        vig.prenhall.com/catalog/

academic/product/1,4096,
0130333514,00.html. Includes

CD-ROM with a complete HTML version of the book.

Sherer:2003:RTS

[She03] Matt Sherer. Real-time signal analysis and Real-Time Linux: Part 1. Dr. Dobb’s Journal of Software Tools, 28 (7):62–65, July 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2003/2003_07/
rtlp1.txt.

Shermerhorn:2006:HOL

[She06] Lee Shermerhorn. HP/OSLO Linux scalability tracking and investigations. In Anonymous [Ano06], page ?? ISBN ???? LCCN ???? URL http://

www.ice.gelato.org/oct06/
pres_pdf/gelato_ICE06oct_
scaltracking_shermerhorn_
hp.pdf.

Shotts:2012:LCL

[Sho12] William E. Shotts, Jr. The Linux command line: a complete introduction. No Starch Press, San Francisco, CA, USA, 2012. ISBN 1-59327-389-4 (paperback). xxix + 446 pp. LCCN QA76.76.O63 S5556 2012.

Shumba:2006:THL

[Shu06] Rose Shumba. Teaching handson Linux host computer security. ACM Journal on Educational Resources in Computing (JERIC), 6(3):5:1–5:??, September 2006. CODEN ???? ISSN 1531-4278.

Sierer:2000:WPC

[Sie00] Brian Sierer. Weighting the pros and cons of the Linux operating system. Research & Development, 42(10):75, October 2000. CODEN REDEEA. ISSN 0746-9179.

Siever:2005:LN

[Sie05] Ellen Siever. Linux in a nutshell. In a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2005. ISBN 0-596-00930-5. xiv + 928 pp. LCCN QA76.76.O63; QA76.76.O63 H453 2005eb; QA76.76.O63 H453 2005; QA76.76.O63 L5459 2003. URL http:/

/www.oreilly.com/catalog/
9780596009304.

Seo:2009:CTR

[SJP^{+}09] Euiseong Seo, Jinkyu Jeong, Seonyeong Park, Jinsoo Kim, and Joonwoon Lee. Catching two rabbits: adaptive real-time support for embedded Linux. Software—Practice and Experience, 39(5):531–550, April 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sery:2001:SLS

[SK01] Paul G. Sery and Mohammed J. Kabi. The SuSE Linux server.

M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2001. ISBN 0-7645-4765-8. xxviii + 609 pp. LCCN QA76.9.C55 S39 2000.

Shankar:2004:COS

[SK04] K. S. Shankar and Helmut Kurth. Certifying open source—the Linux experience. IEEE Security & Privacy, 2(6):28–33, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6028.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6028.
pdf.

Salah:2010:PEC

[SK10] K. Salah and A. Kahtani.

Performance evaluation comparison of Snort NIDS under Linux and Windows Server. Journal of Network and Computer Applications, 33(1):6–15, January 2010. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1084804509001040.

Seong:1996:PSM

[SKP96] Yeong Rak Seong, Tag Gon Kim, and Kyu Ho Park. Packing scheme for mean-filtering of an 8 bit image. Electronics Letters, 32(1):29–30, January 1996. CODEN ELLEAK.

ISSN 0013-5194 (print), 1350-911X (electronic).

Sung:2002:CPE

[SKP^{+}02] Minyoung Sung, Soyoung Kim, Sangsoo Park, Naehyuck Chang, and Heonshik Shin. Comparative performance evaluation of Java threads for embedded applications: Linux Thread vs. Green Thread. Information Processing Letters, 84(4):221–225, November 30, 2002. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sloan:2005:HPL

[Slo05] Joseph D. (Joseph Donald) Sloan. High performance Linux clusters with OSCAR, Rocks, openMosix, and MPI. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-00570-9. xv + 350 pp. LCCN QA76.58; QA76.58 .S56 2005eb; QA76.58 .S56 2005; QA76.58 .S58 2005; QA76.58 .S595 2005. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596005702.

Sidiroglou:2009:AAS

[SLP^{+}09] Stelios Sidiroglou, Oren Laadan, Carlos Perez, Nicolas Viennot, Jason Nieh, and Angelos D. Keromytis. ASSURE: automatic software self-healing using rescue points. ACM SIG-PLAN Notices, 44(3):37–48, March 2009. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Seshadri:2007:STH

[SLQP07] Arvind Seshadri, Mark Luk, Ning Qu, and Adrian Perrig. SecVisor: a tiny hypervisor to provide lifetime kernel code integrity for commodity OSes. Operating Systems Review, 41 (6):335–350, December 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Smith:1999:UCW

[Smi99]  Roderick W. Smith.  Us-

ing Corel WordPerfect 8 for Linux. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-7897-2032-9. xiii + 841 pp. LCCN Z52.5.W65 S6 1999. Edited by Jim Minatel.

Smith:2000:LHH

[Smi00a] Roderick W. Smith. Linux Hardware Handbook. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31918-7. xix + 695 pp. LCCN QA76.76.O63 S58846 2000.

Smith:2000:LNY

[Smi00b] Roderick W. Smith. Linux:

networking for your office. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31792-3. xv + 618 pp. LCCN QA76.76.O63 S5885 2000.

Smith:2001:LSS

[Smi01] Roderick W. Smith. Linux Samba server administration. The Craig Hunt Linux library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2740-1 (paperback). xxiii + 629 pp. LCCN QA76.76.O63 S58847 2001.

Smith:2002:ALN

[Smi02] Roderick W. Smith. Advanced Linux networking. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-77423-2. xviii + 752 pp. LCCN QA76.76.O63 S58845 2002.

Smith:2003:LPT

[Smi03] Roderick W. Smith. Linux power tools. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2003. ISBN 0-7821-4226-5 (paperback). xxiv + 752 pp. LCCN QA76.76.O63 S5886

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0615/2003101647-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0615/2003101647-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0615/2003101647-t.html.

Smith:2005:LNS

[Smi05a] Peter G. Smith. Linux network security. Administrator’s advantage series. Charles River Media, Hingham, MA, USA, 2005. ISBN 1-58450-396-3 (pbk. with CD-ROM). xviii + 541 pp. LCCN TK5105.59 .S59 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip056/2005000312.html.

Smith:2005:LWW

[Smi05b] Roderick W. Smith. Linux in a Windows world. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00758-2. xiv + 478 pp. LCCN QA76.76.O63 S548 2005.

Smith:2012:LE

[Smi12] Roderick W. Smith. Linux essentials. Wiley, New York, NY, USA, 2012. ISBN 1-118-10679-2 (paperback). xxiii + 344 pp. LCCN QA76.76.O63 S5884584

2012. URL http://proquest.
safaribooksonline.com/9781118106792.

Strobel:1997:LUI

[SMM97] Stefan Strobel, Rainer Maurer, and Stefan Middendorf. Linux universe: installation and configuration. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1997. ISBN 0-387-94879-1. viii + 228 pp. LCCN QA76.76.O63 S766 1997. Two CD ROMs. Translation by Robert Bach.

Strong:2009:FST

[SMM^{+}09] Richard Strong, Jayaram Mudigonda, Jeffrey C. Mogul, Nathan Binkert, and Dean Tullsen. Fast switching of threads between cores. Operating Systems Review, 43(2):35– 45, April 2009. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shieh:2008:SAC

[SMS08] Alan Shieh, Andrew C. Myers, and Emin Gün Sirer. A stateless approach to connectionoriented protocols. ACM Transactions on Computer Systems, 26(3):8:1–8:50, September 2008. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Sitsky:1996:IMU

[SMTW96] D. Sitsky, P. Mackerras, A. Tridgell, and D. Walsh. Implementing MPI under AP/ Linux. In Proceedings. Second MPI Developer’s Conference, pages ??–?? (of ix + 207). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7533-0. LCCN QA76.642 .M67 1996.

Sohoni:2001:SMS

[SMXH01] Sohum Sohoni, Rui Min, Zhiyong Xu, and Yiming Hu. A study of memory system performance of multimedia applications. ACM SIGMET-RICS Performance Evaluation Review, 29(1):206–215, June 2001. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Stankovic:2004:MWG

[SNF04] Jan Stankovic, Markus Neteler, and Roberto Flor. Mobile wireless GRASS GIS for handheld computers running GNU/ Linux. Transactions in GIS, 8 (2):225–233, 2004. ISSN 1361-1682 (print), 1467-9671 (electronic).

Saarinen:2022:DRV

[SNM22] Markku-Juhani O. Saarinen, G. Richard Newell, and Ben Marshall. Development of the RISC-V entropy source interface. Journal of Cryptographic Engineering, 12(4): 371–386, November 2022. CO-DEN ???? ISSN 2190-8508 (print), 2190-8516 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s13389-021-00275-6.

Sobell:1997:PGL

[Sob97] Mark G. Sobell. A Practical Guide to Linux. Addison-Wes-ley, Reading, MA, USA, 1997. ISBN 0-201-89549-8. 1216 pp. LCCN QA76.76.O63S5948 1997. US$37.61.

Sobell:2003:PGR

[Sob03] Mark G. Sobell. A Practical Guide to Red Hat Linux 8. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-70313-0. xlvii + 1565 pp. LCCN QA76.76.O63 S59485 2002. Includes two CD-ROMs.

Sobell:2006:PGR

[Sob06] Mark G. Sobell. A practical guide to Red Hat Linux: Fedora Core and Red Hat Enterprise Linux. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2006. ISBN 0-13-228027-2 (paperback). 1168 (est.) pp. LCCN QA76.76.O63 S59485 2006. URL http://

                                        vig.prenhall.com/catalog/
                                        academic/product/0,1144,
                                        0132280272,00.html; http:
                                        //www.loc.gov/catdir/toc/
                                        ecip0613/2006014003.html.

Includes DVD.

Sorfa:2001:OSI

[Sor01] Petr Sorfa. Open source IDEs for Linux/Unix. C/C++ Users Journal, 19(3):8–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Santos:2015:BD

[SP15] Rui Santos and Luis Perestrelo. BeagleBone for dummies. Wiley, New York, NY, USA, 2015. ISBN 1-118-99291-1 (paperback). ???? pp. LCCN ????

Spector:2000:MBC

[Spe00a] David Spector. Managing Beowulf clusters. Journal of Linux Technology, 1(1):18–??, 2000. ISSN 1527-2761.

Spector:2000:BLC

[Spe00b] David (David H. M.) Spector. Building Linux clusters: scaling Linux for scientific and enterprise applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-625-0. xviii + 332 pp. LCCN QA76.76.O63 S6647 2000. Includes CD-ROM with Red Hat Linux clustering software, cluster management tools and scripts, and parallel programming tools.

Soltesz:2007:CBO

[SPF^{+}07] Stephen Soltesz, Herbert Pötzl, Marc E. Fiuczynski, Andy Bavier, and Larry Peterson. Container-based operating system virtualization: a scalable, high-performance alternative to hypervisors. Operating Systems Review, 41(3):275– 287, June 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sprang:2006:XVL

[Spr06] Henning Sprang. Xen: Virtualisierung unter Linux. Open Source Press, München, Germany, 2006. ISBN 3-937514-29-5. 350 pp. LCCN ???? ca. EUR 39.90, EUR 41.35 (AT).

Sprang:2007:XVL

[Spr07] Henning Sprang, editor. Xen:

Virtualisierung unter Linux. (German) [Xen: Virtualization under Linux]. Open Source Press, München, Germany, 2007. ISBN 3-937514-29-5. 350 pp. LCCN ???? URL http:

                                        //deposit.ddb.de/cgi-bin/
                                        dokserv?id=2809360.

Satrapa:1996:LIS

[SR96] Pavel Satrapa and Jiří A. Randus. Linux: Internet server. Neokortex, Praha, Czech Republic, 1996. ISBN 80-902230-0-1 (broz.). 413 pp. LCCN ????

Smart:1999:UCO

[SRBB99] Allan Smart, Erik Ratcliffe, Tim Bird, and David Bandel. Using Caldera OpenLinux, Special Edition. Que Corporation, Indianapolis, IN, USA, 1999. ISBN 0-7897-2058-2. xxii + 1208 pp. LCCN A76.76.O63 U7118 1999. US$39.99.

Santana:2015:FSS

[SRTH15] Ricardo Santana, Raju Rangaswami, Vasily Tarasov, and Dean Hildebrand. A fast and slippery slope for file systems. Operating Systems Review, 49 (2):27–34, December 2015. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Stanfield:2001:LSA

[SS01] Vicki Stanfield and Roderick W. Smith. Linux system administration. The Craig Hunt Linux library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2735-5 (paperback). xxviii + 657 pp. LCCN QA76.76.O63 S7346 2001.

Stanfield:2002:LSA

[SS02] Vicki Stanfield and Roderick W. Smith. Linux system administration. Craig Hunt Linux library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 2002. ISBN 0-7821-4138-2. xxvii + 624 pp. LCCN QA76.76.O63 S7346

2002. URL http://www.loc.
gov/catdir/enhancements/
fy0613/2002106413-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0613/2002106413-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0613/2002106413-t.html.

Schoblick:2005:DGL

[SS05a] Gabriele Schoblick and Robert Schoblick. Debian GNU/Linux. bhv, Bonn, Germany, 2005. ISBN 3-8266-8151-7. 767 pp. LCCN ???? EUR22.95. Includes two CD-ROMs.

Shih:2005:ICA

[SS05b] Timothy K. Shih and Yoshitaka Shibata, editors. 19th International Conference on Advanced Information Networking and Applications: proceedings, AINA 2005, 28–30 March, 2005, Taipei, Taiwan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2249-1 (paperback). ISSN 1550-445X. LCCN TK5105.5 .I5616 2005.

URL  http://ieeexplore.
ieee.org/servlet/opac?punumber=

9746. IEEE Computer Society Order Number P2249.

Sewell:2010:XTR

[SSO^{+}10] Peter Sewell, Susmit Sarkar, Scott Owens, Francesco Zappa Nardelli, and Magnus O. Myreen. x86-TSO: a rigorous and usable programmer’s model for x86 multiprocessors. Communications of the ACM, 53(7):89–97, July 2010. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Souza:1997:EPH

[SSSS97] P. S. Souza, L. J. Senger, M. J. Santana, and R. C. Santana. Evaluating personal high performance computing with PVM on Windows and LINUX environments. Lecture Notes in Computer Science, 1332:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimizu:2003:TLS

[ST03] Naohiko Shimizu and Ken Takatori. A transparent Linux super page kernel for Alpha, Sparc64 and IA32: reducing TLB misses of applications. ACM SIGARCH Computer Architecture News, 31 (1):75–84, March 2003. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Stanik:2004:NNW

[Sta04] John Stanik. News 2.0: New Web provides battery-powered option; there’s nothing like a map; shall I compare thee to a big squeeze?; Linux support: A black hole? ACM Queue: Tomorrow’s Computing Today, 2(8):14, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NLO

[Sta05a] John Stanik. News 2.0: Losing our edge? the real cost of Linux; say no to crackberries. ACM Queue: Tomorrow’s Computing Today, 3 (5):14, June 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NRW

[Sta05b]  John Stanik.  News 2.0:

“ransom-ware” on the loose; adopt IPv6 or die; IBM steps up on Linux training. ACM Queue: Tomorrow’s Computing Today, 3(6):10, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NML

[Sta06] John Stanik. News 2.0: The mobile Linux challenge; Google not into googling; keeping online video legit. ACM Queue: Tomorrow’s Computing Today, 4(8):8, October 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:NGD

[Sta07a] John Stanik. News 2.0: Googlebombs away; DST: This year’s Y2K?; Linux through Windows. ACM Queue: Tomorrow’s Computing Today, 5 (2):7, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

State:2007:REL

[Sta07b] Radu State. Review of ‘Embedded Linux Primer: A Practical Real-world Approach’ by Christopher Hallinan, Prentice Hall PTR, 2006, $44.99, ISBN 0-13-167984-8. ACM Queue: Tomorrow’s Computing Today, 5(3):44, April 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Hal07].

Stephenson:1999:BLW

[Ste99] Neal Stephenson. In the beginning … was the command line. Avon Books, New York, NY, USA, 1999. ISBN 0-380-81593-1. 151 pp. LCCN QA76.76.O63 S7369 1999. US$10.00.

Steil:2000:GOA

[Ste00a]  Michael Steil.  GNOME:

[objektorientiert arbeiten mit X; das GNU Network Object Model Environment optimal einrichten; der gekonnte Umgang mit der Oberfläche; viele Tips und Tricks; Programmieren mit dem GTK+]. Computer-&-Literatur-Verlag, Böblingen, Germany, 2000. ISBN 3-932311-60-4. 606 pp.

LCCN ???? Includes CD-ROM.

Stevens:2000:CPG

[Ste00b] Al Stevens. C programming:

Going undercover. Dr. Dobb’s Journal of Software Tools, 25 (7):113–117, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
cprog.txt.

Sterling:2001:BCCa

[Ste01a] Thomas L. Sterling, editor. Beowulf Cluster Computing with Linux. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-69274-0. xxxiii + 496 pp. LCCN QA76.58 .B46 2002. US$42.95, UK£28.50.

Stevens:2001:CPMa

[Ste01b] Al Stevens. C programming:

A moving target. Dr. Dobb’s Journal of Software Tools, 26 (4):129–130, 132–133, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/.

Stevens:2001:CPA

[Ste01c] Al Stevens. C programming: I almost get a Linux editor and compiler. Dr. Dobb’s Journal of Software Tools, 26(7):113–116, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Stevens:2001:CPG

[Ste01d] Al Stevens. C programming:

It’s good work when you can find it. Dr. Dobb’s Journal of Software Tools, 26(5):121–124, May 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Stevens:2001:CPR

[Ste01e]  Al Stevens.  C program-

ming: Road rage = editors + drivers. Dr. Dobb’s Journal of Software Tools, 26(6):131–134, June 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Stevens:2001:CPS

[Ste01f]  Al Stevens.  C program-

ming: Software development, Linux, and the White House. Dr. Dobb’s Journal of Software Tools, 26(3):124–127, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/.

Stevens:2001:CPYb

[Ste01g]  Al Stevens.  C program-

ming: YAPP: Yet Another Programming Platform. Dr. Dobb’s Journal of Software Tools, 26(10):105–107, 109, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Sterling:2002:BCC

[Ste02] Thomas Lawrence Sterling. Beowulf cluster computing with Linux. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-69274-0. xxxiii + 496 pp. LCCN QA76.58 .B46 2002.

Stones:2000:PLP

[Sto00] Richard Stones. Professional Linux programming. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-301-3. xviii + 1155 pp. LCCN QA76.76.O63 P754 2000.

Stoess:2007:TEU

[Sto07] Jan Stoess. Towards effective user-controlled scheduling for microkernel-based systems. Operating Systems Review, 41 (4):59–68, July 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Strobel:1994:LUJ

[Str94a]  Stefan Strobel.  LINUX:

Unix fur jedermann; der PC als Workstation; Installation Schritt fur Schritt, Unix-Grundlagen und TCP/IP, Editieren mit dem Emacs, DOS-Programme unter LINUX, Textverarbeitung mit LaTeX 2e, Sound, Grafik, Netzwerkspiele. Vogel-Verlag, Postfach 67 40, D-8700 Würzburg, Germany, 1994. ISBN 3-8259-1330-9. 100 pp. LCCN ???? 49.00 DM; 358.00 ÖS; 49.00 Sfr.

Strobel:1994:LUW

[Str94b] Stefan Strobel. Linux: Unleashing the Workstation in Your PC. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1994. ISBN 3-540-94880-5 (Berlin), 0-387-94880-5 (New York). 480 pp. LCCN QA76.5 .S78513 1994. US$29.95.

Stutz:2001:LCT

[Stu01] Michael Stutz. The Linux Cookbook: Tips and Techniques for Everyday Use. Linux Journal Press, San Francisco, CA, USA, 2001. ISBN 1-886411-48-4. 396 pp. LCCN QA76.76.O63 S788 2000. US$29.95. URL

http://dsl.org/cookbook/.

Prepared with TEXinfo.

Stutz:2004:LCT

[Stu04] Michael Stutz. The Linux cookbook: tips and techniques for everyday use. No Starch Press, San Francisco, CA, USA, second edition, 2004. ISBN 1-59327-031-3. xxxiv + 788 pp. LCCN QA76.76.O63 S788 2004eb; QA76.76.O63 S788 2004. URL http:/

/www.oreilly.com/catalog/
9781593270315.

Strobel:1994:LVP

[SU94a] S. Strobel and T. Uhl. LINUX — vom PC zur Workstation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-57383-6.

Strobel:1994:CLK

[SU94b] Stefan Strobel and Thomas Uhl. The complete Linux kit: Linux universe: installation and configuration and Linux: unleashing the workstation in your PC. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-14224-X (set), 0-387-94506-7, 3-540-58077-8, 0-387-58077-8. viii + 152 pp. LCCN QA76.76.O63 S766 1995.

Strobel:1994:LPW

[SU94c] Stefan Strobel and Thomas Uhl. LINUX — vom PC zur Workstation: Grundlagen, Installation und praktischer Einsatz. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58098-0, 3-540-57383-6. xii + 238 pp. LCCN ???? See [SU94d].

Strobel:1994:LWY

[SU94d] Stefan Strobel and Thomas Uhl. Linux: Unleashing the Workstation in Your PC. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58077-8 (Berlin), 0-387-58077-8 (New York). xii + 248 pp. LCCN QA76.5 .S78513 1994. DM49.00, US$29.00. Translation by Robert Bach of German title Linux—vom PC zur Workstation [SU94c]. See also [SU95a].

Strobel:1995:CLK

[SU95a] Stefan Strobel and Thomas Uhl. The Complete Linux Kit. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-14224-X (Berlin), 0-387-14224-X (New York). US$59.95. Package of one CD ROM plus two books, [SU94d, SUM95].

Strobel:1995:LVP

[SU95b] Stefan Strobel and Thomas Uhl. Linux vom PC zur Workstation: Grundlagen, Installation und praktischer Einsatz. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2., uberarb. und erw. aufl. edition, 1995. ISBN 3-540-58097-2, 3-540-59273-3, 3-540-14512-5 (CD-ROM). xiv + 521 pp. LCCN ????

Strobel:1996:CLK

[SU96a] Stefan Strobel and Thomas Uhl. The complete Linux kit: Linux universe and Linux: unleashing the workstation in your PC. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second, release 1.3.x. edition, 1996. ISBN 0-387-14227-4. 156 + xiv + 446 pp. LCCN ????

Strobel:1996:L

[SU96b] Stefan Strobel and Thomas Uhl. Linux. Kossuth Konyvk, Budapest, Hungary, 1996. ISBN 963-09-3896-0. 366 pp. LCCN ???? Hungarian translation by K. Papp Laszlone and Tarjan Gyorgy of [SE96]. See also [SU95a].

Strobel:1996:LWY

[SU96c] Stefan Strobel and Thomas Uhl. Linux — unleashing the workstation in your PC. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second, revised and enhanced edition, 1996. ISBN 0-387-94601-2. xiv + 446 pp. LCCN QA76.5.S78513 1996.

Strobel:1996:LUW

[SU96d] Stefan Strobel and Thomas Uhl. Linux — vom PC zur Workstation (English: Linux: unleashing the workstation in your PC). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second, rev. and enhanced. edition, 1996. ISBN 0-387-94601-2 (paperback). xiv + 446 pp. LCCN QA76.5.S78513 1996.

Strobel:1997:CLK

[SU97] Stefan Strobel and Thomas Uhl. The complete Linux kit. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1997. ISBN 0-387-14237-1. 770 pp. LCCN ???? US$59.95.

Sullivan:2002:SLG

[Sul02] Cary Sullivan, editor. Sair Linux and GNU certification: level II core concepts and practices. Wiley, New York, NY, USA, 2002. ISBN 0-471-40538-8. xvi + 399 pp. LCCN QA76.3 .S253 2002. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/wiley022/2002265408.
html.

Strobel:1995:LUI

[SUM95] Stefan Strobel, Thomas Uhl, and Rainer Maurer. Linux universe: installation and configuration. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94506-7. viii + 152 pp. LCCN QA76.76.O63 S766 1995. US$34.95. Translation by Robert Bach. Includes CD ROM. See also [SU95a].

Strobel:1996:LUI

[SUM96] Stefan Strobel, Thomas Uhl, and Rainer Maurer. Linux universe: installation and configuration. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1996. ISBN 0-387-94600-4 (paperback). viii + 156 pp. LCCN QA76.76.O63S766 1996.

Sun:1999:LFS

[Sun99] Sun Microsystems. Lxrun free software: Run Linux binaries unmodified on Solaris. World-Wide Web document, 1999.

URL http://soldc.sun.com/
articles/lxrun/;   http:
//sun.com/linux/lxrun;
http://sun.com/software/
linux/lxrun/download.html;
 http://www.sunfreeware.
com; http://www.ugcs.caltech.
edu/~steven/lxrun/. “Lxrun

is a freely downloadable utility that executes Linux applications on the Solaris 7 operating environment on Intel platforms. Lxrun is a software layer that sits between Solaris and the Linux Intel binary executable and remaps system calls ‘on the fly’ allowing them to run unmodified on Solaris.” “Lxrun is an emulator for executing Intel Linux a.out and ELF binaries on other types of UNIX running on Intel x86. It was developed originally on and for SCO OpenServer and SCO UnixWare. Lxrun does system call remapping ‘on the fly.’ There isn’t a significant difference between the execution environment required by Linux and SCO binaries. The primary difference is the way in which system calls are handled. In Linux, an int 0x80 instruction is used, which jumps to the system-call-handling portion of the Linux kernel. On SCO systems, int 0x80 causes a SIGSEGV signal. Lxrun intercepts these signals and calls the SCO equivalent of the system call that the Linux program attempted. There is also some mapping of ioctls, various flags, return values, and error codes. The result is that the Linux binary runs–with the help of lxrun– on the host platform with a small (usually negligible) performance penalty.”.

SuSE:2001:SLP

[SuS01] SuSE. SuSE Linux 7.1 Professional: über 2.000 Programme auf CD & DVD; Kernel 2.4, GNU parted, Security, Networking, Development, XFree86 4.0.2, KDE 2.0.1 + Multimedia Tools & Games; [professionelles Linux-Betriebssystem für Inel- & AMD-PCs]. SuSE GmbH, Nürnberg, Germany, 2001. ISBN 3-934678-58-0. ???? pp. LCCN ????

Sutter:2002:ESF

[Sut02] Ed Sutter. Embedded systems firmware demystified. CMP Books, 6600 Silacci Way, Gilroy, CA 95020, USA, 2002. ISBN 1-57820-099-7. xii + 364 pp. LCCN TK7895.E42 S875 2002. Includes CD-ROM.

Stetter:2004:IEZ

[SvBC^{+}04] M. Stetter, J. von Buttlar, P. T. Chan, D. Decker, H. Elfering, P. M. Gioquindo, T. Hess, S. Koerner, A. Kohler, H. Lindner, K. Petri, and M. Zee. IBM eServer z990 improvements in firmware simulation. IBM Journal of Research and Development, 48(3/4):583–594, ???? 2004. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

                                        research.ibm.com/journal/
                                        rd/483/stetter.html;   http:
                                        //www.research.ibm.com/journal/
                                        rd/483/stetter.pdf.

Suzaki:1998:ICT

[SW98] K. Suzaki and D. Walsh. Implementing the combination of time sharing and space sharing on AP/Linux. Lecture Notes in Computer Science, 1459:83– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swaine:1998:SFR

[Swa98]  Michael Swaine.  Swaine’s

flames: The real power behind Linux. Dr. Dobb’s Journal of Software Tools, 23(12): 144, December 1998. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/;
http://www.ddj.com/ddj/
1998/1998_12/#master_top.

Swaine:2000:PPe

[Swa00a] Michael Swaine. Programming paradigms: Oh behave! Dr. Dobb’s Journal of Software Tools, 25(5):99–100, 102, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Swan:2000:TSG

[Swa00b]  Tom Swan.  Tom Swan’s

GNU C++ for Linux. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2153-8. xii + 831 pp. LCCN QA76.73.C153 S93 2000. Includes CD-ROM with Mandrake-Linux 6.0, X Class library for C, and the entire source code from the book.

Sweet:2001:KD

[Swe01] David Sweet. KDE 2.0 Development. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN ???? ???? pp. LCCN ????

Sweetman:2007:SMR

[Swe07] Dominic Sweetman. See MIPS Run. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2007. ISBN 0-12-088421-6. xix + 492 pp. LCCN QA76.9.A73 S88 2007.

Shi:2022:AFL

[SWF^{+}22] Heyuan Shi, Guyu Wang, Ying Fu, Chao Hu, Houbing Song, Jian Dong, Kun Tang, and Kai Liang. Abaci-finder: Linux kernel crash classification through stack trace similarity learning. Journal of Parallel and Distributed Computing, 168(??):70–79, October 2022. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0743731522001320.

Siever:2001:LWS

[SWL^{+}01] Ellen Siever, Matt Welsh, Ben Laurie, Randy Jay Yarger, Alligator Descartes, and Scott Guelich. Linux Web Server CD Bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00208-4. 812 (est.) pp. LCCN ???? US$79.95. URL http:/

/www.oreilly.com/catalog/
linuxwebcdbs.

Sonnenreich:2000:BLA

[SY00] Wes Sonnenreich and Tom Yates. Building Linux and OpenBSD firewalls. Wiley, New York, NY, USA, 2000. ISBN 0-471-35366-3 (paperback). xxi + 362 pp. LCCN QA76.9.A25 S66 1999. US$44.99. URL http:/

/www.wiley.com/compbooks/
sonnenreich.

Shen:2008:HCDa

[SZD^{+}08a] Kai Shen, Ming Zhong, Sandhya Dwarkadas, Chuanpeng Li, Christopher Stewart, and Xiao Zhang. Hardware counter driven on-the-fly request signatures. ACM SIGARCH Computer Architecture News, 36(1): 189–200, March 2008. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Shen:2008:HCDb

[SZD^{+}08b] Kai Shen, Ming Zhong, Sandhya Dwarkadas, Chuanpeng Li, Christopher Stewart, and Xiao Zhang. Hardware counter driven on-the-fly request signatures. Operating Systems Review, 42(2):189–200, March 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Shen:2008:HCDc

[SZD^{+}08c] Kai Shen, Ming Zhong, Sandhya Dwarkadas, Chuanpeng Li, Christopher Stewart, and Xiao Zhang. Hardware counter driven on-the-fly request signatures. ACM SIGPLAN Notices, 43(3):189–200, March 2008. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tackett:1997:UL

[Tac97] Jack Tackett. Using Linux. Que Corporation, Indianapolis, IN, USA, third special edition, 1997. ISBN 0-7897-1132-X (paperback). xxxvi + 774 pp. LCCN QA76.76.O63T28 1997. US$59.99.

Tanenbaum:1992:MOS

[Tan92] Andrew S. Tanenbaum. Modern Operating Systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-098310-1, 0-13-588187-0, 0-13-595752-4 (paperback). xx + 728 pp. LCCN QA76.76.O63 T359 1992.

Tanenbaum:1995:DOS

[Tan95] Andrew S. Tanenbaum. Distributed operating systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1995. ISBN 0-13-219908-4. xvii + 614 pp. LCCN QA76.76.O63 T357 1995.

Tansley:2000:LUS

[Tan00] David S. W. Tansley. Linux and UNIX Shell Programming. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-67472-6. xxiii + 504 pp. LCCN QA76.76.O63 T365 2000. US$44.95. URL http:

//cseng.aw.com/bookpage.
taf?ISBN=0-201-67472-6&ptype=
3017.

Tanenbaum:2001:MOS

[Tan01] Andrew S. Tanenbaum. Modern operating systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2001. ISBN 0-13-031358-0. xxiv + 951 pp. LCCN QA76.76.O63 T359 2001.

Tanenbaum:2008:MOS

[Tan08] Andrew S. Tanenbaum. Modern Operating Systems. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, third edition, 2008. ISBN 0-13-600663-9. xxvii + 1076 pp. LCCN QA76.76.O63 T359 2008.

Tam:2007:TCS

[TAS07] David Tam, Reza Azimi, and Michael Stumm. Thread clustering: sharing-aware scheduling on SMP–CMP–SMT multiprocessors. Operating Systems Review, 41(3):47–58, June 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tauber:1995:CLK

[Tau95] Daniel A. Tauber. The complete Linux kit. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1995. ISBN 0-7821-1669-8. xxiii + 419 pp. LCCN QA76.76.O63T384 1995.

Tauber:1996:LPP

[Tau96] Daniel A. Tauber. LINUX: par la pratique, kit complet. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 2-7361-1612-7. xvii + 308 pp. LCCN ???? 178FFR.

Taylor:1999:LCL

[Tay99a] Grant Taylor. Linux complete: Linux documentation project. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2567-0. xxxviii + 985 pp. LCCN QA76.76.O63 L54583 1999.

Taylor:1999:PBL

[Tay99b] Jeff Taylor. Programmer’s bookshelf: Linux sysadmin on your own. Dr. Dobb’s Journal of Software Tools, 24(7): 127–128, July 1999. CODEN DDJOEB. ISSN 1044-789X.

Taylor:2003:STY

[Tay03]  Dave Taylor.  Sams teach

yourself Unix system administration in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 2003. ISBN 0-672-32398-2. xiii + 508 pp. LCCN QA76.76.O63 T3885 2003. Covers: Unix/Linux/Solaris/Mac OS X.

Tackett:1998:UL

[TB98] Jack Tackett and Steven Forrest Burnett. Using Linux. Que Corporation, Indianapolis, IN, USA, fourth special edition, 1998. ISBN 0-7897-1746-8. xix + 870 pp. LCCN QA76.76.O63T28 1998.

Turk:2005:VLS

[TB05] D. Turk and J. Bausch. Virtual Linux servers under z/VM: Security, performance, and administration issues. IBM Systems Journal, 44(2):341–??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670.

Tackett:1998:L

[TBD98] Jack Tackett, Steven Forrest Burnett, and Raymond Debonne. Linux. Simon & Schuster Macmillan (France), Paris, France, 4e edition, 1998. ISBN 2-7440-0567-3. xxx + 865 pp. LCCN ????

Torrey:2007:CIL

[TCM07] Lisa A. Torrey, Joyce Coleman, and Barton P. Miller. A comparison of interactivity in the Linux 2.6 scheduler and an MLFQ scheduler. Software— Practice and Experience, 37(4): 347–364, April 10, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Templon:1996:EPL

[TD96] Jeffrey A. Templon and Paul F.

Dubois. Evaluation of PC/ Linux systems for use as scientific workstations. Computers in Physics, 10(1):49–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.4822356.

Tranter:1997:SIS

[TD97] Jeff Tranter and Eric Dumas. Son et images sous Linux. O’Reilly France, Paris, France, 1997. ISBN 2-84177-014-1. 380 pp. LCCN ????

Torvalds:2001:JFS

[TD01] Linus Torvalds and David Diamond. Just for Fun: the Story of an Accidental Revolutionary. HarperBusiness, New York, NY, USA, 2001. ISBN 0-06-662072-4 (hardcover), 0-06-662073-2, 1-58799-151-9. xxii + 262 pp. LCCN QA76.2.T67 T67 2001. URL http:

//catdir.loc.gov/catdir/
description/hc043/00054199.
html.

Ts:2003:USF

[TECB03] Jay Ts, Robert Eckstein, and David Collier-Brown. Using Samba: A File & Print Server for Linux, Unix & Mac OS X. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2003. ISBN 0-596-00256-4. xiii + 539 pp. LCCN QA76.9.C55 E38 2003. US$39.95 ; EUR 44.00; SFr 72.90. URL http:/

/www.oreilly.com/catalog/
samba2/.

Terrasa:1998:ERS

[TEGF98] A. Terrasa, A. Espinosa, and A. Garcia-Fornes. Extending RT-Linux to support flexible hard real-time systems with optional components. Lecture Notes in Computer Science, 1474:41–??, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Templon:1996:SPE

[Tem96] Jeffrey A. Templon. Scientific programming: Evaluation of PC/Linux systems for use as scientific workstations. Computers in Physics, 10(1):49–??, ???? 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Teodorovici:2013:BRU

[Teo13] Vasile G. Teodorovici. Book review: Ubuntu unleashed, 2013 edition by Matthew Helmke with Andrew Hudson and Paul Hudson. ACM SIGSOFT Software Engineering Notes, 38 (4):41, July 2013. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tody:1996:PCG

[TF96] D. Tody and M. Fitzpatrick.

PC-IRAF: the choice of a GNU generation. Astron. Soc. Pac.

Conf. Ser. (USA), Astronomical Society of the Pacific Conference Series, 101(????):322– 326, ???? 1996. ISSN 1050-3390.

Tsegaye:2004:CLW

[TF04] Melekam Tsegaye and Richard Foss. A comparison of the Linux and Windows device driver architectures. Operating Systems Review, 38(2):8– 33, April 2004. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tackett:1995:ML

[TG95] Jack Tackett and David Gunter. Le Macmillan: Linux. Simon and Schuster, New York, NY, USA, 3e edition, 1995. ISBN 2-7440-0267-4. xxxi + 812 pp. LCCN ????

Tackett:1996:SEU

[TG96] Jack Tackett, Jr. and David Gunter. Special edition using Linux. Que Corporation, Indianapolis, IN, USA, second edition, 1996. ISBN 0-7897-0742-X. xxxvii + 792 pp. LCCN QA76.76.O63T28 1996.

Tackett:1995:SEU

[TGB95a] Jack Tackett, Jr., David Gunter, and Lance Brown. Special Edition Using Linux. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0100-6. xxvi + 861 pp. LCCN QA76.76.O63T28 1995. US$49.95.

Tackett:1995:UL

[TGB95b] Jack Tackett, Jr., David Gunter, and Lance (Lance Allan) Brown. Using Linux. Que Corporation, Indianapolis, IN, USA, special edition, 1995. ISBN 0-7897-0100-6. xxvi + 861 pp. LCCN QA76.76.O63 T28 1995 Interactive Learning Center. US$49.99, CDN$67.99. CD-ROM includes the complete Linux 32-bit operating system, XFree 86 GUI and source code, plus a complete suite of Internet access tools – Cover.

Tackett:1996:L

[TGBO96] Jack Tackett, David Gunter, Lance (Lance Allan) Brown, and Alfredo Ocampo Rivera, tr. Linux. Prentice Hall HispanoAmericana, Mexico, DF, Mexico / Nueva York, NY, USA, edicion especial edition, 1996. ISBN 968-880-580-7. xxviii + 860 pp. LCCN ????

Tackett:1997:LC

[TGW97] Jack Tackett, David Gunter, and Hua Wan. Linux ta chuan. Tien tzu kung yeh chu pan she, Pei-ching, Peoples’ Republic of China, ti 1 pan edition, 1997. ISBN 7-5053-4814-0. x + 601 pp. LCCN ????

Thiruvathukal:2004:GLN

[Thi04] George K. Thiruvathukal. Gentoo Linux: The next generation of Linux. Computing in Science and Engineering, 6 (5):66–74, September/October 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/cs/2004/05/c5066.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/cs/2004/05/c5066.
                                        pdf.

Thomas:2005:RBC

[Tho05] Marlin Thomas. Review of Building Clustered Linux Systems by Robert W. Lucke, Prentice Hall PTR, 2004, $49.99, ISBN: 0-13-144853-6. ACM Queue: Tomorrow’s Computing Today, 3 (4):60, May 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Thomas:2006:BUL

[Tho06]  Keir Thomas.  Beginning

Ubuntu Linux: from novice to professional. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-627-7. xxxii + 573 pp. LCCN QA76.76.O63 T565 2006.

Throwe:2001:RCF

[Thr01] Tom Throwe. The RHIC computing facility Linux farms. ;login: the USENIX Association newsletter, 26(5): ??, August 2001. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL http://www.
                                        usenix.org/publications/
                                        login/2001-08/pdfs/throwe.
                                        pdf.

TRG:1999:FOS

[Tim99] Timpanogas Research Group, Inc. Free open source FEN-RIS for Linux will be available on June 4, 1999. World-Wide Web document, June

4, 1999. URL http://www.
timpanogas.com/. FENRIS

provides Novel NetWare file system access on other operating systems. The Linux version is available as an open source product; the Windows NT version is licensed.

Timpe:2003:GCJ

[Tim03] R. Timpe. GPIB control with Java on Linux or Windows. Evaluation Engineering, 42(6): 18–23, 2003. CODEN ???? ISSN 0014-3316.

Tsoi:2007:PAS

[TK07] Kelvin Yiu-Lun Tsoi and Yu-Kwong Kwok. Practical algorithms for scheduling video data in a local area network environment. The Journal of Supercomputing, 39(2):131– 147, February 2007. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=39&issue=2&spage=
131.

Taylor:2001:KPS

[TMG01] Don Taylor, Jim Mischel, and Tim Gentry. Kylix power solutions. Coriolis Group Books, Scottsdale, AZ, USA, 2001.

ISBN 1-932111-29-8. xviii + 331 pp. LCCN QA76.76.A65 T415 2001eb. URL http:/

                                        /www.oreilly.com/catalog/
                                        9781932111293.

Tomson:2002:DCR

[Tom02] Phil Tomson. Distributed computing with Ruby. Dr. Dobb’s Journal of Software Tools, 27 (9):16, 18, 22, September 2002. CODEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_09/
                                        taskmas.txt;   http://www.
                                        ddj.com/ftp/2002/2002_09/
                                        taskmas.zip.

Tommasino:2011:HGR

[Tom11] Damian Tommasino. Handson guide to the Red Hat exams: RHCSA and RHCE cert guide and lab manual. Pearson, Indianapolis, IN, USA, 2011. ISBN 0-321-76795-0. ???? pp. LCCN QA76.76.O63 T6494 2011.

Torvalds:1994:LKI

[Tor94] L. Torvalds. Linux kernel implementation. In Anonymous [Ano94f], pages 9–14. ISBN 0-646-20108-5. LCCN TK5105.58.A87 1994.

Torvalds:1999:LE

[Tor99] Linus Torvalds. The Linux edge. Communications of the ACM, 42(4):38–39, April 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/cacm/1999-42-4/
                                        p38-torvalds/.

Toth:1999:LNS

[Tot99] Viktor Toth. Linux: a network solution for your office. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31628-5. xxii + 512 pp. LCCN QA76.76.O63 T6778 1999. Edited by Grace Buechlein and Laura Bulcher.

Toxen:2007:SDS

[Tox07] Bob Toxen. The seven deadly sins of Linux security. ACM Queue: Tomorrow’s Computing Today, 5(4):38–47, May– June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Tozzi:2016:LT

[Toz16] Christopher Tozzi. Linux turns 25. IEEE Spectrum, 53(4): 48–56, April 2016. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Taber:1999:MLS

[TR99] Mark Taber and Randi Roger.

Maximum Linux security: a hacker’s guide to protecting your Linux server and network. Macmillan Computer Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31670-6. xvii + 743 pp. LCCN QA76.9.A25 M387 2000.

Tranter:1996:LMG

[Tra96] Jeff Tranter. Linux Multimedia Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, Fall 1996. ISBN 1-56592-219-0. xx + 363 pp. LCCN QA76.575.T73 1996. US$32.95.

Troan:1996:FSS

[Tro96a] Erik Troan. Free software solutions: Linux 2.0. The X Journal: Computing Technology with the X Window System, 6(2):72, 70, November and December 1996. CODEN XJOUEA. ISSN 1056-7003.

URL http://www.sigs.com/
publications/docs/txjr/9611/
txjr9611.toc.html.

Trocki:1996:PAT

[Tro96b] J. Trocki. PC administration tools: using Linux to manage personal computers. In USENIX [USE96a], pages 187– 192.

Thiruvathukal:2002:WPT

[TSC02] George K. (George Kuriakose) Thiruvathukal, John P. Shafaee, and Thomas W. Christopher. Web programming: techniques for integrating Python, Linux, Apache, and MySQL. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-041065-9. xviii + 745 pp. LCCN QA76.625 .T48 2002.

URL http://www.phptr.com/
ptrbooks/ptr_0130410659.
html.

Trinitis:2001:SBL

[TSEK01] C. Trinitis, M. Schulz, M. Eberl, and W. Karl. SCI-based LINUX PC-clusters as a platform for electromagnetic field calculations. Lecture Notes in Computer Science, 2127:510– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2127/21270510.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2127/21270510.
pdf.

T:2001:TOS

[TT01] Subramania Sharma T. and Matthew J. Thazhuthaveetil. TWLinuX: Operating system support for optimistic parallel discrete event simulation. Lecture Notes in Computer Science, 2228:262–271, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2228/22280262.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2228/22280262.
pdf.

Turnbull:2005:HL

[Tur05] James Turnbull. Hardening Linux. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-444-4 (paperback). xxvii + 552 pp. LCCN QA76.76.D49 T87

                                        2005. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0663/2005280402-d.html.

Tauber:1996:LMD

[TW96] Daniel A. Tauber and Matt Welsh. Linux: mode d’emploi. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 2-7361-2281-X. xvi + 436 pp. LCCN ????

Tan:2007:IBB

[TYKZ07] Lin Tan, Ding Yuan, Gopal Krishna, and Yuanyuan Zhou.

                                        /*icomment:  bugs or bad

comments?*/. Operating Systems Review, 41(6):145–158, December 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tyler:2006:FL

[Tyl06] Chris Tyler. Fedora Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-52682-2. xiii + 639 pp. LCCN QA76.76.O63; QA76.76.O63 T9 2006eb. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596526825.

Ugur:2022:OPF

[UJE^{+}22] Muhammed Ugur, Cheng Jiang, Alex Erf, Tanvir Ahmed Khan, and Baris Kasikci.

One profile fits all: Profileguided Linux kernel optimizations for data center applications. Operating Systems Review, 56(1):26–33, June 2022. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3544497.3544502.

Urgaonkar:2005:AMM

[UPS^{+}05] Bhuvan Urgaonkar, Giovanni Pacifici, Prashant Shenoy, Mike Spreitzer, and Asser Tantawi. An analytical model for multitier Internet services and its applications. ACM SIGMET-RICS Performance Evaluation Review, 33(1):291–302, June 2005. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Urgaonkar:2007:AMM

[UPS^{+}07] Bhuvan Urgaonkar, Giovanni Pacifici, Prashant Shenoy, Mike Spreitzer, and Asser Tantawi. Analytic modeling of multitier Internet applications. ACM Transactions on the Web (TWEB), 1(1):2:1–2:??, May 2007. CODEN ???? ISSN 1559-1131 (print), 1559-114X (electronic).

Hillebrand:1995:MLM

[uRH95] A. Sorgatz und R. Hillebrand.

Mathematik unter Linux: Mu-PAD — Ein Computeralgebra System I. (German) [Mathematics in Linux: MuPAD — A computer algebra system, I].

Linux Magazin, 12/95:11–14, 1995.

Hillebrand:1996:MLM

[uRH96] A. Sorgatz und R. Hillebrand. Mathematik unter Linux: Mu-PAD — Ein Computeralgebra System II. (German) [Mathematics in Linux: MuPAD — A computer algebra system, II]. Linux Magazin, 2/96, 3/96 (Nachdruck):60–67, 1996.

Urgaonkar:2008:ADP

[USC^{+}08] Bhuvan Urgaonkar, Prashant Shenoy, Abhishek Chandra, Pawan Goyal, and Timothy Wood. Agile dynamic provisioning of multi-tier Internet applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3 (1):1:1–1:??, March 2008. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

USENIX:1996:SAC

[USE96a] USENIX, editor. 10th Systems Administration Conference (LISA’96), September 29– October 4, 1996. Chicago, IL. USENIX, Berkeley, CA, USA, 1996.

USENIX:1996:PUA

[USE96b] USENIX Association, editor. Proceedings of the USENIX 1996 annual technical conference: January 22–26, 1996, San Diego, California, USA, USENIX Conference Proceedings 1996. USENIX, Berkeley, CA, USA, 1996. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

USENIX:2000:PAL

[USE00] USENIX, editor. Proceedings of the 4th Annual Linux Showcase and Conference, Atlanta, October 10–14, 2000, Atlanta, Georgia, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-17-0. LCCN ????

URL  http://www.usenix.
org/publications/library/
proceedings/als2000/.

USENIX:2001:PAL

[USE01]  USENIX, editor.  Proceed-

ings of the 5th Annual Linux Showcase and Conference, November 5–10, 2001, Oakland, CA. USENIX, Berkeley, CA, USA, 2001. ISBN

????  LCCN ????  URL

http://www.linuxshowcase.
org/tech.html.

Urgaonkar:2009:ROA

[USR09] Bhuvan Urgaonkar, Prashant Shenoy, and Timothy Roscoe. Resource overbooking and application profiling in a shared Internet hosting platform. ACM Transactions on Internet Technology (TOIT), 9(1):1:1– 1:??, February 2009. CODEN ???? ISSN 1533-5399 (print), 1557-6051 (electronic).

Vadala:2000:RLB

[Vad00] Derek Vadala. RAID on Linux:

Bringing increased reliability and disk throughput to an open source server. Journal of Linux Technology, 1(2):25–32, 2000. ISSN 1527-2761.

Vadala:2002:MRL

[Vad02] Derek Vadala. Managing RAID on Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 1-56592-730-3. 304 (est.) pp. LCCN ???? US$39.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        mraidlinux.

VanHensbergen:2006:PRP

[Van06] Eric Van Hensbergen. P.R.O.S.E.: partitioned reliable operating system environment. Operating Systems Review, 40(2):12– 15, April 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Varhol:1994:TOS

[Var94] Peter D. Varhol. Trends in operating system design. Dr. Dobb’s Journal of Software Tools, 19(5):18–??, May 1994. CODEN DDJOEB. ISSN 1044-789X.

Vaamonde:2001:ISG

[VD01] Fernández Vaamonde and Manuel David. Implantación de un sistema de gestión centralizada de paquetes deb para su uso en configuraciones Debian GNU/Linux. (Spanish) [Implantation of a system of centralized management of deb packages for its use in configurations of Debian GNU/Linux ]. Technical report, Trabajos fin de carrera de la Facultade de Informática de A Coruña. Departamento de Electrónica e Sistemas, Facultade de Informática Universidade da Coruña, Coruña, Spain, 2001. 111 pp. Final engineering project. Includes one CD-ROM.

vanDoorn:2006:HVT

[vD06] Leendert van Doorn. Hardware virtualization trends. In ACM [ACM06], page 45. ISBN 1-59593-332-6 (??invalid ISBN??). LCCN QA76.9.V4.

Villeneuve:1999:LWN

[VE99] Arnold Villeneuve and Phillip Edwards. Linux and Windows NT. UNIX tools. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 1999. ISBN 0-07-134983-9. 640 pp. LCCN ????

Venkateswaran:2008:ELD

[Ven08] Sreekrishnan Venkateswaran.

Essential Linux device drivers. Prentice Hall open source software development series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2008. ISBN 0-13-239655-6 (hardback). xxx + 714 pp. LCCN QA76.76.D49 V35 2008. URL http:

//www.loc.gov/catdir/toc/
ecip087/2008000249.html.

VanHensbergen:2008:HAR

[VFMM08] Eric Van Hensbergen, Charles Forsyth, Jim McKie, and Ron Minnich. Holistic aggregate resource environment. Operating Systems Review, 42(1):85– 91, January 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

vonHagen:2002:LF

[vH02] William von Hagen. Linux Filesystems. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32272-2. xii + 555 pp. LCCN QA76.76.O63 V66 200.

vonHagen:2006:ULB

[vH06] William von Hagen. Ubuntu Linux bible. Wiley, New York, NY, USA, 2006. ISBN 0-470-03899-3 (paperback). xxvi + 904 pp. LCCN QA76.76.O63 V665 2007.

vonHagen:2007:ULB

[vH07] William von Hagen. Ubuntu Linux bible. Wiley, New York, NY, USA, 2007. ISBN 0-470-03899-3, 0-470-12454-7 (electronic book). xxvi + 904 pp. LCCN QA76.76.O63 V685 2007eb.

VonHagen:2006:LSH

[VJ06] William Von Hagen and Brian K. (Brian Kenneth) Jones. Linux server hacks. Vol. 2. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10082-5. xx + 456 pp. LCCN QA76.76.O63; QA76.76.O63eb. URL http:/

/www.oreilly.com/catalog/
9780596100827.

Volkerding:1996:LP

[VJR96] Patrick Volkerding, Eric F.

Johnson, and Kevin Reichard. Linux programming. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1996. ISBN 1-55828-507-5. ???? pp. LCCN QA76.76.O63 V65 1996. US$34.95. Includes CD-ROM.

Venton:2005:ULH

[VMKB05] T. Venton, M. Miller, R. Kalla, and A. Blanchard. Using Linux for hardware bring up, development, and manufacturing. IBM Systems Journal, 44 (2):319–??, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/venton.pdf.

Volkerding:1997:SL

[Vol97] Patrick Volkerding. Slackware Linux, 1997. ISBN 1-57176-151-9. Includes four CD-ROMs and installation pamphlet.

Volkerding:2000:LSC

[VR00] Patrick Volkerding and Kevin Reichard. Linux system commands. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2000. ISBN 0-7645-4669-4. xiv + 462 pp. LCCN QA76.76.O63 V463

Vrenios:2003:LCA

[Vre03] Alex Vrenios. Linux cluster architecture. SAMS Publishing, Indianapolis, IN, USA, 2003. ISBN 0-672-32368-0. x + 247 pp. LCCN QA76.58 .V74 2002.

Volkerding:1997:LPE

[VRFJ97a] Patrick Volkerding, Kevin Reichard, and Eric Foster-Johnson. Linux in Plain English. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-542-3. iv + 571 pp. LCCN QA76.76.O63V645 1997. US$14.95, CDN$20.95.

URL http://www.mispress.
com/.

Volkerding:1997:LIS

[VRFJ97b] Patrick Volkerding, Kevin Reichard, and Eric Foster-Johnson. The Linux Internet Server. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-545-8. xxi + 530 pp. LCCN QA76.9.C55R444 1997. US$39.95, CDN$55.95. URL

http://www.mispress.com/.

Includes CD-ROM.

Volkerding:1997:LP

[VRFJ97c] Patrick Volkerding, Kevin Reichard, and Eric Foster-Johnson. Linux Programming. The MIS:Press slackware series. MIS Press, P.

O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-507-5. xxi + 374 pp. LCCN QA76.76.O63V65 1997. US$39.95, CDN$55.95. URL

http://www.mispress.com/.

Includes CD-ROM.

Volkerding:1998:LCI

[VRFJ98] Patrick Volkerding, Kevin Reichard, and Eric Foster-Johnson. LINUX configuration and installation. M and T Books slackware series. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, fourth edition, 1998. ISBN 0-7645-7005-6. xxxii + 554 pp. LCCN QA76.76.O63 V64 1998.

Volkerding:1995:LCI

[VRJ95] Patrick Volkerding, Kevin Reichard, and Eric F. Johnson. Linux configuration and installation. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1995. ISBN 1-55828-426-5. xv + 463 pp. LCCN QA76.76.O63 V64 1995.

Volkerding:1996:LCI

[VRJ96] Patrick Volkerding, Kevin Reichard, and Eric F. Johnson. Linux configuration and installation. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1996. ISBN 1-55828-492-3. ???? pp. LCCN QA76.76.O63 V64 1996.

US$39.95. URL http://www.

                                        mispress.com/. Includes two

CD-ROMs.

Volkerding:1997:LCI

[VRJ97] Patrick Volkerding, Kevin Reichard, and Eric F. Johnson. Linux Configuration and Installation. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, third edition, 1997. ISBN 1-55828-566-0. xxix + 512 pp. LCCN QA76.76.O63V64 1997. US$39.95, CDN$55.95.

                                        URL  http://smartbooks.
                                        com/bw708linuxconfg.htm;
                                         http://www.mispress.com/.

Includes two CD-ROMs.

Vallee:2006:OTX

[VS06] Geoffroy Vallee and Stephen L. Scott. OSCAR testing with Xen. In IEEE [IEE06], pages 43–? ISBN 0-7695-2582-2. ISSN 1550-5243. LCCN QA76.88. ACM product number E2582.

Valente:2023:HVH

[VTS^{+}23] Luca Valente, Yvan Tortorella, Mattia Sinigaglia, Giuseppe Tagliavini, Alessandro Capotondi, Luca Benini, and Davide Rossi. HULK-V: a heterogeneous ultra-low-power Linux capable RISC-V SoC. In IEEE, editor, 2023 Design, Automation & Test in Europe Conference & Exhibition (DATE), pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2023.

Welsh:1995:LBG

[W^{+}95] Matt Welsh et al. The Linux bible: the GNU testament. Yggdrasil Computing, Inc., Berkeley, CA, USA, third edition, 1995. ISBN 1-883601-12-6. 1596 pp. LCCN QA 76.76 O63 L56 1995.

WalnutCreek:1995:SL

[Wal95] Walnut Creek CDROM. Slackware Linux, 1995. ISBN 1-57176-117-9. Two computer laser optical discs.

Wallich:1999:CVB

[Wal99] Paul Wallich. Cyber view: The best things in cyberspace are free. Scientific American, 280 (3):44, March 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Discusses the U.S. Justice Department lawsuit against Microsoft, and the Linux phenomenon.

Wall:2000:LPE

[Wal00] Kurt Wall. Linux Programming by Example. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2215-1. xvii + 533 pp. LCCN QA76.76.O63 W357 2000. US$24.99, CDN$37.95, UK£17.99.

Wall:2001:LPU

[Wal01a] Kurt Wall. Linux programming unleashed. SAMS Publishing, Indianapolis, IN, USA, second edition, 2001. ISBN 0-672-32021-5. ixi + 886 pp. LCCN QA76.76.O63 W3573 2001.

Walters:2001:EGC

[Wal01b] E. Garrison Walters. The essential guide to computing: the story of information technology. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-019469-7. xxv + 499 pp. LCCN QA76.6 .W335 2001.

Walton:2001:LSP

[Wal01c] Sean Walton. Linux socket programming. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-31935-7 (paperback). xv + 533 pp. LCCN QA76.76.O63 W358 2001.

Wang:2002:DIR

[Wan02] Yu-Chung Wang. Design and implementation of RED-Linux. Thesis (ph.d.), Electrical and Computer Engineering, University of California, Irvine, Irvine, CA, USA, 2002.

Wang:2011:ML

[Wan11]  Paul S. Wang.  Master-

ing Linux. Chapman and Hall/CRC, Boca Raton, FL, USA, 2011. ISBN 1-4398-0686-1 (paperback). xviii + 421 pp. LCCN QA76.76.O63 W365143 2011.

Ward:2000:LPS

[War00] Brian Ward. Linux Problem Solver. No Starch Press, San Francisco, CA, USA, 2000.

ISBN 1-886411-35-2. xiii + 283 pp. LCCN QA76.76.O63 W3655 2000. US$34.95.

Ward:2001:QPL

[War01] Patrick Ward. Qt programming for Linux and Windows 2000. Hewlett–Packard professional books. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-027001-6. xvii + 280 pp. LCCN QA76.9.U83 W37 2001.

Ward:2004:HLW

[War04] Brian Ward. How Linux works:

what every super-user should know. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-035-6. xvi + 347 + 4 pp. LCCN QA76.76.O63 W3654 2004eb; QA76.76.O63 W3654 2004. URL http:/

/www.oreilly.com/catalog/
9781593270353.

Ward:2015:HLW

[War15]  Brian Ward.  How Linux

works: what every superuser should know. No Starch Press, San Francisco, CA, USA, second edition, 2015. ISBN 1-59327-567-6 (paperback). xxii + 366 pp. LCCN QA76.76.O63 W3654

2015. URL http://proquest.
safaribooksonline.com/9781457185519.

Wayner:2000:FAH

[Way00] Peter Wayner. Free for all:

how Linux and the free software movement undercut the high-tech titans. Harper Business, New York, NY, USA, 2000. ISBN 066620503 (??invalid ISBN??). 340 pp. LCCN QA76.76.O63 W394 2000.

Wilding:2006:SSL

[WB06] Mark Wilding and Dan Behman. Self-service Linux: mastering the art of problem determination. Bruce Perens’ Open Source series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2006. ISBN 0-13-147751-X. xxii + 431 pp. LCCN QA76.76.O63 W52 2006.

Worsley:2001:GSY

[WBS01] John Worsley, Andrew Brookins, and Kai Staats. Getting started with Yellow Dog Linux. Open-Docs, Salem, OR, USA, 2001. ISBN 0-9700330-3-6. 267 pp. LCCN QA76.76.O63; QA 76.76 .O63W67 2001 SCAR.

Wang:2022:DII

[WC22] Shie-Yuan Wang and Jen-Chieh Chang. Design and implementation of an intrusion detection system by using Extended BPF in the Linux kernel. Journal of Network and Computer Applications, 198(??):??, February 2022. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1084804521002769.

Welsh:1995:SL

[WCC95] Matt Welsh, Remy Card, pref, and Rene Cougnenc, trad.

Le Système Linux. Editions O’Reilly International Thomson, Paris, France, 1995. ISBN 2-84177-002-8. xxvii + 629 pp. LCCN ????

Wright:2006:VUS

[WDG^{+}06] Charles P. Wright, Jay Dave, Puja Gupta, Harikesavan Krishnan, David P. Quigley, Erez Zadok, and Mohammad Nayyer Zubair. Versatility and Unix semantics in namespace unification. ACM Transactions on Storage, 2(1):74–105, February 2006. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Weeks:2004:LU

[WDJ04] Roger Weeks, Edd Dumbill, and Brian Jepson. Linux unwired. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00583-0. xii + 297 pp. LCCN QA76.76.O63 W433 2004. URL http:/

/www.oreilly.com/catalog/
9780596005832.

Welsh:1999:RL

[WDK99] Matt Welsh, Matthias Kalle Dalheimer, and Lar Kaufman. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 1999. ISBN 1-56592-469-X. xix + 730 pp. LCCN QA76.76.O63 W465 1999. US$34.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        runux2.

Wehrli:2001:LEP

[Weh01] Rob Wehrli. Linux Embedded Programming. CMP Books, 6600 Silacci Way, Gilroy, CA 95020, USA, 2001. ISBN 1-57820-085-7. 352 (est.) pp. LCCN ???? US$39.95. URL

                                        http://www.cmpbooks.com/.

Weinstabl:2004:PAE

[Wei04] Paul Weinstabl. PostgreSQL: [Administration und Einsatz ; für Linux und Windows]. Computer and Literatur. Computerund Literatur-Verlag, Böblingen, Germany, 2004. ISBN 3-936546-22-3. 589 pp. LCCN ???? EUR 49.90.

Welsh:1994:LIG

[Wel94] Matt Welsh. LINUX installation and getting started. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1994. ISBN 0-916151-71-9 (hardcover), 0-916151-77-8 (paperback). xviii + 231 pp. LCCN ????

Welsh:1995:ILK

[Wel95a] Matt Welsh. Implementing loadable kernel modules for Linux. Dr. Dobb’s Journal of Software Tools, 20(5):18–20, 22, 24, 96, April 1995. CODEN DDJOEB. ISSN 1044-789X.

Welsh:1995:LIG

[Wel95b] Matt Welsh. Linux installation and getting started. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1995. ISBN 0-916151-76-X. xvi + 246 pp. LCCN ????

Wells:1998:LWS

[Wel98] Nicholas Wells. Linux Web server toolkit. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-3167-0. xxvi + 502 pp. LCCN QA76.76.O63W464 1998.

Wells:1999:STYb

[Wel99a] Nicholas Wells. Sams teach yourself KDE in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31608-0. xiii + 457 pp. LCCN QA76.9.U83 W48 1999. Includes CD-ROM.

Wells:1999:LLS

[Wel99b] Nicholas D. Wells. Learning Linux system administration, 1999. ISBN 0-7400-2097-8 (level 1), 0-7400-2098-6 (level 2), 0-7400-2099-4 (level 3), 0-7400-2100-1 (level 4), 0-7400-2101-X (Level 5).

Wells:1999:LDK

[Wel99c] Nicholas D. Wells. Linux! I Didn’t Know You Could Do That. . .. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2612-X. xiv + 283 pp. LCCN QA76.76.O63 W4638 2000. US$19.99; CDN$29.95; UK £14.99. URL

http://scooter.sybex.com/
sybexbooks.nsf/2604971535a28b098825693d0053081b/
f68011722d45db5c8825693d0057f9a6?
OpenDocument&Highlight=0,
0-7821-2612-X.

Werden:2010:BRT

[Wer10] David A. Werden. Book review: TCP/IP architecture, design, and implementation in Linux by Sameer Seth and M. Ajaykumar Venkatesulu. ACM SIGSOFT Software Engineering Notes, 35(5):57, September 2010. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ward:2003:VWH

[WF03] Brian Ward and Gerhard Franken. VMware Workstation: [das Handbuch; Installation, Konfiguration, Anwendung und Troubleshooting; Gast-Systeme: Windows, Linux, BSD, Novell NetWare, Solaris, FreeDOS und Oberon; virtuelle Netzwerke, Netzwerkkonfiguration und -dienste]. mitp-Verl., Landsberg, Germany, 2003. ISBN 3-8266-0964-6. 336 pp. LCCN ???? EUR 32.00 (DE). URL http:

//www.gbv.de/du/services/
agi/FCC0A57071BE8695C125704A0029797F/
FLMA122525.

Wampler:2008:NBM

[WG08] Doug Wampler and James H. Graham. A normality based method for detecting kernel rootkits. Operating Systems Review, 42(3):59–64, April 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wilde:2018:MTV

[WG18] Evan Wilde and Daniel German. Merge-Tree: Visualizing the integration of commits into Linux. Journal of Software: Evolution and Process, 30(2):e1936:1–e1936:??, February 2018. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Wiegand:1993:CDL

[Wie93] J. Wiegand. The cooperative development of Linux. In IEEE [IEE93], pages 386–390. ISBN 0-7803-1466-2, 0-7803-1465-4, 0-7803-1467-0. LCCN T 10.5 I54 1993. IEEE catalog number 93CH3367-0.

Wielsch:1996:L

[Wie96] M. Michael Wielsch. Linux. PC Poche. Edition Micro Application, Paris, France, 1996. ISBN 2-7429-0767-X. 491 pp. LCCN ????

Wielsch:1997:L

[Wie97] Michael Wielsch. Linux. Grand livre. Edition Micro Application, Paris, France, 2e edition, 1997. ISBN 2-7429-0619-3, 2-7429-1127-8. 1002 pp. LCCN ????

Wielsch:1999:Lb

[Wie99a] Michael Wielsch. Linux. PC 100 pour 100 pratique; 2500.

Edition Micro Application, Paris, France, 1999. ISBN 2-7429-1500-1. 509 pp. LCCN ????

Wielsch:1999:LTD

[Wie99b] Michael Wielsch. Linux toutes distributions. PC Poche. DATA-Becker, ????, 1999. ISBN 2-7429-1318-1. 524 pp. LCCN ????

Williams:1998:MLU

[Wil98] James D. Williams. A methodology for Linux as a user process based on Solaris Minix on the SPARC architecture. Thesis (M.S.), New Mexico State University, as Cruces, NM 88003-8001, USA, 1998. xiii + 141 pp.

Wisniewski:2003:LOI

[Wis03] John Robert Wisniewski. Linux and OpenVMS interoperability: tricks for old dogs, new dogs, and hot dogs with open systems. Compact solutions. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2003. ISBN 1-55558-267-2. xiii + 198 pp. LCCN QA76.76.O63 W584 2003.

URL http://www.loc.gov/
catdir/description/els041/
2003276793.html;   http:/
/www.loc.gov/catdir/toc/
els041/2003276793.html.

Wang:2007:DAS

[WJS07] Haining Wang, Cheng Jin, and Kang G. Shin. Defense against spoofed IP traffic using hopcount filtering. IEEE/ACM Transactions on Networking, 15(1):40–53, February 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Wu:2007:EPM

[WJY^{+}07] Wei Wu, Lingling Jin, Jun Yang, Pu Liu, and Sheldon X.-D. Tan. Efficient power modeling and software thermal sensing for runtime temperature monitoring. ACM Transactions on Design Automation of Electronic Systems, 12(3): 26:1–26:??, August 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Welsh:1995:RL

[WK95] Matt Welsh and Lar Kaufman.

Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, February 1995. ISBN 1-56592-100-3. xvii + 575 pp. LCCN QA76.76.063 .W3 1995. US$24.95.

Welsh:1996:RL

[WK96] Matt Welsh and Lar Kaufman.

Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, Summer 1996. ISBN 1-56592-151-8. xviii + 630 pp. LCCN QA76.76.O63W465 1996. US$29.95.

Welsh:1997:SL

[WKC^{+}97] Matt Welsh, Lar Kaufman, René Cougnenc, Manuel Makarevitch, and Nat Makarevitch. Le Système Linux. O’Reilly France, Paris, France, 2e rev. et augm. edition, 1997. ISBN 2-84177-033-8. xix + 653 pp. LCCN ????

Welsh:1997:PLO

[WKD97] Matt Welsh, Lar Kaufman, and Miroslav Dressler. Pouzivame Linux: operacni system pro osobni pocitace kompatibilni s Unixem. Computer Press, Praha, Czech Republic, 2. vyd.

[sic] edition, 1997. ISBN 80-7226-001-4 (vaz.). 612 pp. LCCN ????

Welsh:2002:RL

[WKDD02] Matt Welsh, Lar Kaufman, Terry Dawson, and Matthias Kalle Dalheimer. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2002. ISBN 0-596-00272-6. xviii + 672 pp. LCCN QA76.76.O63 R855 2003. US$44.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        runux4.

Walsh:2003:RL

[WKDD03] Matt Walsh, Lar Kaufman, Terry Dawson, and Matthias Kalle Dalheimer, editors. Running Linux. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2003. ISBN 0-596-00272-6. xviii + 672 pp. LCCN QA76.76.O63 R855 2003. US$44.95. URL http:/

/www.oreilly.com/catalog/
9780596002725;   http://
www.oreilly.com/catalog/
runux4.

Welsh:2001:LWS

[WLS^{+}01] Matt Welsh, Ben Laurie, Ellen Siever, Randy Jay Yarger, Alligator Descartes, and Scott Guelich. The Linux Web server CD bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00208-4. LCCN Disk 2057 Protect. CD-ROM with Running Linux, 3rd ed. – Linux in a nutshell, 3rd ed. – CGI programming with Perl, 2nd ed.

– Apache: the definitive guide, 2nd ed. – MySQL and mSQL – Programming the Perl DBI.

Wurmsdobler:2000:MMR

[WM00] Peter Wurmsdobler and Nicholas McGuire. MiniRTL: A minimal real-time Linux. Dr. Dobb’s Journal of Software Tools, 25 (12):48, 50, 52, 54, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2000/2000_12/
                                        minirtl.txt.

Wang:2015:SDB

[WN15] Yongge Wang and Tony Nicol. On statistical distance based testing of pseudo random sequences and experiments with PHP and Debian OpenSSL. Computers & Security, 53 (??):44–64, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167404815000693.

Woehr:2002:PBE

[Woe02] Jack J. Woehr. Programmer’s bookshelf: Embedded systems programming. Dr. Dobb’s Journal of Software Tools, 27 (10):86, October 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/.

Woehr:2003:PBL

[Woe03] Jack J. Woehr. Programmer’s bookshelf: Linux, XP, and everything in between. Dr. Dobb’s Journal of Software Tools, 28(3):69, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/documents/s=7826/
                                        ddj0303j/.

Wolfe:2002:OSP

[Wol02] A. Wolfe. An open-source platform to tackle scientific computing. IEEE Spectrum, 39 (6):22–24, June 2002. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolfe:2004:TSD

[Wol04a] Alex Wolfe. Toolkit: Samba does Windows-to-Linux dance. ACM Queue: Tomorrow’s Computing Today, 2(5):18–21, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2004:GTC

[Wol04b] Alexander Wolfe. Toolkit: Grid tools: Coming to a cluster near you. ACM Queue: Tomorrow’s Computing Today, 2(4):20–23, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

WorkGroupSolutions:1998:LE

[Wor98] WorkGroup Solutions, Inc.

Linux encyclopedia. Work-Group Solutions, Aurora, CO, USA, sixth edition, 1998. ISBN 0-9644309-2-4 (paperback). xxvi + 1631 pp. LCCN ????

Wielsch:1999:La

[WPL^{+}99] Michael Wielsch, Jens Prahm, François Ligier, traducteur, Pierre M. Wolf, and Serge Springinsfeld. Linux. La Bible; 2428. Edition Micro Application, Paris, France, 1999. ISBN 2-7429-1428-5. 1408 1 disquette pp. LCCN ????

Wu:2013:SFS

[WQR13] Xiaojian Wu, Sheng Qiu, and A. L. Narasimha Reddy. SCMFS: a file system for storage class memory and its extensions. ACM Transactions on Storage, 9(3):7:1–7:??, August 2013. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Wisniewski:2003:EUS

[WR03] Robert W. Wisniewski and Bryan Rosenburg. Efficient, unified, and scalable performance monitoring for multiprocessor operating systems. In ACM [ACM03a], page ?? ISBN 1-58113-695-1. LCCN

????  URL http://www.
sc-conference.org/sc2003/
inter_cal/inter_cal_detail.
php?eventid=10721#0; http:
//www.sc-conference.org/
sc2003/paperpdfs/pap121.
pdf.

Witchel:2005:MMI

[WRA05] Emmett Witchel, Junghwan Rhee, and Krste Asanović. Mondrix: memory isolation for Linux using Mondrian memory protection. Operating Systems Review, 39(5):31–44, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wright:2000:BGG

[Wri00]  Peter Wright.  Beginning

GTK+/GNOME Programming. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-381-1. xi + 613 pp. LCCN QA76.9.U83 W754 2000. US$39.99.

Wall:1999:LSL

[WSL^{+}99] David A. E. Wall, Roger Sessions, Marty Leisner, Michael Stiefel, Robin F. Goldsmith, and Suresh Peram. Letters: Sounding off on Linux; (re)building the community; real success stories; improving performance by improving design. IEEE Software, 16(3): 8–12, May/June 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1999/pdf/
s3008.pdf.

Welsh:1995:DL

[Wt95a] Matt Welsh and the Linux Documentation Project. DRX. Linux. Linux System Laboratory, Chesterfield, MI, USA, 1995. 1176 pp. US$49.95.

Welsh:1995:LGS

[Wt95b] Matt Welsh and the Linux Documentation Project. Linux: Getting Started. Linux System Laboratory, Chesterfield, MI, USA, 1995. 565 pp. US$59.95. Includes CD-ROM.

Wells:1999:STYa

[WT99] Nicholas Wells and R. Dean Taylor. Sams’ teach yourself StarOffice for Linux in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31412-6 (paperback). xvi + 505 pp. LCCN QA76.73.I57W4637 1999.

Wong:2008:TAF

[WTKW08] Chee Siang Wong, Ian Tan, Rosalind Deena Kumari, and Fun Wey. Towards achieving fairness in the Linux scheduler. Operating Systems Review, 42(5):34–43, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wunschiers:2013:CBP

[Wün13] Röbbe Wünschiers. Computational biology: a practical introduction to biodata processing and analysis with Linux, MySQL, and R. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2013. ISBN 3-642-34748-7, 3-642-34749-5 (e-book). xxix + 449 pp. LCCN QH324.2 .W86 2013.

Wookey:2001:GAD

[WW01a] Wookey and Paul Webb. Guide to ARMLinux for Developers. Aleph One, ????, 2001. ISBN ???? ???? pp. LCCN ????

URL http://www.aleph1.co.
uk/armlinux/thebook.html.

Wookey:2001:GAR

[WW01b] Wookey and Paul Webb. Guide to ARMLinux for RISCOS machines. Aleph One, ????, 2001. ISBN ???? ???? pp. LCCN ???? URL

http://www.aleph1.co.uk/
armlinux/thebook.html.

Williamson:2002:CCA

[WW02] Carey Williamson and Qian Wu. A case for context-aware TCP/IP. ACM SIGMETRICS Performance Evaluation Review, 29(4):11–23, March 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Witherspoon:1997:LD

[WWH97] Craig Witherspoon, Coletta Witherspoon, and Jon Hall. Linux for dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-0275-1. xii + 360 pp. LCCN QA76.76.O63 W59 1997.

Witherspoon:1998:LPN

[WWH98] Craig Witherspoon, Coletta Witherspoon, and Jon Hall. Linux pour les nuls. Pour les nuls. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1998. ISBN 2-7361-3138-X, 2-7361-2989-X. xvi + 333 pp. LCCN ????

Wielsch:1998:LAP

[WWS98] Michael Wielsch, Michael Weber, and Matthias Sylvester. Linux: administration et programmation. La Bible. Edition Micro Application, Paris, France, 1998. ISBN 2-7429-1012-3. 943 pp. LCCN ????

Wall:1999:LPU

[WWW99] Kurt Wall, Mark Watson, and Mark Whitis. Linux programming unleashed. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-672-31607-2. xxiv + 818 pp. LCCN QA76.76.O63 W3573 1999. Edited by Brian Gill and Ron Gallagher.

Wu:2008:DNL

[WXX08] Fengguang Wu, Hongsheng Xi, and Chenfeng Xu. On the design of a new Linux readahead framework. Operating Systems Review, 42(5):75–84, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Welsh:1994:LBG

[WY94] Matt Welsh and Yggdrasil Computing Inc. The Linux bible: the GNU testament. Yggdrasil Computing, Inc., Berkeley, CA, USA, second expanded edition, 1994. ISBN 1-883601-10-X (paperback). 1176 pp. LCCN ????

Wylde:1995:CLG

[Wyl95] R. Wylde. Creating a Linux gateway. Personal computer world, 18(7):534–??, ???? 1995. CODEN PCWODU. ISSN 0142-0232.

Wang:2013:ELK

[WYW^{+}13] Lei Wang, Pengzhi Yu, Zheng Wang, Chen Yang, and Qiang Ye. On the evolution of Linux kernels: a complex network perspective. Journal of Software: Evolution and Process, 25(5):439–458, May 2013. CO-DEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Xing:2022:DDG

[XCS^{+}22] Yunlong Xing, Jiahao Cao, Kun Sun, Fei Yan, and Shengye Wan. The devil is in the detail: Generating system call whitelist for Linux seccomp. Future Generation Computer Systems, 135(??):105–113, October 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X2200139X.

Xian:2008:CAS

[XSaJ08] Feng Xian, Witawas Srisa-an, and Hong Jiang. Contentionaware scheduler: unlocking execution parallelism in multithreaded Java programs. ACM SIGPLAN Notices, 43(10):163– 180, September 2008. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Xia:2010:ITA

[XZY^{+}10] Liang Xia, Yongxin Zhu, Jun Yang, Jingwei Ye, and Zonghua Gu. Implementing a thermalaware scheduler in Linux kernel on a multi-core processor. The Computer Journal, 53(7):895–903, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

                                        //comjnl.oxfordjournals.
                                        org/cgi/content/abstract/
                                        53/7/895;   http://comjnl.
                                        oxfordjournals.org/cgi/reprint/
                                        53/7/895.

Yager:1996:LM

[Yag96] T. Yager. Linux matters. BYTE Magazine, 21(2):123– 124, 126–128, February 1996. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Yaghmour:2003:BEL

[Yag03]  Karim Yaghmour.  Build-

ing Embedded Linux Systems. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00222-X. xxii + 391 pp. LCCN QA76.76.O63 Y35 2003. US$44.95, CAN$69.95, UK£31.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        belinuxsys/.

Yeoh:2005:BAL

[Yeo05] C. Yeoh. Building applications for the Linux Standard Base. IBM Systems Journal, 44(2): 369–??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670.

Yggdrasil:1993:YLG

[Ygg93] Yggdrasil Computing, Inc. Yggdrasil Linux/GNU/ X operating system, 1993. ISSN 1069-3955. 1 computer laser optical disk.

Yggdrasil:1994:YLG

[Ygg94] Yggdrasil Linux/GNU/X operating system, 1994. ISSN 1069-3955. Yggdrasil Computing, Inc., Berkeley, CA, USA. CD-ROM. Updated quarterly.

Yoder:2015:BCS

[YK15] Mark A. Yoder and Jason Kridner. BeagleBone Cookbook: Software and Hardware Problems and Solutions. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-0539-5. ???? pp. LCCN ????

Yanagiuchi:2002:LJI

[YKS^{+}02] S. Yanagiuchi, T. Kiyohara, N. Shiraishi, K. Mori, and M. Ohkita. Linux/Java implemented personal mobile tool SL-A300. Sharp Technical Journal, 84:31–32, 2002. CO-DEN ???? ISSN 0285-0362.

Yalta:2008:GLP

[YL08] A. Talha Yalta and Riccardo Lucchetti. The GNU/Linux platform and freedom respecting software for economists. Journal of Applied Econometrics, 23(2):279–286, 2008. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

York:1999:BLC

[Yor99] Dan York. Building a Linux certification program. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CODEN LOGNEM. ISSN 1044-6397. URL http://www.

usenix.org/publications/
login/1999-6/features/linuxcert.
html.

Yu:2004:CCC

[YSCO04] L. Yu, S. R. Schach, K. Chen, and J. Offutt. Categorization of common coupling and its application to the maintainability of the Linux kernel. IEEE Transactions on Software Engineering, 30(10): 694–706, October 2004. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=1339279.

Yu:2007:UCC

[Yu07]  Liguo Yu.  Understanding

component co-evolution with a study on Linux. Empirical Software Engineering, 12 (2):123–141, April 2007. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/accesspage/
article/10.1007/s10664-006-
9000-x.

Yin:2007:BAA

[YW07] Heng Yin and Haining Wang. Building an application-aware IPsec policy system. IEEE/ ACM Transactions on Networking, 15(6):1502–1513, December 2007. CODEN IEANEP. ISSN 1063-6692 (print), 1558-2566 (electronic).

Yu:1996:EDC

[YY96] Kyoung-Sang Yu and Hyuck Yoo. Elimination of data copying for multimedia communication. In Proceedings. Third International Workshop on Real-Time Computing Systems and Applications: October 30 – November 1, 1996, Seoul, Korea, pages ??–?? (of xi + 328). ????, ????, 1996. ISBN 0-8186-7626-4. LCCN QA76.54.I54 1996. IEEE catalog number 96TB100071.

Zack:2001:DUG

[Zac01] Randy Zack. Debugging under GNU/Linux. C/C++ Users Journal, 19(2):??, February 2001. CODEN CCUJEX. ISSN 1075-2838.

Zadok:2001:LNA

[Zad01] Erez Zadok. Linux NFS and Automounter administration. The Craig Hunt Linux library. Sybex, Inc., 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2739-8. xlvi + 638 pp. LCCN QA76.76.O63 Z34 2001.

Zaruba:2019:CAC

[ZB19] Florian Zaruba and Luca Benini. The cost of applicationclass processing: Energy and performance analysis of a Linux-ready 1.7-GHz 64-Bit RISC-V core in 22-nm FDSOI technology. IEEE Transactions on Very Large Scale Integration (VLSI) Systems, 27(11):2629– 2640, 2019. CODEN IEVSE9. ISSN 1063-8210 (print), 1557-9999 (electronic).

Zhang:2005:FVM

[ZBG^{+}05] Yuting Zhang, Azer Bestavros, Mina Guirguis, Ibrahim Matta, and Richard West. Friendly virtual machines: leveraging a feedback-control model for application adaptation. In ACM [ACM05b], pages 2–12. ISBN 1-59593-047-7. LCCN QA76.9.V5 I575 2005. URL http:

//www.loc.gov/catdir/toc/
fy0611/2006530661.html. ACM

order number 548059.

Ziegler:2002:LF

[ZC02] Robert L. (Robert Loren) Ziegler and Carl B. Constantine. Linux firewalls. New Riders Publishing, Carmel, IN, USA, second edition, 2002. ISBN 0-7357-1099-6. xxiii + 562 pp. LCCN QA76.9.A25Z54 2002.

Zwicky:2000:BIF

[ZCC00] Elizabeth D. Zwicky, Simon Cooper, and D. Brent Chapman. Building Internet firewalls. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2000.

ISBN 1-56592-871-7. xxi + 869 pp. LCCN TK5105.59 .Z85 2000.

Zemla:1994:WTC

[Zem94] A. Zemla. Wavelet transforms computing on PVM. In Dongarra and Wasniewski [DW94], pages 534–546. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). LCCN QA76.58 .P35 1994. DM104.00.

Ziccardi:2001:DAN

[Zic01] Giovanni Ziccardi. Il diritto d’autore nell’era digitale: evoluzione tecnologica e copyright: Internet, mp3, DivX;-), open source, Gnu/Linux, free software, mezzi di protezione. Prima ed., Milan, Italy, 2001. ISBN 88-324-4459-3. xxi + 343 pp. LCCN KE1809 .Z53 2001.

Ziegler:2000:LF

[Zie00] Robert L. (Robert Loren)

Ziegler. Linux firewalls. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 0-7357-0900-

9. xvi + 470 pp. LCCN TK5105.59 .Z54 2000.

Ziskovsky:1993:CCE

[ZS93] J. Ziskovsky and B. Sullivan.

Creating a cost effective news and email interface to the Internet. In ACM [ACM93], pages 189–193 (vol. 1). ISBN 0-89791-631-X. LCCN QA74.A26 1993.

Zee:2007:ISZ

[ZST^{+}07] M. Zee, J. W. Stevens, B. L.

Thompson, J. A. Fowler, J. Goldman, P. T. Chan, and T. P. McSweeney. IBM System z9 open systems adapter for communication controller for Linux. IBM Journal of Research and Development, 51 (1/2):119–??, January /March 2007. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

                                        research.ibm.com/journal/
                                        rd/511/zee.html.

Zhou:2010:PAT

[ZYCZ10] Xiuyi Zhou, Jun Yang, Marek Chrobak, and Youtao Zhang. Performance-aware thermal management via task scheduling. ACM Transactions on Architecture and Code Optimization, 7(1):5:1–5:??, April 2010. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic).